A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.02

Title word cross-reference

(0 ≤ m < k) [PP87]. (l + 1) [Shi82c]. (M, S) [ER80]. (N − 1) [GHO86]. (n − l) [BM81b]. (Σ + Π) [SN89a]. 1 [HY86, SD83, SD84a, Wil88b]. 2 [Kul85, SD83]. 2^{6} [Gup88a]. 2^{M} [Shi82c, GC84, KM84c]. 2^{n} [EHA84, Kuw89]. 2^{n−p} [MCN81]. 2l + 1 [Kuw89, Shi82c]. 2 × 2 [Bur81, Hab89, Mar89a, SG84c, TO85]. 2 × 3 [GR81]. 2 × j [Wal87]. 3 [BV89]. 3.1 [Gup88a]. 3^{m} [Kuw83]. 3 × 3 [BV89]. ^{2} [ER83]. _{1} [RMW86]. _{K} [Gup82]. _{p} [Nyq83]. _{r} [KK81]. A [SS87b]. α [AR83, Pil85]. AR(1) [And88, And89a, DD85b, Mag89, Rus88]. AR(2) [SM87]. ARMA(P, Q) [Hie85]. C [Gup88d, Arb85, HB86]. C_{p} [SCT83]. cX̄ + d [Roi87]. χ^{2} [BDF^{+}84]. D [MH85, SS87b, Czi88, PG83]. D^{−1} [KSR88]. d_{o} [RS80]. Δ [Bof84]. F [BDF^{+}84, Dey89, HF80, Joh86b, Lam86, Lec85, Mos87, MSW89, PL83, SD84b, SB82a, YM85, ZG80, MT85, SFR84]. F {N_{1}} = ∞ [McC82]. G [Rao81b, SD83, JB84]. Γ [HT83b]. h [JB84]. i × j [Mic85]. i × j × 2 [Chu83]. i × j × k [SS86d]. K [PK85a, AR83, AH86a, BCDE85, BG83, CS87a, HK87, KP81, LC85, PP87, PTA89, SJ87a, Shi80, Som85, Som89a, SM82b, Won82]. k(≥ 2) [Sin85]. K = 2^{N−1} − 1 [Smi84]. k = 3 [WT83]. L [Par82, Sha89]. L_{1} [Far85, SKG80, ST86a, SK87b]. L_{∞} [Far85]. L_{P} [GM85a, GM85b, GdT87]. LU [DL88]. M [Fra83, Hol87, JP87, Kaf82, KY81, SS85b, Sil88, Kap86, Par81, PP87, SS88b, Str84]. MA(2) [SM87]. μ^{y} − μ^{x} [ESP84]. N [Smi84, Blu87, BM81b, GHO86, HW88, MT83b, Old87, PP81, Par81, Ped85a, Ped85b, SD84a, Sat88, Som89a, Sun88]. n = 3(1)14 [Ima87]. O [DL88]. P [Bri87, Som85, BG83, Ber89, GM85a, GM85b]. p(l, k) [KY82]. P (X > Y) [SS81a]. p = 3(1)10 [WT83]. Pr(X_{1} < Y < X_{2}) [Sin80b]. Pr(y, x) [Beg80]. P r{y, x} [Beg81b]. P {X > Y } [DS81]. q [Dun81, Tam86]. R [HJ83, SS86b, aS89b, BR85, Gap84, GM83b, MT83a, Sat88, Tuc82]. R^{2} [Sri83]. r × c [MP86]. s [BR85, Tuc82, YKY85]. σ^{2} [MJ82]. Σ_{1}^{−1}Σ_{2} [Dun86]. sn [StR82]. t [AH84, BB89b, BM86b, CS87a, Cho83a, HN84, JG85, KS82b, Kul85, LP88, LM84c, MT85, MT83a, MSL86, MSW89, PYO82, Som84, SH89, SA86, Won88, YKY85, YG82, vdMN87]. T^{2} [SA82, Sri83, TB88c, TB88b]. U [DK83, Gho81c, Lee85, Muk81, OM80, Sen84, Wit88, YM86, Yam87, YM89]. V [Kuw83]. W [Abd89, Sam85]. x [GRY81]. X^{2} [Chu84, KS82b]. x ≤ y [Smi83]. Z > X + Y [ZTJ86].

-ary [PP81]. -Ball [Ped85b, MT83b, Ped85a]. -best [BB89b]. -binomial [Dun81]. -charts [GRY81]. -class [SM82b]. -correct [Bof84]. -dependent [SS88b]. -dimensional [Old87]. -distribution [YM85]. -estimates [Kaf82]. -estimation [SS86b]. -estimators [Abd89, Fra83, HJ83, Hol87, JP87, KY81, SS85b, aS89b, Sil88]. -exponential [Sam85]. -factor [Shi82c]. -inverse [Rao81b]. -Laplace [Pil85]. -levels [Som85]. -matrix [Gup88d]. -means [HB86, Won82]. -norm [GM85a, GM85b, GdT87]. -optimal [Czi88, PG83, HT83b, ER80, MH85]. -optimality [SS87b]. -out-of- [BG83, SD83, SD84a]. -Partially [KSR88]. -population [AH86a]. -quantile [AR83, Tam86]. -ratios [Joh86b]. -ray [BV89]. -sample [Shi80, Som85, BCDE85]. -samples [Par81]. -squares [SFR84]. -statistic [DL88, Lee85, Muk81]. -statistics [DK83, DL88, Gho81c, OM80, Par82, Sen84, Sha89, Wit88, YM86, Yam87, YM89, ZG80]. -symbol [YKY85]. -tail [Som84]. -test [PYO82, Lec85, SA82]. -tests [SD84b]. -th [Sat88]. -type [Lam86]. -values [Bri87]. -variables [YG82]. -variate [Sun88]. -variates [SB82a]. -wise [BM81b].

1 [Gup82, Ped85a, PH84].

2SAE [SA89].

7 [Pet82].

a-priori [OP84]. absence [KSC81]. absenteeism [Poc82]. absolute [Die84a, Kea83]. absolutely [GMM84]. absorption [Dun81]. abundance [OW86]. abuse [KK86a]. Accelerated [KB82a, BS89, CEA85, GM84, KB81]. acceptance [CB81, KB84, Sch85a]. accessory [Llo87b]. Accuracy [Owe85b, Fal86]. Accurate [Spr89, Wit89a]. Acknowledgement [KJS^{+}84]. Acknowledgment [SG86b]. acreage [SG84a]. adaptation [BH84]. Adaptive [EH83b, Kap86, Bra81, BYD85, BD86c, HT81, Hog82, HL84, KK80, KC84b, Kha84, Loh84a, McC82, MN89, Par82, PM89b, Rub86a, WR87]. Additional [MAT82, SD87a, SA85, Dah88, KR82, NDM84, SCG85]. Additive [Gup87a, GM85a, Nar83]. adequacy [AV88, KS85, SND84]. adequate [Rad85]. adjusted [BR86, KW84, Mic87, SD87a, Sha88a]. adjusting [ME86]. Adjustments [WK86]. Admissibility [MB89, SG84b, Sri87d, MC84, Roi87, SAN89a]. Admissible [CZ87, Dah88, MRS84, Mat85, Lis88]. admit [Sob87]. Advances [Sri89]. adverse [Can84]. aetiology [Cla82]. affinities [SS88d]. After [Kun83, JP86, JP88c, LSS85a, LSS85b, LO89, Oht87a, Oht88, SK82, SS83]. against [AB89, Bar83, EZ89, Goo89, KI88, Kle83, Koc85, KS84, Kul85, Lin89, Man89, Mar80, MT83a, SQ81, SYT89, aS82, SK86]. Age [KJS^{+}83, LP88]. Age-period-cohort [KJS^{+}83]. aged [Fu80]. agglomerative [DW87b]. aggregates [Fin85, WA81]. aggregation [Ven80]. aging [Kle83]. agreement [Fla87, IB89]. agricultural [Fei84, HH84]. ahead [Har82]. Aiala [KJ90]. algorithm [AH88, tFAH84, BA87, BH84, Boh83, Din87a, Din87b, FW86, KC87a, KC87b, KC87c, MP86, Sil88]. Algorithms [Pet82, GdT87]. alias [Kuw83, Shi82c]. Allocating [Bar81]. allocation [AS84a, DS87, DF89, PG83, VB80, Wri83, Wri88a, YS84]. allowing [JK84]. Almost [Goo89, SIU85, Hos87b, Oht86]. Alternative [BS88, Dav85b, KO84, MN89, AB89, Ali89a, Ang83b, Ber86, Cal87, HL87, Hei88, Hor81, KH86, Mat81, Mon87]. Alternatives [FH87, AJK89, Bar83, Bas89, BJ83, Ber82, BS83, BS84, Die89, Fer87, Goo89, Gov86, Isa86, KI88, Ken81, KR84, KS84, Kul85, Lin89, Man89, Mar80, MT83a, Mar84b, Miy87, SQ81, zS88, Shi88, Shi80, aS89a, SK86, SSS87]. always [Ber81b]. amendments [Ban80b]. American [Net85]. among [BM89, BA88, KR89b, McC85, Pri82, Sha86, ZF87]. amplitude [Sho82b]. analogues [DPS81]. Analyses [SKSF83, Atk82, GgT86, NS86]. Analysis [CM86a, Coh82, Har86, KC84a, Kuw89, LPW82, PK81, Rat89, SC85, SL82b, Ach84, AB86, AD83, BM82, BM85b, BLW87, Bha82, Bra81, BH80, Byr88, CG87, Che83b, Chr88, CEA85, Cla82, Coo85, CVDS85, ĆM89, Dav84, Dav89, DK89, Dre85, Fic84, Fra80, GG86, GM80a, GR83, GM83a, Gil82, GBS83, GR89, GW84, Gun84, Gun83, Gup84, Hoc88, IC80b, JBF^{+}82, JN89, Kay82, Kha83a, Khu82, KKC82, KY82, KJS^{+}83, LS80, Lec85, Lem80, LS81, Lis84, Mar80, Mat81, Men81b, MA80, Miy87, Nel87, Nel82, OA87, OAH89, PH80, PL82, PL86, Pen85a, Pen85b, PVB86, Pou83, PK85b, RK81, Rad84, Rad85, SD88, Saw87, Sha80, Shi85, SG88b, Spa88a, Sri87c, SL89b, Tan88, Tan89, TO89, TL89b, Tra83a, Tri88, Tsu85, Vos88, Wan83, Wan89a]. analysis [WL80, Wu88]. analystcs [SG84c]. Analytic [FV89]. analyze [DeP83]. analyzing [GSS87, MI82, Rav81, Spa85, Spa86]. ANCOVA [BF85, OA87]. and/or [KR84]. Anderson [Boo82, BWMD83, Fri87, Leu88a]. Andrews [KP84]. angular [KP89a]. animal [Cru82]. anisotropy [Pro84]. annotated [Die84a]. ANOCOVA [Man89]. Anova [RS84a, Adk86, BM81a, Man89, Mar82, Nai82, Saw88, SN82]. antitonicity [Nor89]. AOQL [SV89]. appear [Tan89]. applicability [JC87]. Application [Ang89a, Dun86, BE88, CL83, ĆM89, Czi89a, DL84, FBV84, Hai88, HGB89, HS83, HZ82, IC80b, Joe88, JP87, KJ83, LB84, MS81c, PH80, Pou83, RJ84, Rus85, Sau82, Saw87, Sen81, SU86, Ste89b, SB89b, Tam86, Vog89, WK81, WYF87]. Applications [Har86, Som89a, BA88, CS88b, CP82, CO84, Chr86, Cru82, EN80a, Fra83, GP82, Gro82, HES84, Kai88, KHB^{+}88, KPS88, Kum81, LR87, LWA80, tL84, MR86, Ng89a, Sob89a, TC81, TGG86, TG87, TT89b, VJ88, Xek83b, Yas81a, eL82a]. applied [JN89, NS84, Sha80, Wan86]. appraisal [Chm81]. approach [AB89, Ali89a, Alw86, Ban80a, Bjø81, BR86, BSC80, BA84, Che87a, CT89, Con85, CW87, CB81, Dav88b, Dre85, DC84b, Ebr86, EKR86, Fon89, GG86, GK88b, Gia86, God85, Gra89, HG86, Haw89b, HL85, JI84, JPD82, Kab83, Kap85, Kok85, Lai84, LM84b, Men81b, MSL86, NDM84, Obr82, Pap89, PS87, Pur86, Rav81, Sam88, Sam89, SD86, TB88a, TT88, TTB88, Tra83b, TG86, Ven80, WNK80, dWM85, dR89]. approaches [Chr88, GM80a, HN84, SG84c]. appropriateness [KW83]. Approximate [Cha83b, CM82, DFG87b, Eub84, How85, PU86, RM88, SB82a, SL87, SL89b, Wil86, AG89b, HT81, KY82, LWA89, MSW89, Som85]. approximated [LSB89]. Approximating [GN86b, MR88, She86, JS82, TG83]. Approximation [Whi89a, Anb84, Bet87, BE88, Bri87, CO84, DW89b, Fal86, Hie85, IM88, KS85, Kul88, Lac84, LT85, LM84c, Obr82, Ojo88, PM85, RS81, Spa86, SB89b, Wan87c, Wef87, YG82]. Approximations [AH84, Bra88, ID80, AC84, Dav85b, Fra88, Gho80, Mag89, SA85]. AQL [SV89]. aqueous [AD83]. arbitrarily [LP85, MP82c, Nic89]. Arbitrary [Sea80, BLW87, MBA88, SD87a, Siv88, Som84, SKN84, gW82]. Archimedian [Bil89]. Area [Ano83x, Fu80, Som84]. areal [LS80, Lac84]. areal-fraction [LS80]. Areas [Ano83y, GG81, WS84]. ARIMA [Liu89, Wei81a]. arising [KJ83, PX86, PR84, SL84, XP89, eL82a]. ARMA [BR85, NN82b, Pei86a, Pei88, RG88, SB85, SN89a, Tuc82]. arrangements [HK89]. arrays [BR85, Cho83a, KM84c, Sri87a, yS89, Tuc82, YKY85]. ary [PP81]. Aspects [FV87, CC82, KJ82, Nor89]. assay [AD83]. assays [Gup88c, SD87b]. Assessing [Koz86a, KJS^{+}83]. assessment [IC80b, JC87]. Assignment [SKN84, IV89]. assignments [Mor87]. associated [AK86b, Den89, JP88c, Mar85, NN82a, SS86d, Zac81b]. Association [AK83, AG83, SY87, GR81, Gil82, Gro84, KP89a, Kha87a, LSB89, Lin87c, WC87, Wol80, Net85]. assuming [PW81, PL86]. Assumption [SB82b, Nel87, PYO82]. assumptions [MS81d]. Asymmetric [Kha88a, Bol89b, Cha89c, KRS83, Mac82, Vos86]. Asymptotic [Bar88b, Bjø85, Blu85, CS83, Che86a, FKN83, Gim87, GN88, HJ83, Kim86, KY81, KB82b, KN83a, KN83b, LS83a, Lai89, LS83b, LWA80, Mag87, Mar85, NNQ84, OM80, PS85a, PS86, SS85a, Shi85, SW84b, SC86, Tah89, TW81, VJ88, Won82, Won88, YSD85, Yum85, Ang83b, CS88a, CW80, DN80, Del81, De 89, DM87, FG88, GM83a, GM85b, Gov89, Gov86, GL87, HB86, Hus88, JC84, KL84, Kok85, KT85, Koz86b, KN83c, Li87, MS81a, Pat85, Phi80, RM88, SMS84, Suz83, Woo80, Yu85]. Asymptotically [Che84b, Neu88, BW82, LK83, PG82, Sin85, Wit87b]. Asymptotics [JP87, Wit88]. Attaining [Obr82]. attribute [Kha84]. attributes [Bla84, HS82, SV89]. augmented [Cer89, SRP82]. Augmenting [Gov86]. Author [Ano83a]. autocorrelated [BBH82, SBC81]. autocorrelation [Bar84, Kan84, Mit88, SM87, SdB89, Zei89]. autocorrelations [DN80]. Autocovariance [NT87, Mit88, SI87]. Autocovariances [AR87]. Automatic [Hur87]. autoregessive [Mit88]. autoregression [KM87]. autoregressions [SL89a]. Autoregressive [AR87, TA83, BS86, BS87, Coo85, Coo87, EK85, Has80, LM84b, SL89a, Wil88b]. autoregressive- [Wil88b]. auxiliary [Hor85b, Hun89, Sin82, SKC83, SSK89]. availability [Ang86, EG88]. Average [AC85, NT87, AC84, DN80, Ete84, HH80, Koc85, LT85, Lin89, Mit88, SL88, SL89a]. averages [SB84a]. axiom [Vog87].

Backshift [Fin85]. Backshift-Operator [Fin85]. Bad [Kle87a]. Bahadur [FP84, GS81b]. Bailey [KS87]. Balakrishnan [O.’89]. balance [CH87, Gup87c, Gup88d, Hem80, Kag88, MB84a, Swa80, Swa82]. Balanced [WT83, AK86a, AM87, BK85c, BG88b, Cho83a, DK85, DSS81, Goo87, KC84a, Kag80, Kag81, KSR88, Khu82, Kun84, Kuw83, KM84c, Kuw89, Mag80, MP82d, PP88a, PK85b, Shi82c, Swa82, WK82, YKY85]. balancing [Pre89]. Ball [Ped85b, MT83b, Ped85a]. balls [IM88, Sob89b]. Ban [BB83]. Bands [Spu83b, Eva82, KM82]. Bandwidth [BT82]. Barlett [Sen87]. Barr [KJ90]. Bartlett [CM82, DN80]. basal [Fu80]. Based [SY87, WM87, AD80, Ang83a, Ang86, BE80, BK85a, Bal89, Bro82, CF89a, CW88, CH88, Cho80, DS81, DH89, ESP84, FY86, Fer87, FB82, FB89, GHK85, GP86, GR87c, GS80, Han80, HRR88, HR81, JV87, JP88c, JKK86, KL80, KP89a, Kea83, KM84a, KM88, Kha89a, Kle83, KB84, KKB87, Koz83, LaR84, LaR88, Leu88a, LR82, Llo87b, MF83, Mar84c, MR88, MP82c, Meh82, MDH81, MH88, MS81d, Nay85, NJ89, Nic89, PU86, PS86, Pre80, RS81, RMW86, SS85a, SR88b, Sat88, SG85a, aS82, Sic86, SIU85, Sin85, SS81b, Spa85, Spr89, SA81, Sta86, SG85b, Swa83, TB88a, TB84a, TB88c, TG89, TB89, Tit83, TT85, Von83, Wei81b, WS87, WJT89, Wha87, Wij81, Wu82, WE80]. baseline [Pat83, WF88]. batch [RB83b]. Bathtub [RR88a]. Battese [WS84]. Bayes [AM86, Alb81, Alb87, Ber89, CW80, CB82, Chu82, CL83, CB81, DP83, DB86, Dav87, EZ81, Geo86, Gho81a, Gho85, GLT89, GR82, Gru85, Güi82, Gut82, HW88, How85, HW89a, HW89b, Kar89a, Kar89b, LP88, Li82, Li87, Mar89a, MG87, PR86, RA80, SB84b, SG84b, SB82b, TZ85, TJZ88, Tiw88, TCJ89, TL89a, TT89a, TL89b, VS81, WH87, Wei81b, Woo80, Wri83, Yam86, Yam87, Yu86, ZS81, Zac81a, ZTJ86, Zeh85, dWM85]. Bayes-type [CB82]. Bayesian [ARP86, Abe80, Ach84, AB89, Alb85, AB86, AR81, Bar88b, BYD85, BD86c, Bru80, Cha89a, Coo85, Coo87, DG81, DGG89, DLP85, EN80b, ER83, Fon89, GG86, GA89, GK88b, Gho89b, Gun84, Gun85, HN84, HK83, JI84, KP89b, Kok85, Lec85, LM84b, ML83, Men81a, NH87, Nig89, NM88, OSG81, Owe85b, PW81, Pep88, PS85b, Rad84, RB83a, RH88, SB84a, Sin80a, SW83a, Sin83, SG88b, SG84c, SSS^{+}85, Smo83, SG86a, Spa88a, Ste86b, Str84, TB88a, Tra83a, Tra83b, Tra85, Tsu85, WT89, WS82, dlHRR88]. BCB [Rau87]. be [Fla85, HSC80, Roj86, SR88c]. Beale [SDCB83]. BEC [Ang89a]. behavior [BH84, BG84, DK81, GR81, Gel87, KY81, PS85a]. behaviour [Cer89, Hus88, KS82a, KS82b]. Behrens [AG89b, BR86, NvdM86, RF84b, SD86]. Beliefs [AG83]. bell [PG84]. belonging [Sam88]. beneficial [Can84]. Berkson [Dav85b]. Bernoulli [Ber89, GC80, GQ84, Jen83, KB80, PH88, San87, San89b, Tar86b]. Berry [CR89]. Bessel [De 81]. Best [BM87a, Cha82b, CS87a, Gro82, Kuw83, BB89b, BM86b, CV89a, EK88, HM81, Hsu85, HT83b, KL84, KP80b, Roj85, Sch80]. beta [Dan88, Kot89b, MN89, Ojo88, Ped85a, Ped86, PF84, Rau87, Wri88a]. beta-binomial [Dan88]. beta-correlated [Rau87]. better [BE84, Ber81a, dSBPR84, Che80, CP84, DK83, EZ89, EN81, Kum83, Ris85, Sha88a, WK81]. between [BR85, BP83, Cru82, Dal81, ESP84, Gil81, GC80, Goo87, GHO86, HG80, IB89, Joh86b, KR86, Kha87a, KN84, Lin87c, MM88, MS81b, MR86, Nan89, OS88, PKN89, Rao81a, Sat88, SA89, Shi83a, Sza85, Ter88, TT84, TG87, Wol80, YS81]. Bhattacharyya [GS85b]. Bias [Mag89, Wit87a, Bre82, Cha82b, GgT86, Mic85, Phi80, SN82]. biased [Ber81b, Cna85, FR88, KD89, Mas86, Tre80, TT84, TG87, gW82]. biases [Buo89, Lai89]. biasing [Kad81]. bib [WEA80]. bibliography [Die84a, Fra87, HK89, SKK85, STSM89, Wil81]. Bilinear [Liu85, NS86, TC81]. binary [Cru82, Dav84, KC84a, KK83, Moo88, PST87, Pro84, SC85]. Bingham [Woo88]. binned [SS85c]. Binomial [Bur81, BS80a, Che84a, Con89, Dan88, Dun81, DC84b, Gho80, Gut82, HN86, HW88, Har83, HT83a, Jel87, Kem86a, Kim88a, Lem81, Ng89a, O’C87a, OL86, Pau85, Rau87, Ris85, Som85, Tea84b, WH87, Yan89a]. bioassay [Anb84, CHW87]. bioavailability [AH83, Fri87]. biological [AD83, WYF87]. Biopharmaceutical [HES84]. Biostatistics [Cor82, Har82]. bipartite [Swa82]. birth [BE88, PM82]. Bivariate [DP83, HH86a, HW87b, Pap86, SlR87, SJ87b, AT84, Bar80, BE84, Bil89, CO84, Con85, DW89b, EZ89, Eva82, FS80, GMM84, HN86, HW87a, Hut84, Kel86, KR84, KN87, KK81, KS82a, KS82b, Koc86, KD82, KM80, MY85, Nel86, PL88, PH89, Smi83, SH84, SH89, TB88b, Tik88a, TG89, TGB89, Wei81b, Won82, eL82a]. bivariates [MT85]. biweight [Kaf82]. blackjack [ES88]. block [AK86a, BK85c, BM85b, Ber82, Bha82, BS84, BW88, DSS81, Fei88, GSS87, Gup88d, HT83b, MI82, Per81, SC85, SD87b, Stu88, Whi89b, WT83, YB83]. blocking [Lor84, Lun82, Per81]. blocks [Che83b, FY86, Fer87, SQ81]. Board [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano82u, Ano82v, Ano82w, Ano82x, Ano82y, Ano83x, Ano83y, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano83u]. Board [Ano83v, Ano83w, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Owe81c, PB83c]. Bock [LS81]. Bonferroni [Bri87, SH89]. Boolean [Dav87]. Bootstrap [Sri87b, Swa83, Ang89b, CS88a, Fal86, Sch87, Ste86b, Swa86b, Swa86a, Wan89b]. Bootstrapping [HS86, Jou88, Sha89, Sho82a, Gu87]. borrowing [Hor85b]. Bose [WYF87]. Both [Ala87, CS88a, EZ81, LY89, Sel89]. bound [CR89, Cho83a, Gup88a, Rob89, Sch83]. boundaries [BA87]. Boundary [Joh84, De 81, Yu85]. Bounded [AS84b, Dor89, kL82b, DD85b, Fra83, Hay89, Yu86]. Bounded-leverage [Dor89]. Bounds [Arn88, Kad81, Kor85, WW84, CB87, CB88, CV89a, CM82, Dav88a, DJD89, Gro82, KPS83, MP82a, NH87, Nig88, PS83, SS86c, Som84, WE80]. Box [Sar85, SN82]. Box-cox [Sar85]. branching [Des84, JB86, Pur86]. breakdown [JP87, Sim87]. bridges [Kul88]. broad [Rub86b, Spa86]. Brownian [Kul88]. BTIB [WT83]. Buckley [WS87]. Burr [ER83, Nig88]. Busy [RB83b].

C [KK85]. Calculations [Lev83, FD85]. calculus [HK89]. calibration [DFG87a, Fox89, LY89, Per87]. Caliper [Ray83]. CALS [KW84]. can [Kal84, SR88c]. Canada [BM85c]. cancer [PH84]. canonical [Kha89b, WC87]. capability [CO89]. capital [Mad85]. Capture [WM87, SMS84, VS86]. Capture-Recapture [WM87]. carcinocenicity [Cru82]. carcinogens [Rao86]. Carlo [Ber86, DD85a, Fri81, Kul88, Mar85]. carriers [Mil82a]. case [AH86a, AH86b, BKZ88, BM86a, Bet87, Cha83b, DS82, GG88, Isa81, JK85, LY89, MB84a, Mil82a, MM89a, MS80b, NNQ84, Pad80, Rao81b, Rao86, Ray83, SS83, SA85, Tse89, WS84, Wef87, WT89]. case-control [Rao86, Ray83]. cases [Din87a, Din87b, GS81b, KC87a, KC87c, MC85]. CATANOVA [AL80, AL82b]. Categorical [BB83, SKSF83, SA81, Bha82, Bou82, Dre85, GA89, LS81, MA83, MA84, WL82]. categories [Gil82]. categorized [SG84c]. category [BA87]. Cauchy [Bal87c, SHB85]. cause [Din87a, Din87b, Ebr86, KC87a, KC87b, KC87c]. causes [KB81, KB82a, KC87b]. Cell [Lev83, BSC80, Cau84, CB82, Hoc88, Kan88b, KJS^{+}83, MP85, MS85, WJT89]. cells [BB89b, Mur89, PM85]. Censored [Sin83, Win87, AB89, BT83, Bar86, CM86b, Che84b, Che87b, CEA85, CT89, Ebr84, EL82b, ER83, FW86, FH81, FH84, Fro87, GHK85, GM83a, Ger88, GR87b, GSS87, Kim86, KRB89, LaR88, LPW82, LP85, MP82c, Meh82, MC87b, Mus89, MS81d, Neu88, Nic89, PM84, PS80, PR85, SR88a, Sch84b, Sel89, SJ87a, SJ89a, Sin85, Sir84, TB88a, TG89, Tit83, VJ88, Wei89, WAS88, Wu82, WE80]. censoring [Cha88a, CB87, CB88, Ebr85, GM88, Kim88b, KF89, LP87c, ME86, PM89a, Sir86, Spu83a, TJZ88, Wei80, YAH89]. censorship [GV89a, LP87d, PS85a, PR88]. centers [HB86]. central [Ang89b, HLN88, NNQ84, OL86, Poc82, TG83]. centrally [Kar88]. Certain [Joa83, Cha86a, DJD89, Gal80, Gho85, MS86b, MS80a, Ng89a, PX86, RK81, Rad85, Rao87, Sil86]. Chain [SSK89, CF86, WS89]. chains [And89b, BD86a, YS82b, YS82a, ZS89]. challenges [Net85]. change [AB89, CF86, Ebr86, GS87, MZ88, MB84b, MB85, RG89, Sak88, SW88, Yao86]. change-over [GS87]. change-point [AB89, Ebr86, SW88, Yao86]. changepoint [Sch83]. changes [Haw89a, MSB85, SBC81, Sch82]. changing [ML83, MB85, Ng89b]. characteristic [BP83, DK81, Hol89a, MCR88, PF84]. characteristics [Scl80]. Characterization [ARS86, Gup87a, JR82, KA87a, KA80, Nis89, PT86, YN86, AT84, GG83, Kan84, Mar88b, RJ84]. Characterizations [DL88, Nar83, Sta86, Ahs87, Gib84, Gor87, Tal86]. characterizing [Tal80]. characters [KKB87, SSK89]. Charts [SM87, CAR89, GRY81]. Chebyshev [Bha87]. check [AV88]. chemical [Kag88, Swa82]. Chen [Din87a, Din87b]. Chernoff [TT83]. Chernoff-type [TT83]. Chi [Ala87, BB83, Gim87, KW86, MS81c, BA87, BWD86, CARO84, Dav85b, HH86a, Hos86, McC85]. Chi-Square [BB83, Gim87, KW86, MS81c, BWD86, CARO84, HH86a, Hos86, McC85]. chi-squared [BA87, Dav85b]. chief [SK88, SSKK88]. children [Rao81a]. Chinese [Lin86]. Chipman [MC85]. choice [Dey83a, GR83, GM85a, PRR86, RM89, SA89, SL83]. Choosing [EKR86, GC80]. chosen [Che88]. Circular [Mag80, DM87, Kun84, MH86, Wol80]. claim [Har85b]. class [AK86a, AS84a, Ali89b, BF85, DM89, EN81, FH81, HT81, KP85, KI88, Kam88, Koo85, Kum83, Kuw83, MB82, Ng89a, Pre80, RS80, RS88, Rub86b, SM86a, Sho84, Spa86, SKN84, SM82b, Tri88, Urz88, Wu89]. classes [Abe80, KSR88, MS82, Rob89]. classical [Vos86]. Classification [Bha86, LS83b, BS87, BG85, BMC86, DS80, DS82, GM80a, HG86, KC84a, KP89b, LS83c, Leu88a, Leu88b, Mar84a, MM89a, MAT82, TB84a, TB89, VW83, Wan86, YO86]. Classifications [AK83, Gil82, Gup84, KK83]. classifiers [Kok85]. Classifying [KP82, PP87, Dun82a]. clinical [Ach84, Anb84, AH83, Can84, CLl86, DL84, Gra87, Lai84, LDH84, PH80, PK81, TA83]. Closed [BK84, Bha86, ES84, Fu80, LK83]. closure [Bev84]. Cluster [BM85b, GM80a, HB86, KW86, XP89]. clustered [GWBG87]. clustering [DW87b, Men81a, Pro84, TT83, Wil86]. clusters [GN83, Sin88, Won82]. co [Dey83a]. co-ordinates [Dey83a]. coal [LM84a]. coefficient [CY86, Ebr84, Fra88, Gad83, GHO86, Gup87b, HI89, Ima87, IB89, Iwa89, JS82, Ken81, KS82a, KS82b, LSB89, Lee81, MSL86, Pau88, Sam84, Sat88, Sha80, SA85, Sin86, SL83, SH84, SL85, TCJ89, Tsu85, Via82, YI88]. coefficients [AM82, Ali89b, Chi81, Haf86, Hos87b, Hos88, Li88, Oht88, PS83, Rao81b, Rie80, RU89, Str84, TG83, TSM86, Ven80, ZG80]. coherent [CV89b]. cohort [Bre82, KJS^{+}83, RM88]. collapsibility [Dav86]. collapsing [PG89]. collinear [Mas86]. Collinearity [MP89, Wal89]. collinearity- [Wal89]. column [ER80, Joe88, Sha88a, WK86]. combination [AO82, DK83, Sch89, TG83]. combinations [Buo85, CVV82, Mat83, SSS87]. Combinatoric [DS80, Dun82a, DS82]. Combined [Via82, Sha88a]. Combining [BP84, JP87, NO85, PS86, TO85]. Comment [Pag81, Kle87a]. Comments [Atw80, Eas80, KM84a, Law80, MB80, MSVS80, Ney80b, Smi80, Bal87d, BF85, Gab80, IC80a, KJ90]. commodities [FV89]. Common [AT81, AS84b, Bha81, CHS86, GM83b, Hei88, Kha83b, Kor85, Kub89, LP87a, MS81c, Oht87a, Pau88, SM82a]. Communications [Chh84, SK88, SSKK88, MSVS80, Smi80, Ano83a, Ano83z, Ney80b]. commutative [El 81, Elp89]. comovements [YS81]. Comparative [AB87, AH83, BT86, Bro87, Fri87, KO84]. Comparing [AP86a, Huf85, BS80b, EH83b, KW86, LP81, Maj88, Mat81, Sin80a, SAH87, Tan82, Tar86b, WT83]. comparisions [Mag87]. Comparison [Bjø82, BCK85, CS88a, CK87, Dun82a, KT84, KR84, Kha87b, LP87a, LS83b, Sch85b, Sir86, Spu83a, SC89, SD87c, ZG80, AC84, BK86, Ban80a, CDS84, DD85a, DW87b, Dun86, EHA84, Fon89, FR88, Fra89, Gad83, GM80a, GBS83, Gru85, Gun85, HH86b, HT81, Hei88, Lam89, MB84a, ME86, NJ89, SQ82, SG87, Sim84, Sri89, Ter81, TG87, Tse89]. Comparisons [Hel88a, Pri82, SC80, SB82b, Tre83, Bal87c, Bas89, BG88a, CW87, DW87a, DW89a, Dav88b, Dun82b, EK88, FP84, Hsu85, KPS88, LY89, MA83, MAT82, Sad82, Som85, Ter82, Ter86, Ter88, TT86, Tre80, öT85, Von83]. compartmental [PM82, PM83]. compendium [Koz86a]. compensation [GH84]. competing [BDK89, BG80, Ebr86, GW84, KB81, KB82a, RMW86, RM88, Rat89]. competitors [PG84, SG86c, Yan88]. complementary [BG80, CH87, Nan89]. Complete [RS87, Abe80, AH86a, AH86b, Ang89c, BM85b, BS84, FY86, HT83b, Per81, SQ81, SR88a, SFR84, SC85, WAS88, WT83, WS82, WL80]. complete-data [AH86a, AH86b]. completely [Bra88, VB80, Von83, WEA80]. completeness [YI88]. Complex [KL86, CB85, FKN83, NNQ84]. component [AJG82, BP84, BC88, Cer89, Dav87, Hsu81, MS86b, NO85, Obr82, Sch85b, Soo88, Tan88]. components [AM87, BKZ88, BG88b, Coh80, Czi89b, EG88, GD88, Gie83, Kap82a, Lai85, LMP88, MAT82, Oks88, PRPR88, RB83a, Rao82, RP88, Ros89, SKK85, Sch86b, Sch86c, SJ89b, Som88, Tan88, Ver88a, Ver88b, WEA80, WG81, WK82]. composed [Lai85]. Composite [Gue87, BP87, Goo89, SN89a]. composition [JRF80]. compound [AM88, AR81, HG86, SJ87b]. compounded [MR83]. compounding [WCK85]. Computation [Sha81, ST86b, Rao81b, eL82a]. compute [Gov89]. Computed [Gim87, Sil88]. Computer [Rod85, IC80b, MS87, Wha87]. computerized [Dou86]. Computing [MG87, Ram87, Fra80, Gie83, Mon87, SR88c]. concatenated [JP88a]. concentration [BM87b]. concentric [AT81]. concept [Pre80]. concepts [BT81b]. concerned [MSVS80, Ney80b, Smi80]. concerning [Har85b, MS80a, MT84, PH89, Sch86c, Seh88, SL82a, Vog87]. concomitant [Mos89]. Concordance [DP83, FP89a, FV87]. concurrence [CP82]. condensed [SKC83]. condition [GC84, KL84, Roj86, Sta86, Tal86]. Conditional [AGR89, Fox89, PL88, PR85, VS86, AT84, Ahm82, Bon88, DFG87a, DFG87b, Han80, HRR88, Llo87b, Mar89b, PKG87, SU86, Sob89a, Wal84a, Yan88]. Conditionally [Rus85, Nay88, Sri87d]. conditioned [GdT87]. Conditions [McC82, PK85a, Gal80, HZ83, Hos87b, KP89c, Sil86, SS88d]. Confidence [Ars87, AJG82, BG85, BMC86, CH86, Eva82, Fla87, FGV89, LMP88, Pat81, Sob87, Spu83b, Wal87, WG81, Wei81c, Wei82, Wij81, WM87, YS82b, YS82a, AS84b, Asr85, Bis80, Bof85, BG88b, CB87, CB88, CV89a, Chi81, CHS86, EG88, JH88, JK85, JKK86, KN85, Kab83, KB83, KL80, KM82, Kon81, Kor87, Mar82, Mil82b, MM88, Muk81, Owe85b, PS83, PRPR88, RMW86, RH88, Sau82, Sch83, Sch87, aS89b, Sob89a, SD85, SvW82, Tah89, Tub80, Ver88a, Wie85, Wit89a, ZLMK82]. confidence-interval [JH88]. configuration [Che86b, MS80a, Seh88]. conflict [SS87b]. confounded [Lin86, Lin87b]. Confounding [Lin87a, Bed89a, Bed89b, Vos86]. conjecture [BDF^{+}84, MS80a, Seh88]. conjectures [RF84a]. Conjugate [DG81, Alb88, SJM86]. conjugates [ARP86]. connected [Gup88a, KN83a, Nag80, Pro86, Pur86, RS88]. connectedness [MS85, Mur89, Rus80]. connectivity [DL86]. consensus [Owe85b]. consequences [DK85]. conservative [Sch83]. considerations [Jai81, Lai84]. Consistency [Gal82, Sha88b, Bjø85, Gup86, LP84, Sha86]. consistent [KW84]. Constant [AB89, Ban80a, Iwa89, Ket82, NM88, YI88]. constant-utility [Ban80a]. constrain [PH88]. Constrained [Lee87a, Bon88, CB82, Far88, Gru85, MS80b, RLN86, Smi85]. constraints [Cau83, Cau84, Cho83a, DF89, ES84, ES87, Far86, Far87, FG88, PP88b, Sch85c, SL89b, YKY85]. construct [Hut84, KP84]. Constructing [Ray82, KM84b, Lin86, Lin87b, Sau82]. Construction [KP87, Muk86, Cha89b, Dal88, Kab83, Swa86b, WNK80, YB83]. Constructions [Kag81, PK85b, BK85c]. contagion [Wal84b]. contain [Hab89]. containing [Bal88]. contains [Che80]. contamination [Gat89, KRS83, PC82]. contaminations [Siv88]. contemporal [WA81]. content [AT81]. contents [Ped86]. Contextual [KP89b]. Contingency [AG83, DG83, Mic83, Alb87, AL80, AL82b, Bjø82, BWD86, Cau83, Cau84, CQ82, Chr86, Chu84, CGO85, Coh80, Dav86, Gen88, Gro84, Güi82, Hab85, Hab89, Hos86, Joe88, Kha83a, Mar89a, MP86, SG84c, SS86d]. Continuous [BMA88, BA84, HW87b, AM84a, And89b, BD86a, CA81, Fra89, GMM84, HW87a, MBA88, PC82, Raf84, Ray83, Wol80, YSTM88]. continuously [Rub86a]. contour [Fic84]. contours [DW89b, Sau82]. Contrast [OAH89]. contrasts [KB83, Mar82, Tse89]. control [Alw86, Ber81a, BM85c, BD86c, Che80, CP84, Chr88, DW89a, DS87, EK88, Jac85, Kat80, KP82, KP84, LC82, Luc85, Lyb85, Maj88, Nat85, PP87, Rao86, Ray83, Rod85, Sch85a, SM85, SM89, Sri87a, Tse88, WT83]. control-multivariate [Alw86]. controlling [Muk86]. controversy [Pie83]. Convergence [And83, Boh83, Muk81, Tiw88, Che88, Fin84, Fra88, Mus89]. Convex [Sch86a, Bha86, MT83b]. convexity [EH83a]. convolution [Fin84, Har85b]. coordinate [dR89]. coordinate-free [dR89]. corner [KT85, PD84]. correct [BE80, Bof84, Jen83]. Corrected [SD84b]. Correcting [Ste89a]. Correctings [Tar86a]. Correction [Wan09, SN82]. Corrections [Ban80b]. Correlated [PL83, Ali87, Fei82, GS85a, GS87, Hab85, Hol89b, IS85, LS83c, Mar84d, PL82, Rau87, Tub86]. Correlation [Rao81a, SL82b, AC85, BMC86, CD89, Fra88, GR83, Goo87, GHO86, Gup87b, GN87, Ima87, IS87, Joh86b, Ken81, KS82a, KS82b, LP87b, MC86, Pau88, PL86, Rao81b, Rav89, Rie80, SQ82, SR88b, Sat88, SND84, SM86a, Sha86, Shi82b, Shi82a, SW84a, Six81, SL80b, SN82, SL83, Sri83, SH84, SL85, TCJ89, Tub80, Via82, ZG80]. correlations [JP88a, Kha89b, Rie82, WC87]. corresponding [BB89a, BH84, KP87]. cost [Ber85, ESP84]. Costello [Dav88b]. costs [Ham89, HL85]. could [Fla85]. counterexample [Har85b]. Counterexamples [BH80, RF84a]. counting [BJ88, Wec85]. county [WS84]. Covariance [SL82b, tFAH84, AH86a, AH86b, BKZ88, BF86, Bed88, BLW87, Byr88, El 81, Fra80, GS85b, GR89, GN87, HH86b, Hie85, HF80, Kan88a, Kay89, Mag89, Mat81, Nag85, O.’89, PL86, Pei88, Ray85, Rei84, RS89, Sak87, SD87a, SOL81, Smi85, Sza85, TB85, TGB89, Wie85, Yas81b, YO86]. Covariances [Shi83a, MM86, Shi83b]. Covariate [LS83b, LS83c, Leu88b]. covariates [Arb85, CEA85, Eri83, ME86, Per82, Saw87, SF85]. coverage [Shi87]. cox [Sar85]. Cramér [Ang83b, BWMD83, CVV82, PD81, WE80]. criminology [LB84]. Criteria [BB83, AOA89, Bas89]. Criterion [Kha87b, DBO83, EKR86, GN87, Kab86, Lis88, Man80, Mik85, Pri82, RRM86, Sam89]. Critical [Miy87, CM82, ID80, Nel82, Sch86b]. criticisms [Far83]. crop [CHLL84, Hey84, SG84a, WS84]. Cross [AK83, Loy87, AJG82, CM86a, DGG89, Gil82, Hur87, JP88a, KK83, MC86, MC87a, RM89, Tut89].

Cross-Classifications [AK83, Gil82, KK83]. cross-correlations [JP88a]. cross-effects [CM86a]. cross-entropy [DGG89]. cross-lagged [CM86a, MC86, MC87a]. cross-validation [Hur87, Tut89]. Cross-validatory [Loy87, RM89]. crosses [KP89a]. Crossing [De 81, MCR88]. crossings [Yu85]. crossover [PH80, Pat83, WF88]. crude [Mic87]. cube [LL82, Lun82]. cubic [IV89, Li88]. Cumulants [TC81]. Cumulative [Luc85, CARO84, Iwa86a, Mal80, Pat85, SH89]. curtailed [Eis81, KSJ86]. curve [Bra81, EH83a, KD83, Kub89, Kul87b, LC87, OS86, PB85, Tam86, WL80]. curves [KS81a, KP84, MCR88, Tam86, Zeh85]. customary [Sri87a]. cusum [LC82, Luc85, Zac81b, Alw86]. cutoffs [PKG87]. cyclic [DL86, Gup87c, mN83b].

damage [Nar83]. Darling [Boo82, BWMD83]. Data [BB83, Gim87, Kor87, PL83, SKSF83, Sin83, Wan09, Win87, AB89, AP86a, Ahm82, AH86a, AH86b, Ang86, Arc82, BT83, Bar89, Bha82, Blu85, Blu87, Bou82, Bra81, CB85, CM86b, Che84b, Che87b, CC88, CEA85, CT89, Coh82, Dan88, Dav87, Dav84, DeP83, DS86b, Dre85, DH89, EL82b, Fei82, FW86, FH81, FH84, FS86, GHK85, GM83a, Gen88, GA89, GWBG87, GR87b, GW84, GSS87, Hai88, HGB89, Ham89, Hem80, Hey84, HZ82, HT83a, Hoc81, HSC80, Hol89b, HH84, JC87, JL89, KC84a, KO82, KK83, KN87, Khu82, KP89b, KKC82, Krä85, KJS^{+}83, LaR88, LPW82, Lee87b, LC87, LS81, Mar84d, Mar89b, McC87, MP82c, MA83, Men81a, MA84, MI82, MC87b, MS81d, Nan89, Neu88, Nic89, PM84, Pad80, PK81, PRR86, PL82, Poc82, Pro84]. data [Pro88, Rav81, Sch84a, Sch84b, SS85c, Sea80, Sea88, SM86a, SC85, Sir84, SG84c, Som85, Spa85, Sri85, SC86, Ste89a, Tit83, TYM86, VJ88, WK82, Wei89, WT89, WL82, WL80, WE80]. Data-randomized [Kor87]. date [Blu88]. dates [KY82]. Davis [Ste89c, YSD85]. Deal [Kor85]. death [BE88, Din87a, Din87b, Ebr86, KC87a, KC87b, KC87c, PM82, RMW86]. decaying [PR84]. deciding [ESP84]. Decision [Pal88, Bjø81, Bof84, ESP84, HG86, Hal84, HT83b, Men81b, Ven80]. Decision-theoretic [Pal88, Bjø81, Men81b]. decisions [PO84b, PO87, PP87]. decomposition [Chu84, WK82]. decompositions [Tom87]. decreasing [MP82c, Nic89]. deduced [Rob89]. defect [JK84]. defective [BCK85]. defectives [CH87, KJ83, WR87]. defects [HL85]. Deficiencies [YM89]. Deficiency [Gho89a]. defined [Sch84a, Spa86]. definite [AP86b, Kik83, LP89]. definiteness [Smi85]. definition [Gun83]. Definitions [Hel88b, Hel88a]. degenerate [Umb81]. Degradable [ENP84]. degree [Gil81, LT84, OS86, SA89, YM89, dlHRR88]. degrees [Som84, SN82, Vos88]. delay [Pra83, ZF87]. demand [BP84, PG84]. Dempster [Sch89]. Densities [HW87b, AT84, HW87a, IB89, MP82c, Nis89, Sch85c, SB84b, Swa87, Won88]. Density [KF89, Win87, BT82, CARO84, ĆM89, Elp83, FR88, Jou88, Kam88, KV80, KLV85, KRB89, LP85, MT82, Nic89, PM84, Ped86, Pou83, SS85c, Swa86b, Tit83]. departure [Tan82]. Dependence [HW87a, HW87b, Bil89, BT81b, CK83a, Ebr82, Gho81b, Gil81, KS81b, KB80, Wol81, dlHRR88]. dependent [Adk86, Bal87d, Bjø82, Bor85, CLMFC87, Cha82a, Gal88, Hig82, Lai85, MR88, Mus89, Pei86a, Pei88, Riv88, SD88, SS88b, Yu88, dlHRR88]. depletion [EN80a]. Derivation [Cha86a, Mit88, GW84, Sam85]. derivations [Feu87]. derivative [KV80]. derivatives [Hol89b]. derived [AD83, KM84c, LS83a, Ros89, YM85]. deriving [Mos83a]. Description [LL82]. Design [KM81, Ach84, BK85c, BM85b, Box85, BS84, Bra88, Bro87, BW88, Cha89a, Cha89b, Cha87, Czi89b, DK85, EHA84, Gab87, Gho83a, GC80, HH86b, HT83b, Kap82a, Kar88, Kat80, Kee80, Kha84, KP86, Kos87, KS84, KM87, Lai84, Lor84, LL82, McC80, MS87, MM83, Obr82, PH80, Pat83, Per81, PH84, Rat89, SC85, VB80, WF88, WEA80]. Designs [Cru82, Gup88c, Kun84, AK86a, AH84, Bal87d, Bha82, Bla84, Bou82, BV89, Cha89c, CHLL84, Czi88, DW87a, DW89a, DL86, DSS81, ER80, Far83, Fei88, GP86, Gho82, Gho87b, GS85a, GS87, Goo87, GSS87, Gup84, Gup86, Gup87c, Gup88a, Gup88d, Gup88b, HES84, HF80, Joh85, JRF80, Kag80, Kag81, KP85, Kag88, KSR88, KM84b, Kle87a, Kle87b, KP87, KM87, Kuw83, KM84c, Kuw89, LR87, LDH84, Lee87a, Lin86, Lin87b, Lun82, Mag80, Maj88, MB84a, MK82, MP82d, MH85, Muk86, MH86, Muk87a, MB89, MM89b, MAT82, OP86, Odh86, OM87, OP84, PP88a, PP81, PO84a, PO84b, PL82, PK85b, RS80, RS84b, Rus80, SRP82, Sha88a, Shi82c, SD87b, SW83b, Sri87a, Str88, Stu88, Swa80, Swa82, Tri88, Ver88a, Von83, Vos88, Wan83, WR87, WK81, Whi89b]. designs [WT83, YB83, Yeh88]. Detecting [CH87, Jai81, MC86, Sak88, AB87, MYT87, Sch82, Spa88b]. Detection [Bal89, Bha89, FP89a, KHB^{+}88, KN87, MZ88, Nae86, Nai89, PB87, Wal89, BT86, JP81b, JPD82, Kim88a, NDM84, PCP87, Pro84, RG89, Sim84, Spr89]. determinant [Hie85]. determinants [HLN88]. determination [Bla84, Rao86, RRM86]. determinations [Che84a]. Determining [SL80b]. detour [MS88, Sen84]. deviance [AAY86]. deviate [SYT89]. deviation [CVV82, Wit89b]. DFR [Bev84]. DFR-mixture [Bev84]. DGP [SU86]. diagnosing [CGO85]. diagnosis [Bra84, Tea84a, eL82a]. diagnostic [Atk82, RC85, Wei89, WL82]. Diagnostics [PP88b, Gra89, Old87, Wan87c]. Diagonal [Tom87]. diagonals [Tom87]. diagonals-parameter [Tom87]. diallel [KP89a]. dichotomous [Alb85, BV89, Fra89, Gun84, Smo83]. differ [JC87]. Difference [SY87, YS81, AP86b, BC88, Cha84, Kha89a, LP89, Mic87, RD88, YG82]. differences [Con89, OAH89, SD87c, Wil88b]. different [Cha80, GM83b]. differential [Gad83]. difficult [Har82]. diffusion [BE88]. dilemma [HP83]. Dimensional [AK83, Hos86, Old87, Scl80]. dimensionality [BM82, Kok85, NS84, Sam89]. dimensions [GS85a, KP86, Kos87, MC85, PK85a]. Dinse [KC87b, KC87c]. Direct [Khu82]. direction [Gun83]. Directional [Mar81a]. Dirichlet [CD87, CC85, Dan88, Den89, Tiw88, Wil86, Yam86, Yam87, Yas81a]. Dirichlet-multinomial [CD87, CC85, Dan88]. discovering [TT89a]. discrepancies [Kep86]. Discrete [Cha86b, Kha84, AM81, Alb88, ARS86, Bal86, Bar89, BA84, CB85, Cha84, Cha86a, Che86a, Dah86, Dey83b, Gil81, GCH83, Hug80, Kep86, Kha83c, KS82a, KK86b, Kor87, PL88, PT86, PRR86, SLH83, Sta86, Tal86, Tut89, VS81, Xek83a, ZS89]. Discriminant [KD83, AA81, BK85b, BCL80, Chi81, ĆM89, GR89, Haf86, Leu89, Men81b, NS85, Rad85]. Discriminants [LS83b]. discriminating [Cru82]. Discrimination [Nan89, Wu88, Bor83, CI80, Hor80, Job87, LWA89, Mar84a, NR83, Tut89]. Discussion [Din87a, Din87b, KC87b, KC87c, OEJ^{+}85]. disease [Cla82, Kha87a, Wol81]. Dispersion [Kha87b, Cha80, DLP85, Gho87b, PM88, Sie83, Wec85]. displacement [CPW88]. display [Coh80]. displays [Obe84]. Distance [Kam88, BP83, Cha81, Kan88a, KG87, LR82, PG84, Rad84, Ros89, Sch86a, TT84, TT89b, Wil81]. distances [CV89a]. distinct [Bed87]. Distributed [Sin83, AM84b, BB89a, GHO86, Mar88a, Ped85a, Ped85b, Sch87, SH82, Spa85, TSM86]. Distribution [AR87, AO82, Ali87, Gim87, Gov89, Joa83, KD89, OW80, Roh87, Smi82, SD85, Win87, AC82, AWS87, Ahs87, AH84, AC84, Aly87, AK86b, AH86a, AH86b, Ang83b, Ang83a, Asr85, BDK89, BE83, BM86a, BM87a, Bal89, BT83, BF86, Beg81a, BHL82, BLW87, Ber81b, BB89b, Blu87, Blu88, BCK85, BM88, CS83, CD87, CS88a, CS81, CP84, CV89a, Cho83b, CO84, Cho88, CC85, CW82, CS84b, CF88, Con88, CF89b, CW87, CA81, DW87a, DW89a, Dan88, Dav88b, DeP83, Del81, Den89, Dey89, DLP85, Dix89, Dun81, DC84a, Ebr84, Ebr85, Elp83, EK82, ER83, FCP^{+}82, Fal86, FKN83, FH84, Fra88, FS80, Gat86, Gho81a, Gho80, GLT89, Gle82, GM85b, Gor87, GL87, GN87, GN88, GV89b, GCH83, HG86, HN86, HD89b, HD89a, Hil85, HI89, How85, HW89a]. distribution [HW89b, HLN88, HF88, Hut84, ISS82, Iwa86a, Iwa86b, Iwa89, JG85, JC84, JS82, KS85, KI88, Kan82, Kap86, KM84a, Kel86, Kem86b, Kim85, KA80, KK81, KS82b, Koc86, KD82, KB82b, Koz83, Koz86b, KN83a, KN83b, KN83c, KC80, LK83, kL82b, Lam86, LSB89, Lee87b, Lem81, LWA80, Llo87b, LM84c, Mal80, MT85, MBA88, MR88, Mee88, MG87, MD89, Moo89, NN82b, NCLN82, NNQ84, NS85, Nig89, NM88, OL86, OSH88, OW86, PS80, PM88, PM89a, PR84, Par81, PM83, Pat85, Pau85, PH89, PG82, PM85, PR86, Pro86, Rad82, Raf84, RG84, RJ84, Rao86, RD88, Rau82, Rau87, RR87, Riv88, Rob81, SHB85, Sam84, Sam85, SL80a, SS81a, Sch83, Sen80, SI85, Sha81, Sha84, She86, Sho82c, Sho82b, Sic86, SA85, SGP83, SG88b, Sir84]. distribution [Smi83, SMS84, SW84b, Spa88b, SS88c, SA82, Sri83, Sri89, SJ88, SH89, SN89b, SB89b, SvW82, Swa88, TP86, TS81b, TG85, Umb81, VI89, WSR84, Wal84a, Wal84b, WA84, Wil86, Win82, Wri83, WH88, Xek83b, Xek86, Yan88, Yan89a, YB82, YG82, YM85, Zac81b, dlHRR88]. Distribution-free [OW80, Smi82, SD85, BDK89, GL87, HG86, Lee87b, RD88, Sch83, VI89]. Distributional [KK86a, Kap82b, KJ82, Nag80]. Distributions [BMA88, CK83b, DP83, KA87b, MT85, NN82a, Pan83, ARP86, AM81, AS84a, AM87, ARS86, Amo84, Ang89a, BK85b, Bal86, Bal87c, Bar80, BG80, Beg80, Beg81b, Bra88, BWMD83, Cha84, Cha86a, Cha86b, Che86a, CK83a, CO89, CD89, CV89b, Con81, CHS86, Dah86, De 89, EA89, EN81, EP85, EkBL88, FKN83, Fra89, Fro89, Gel83, Gho81d, Gro86, HP87, Har83, Har85b, HH86a, HH89, IB89, Isa81, JSB89, JR82, Joh82, JKS80, JB84, Kan88a, Kep86, KA87a, Kha83c, KS81b, Kim88a, KL84, KK86b, Kor87, Kor85, KJ83, KL86, Kul88, Kum80, Kum81, LT85, Loy89, Mac82, MT82, Mar84c, McD87, Men88, MS86b, MS81a, MDH81, MM88, Mos83a, Nag85, NCLN82, Nay88, Nel86, Ng89a, Ng89b, Oht87a, PX86, Pap86, PL88, PT86, PRR86]. distributions [PG89, PKN89, PTA89, Pil85, RR88a, RR88b, RM88, RB88, RS88, Rub86b, Rus85, SR88a, Saw88, SS88a, SYT89, SJ87a, SJ89a, SI81, SJM86, Sin85, SJ89b, Sir86, Six81, SL84, SC86, Sta86, SJ87b, SS86d, Tal86, Tal80, TS82, Tra85, TGG86, TG86, TG87, Urz88, VJ88, Wei81b, Wri88a, Xek83a, XP89, YM89, Yas81a, Yat86, dWM85, vdMN87]. disturbance [Oht88]. disturbances [KR86, KO84]. diversity [CG88, Gil89, Nay85]. Divisia [FV89, Vog86, Vog89]. divisible [Gup88c]. Do [Hab89, Huf85, Kal84]. domain [Hur87]. dominance [BHL82, DS85]. domination [DB86, KT86]. done [Huf85]. Dorfman [HJK89, JKR89, KJG88]. Dorman [KSJ86]. Dorman-type [KSJ86]. Dosage [AD83, EZ81]. Dosage-response [AD83]. dose [Cru82, Wal87]. Double [SV89, SM89, Bal87c, HN84, MDH81, Sho82c, SN89b]. double- [HN84]. doubly [Ebr84, SJ87a, SJ89a, YM85]. Douglas [Dav88b]. down [Ima87, Kab84, MS88]. drawn [Par81]. drug [BV89, KK86a]. dry [Llo81]. dual [Isa86, IS87, SG87]. duality [And89b, OS88]. due [BDK89, Mar89b]. dummy [KD83]. Durbin [She86, Zei89]. Dutter [Sil88]. dynamic [BYD85, Gov89, GH84, LT80, TP85, Wil88a]. dynamics [RG89].

each [Bla84, Ebr86, Wri83]. earth [Nar81]. econometric [SU86]. econometrics [Oks88]. economic [Lai84]. Economically [CA86]. EDIOOq [Tam86]. editor [Ney80a, Zer81, SK88, SSKK88]. Editor-in-chief [SK88, SSKK88]. Editorial [Ano86a, Ano88a, Ano89a, GS81a, Owe81a, Owe81b, Owe85a, PB83a, PB83b, PB84a, PB84b, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano82u, Ano82v, Ano82w, Ano82x, Ano82y, Ano83x, Ano83y, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l]. Editorial [Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j]. Editorial [Ano89k, Ano89l, Ano89m, Owe81c, PB83c]. editors [OEJ^{+}85]. education [God85]. Effect [MK82, ME86, Spa88a, Spa89, Wei81a, AB86, BFR80, Bre82, BCL80, CH86, DK89, Fla85, GgT86, GC84, MYT87, NS83, Per81, Roj85, Tub80, Wal87, Wil86]. Effects [KSJ86, Rei84, SL82b, AM87, AH88, Ben89, Can84, CM86a, FY86, Gho87b, Gie83, Goo80, Gun83, HH86b, HJK89, IS87, JH88, JKR89, KS81a, Kun84, LR87, ML88, Muk87b, Mur89, RS84a, Rat89, SR88c, SB89a, Sha86, Shi85, aS89b, SL80b, TT88, TTB88]. Efficiencies [Rub86b, GS85a, HJ83, Muk86, Tlk83, WW84]. Efficiency [DSS81, FY86, Fer87, Gup86, KK80, KS87, Pra89a, Pra89b, Von83, YAH89, AL82a, CDS84, ES87, FP84, Gov86, JC84, JP87, Kag80, Kag81, KSR88, Kar88, Koo85, MP82d, PS86, PK85b, SS85a, SG87]. efficiency-balanced [Kag80, Kag81, KSR88, MP82d, PK85b]. Efficiency-consistency [Gup86]. Efficient [AM88, BJ88, JB86, KDK85, mN83b, Gie83, GC80, KK84, LK83, SL88]. Eggenberger [JR82, Wal84b]. eigenvalue [Tan89]. eigenvalues [Dun86, Nag85]. Einstein [WYF87]. Elemental [Far88]. elements [Par81]. elicitation [CD87]. eliminating [HP82]. elimination [HM81, RS80]. Ellipsoidal [Wie85, SL89b]. elliptical [CD89]. elliptically [KL86]. embedding [Loh84a]. Empiric [WH87]. Empirical [Alb87, Chu82, Gho81a, JC87, Mar89a, PR86, TZ85, TL89a, TT86, VS81, AM86, CB81, Dor88, FB82, GgT86, Gho85, GLT89, Gut82, HS86, Kar89a, Kar89b, Koz83, LP88, Lem80, Li82, Li87, SS88c, TJZ88, TL89b, Yu86, ZTJ86, dWM85]. empty [MP85, PM85]. emulation [Fic84]. endogenous [SA85]. engineering [Tag85]. enough [Nar81]. entropy [DGG89, Kap89, MR86, Nay85, SA81, Urz88]. Environmental [Pra83]. enzyme [CP82]. epidemic [ML88]. epidemiologic [KKC82, RMW86]. epidemiology [Cor82, Kum81]. Equal [Jen83, SB82b, BG85, DF89, PO84b, SKN84]. Equality [SS88d, BDK89, BW82, Cha88a, FY86, GN87, HK87, KT84, KSC81, KN83b, McC87, NCLN82, Nay88, O.’89, Oht87a, SJ89a, Sin85, Tik82, TB84b, TB85, TGB89, VI89]. equally [GQ84, MS82]. equation [BC88, Krä85, Lee87b, Lee88, Li88, Phi80, ST86b, TP85, Vog89]. equations [Li88, Llo87a, Tal86, Tsu85, Xie89]. equicorrelated [TYM86]. equicorrelation [MYT87]. equiprobability [DW89b]. Equivalence [Tun85, AH83, FF86, Fri87, Ger88, YSD85]. equivalent [Hen88]. Equivariant [Bol89a, HI89, JS82, Roj85, Sha83, ZS81, Zac81a]. equivocal [Din87a, Din87b, KC87a, KC87b, KC87c]. Erratum [Ano87m, Pra91]. Error [AR87, Job87, LT82, Wan89a, Ang83a, BM87b, CM82, ES87, Fic84, Gal82, GD85, Gel87, IS87, JK84, KM84a, Kem86b, Lee87c, LY89, LWA89, LP87d, Lis88, MC86, Men88, MS86b, MA80, Mit81, PH84, PRPR88, Pri82, SJ89b, Ste89a, Ste89b, SJ88, SA86, Swa88, Tre80, öT85, Val86, Vos88, WH88]. Errors [AK86b, Buo89, KW84, KJ82, CC82, Cha83b, Fei82, FP86, GM85a, HG80, Hig82, HT83a, Iac83, JKW89, Ket82, KW83, Kon81, KSJ86, KJG88, Mag89, OM87, OP87, PO84b, PO87, SBC81, SNR84, ST86b, SKN84]. Errors-in-variables [Buo89, CC82, Ket82, KW83]. especially [GS85b]. Esséen [CR89]. essentially [Abe80]. Estimability [MS85, Tun85]. estimable [Yam86, Yam87]. estimate [AM86, Bar88b, Hey84, Hoc81, MAT82]. estimated [CO89, GR81, GRY81, KKOV87, Kul87a, Mil82b, SG84a, TSM86]. Estimates [Kha87b, NT87, tFAH84, BT82, Bru80, CH83, EP85, Gal82, Gal80, Gie83, Jou88, Kaf82, MR88, MG87, Mic85, SD87a, Sha88a, SGP83, SG84b, Spa85, Spa89, Str84, Tiw88, Tut89, VW83, WS84, Wan87a, Wit87b, Yam86, Yam87]. Estimating [AUSH83, Bed88, Blu87, Blu88, BR82, ĆM89, DMH81, GL89, GM83b, Hai88, Haw89a, KM80, Kub89, MD89, Rad84, RHB87, Sag83, SL89a, VSD88, Wal80, WB80, BHW81, Cha89a, Gil89, HP87, Han80, HI89, HT83a, Hsi81, JP81a, Kag88, Ket82, Lee85, Llo87a, Oht87a, Rao82, SS81a, Sim84, Swa80, Swa87, Swa88, TG86, dWM85]. Estimation [Alb88, Amo84, Bas89, Beg80, BB83, Bha81, BS80a, BP83, BSC80, Cau83, Cau84, CC88, DB86, DS81, Dix89, DC84a, DH89, GD85, Gel83, GG81, GR87b, Gle82, Gup87b, GG88, HA83, HD89b, HD89a, Isa81, Jel87, JP86, JN87, Jos88, Kha83b, KG87, KV80, KD82, KLV85, Lee81, Lee88, LWA89, MJ82, Mee88, MDH81, MC87b, NT87, PS81, PNA85, Pau88, PH89, PR88, Rau82, Rob81, SR88a, Sam84, SD83, SD84a, SM86a, SJ87a, SOL81, Sin88, STSM89, SW83a, SM82a, SA86, VS88, Ven88, Wan09, Wei80, Wil88a, Win87, AC82, AM84a, AM88, AM87, Alb81, Alb87, AH86a, AH86b, Ang83a, AGR89, Arc82, BJ88, BM87a, Beg81b, Bel88, Ber89, BC88, Bon88, Boo82, BM88, BYD85, BG88b, CG88, CW88, CW80, CQ82, CL89, CC85, CS84b, CF88, CF89b]. estimation [Coo87, Dah86, Dan88, DGG89, Des84, Dey83a, DS86a, Dey89, DLP85, Die84a, Dou86, Ebr84, Ebr85, EH87, Elp83, EK82, EL82b, EkBL88, EB89, Far85, Fei84, FH84, Fra83, FR88, Fri81, Fro89, GgT86, Gat86, GK88a, GD88, GA89, GK88b, Gho81a, Gho89b, Gho81c, Gho85, GN86a, GM88, GLT89, GV89a, GM85a, GM85b, GdT87, GR89, GH84, Gup82, Gut82, Haf86, HGB89, HN86, HW88, Has80, Hen88, Hil85, Hol89b, How85, IB89, IS87, Iso89, Iwa86a, JPD82, JB86, JP88c, KH81, Kai83, Kal81, Kan82, Kap85, Kap86, Kap89, KW84, Kar89a, Kem86b, Kep86, KW83, KS81a, KB80, KP81, KDK85, KF89, KP87, Kou81a, Krä85, KR88, KRB89, Kul87b, KC80, KS88, LaR84, LP88, LSS85a, LM84a, LP87b, LK81, LM86, LR82, LT80, Loh84b]. estimation [LP85, MP82a, MS82, Mar89a, Mar88a, MRS84, Mat85, MN89, MP82c, MS86b, MM89a, MS80b, Moo89, MD87, Muk88, MS81d, PM84, Pal88, PS80, PL88, PD81, PG84, PG89, PB85, Pep89, Pon83, PH85, PR86, Rei84, RU89, Rob89, SL88, SHB85, SS86b, SS86a, Sam88, San89b, Sch86a, SS85c, Sel89, Sen80, Sen86, SB89a, SS89, Sha83, SS88b, SI81, Shi83b, SI87, SB84b, Sho82b, SW84a, SJM86, Sic86, SG84a, Sie83, Sin80b, Sin86, SU86, Smi85, Sob89b, ST86a, Spr82, SL87, SL89b, SM86b, Ste89b, Suz83, Tea84b, Tik88a, Tit83, TZ85, TJZ88, TCJ89, TL89a, TT89a, Val86, VS81, VS89, WEA80, WH87, Wan89a, Wei81b, WAS88, Wil81, WS82, Wil88b, Win82, Wri83, Yao86, ZTJ86, Zeh85, dR88]. estimations [Cha83b, FG81]. Estimator [Fen83, Ama88, Bed87, BWD86, BU84, Bre82, Cha83a, CR89, CDS84, Cha88b, Che88, Che87b, Dah87, Dav85b, Dey83b, DBO83, ES87, Far84, Feu87, Fir87, GM83a, Geo86, HI89, Hos88, Isa86, Iwa86b, JS82, KL82a, KK80, KC84b, Kim86, Kon81, KPS83, Kor85, Kub88, KSC81, KS89, LK83, Lai89, LT82, LP87c, Mit81, MC85, Nyq83, Nyq86, Oht86, PW81, PU86, PS85a, PNS89, Per87, Phi80, Pli87, Rad82, RR87, Roj85, Roj86, SRW84, SA89, SA85, Sir84, SW84b, SDCB83, SSK89, Ste89c, Tre88, Wan86, Wan88, YSD85, Yum85, dPV87]. Estimators [Joa83, Kea83, Kun83, Sin83, Tre83, Abd89, Ali89b, AM84b, BK85a, BHL82, Ber81b, Blu85, BCK85, BR82, Cha82b, CB82, Che83a, CW82, CL83, DD85b, DM80, Dor89, DS89, DK81, ESP84, Far88, Fra83, Fra89, FB89, Gad83, Ger88, GR82, Gru85, HY86, HJ83, HB86, Hol87, Hos87b, HW89a, HW89b, ISS82, IS86, JV87, JK84, JC84, JP87, Kap82a, KM84a, KT86, KY81, KB84, Koo85, KB82b, Kum80, LaR84, LS83a, Lac84, LP87a, Lee87c, Lem81, LlHK82, Li82, Li87, LO89, LP87d, Lis88, Loh84a, Mag87, Mag89, MR83, MC84, NM89, Off86, OSH88, Pad80, PM88, PK86, Pra89c, Pra91, Pri82, RS84a, Rao87, RHB87, Rob89, SS85b, SS87a, SS86c, Sau82, SW88, Sch85c, SH82, SI85, SG87, Sha88b, Shi83a, aS89b, Sho84, Sil86, Sil88]. estimators [SIU85, Spa86, SKG80, SC89, SD87c, SN89b, Swa83, Swa86b, SM82b, Ter81, Ter82, Ter86, Ter88, TB82, TB84a, TB88c, TG89, TB89, Tlk83, Tre80, TT84, öT85, UV84, Ver88b, Via82, WS84, gW82, WK81, WE80, Yu86, ZS81, Zac81a]. ethical [Lai84]. Euclidean [TT84]. Evaluating [BHL82, GR83, SH89, DBO83, WL82]. Evaluation [SQ80, Bel88, LL82, Nyq86, WS84]. even [PL88]. event [RA80]. events [BK84, LP81, RG89]. evidence [KW83]. Exact [Bon88, CT83, CB87, CB88, Eri83, Fir87, Hei88, IV89, JKK86, Mal80, Mar84d, Nel82, Per82, Pou83, Rus88, Spu83b, SS86d, Tlk83, VI89, Ver88a, Asr85, DW87a, DW89a, Ima87, Joe88, Lee87c, Lee88, Li88, MP86, Nic89, Ped86, Som85, Tri88]. Examination [Smi84, Far83, TSM86]. examining [MY85]. example [BM81b, Poc82]. Examples [Roj85, CI80]. excess [Spr89]. Exchangeability [Tea84a, PS85b, Soo88]. exchangeable [Bal87a, Jos88, Tea84b]. executive [GH84]. existence [GC84, Stu88, SM82b, YB83]. existing [Ber81a]. Expanded [Bof85, Bri87]. expansion [DA80, GN88]. expansions [Blu85, KN83c, NNQ84, Tah89, TW81]. expectation [dSBPR84, EZ89, Gun86, Yu85]. expectations [Gro82]. expected [Arn88, Fin82, Nis88, Nis89, SQ80, Zac81b]. Experience [Ste86b]. experiment [Bar83, BS80c, Gab80, MSVS80, Smi80]. Experimental [Bal87d, Box85, Czi89b, GS85a, Czi88, LL82, McC80, PL82, Tri88]. experimentation [Kle86, Sri87a]. Experiments [PL83, BD87, Bed89a, Bed89b, Ber82, BV89, Che83b, Cru82, Czi88, EHA84, Fre83, Gho87b, GC80, GBS83, Kee80, KS87, Lin86, Lin87a, Lin87b, Lor84, MS87, MI82, MM83, MM89b, Ney80b, Ng89a, Odh87, PNA85, PK81, Pre89, SF85, SL80b, Smi82, Vos86, Wah81]. explicitly [Spa86]. exploration [Dav81]. explosive [Kul87b]. exponent [GM85a, GM85b]. Exponential [Beg81a, PS85a, AWS87, Ahs87, AO82, Ami89, And88, BM86a, Bal87c, Bal89, BL84, Cha83b, CB87, CB88, CV89a, CK83b, CK87, Cho88, CV89b, CHS86, Dey83b, DK81, EH87, EkBL88, FV89, Fro89, Gat86, Gok89, GMM84, GG88, Ham89, HK87, Haw89a, JP86, Jos88, Kar89b, Kha83c, Kim86, Kim88b, KN83a, KN83b, Lai85, Lia88, Loh84a, Mat83, MDH81, MC87b, Muk88, NN82a, NCLN82, Nay88, PS80, Pap89, Pep88, PH85, Raf84, Rau82, Sam84, Sam85, Sam88, SS81a, Sat88, Sen80, SJ87a, SJ89a, Sin85, SL82a, SN89b, SvW82, Tra85, VS81, Vog86, Wei81b, Zac81a]. exponentiality [Ang89c, CDS82, Koc85, Lin89, Pat88, SK86, Spu84]. Exponentially [Sin83, GHO86]. exponentials [GM88]. exposures [Rao86]. expression [Pat85]. Extended [DW87a, DW89a, TT85, MB82]. Extending [MC85, LS81]. Extension [SlR87, Ste89c, BE82, Che83b, Som84, SG86a, YB82]. Extensions [EN80a, Nai82]. extinction [EN80a]. extraction [Dav87]. Extrema [tL84, SG85a]. Extreme [Joe88, Ang89a, CW87, Dav88b, FP89a, FKML89, Gal88, HY86, SYT89, WJT89]. extreme-value [HY86]. Extremes [FKML89].

faces [TT83]. factor [BM89, BD87, Bed89a, Bed89b, Gho87b, Mor87, PH85, aS89b, Shi82c, Smi84, SN82, TO89, WG81]. Factorial [PL83, Ban80a, BV89, EHA84, Fed87, Gup87c, Gup88a, HK89, JRF80, KS87, Kuw83, KM84c, Kuw89, Lin86, Lin87a, Lin87b, Muk86, MB89, StR82, SRP82, Shi82c, SL80b, Smi82, Vos86, Vos88]. factorials [Cha89c, GC84, MCN81]. factorization [FS86, Pou83]. factors [AG89a, BM89, Fed87, Güi82, Kle87a, Smi84]. Failure [MR87, AOA89, BK86, BDK89, BS89, Bor85, CEA85, Dav87, DS86b, DK81, GM84, KB81, KB82a, Koc85, Lin89, PW81, PU86, Pap89, SJM86, SW83a]. failures [GR87c, San87]. familiar [Dav85a]. families [Ami89, CV89b, Dey83b, EkBL88, Fou81, FKML89, Gat89, Kar89b, KR88, MC87b, Sel89, SL84]. family [Ang89a, Arb85, Beg81a, BG84, FKML88, Haw89a, Joh82, JB84, Kha83c, KL84, Koo85, Nel86, Rao81a, SR88a, Sam88, SG84b, VS81, Woo88, Wri88a, YI88]. fatalities [LM84a]. fatigue [CLMFC87, KC84b]. fault [GL89]. faulty [HJK89, JKR89, KJ83, JKS80]. favorable [Che86b, MS80a, Seh88]. Feature [YO86]. Feature-subset [YO86]. fi.S [Sea80]. Fibonacci [TP86]. fibres [FR88]. fiducial [LT85]. field [PNA85]. Fifth [NP88]. filter [Coo87, SOL81, SW84b]. filtering [Liu89, Rus85]. filters [Fin84]. fingerprint [Scl80]. Finite [SDCB83, AM84a, AR83, BD86a, Bol87, Bol89a, Bol89b, Cha84, Fou81, Gho81c, Har83, Hun89, JL89, KM87, MC87b, Off86, Par81, PS85a, PS83, Pra89c, Pra91, RR87, SS88d, Ste89c, TL89a, Val86, YSTM88, ZS81, Zac81a]. firms [GH84]. first [And86, Bha87, EK85, Gab80, GN86b, Has80, Iwa86a, KM87, Mar89b, Sat88, Shi83b]. first-order [Has80, KM87, Mar89b]. Fisher [AG89b, BR86, CG81, Joe88, MP86, NvdM86, PS86, RF84b, SD86, Woo88]. Fisherian [MS88]. Fit [Gim87, AAY86, BA87, BT83, Bar86, Bro82, Cal87, EP85, Fre84, FGP88, Hor85a, Hos87a, HL80b, Kan82, KK86b, Koz89, Mar84c, McC85, Meh82, NJ84, NJ89, Pie83, RB88, Rig89, Sch85b, SK87a, Tik88b, Utt82, VJ88, Wei89, WJT89, Wil86]. Fitting [CM86b, Bra81, Joh82, SL89a]. five [Gap84]. fix [BHW81]. Fixed [Ala87, SvW82, Ang86, Bis80, Boh83, CHS86, Gie83, Goo80, JH88, KS81a, Mur89, Pei86b, RH88, SR88c, VB80]. fixed-effects [JH88]. fixed-width [Bis80]. flexible [LDH84]. Florida [Gap84]. fold [BG85, MM89a]. folded [LK81]. following [AGR89, San89b, Wan88]. follows [Ebr86]. forced [NR83]. Fordham [Vin84]. Forecast [AL82a]. Forecasting [WA81, Gue87, LM84b, ML83, MB85, Pou83, WH88, ZF88]. forecasts [SB84a]. form [ES84, KA80, LK83, NM88, SM82b, Won89]. formal [Geo86]. forms [BS88, Cha82a, CT83, HGO85, tL84, Lis84, Mac82, MW85, SA82, SB89b, TC81, VJ88]. formula [DN80, KY82, Mal80, SS88c]. formulas [RM88]. Formulation [CS87a, CS87b]. forward [PKG87]. four [Bha87, Czi89b, Kos87, PVB86, SD86]. four-moment [SD86]. fourth [KP86, KP87, Kos87, Muk87a, PK85a]. FPE [Che88]. fraction [BCK85, LS80, Lac84]. fractional [JRF80, Kuw83, KM84c, Kuw89, MCN81, SRP82, Shi82c]. fractions [Lun82, StR82]. free [BDK89, GL87, HG86, KI88, Lee87b, OW80, RD88, Sch83, Smi82, Spa88b, SD85, Stu88, VI89, Whi89b, YB83, dR89]. freedom [Som84, SN82, Vos88]. freezing [KY82]. frequencies [Bar83, DH89, JK84, WJT89]. Frequency [Lev83, Hur87, Sic86]. frequently [KC84b]. Friedman [ID80, Shi85]. friendly [Ber85, Dou86]. fulfilment [JKW89]. full [GR82, Khu80, MS80b, MS81b]. Fuller [WS84]. Function [Ber84, Fen83, NT87, SKSF83, Sin83, And83, Arb85, BK85b, BE84, Bed88, BCL80, BM87b, Gho81a, Gho89a, Gho89b, GLT89, GM85b, Gup80, Hil85, Hol89a, Hol89b, HW87a, Iwa86a, JP83b, Kan84, KM84a, KR89a, KT85, KD82, Koz83, KLV85, KP84, LaR84, LaR88, LP88, Leu89, LP84, LM86, LH82, Llo87b, Mal80, MT82, MR88, MP85, Mit88, MS81d, Ojo88, Rad85, Roi87, Roj85, SAN89a, Sam88, SND84, SI87, Sie83, SK87a, Spa85, SS88c, Ste89b, SH89, Swa86b, Swa88, TJZ88, Win82]. Functional [Tal86, Abd89, CC82, DFG87b, MR86, NV84, Pat81, Pro86, Yum85]. functionals [TZ85]. Functions [Joa83, AA81, Ang83a, Bol89b, BS80b, Bru80, BP83, Che83a, Ebr85, Gle82, Hel88b, Kam88, Kan88a, KV80, KDK85, Kot89b, KD83, MR83, NS85, Nay85, PRPR88, RK81, Rad85, SGP83, SW89, SM87, Tan89, Xek83a, dR88]. funtionals [Gho85]. Further [Din87b, GgT86, KW83, KN83c, MP82b, MT84, Rus80, Far87, FS80, KC87c, Krä85, LP89, Nor89]. future [Cho88, JC87]. fuzzy [HB86].

g [SD84a]. G. [KC87b, KC87c]. gain [Coo87, Gil89, KR82]. game [Pur86]. Gamma [SlR87, AAY86, BE80, Bar86, Bar81, Bel88, Ber81b, DB86, Dix89, Gho81a, HD89b, IS85, LT85, Pep89, Rob81, Tub86, VS88, VS89]. gammaization [Cha82a]. gap [KA87a]. Gauss [Cha82b, Yas81b]. Gaussian [Amo84, AK86b, BM88, GN86b, HI89, How85, Iwa89, JC84, Koc86, Mar84d, Mar89b, NNQ84, OW86, PM88, RC85, Rus85, SI85, Sha81, SJ88]. Geary [SG82]. Gehan [GM84]. gene [DH89]. gene/haplotype [DH89]. General [Dav88a, Fra83, SS88c, AH86a, AH86b, BM85a, Buo85, Cha89c, Cha88b, CEA85, CA81, Hel88a, HZ82, Kal81, Kel86, Lam86, LT84, Mit81, Mos83a, Mos87, PL86, RB83a, Rie82, Rob89, SD84b, Smi85, Spu82, Sri87a, Sun88, SB89b, Tse89, WL82, Won89]. generalisation [EK85]. Generalization [TB88b, GS85b, Ier85, Kha89a, ND81, Neu89, Pau85, RR88b, Rau87, Sho82c, Sin87, TG85]. Generalizations [PC82, Sen87, CS88b, Die89, Vos86, vdMN87]. Generalized [Bel88, DR86, GE88, HG83, Kha87b, LlHK82, Mos87, Tre80, Vos88, Ama88, AH88, Ber86, BE88, Cau84, Cha84, Cha86a, CS84b, CF88, Con88, Con89, CF89b, CL83, DL86, EK82, FKML88, GWBG87, GS83, Gia86, Gie83, GS85b, Gro86, Gup87c, HG80, HZ82, Hos88, HF88, JR82, KL82a, KP89a, KPS83, LW85, LT85, LP87c, LP89, MP89, McD87, MSL86, Nor82, Oht86, OL86, Pal88, PX86, PT86, Ped83, Pep89, Pri82, Rei84, Rus88, RF84b, Sch84a, Sha89, SI87, SG84b, SKG80, Sri89, Ste89a, TG86, Wan87a, Wil88a, Wit84, Xek86]. generate [Hal84]. generated [EkBL88, SL84]. Generating [Gup87c, Pan83, And83, El 81, Kem86a, MR88, MP85]. generator [WT83]. genetic [DH89]. genotype [Kha87a]. Geometric [DS87, Con85, DF89, Har83, Iwa86b, KA80, RR88b, Yan88, dWM85]. geometrical [Sam85]. geometry [Loh85, MS80b, Old87]. geosciences [Mar81a]. geostatistics [Mar81b]. Gin [SY87]. GINI [Moo89, CG88, LSB89, Rav89]. given [ER83, Smi83]. Glejser [HL80a]. GLMS [PP88b]. Globally [StR82]. GLS [DM80]. GMANOVA [CE85, Kab86, Mur89]. goal [Gia86]. goals [AD80]. Goldfeld [Hor81]. Gompertz [Pra83]. good [Bjø85, GK81, HH82, Lia88, Pli87, Tre88]. Goodness [Gim87, Hor85a, HL80b, KK86b, Mar84c, AAY86, BA87, BT83, Bro82, Cal87, Fre84, FGP88, Hos87a, Kan82, Koz89, McC85, Meh82, RB88, Rig89, Sch85b, SK87a, Tik88b, WJT89]. Goodness-Of-Fit [Gim87, Mar84c, BA87, Kan82, Koz89, Rig89, WJT89]. GPSD [Gup87a]. grading [KJ82, KJ83]. Graff [KJG88]. graph [Gal88]. graphical [Coh80, LP81]. graphs [Czi89c]. Graybill [El 82, Kor85]. greenwood [Bre82]. Gregor [Pie83]. grid [KM88]. GRLS [Yum85]. Group [BM87b, PST87, Pet82, FV87, Gup88c, HES84, Hoc81, HJK89, JKR89, Kle87b, LC87, LWA89, MB84a, MS87, OP86, Odh86, Odh87, OM87, OP87, OP84, PO84a, PO84b, PO87, Wat87, WR87, ZF87]. group-screening [Kle87b, OP84, PO84a, PO84b, PO87]. Group-testing [PST87]. Grouped [Wan09, Arc82, Bar83, Blu87, Blu88, CC88, Pad80, Tit83, Wu88]. grouping [Mor87]. groups [Kul85, MT83a, Nel87, Sri87c]. Growth [WL80, Fu80, KS81a, KD83, Kub89, Kul87b, LC87, OS86, PB85]. guidelines [DeM84, Wha87]. Gumbel [EZ89].

Hadamard [SFR84]. Haenszel [Per85]. Half [SS89]. Half-sample [SS89]. halving [PST87]. Handling [Din87a, Din87b, KC87a, KC87b, KC87c]. haplotype [DH89]. happens [Gu87]. Harmonic [TT88, TTB88]. Harrell [Ste89c, YSD85]. Hartley [Sen87]. hash [Ram87]. Hauck [Fri87]. having [Bed88, Den89, Gil82, Kul85]. Hazard [Xek83a, AB89, BM86a, Che87b, Ebr86, GR87b, Gho89a, Gho89b, Iwa86a, Kay82, MS81d, Tak88, Yao86, YAH89]. hazards [CB85, CEA85, GV89a]. health [Hut84, Ney80c, RG89]. heaviness [CS81]. heavy [SYT89]. heavy-tailed [SYT89]. Heckman [Lai89]. Hellinger [GS85b]. helping [Nar81]. Hermite [Pat85]. heterogeneity [HF80, RS80]. heterogeneous [Ama88, KB83, SJ89b, Ter86]. heteroscedastic [Dor88, Nel87, TT88]. Heteroscedasticity [Hor81, CL89, HL80a, HN84, KO84, Lis84, McC86, Sar85, Tan82, TTB88, TP85]. heteroskedasticity [SA89]. hierarchal [KJ82]. hierarchical [Bas89, KJG88, Pap89, Pep89]. high [Czi89c, JP87, Zei89]. Higher [EK85, KA87a, Sri87a, ZF88]. higher-order [ZF88]. hiked [Kle88]. histogram [Kan88b]. histograms [Dal88, SKN84]. History [Bar88c]. histospline [Wah81]. hitting [GL89]. HNBUE [SK86]. hoc [SF85]. Hodges [Pad80]. Hoerl [Pag81]. Homogeneity [Tar88, CM82, Del81, GL81, KI88, Ken81, KS84, Kul85, LSS85b, Man89, Mar80, MT83a, MC87a, PYO82, RS89, Sen87, Sha84, Shi88, aS82, TO85, Tom87]. Homogeneous [GP86, Tre83, KR89a, Ter82]. Hotelling [Bet87, SA82, Sri83, TB88c]. Hotelling-type [TB88c]. household [GG81]. Huber [Nyq86, Sil88]. Huberizing [BCL80]. Huddleston [WS84]. hull [MT83b]. hybrid [CB87, CB88, MP86]. hydrology [Llo81]. hypercubic [Cha89b]. hypergeometric [WR87]. hypersimplex [AT81]. hypersphere [AT81]. hypertension [eL82a]. Hypotheses [HSC80, Aic83, CEA85, El 81, Fla85, Fou81, Goo80, Hel88b, HS83, Kle88, Mos89, Pei86b, Sea80, SOL81, Spu82, Vad83, dR89]. Hypothesis [DG83, LC87, NV84, RLN86, BSC80, FH81, Hel88a, Jel87, KD89, Lam86, MS80b, Oht88, Owe85b, PH89, Pep88, PD84, Sak87, SHB85, SC80, SR88c, WCK85]. hypothetical [Sch85b].

i.i.d [Ete84, Wef87]. Ian [SB82b]. Identifiability [BG80, BG83, Tun85]. Identification [KT85, LH82, Liu89, SN89a, BR85, GM80b, JP88a, mN83b]. identifying [WR87]. Identities [BMA88, BB89a, Nis88, BM85a, Bal87b, Bal88, MBA88]. if [Li88]. ignorance [MC85]. II [BMA88, DK89, GM85b, GdT87, Har86, KRB89, KM87, LaR88, Meh82, MP82d, PM89a, SR88a, Sel89, Sin85, Spu83a, TG89, Wei82, YS82a]. iid [GHO86]. III [MCN81, Pad80, SA82, SH89]. ill [GdT87]. ill-conditioned [GdT87]. illustration [BP84, KJS^{+}83]. imbalance [DK89]. immigration [JB86, PM82]. imperfect [Tra83a]. implementation [SSS^{+}85]. implications [Win82]. implicit [Spa85]. implicitly [Spa86]. importance [Gho83b]. imprecisely [Spa89]. Improper [BU84]. Improved [AAY86, Bis80, BM88, GD88, Hos86, Hos87a, MR83, SN89b, Wan87c, Kap89, PM88, Pra89c, Pra91, SG82, SC89]. Improvement [UV84, Box85, KPS83]. improvements [Kar89b]. improving [Gov86, Net85, TT85]. imputation [KK84]. Imputed [Gim87]. IMSE [Bar88b]. inaccuracy [PT86]. inadequacy [Ali89a, Sri87a]. Inadmissibility [Rad82, PNS89]. incidence [mN83b]. included [SA85]. Incomplete [IS86, TGG86, AK86a, BK85c, BWD86, DS86b, Ebr85, Fei88, Gup88b, Ham89, KD82, Kot89b, Lee87b, LC87, Mar84d, Mar89b, Ojo88, PK81, SM86a, SD87b, WT89, WT83, WS82, WL80, Zeh85]. incomplete-data [LC87]. incompletely [SG84c]. Incorporating [Lai84, Sch85c, Arb85, DC84b, Ste86b]. incorrect [PP87]. increasing [AOA89, BM86a, PW81]. incurred [Woo80]. independence [Ahm82, ARS86, BL84, BM81b, FR83, Hos86, Joe88, KR86, KK85, KN84, NN82b, PH89, Pro88, SLH83, SC86, WS89, Wu82]. Independent [Ete87, Bar81, Bur81, DR82, GN86a, HH86a, HLN88, KSC81, LM86, Mat83, MS86b, MM88, Nay88, Sat88, Spu83a, WCK85, ZG80]. indeterminate [Gio88]. index [Ano83a, Ano83z, Ban80a, CG88, Che88, FV89, GS85b, Moo89, Vog86, Vog89]. indexed [SM85, SV89, SM89]. indices [CO89, Kha88a, Kha88b, Vog87]. indifference [Gia86]. indistinguishable [IM88]. individual [FBV84, NS83]. induced [Sen81]. industrial [Dav85a]. industry [Lai84]. Inefficiency [Slu86]. Inequalities [Smi83, Dav88a, Ete84, Kot89b, Nis88, RLN86, SL80a, YSTM88]. Inequality [Kha88b, And86, Bha87, ES84, Far86, Far87, PP88b, SS88d]. Inference [BS86, DP83, HRR88, MH88, Roc82, SB85, ARP86, BF85, CY86, Cho83b, DG81, DFG87a, DFG87b, GR87a, Gho83a, Gra87, Hau89, HK83, Hun89, Owe85b, RB83a, RD88, RC85, Sin80a, Smo83, Som89a, ST86a, Tra85, Umb81, VS86, WT89]. Inferences [WS87, ER83, Hog82, HP80, KY82, SB84a, Slu86, TT85]. inferential [Kap83]. inferior [HP82]. inferring [dlHRR88]. infinite [EP85]. inflection [SM85, SM89]. Influence [RK81, Rad85, Tan89, CPW88, DD85b, FCP^{+}82, Fra83, Gra89, Sim87, Tan88, Wan87b]. Influential [Gho83a, Gho87a, Pen85a, Pen85b, Wal89]. Information [Gho82, Mar89b, Sam89, Dah88, Dav87, DC84b, EA89, EC83, GP86, Hab89, Lee88, Loh85, NDM84, PP81, Pre80, Rao82, Sin82, STJ88, SCG85, SG86a, Ste86b, Tsu85, WK86, Whi89a, ZFP89]. information-theoretic [DC84b]. informative [CV89b, Nig89]. initial [Fla85, SOL81]. inner [Kan88a]. input [LH82, Ven80]. input-output [Ven80]. inspecting [Min80]. Inspection [KJG88, SNR84, Abe80, HJK89, JKS80, JK84, JKR89, KJ82, KJ83, KSJ86, Tra83a]. instrumental [RG88]. Integer [Dah86, Zac81b]. Integer-parameter [Dah86]. integral [eL82a]. integrals [Yas81a]. Integrated [CS87a, DS87, HG80, Swa88]. integration [FV89, Vog89]. intensity [BS80b]. intention [CF86]. inter [WK86]. inter-column [WK86]. inter-row [WK86]. Interaction [DG83, BV89, Che87a, Chu82, Chu83, CGO85, KR89b, KJS^{+}83, MJ82, Mar85, Muk86, Sch86b]. interactions [BM89, Shi82c]. interblock [PP81]. intercept [JP88b, RP88]. intercepts [LSS85a, LSS85b]. Interchangeability [ST83]. interclass [Shi82a]. interdependent [BHW81]. internal [Gra89]. International [You85]. Interpenetrating [Bed87, SIU85]. Interpretation [PVB86]. interquantile [SD87c]. interrater [Fla87]. interrelated [PH88]. Interruptible [RS84b]. intersection [BS83, BS84, NS84]. Interval [KP81, LK81, MM89a, AS84b, AGR89, Asr85, BG88b, CL89, ES87, FW86, JH88, Kai83, Mee88, Mil82b, RH88, SI85]. Intervals [WM87, Ami89, AJG82, Bis80, Bof85, BG85, BMC86, CAR89, CO86, Cho88, CHS86, Fla87, FGV89, JP88b, Kab83, KB83, Kon81, Kor87, LMP88, Mar82, MM88, Muk81, Pat81, Pat89, PS83, PRPR88, RMW86, Sch84a, Shi87, Sob87, Sob89a, SKN84, SD85, SvW82, Ver88a, Wal87, WG81, Wei81c, Wei82, Wit89a, YS82b, YS82a]. intra [PP81, SM86a]. intra- [PP81]. Intraclass [Mic83, SL82b, BMC86, GN87, Kha83a, NM88, Shi82b, SW84a, SL80b, ZG80]. introduction [Koc86, Llo81, STSM89, Vin84]. invariance [Kab84, OM80, Sen81]. Invariant [YM86, CS83, CV89b, Rob81, Sob87, Tiw88, Yam86, Yam87]. inventory [EN80a, Kat80, Kot89a, PR84]. Inverse [NT87, Amo84, AK86b, BS80a, BM88, HI89, How85, Iwa89, JC84, Kan84, Kik83, Koc86, KPS83, Nor89, OW86, PM88, Per87, PTA89, Rao81b, SI85, Sha81, SJ88, Tah89, XP89, Yan89a]. inverses [AP86b, LP89]. investigations [Fed87]. Involving [Lev83, Pan83, BM85a, Bal87b, GCH83, RMW86, SG88b, Yat86, YO86]. isn’t [Huf85]. Isolating [Bre82]. isomorphism [Joh85]. isomorphisms [Lin87a]. isopleths [Sag83]. Isotonic [Cho83b, FW86, Lia88]. isotropic [Sch85b, Sch86c]. Israeli [Gab80]. issue [MSVS80, Ney80b, Smi80]. Item [GR87c]. iterated [DM80]. iterative [KY81]. IV [Kle87a].

J [Din87a, Din87b]. jackknife [Fra87, Nag85, Sha88b]. jackknifed [Phi80]. Jackknifing [GM83a, KKOV87, Sen86]. James [LlHK82, Li82, SS86c, Sen84, WS87]. Japan [Tag85]. Jerzy [MSVS80, Smi80]. Jewls [KKOV87]. Joint [Rao86, BM81b, FP84, Fal86, LWA80, RM88, SS85a, SHB85]. joint-tests [SS85a]. Jonckheere [Die89]. journal [OEJ^{+}85]. Jury [Sol83]. justification [Phi80].

Kageyama [Joh86a]. Kalman [Rus85, SOL81, SW84b]. Kaplan [Ama88, CR89, GM83a, Ger88]. kappa [Fla87]. Kendall [AC84, DPS81, FR83, SR88b]. Kennard [Pag81]. Kernel [Hil85, SS85c, Tit83, BT82, Joh84, Tut89]. Kernel-based [Tit83]. kernels [Swa87, Swa88]. kind [HF88, Iwa86a]. kinetics [CP82]. knockout [HH80]. knowledge [Sin86]. known [BT81a, Bha87, Ebr84, GHK85, HI89, JS82, Lee81, Sam84, SD87a, Sha80, Spa89, Tar86b]. Kodel [Din87a]. Kodell [Din87b]. Kolmogorov [Bri87, Cal87, FGP88, JK85, KH86, LR82]. Kruskal [RF84b, Shi85]. Kullback [EA89, Loh85]. Kurtosis [Bal87c, GM85b, Gro86, Koz86b, MF83].

L [Din87a, Din87b, Nyq83]. laboratory [Fed87]. lack [KHB^{+}88, NJ84, NJ89, Utt82]. lag [AM84b, Mar88a, TSM86]. lagged [CM86a, MC86, MC87a, ZF87]. Lagrangian [Con81, Kum81, Sha84]. Laguerre [GCH83]. Lake [KY82]. lambda [FKML88, Gro86, Mac82]. LANDSAT [CHLL84, WS84]. Laplace [Asr85, Pil85]. Large [CY86, Cha88a, Iso84, Kou81a, Kul87b, Kul87a, LY89, Mic85, RC85, CVV82, CG87, Cha83a, Coh82, Gil89, GdT87, GH84, Loy89, PS87]. Large-sample [Kou81a, LY89]. Larger [Ber84, BE82, CF88]. largest [AR83, BK84, Bis80, Cha81, CL89, GHO86, GS80, Hay89, KP81, PF84, Yan89b]. last [MM89a]. last-stage [MM89a]. Latent [Mas86, BHW81, Gio88]. latin [Joh85]. lattice [Pro84]. law [Rig89]. Lawless [Fir87]. Lawley [Bet87]. laws [Kou81a, Kou81b, KB82b]. layout [DW87a, DW89a, FY86, GL87, aS82]. leading [Con88]. Least [BC88, Bof88, Die84a, PM89b, Ver88b, BHW81, BU84, Bru80, CM86b, Che86b, CH83, Dre85, EK82, ES84, ES87, Far84, Fri81, GgT86, Gie83, Gru85, KW84, KKOV87, Krä85, LSS85a, MS80a, MC84, PS87, Phi80, SS86a, SS87a, San87, Seh88, SB84b]. least-failures [San87]. least-squares [CH83, Fri81]. leaving [LWA89]. leaving-one-out [LWA89]. Lebesgue [Kar89b]. Lebesgue-exponential [Kar89b]. Lehmann [Gov86, Pad80]. Leibler [EA89, Loh85]. lemma [Rao81b]. length [AS84b, Ang86, Cna85, FR88, TG87]. length-biased [FR88, TG87]. less [GR82, Khu80]. Letter [Ney80a, Zer81]. level [Fri87, RS84b, Shi88, SQ80, Sob89a]. levels [AAST88, BF81, KS87, Som85, SS88c, Tah89, VB80]. leverage [Dor89]. Levinson [BR85]. Life [Sin83, BS89, Bre82, CW80, DK81, Ebr85, GR87b, Gup80, GG83, JP83b, JP83a, Kha84, KB81, KB82a, NS83, SK87a, SG88b, SL82a, Spu83a, WNK80, Xek83a]. life-testing [SG88b]. lifetime [Bro87, CLMFC87, EN80b, NH87, Nig89]. lifetimes [Fro87, PS85a, Sri89]. like [HZ82, SM82b]. Likelihood [BD86a, El 81, Pro88, Sak87, SJ89a, tFAH84, Arc82, Ber81b, Blu85, BR86, BA84, CB85, CM86b, Cho80, CC85, CW82, CS84b, CF88, CPW88, Dah87, Des84, Dou86, DS89, Dun82a, Ebr84, EB89, Feu87, Fin82, FS86, GK88a, GM88, Gro84, GN87, GN88, HGB89, HN86, Has80, Hos86, Hos87a, Isa81, JC84, Kab86, Kep86, Kim86, Kon81, KR89b, LT85, LP84, Llo87a, LP85, Mag87, Mar84d, MP82c, Mic85, Min80, MS81d, Nis89, Obe84, Rad82, RV82, RR87, Sau82, SB89a, SW89, Slu86, Smi85, SS81b, Spa85, Spr82, TO89, Tik88a, TB88c, TG89, VS86, WEA80, Wan87a, Wef87, Wil88b, Win82, Yao86]. likelihood-ratio [BR86]. likely [GQ84, MS82]. Lilliefors [TYM86]. limit [Ang89b, EG88, PU86, Poc82, Wan87a, Wri88b]. limitation [BH86]. Limited [Tsu85, Roc82]. limiting [Che86a, Nag85]. limits [BK85a, FB89, Pat86, RC85, Sch87, SI85, Tiw88, ZLMK82]. Lindstrom [Som88]. line [GN83, Gup88c, LS83a, SD87b, Vog86]. Linear [BW88, CW88, DG83, Die84b, DFG87a, FH87, Iwa89, Man89, SKSF83, SJ88, Tre83, Won89, vdMN87, Abd89, AO82, AH88, AH86a, AH86b, BK85b, BM86a, BM87a, BT86, Bha86, Bon88, BS83, BR82, BYD85, BD86c, BSC80, Buo85, CVV82, CM86b, Cha82b, CP82, Con85, CVDS85, Czi89c, DK83, DFG87b, DR82, Dor88, EZ81, Elp89, ES84, EB89, Far86, Far87, Far88, Fei82, Fic84, Fin84, FP89b, Fra83, GD85, Gel87, Gho87a, Gho87b, Goo80, Gov89, GR82, Gut82, Hab85, Haf86, Haw89b, Hel88a, Hel88b, Hen88, Hig82, Hoc88, Hur87, JH88, JP88b, KG89, KH81, KB83, Kal81, Kha83a, Kha83b, Kha88b, Kle88, KY81, KD89, Lam86, LT84, Leu89, LP84, Lis88, Loy87, MP89, Mag89, ML83, Mar84b, Mat83, MRS84, Mat85, Mit81]. linear [MN83a, MB84b, MS80b, MS81b, Mos89, Mos87, Nae86, NV84, Nai89, NS85, Nor82, NM88, Nyq83, Oht87b, Oht88, PNA85, Pat81, PP88b, PNS89, Pei86b, Pep89, Per87, Pro86, RB83a, RP88, RLN86, Ray85, Rei84, SS87a, SD84b, Sch80, Sea80, SH82, SR88c, SB84b, Sie83, ST86a, Spu82, SA82, SL87, SL89b, Ste89a, Stu88, SK87b, Sza85, Tah89, TW81, TG83, Ter81, Ter82, Ter86, Tom87, TS88, öT85, VI89, Wal80, gW82, Wil88a, Yu88, Zeh85, dR88]. linear-exponential [BM86a]. linear-linear [GD85]. linear-structure [AH86a, AH86b]. linearity [EH83a, Nar83]. linearly [CB82]. lines [AB87, LSS85a, OS88, RD88, Tan82]. LINEX [Roi87, SAN89a]. Link [HS82, SNR84]. linkage [BE82, Wha87]. LISREL [Gio88]. lit [MB84b]. literature [Gla88, Tam86]. livelihood [BE80]. Locally [GV89b, Jan88, Rie80, CS83, Kim88a, Mar88b, MS81a]. located [Kar88]. Location [EL82b, BM87a, Bha81, Bof84, BWMD83, CW88, CV89b, CHS86, Die84b, DM89, EH87, Far88, Gat89, GS80, HJ83, HK87, Hol87, JKK86, Kap86, KR84, Kha83b, Leu89, LR82, Loh84a, Nai82, ND81, RP89, Rub86b, SK82, SS83, SS85a, SHB85, SS85b, SR88a, Sam84, Sch82, SG85a, SG88a, SJ89a, aS89a, Sim84, SSS87, SM86b, SN89b, SvW82]. location-scale [CV89b, DM89, Nai82, SR88a, aS89a, SSS87]. location/scale [Gat89]. location/two [Kot89a]. Loewner [Nis88]. Log [Hab85, Kha83a, BM87a, Bon88, CB82, EB89, Fin82, Haf86, KK80, Mar89a, Mic87, Nan89, Nis89, RG84, TG85, Wef87]. log-likelihood [Wef87]. Log-linear [Hab85, Kha83a, Bon88, EB89]. log-linearly [CB82]. log-log [Nan89]. log-logistic [BM87a, RG84]. log-odds [Haf86]. log-series [TG85]. log-zero-Poisson [KK80]. Logarithmic [Pan83, BFR80, LK81, RR88b, RJ84]. Logistic [KKC82, AA81, Aly87, BM87a, Bro82, CB85, DS89, GWBG87, HL80b, HW89a, HW89b, LWA89, LS81, LM81, Mal80, Nan89, RG84, Sam89, SRW84]. logit [Bon85, Chu83, Dav85b, FP86, GgT86, Moo88]. logits [DS86a]. lognormal [AAST88, Hol89a, ISS82, Kan82, LK83, Mar84c, Moo89, SI81, SI87, Suz83]. long [Arb85, BS88, Can84, LDH84, PG84, SL89a]. long-term [Arb85, Can84, LDH84]. longitudinal [SM86a, WL80]. look [CG81, Har84]. loss [Bol89b, Kea83, Kub89, LM86, Mar89b, MR83, MS86b, RS87, Roi87, Roj85, SAN89a]. low [Ber85]. low-cost [Ber85]. Lower [EG88, PS83, Rob89, WE80]. Löwner [Nor89]. Löwner-ordering [Nor89]. LR [AH86a]. LRT [NN82a, NCLN82]. lying [Mee88].

M [FW86, Gup82]. machines [Sri89]. Madden [Som88]. magnitude [Mic87]. magnitudinal [CG87]. Mahalanobis [Cha81, KG87, Rad84]. main [GC84, aS89b]. Making [RM89, PP87]. management [GH84]. manifolds [Bha86]. Mann [DS81, Hoc81]. MANOCOVA [Cha80, PS85b]. MANOVA [Ber86, CE85, Kab84, MS88, Mur89, PS85b, Saw87, aS89b]. Mantel [Per85]. Manual [Pet82]. many [CK87, Gup84, NR83]. maps [Mar84a, Wal80]. Mardia [Koz86b]. marginal [Ahm82, Cau83, Cau84, Hab85, Hab89, Llo87a, SU86, Tom87]. marginals [Nel86, Smi83]. Margins [Ala87]. Mark [Pet82]. marker [DH89]. market [MN89]. Markoff [Cha82b]. Markov [AM84a, And89b, BK82, BD86a, BD86b, CF86, Iso89, JR82, JB86, MC89, WS89, Yas81b, YS82b, YS82a]. Markovian [CA86]. MARL [MC85]. martingale [Kok85]. Mas [Isa86]. Masking [Ben89]. mass [JL89]. mass-size [JL89]. matched [Eri83, LPW82, MB82, OSG81, Per82, Per85, PR85]. matching [IS87, Ray83]. material [KC84b]. matric [JG85, vdMN87]. matric- [vdMN87]. matrices [AP86b, Cha80, GN87, Kan88a, Kay89, LP89, Nis88, O.’89, RS89, Sak87, SFR84, Shi82c, TG83, TB85, TGB89, YO86]. matrix [AH86a, AH86b, BF86, Dey89, DLP85, GR89, Gup88d, Kik83, Mat80, MP82b, MB84b, Nag85, PF84, Rao81b, Rei84, Smi85, öT85, Wie85, Yas81b]. maximal [AH84, Ete84, Kab84, Rob81]. Maximam [Arc82]. Maximum [tFAH84, CB85, CS84b, CF88, Des84, Ebr84, EB89, GK88a, GM88, HGB89, HN86, Kon81, Lev83, MP82c, MS81d, Obe84, SB89a, Smi85, WEA80, Wil88b, Yao86, Arn88, Ber81b, Blu85, BA84, CM86b, CC85, CW82, Dah87, Dou86, DS89, Feu87, GWM89, Has80, JC84, Kel86, Kep86, Kim86, Kul88, LP84, LP85, Mag87, Mar84d, MR88, Mic85, Rad82, RV82, RR87, Sau82, Spa85, Spr82, TO89, Tik88a, TB88c, TG89, Urz88, VS86, Wan87a, YKY85]. maximum-entropy [Urz88]. MBGDD [Muk87b]. Mean [AR87, ES87, Gel87, SY87, Swa88, öT85, AWS87, AP86a, AS84b, Alb88, Ali87, tFAH84, AH86a, AH86b, Ang89b, AGR89, BF86, Bis80, Cha83b, CG81, CL89, CF88, CL83, Dah88, DD85b, Ebr84, Fin84, Gat86, GR87b, GLT89, GWM89, Gro82, Gup80, GV89b, GM83b, Han80, Hay89, HG80, Hin86, HI89, Hsi81, JS82, KN85, Kem86b, Kha89a, KR89a, Kor85, Kub88, Lee81, Lee87c, LWA80, LP87d, Lis88, Man80, Mar84b, MG87, Mit81, Muk81, Off86, Oht87a, PG82, Pra89c, Pra91, PR86, Pri82, Rau82, Sen80, SB89a, Sha80, SOL81, Sim87, Sin86, SK87a, SM82a, Smi85, Ste89b, Sza85, TT88, TTB88, TS82, TGB89, TL89a, Tre80, VS88, VSD88, VS89, Ven88, WSR84, Wal84a, Wit89b, Wri83, WH88]. Means [Fen83, Kun83, BD87, BG88a, BW82, DD85a, Dey83a, Fon89, Gho81c, GN86a, Hai88, HL87, HB86, Hoc88, KB83, Kap82a, Kel86, Ken81, KSC81, Kul85, KS89, LM86, Mar80, MT83a, MS85, Nel82, Pep89, Pli87, RLN86, Rob81, SQ82, Shi83a, Sin80a, SW89, SCG85, SAH87, SSK89, TS81a, Tik82, TL89b, Tre88, Won82, dPV87]. Measure [Kha87a, SY87, AH84, CS81, EA89, Fla87, IB89, Koz86b, LP87a, PT86, Rav89, Ros89, SQ82, Shi82a, SG86a]. measured [Gil81]. Measurement [Gro86, MA80, MC89, GD85, Mad85, Men88, Ste89a, Ste89b, Vog89]. measurements [LY89, Mag80, Maj88, PK81, Pat83, Str88, WF88]. Measures [Lin87c, AP86a, CPW88, Fla89, FB82, GR81, HF80, JSB89, Kam88, Nay85, PM88, PH80, RMW86, Shi82b, STJ88, Siv88, SS88d, TZ85, Tiw88, Von83, Wan83, WC87, ZFP89]. Measuring [SA89, Gho87b, Gil81]. median [DM87, Mus89, SW86a, ST86a, YSD85]. medians [Cha88a, GS80, LM81]. medical [Tea84a, WL82]. Meier [Ama88, CR89, GM83a, Ger88]. Mem [Val86]. Members [Ano83x, Ano83y, LT85]. Mendel [Pie83]. Mendota [KY82]. merging [Duf83]. method [AAST88, BHW81, CI80, DW89b, Fra87, Fri81, GdT87, Gov89, Hsu85, Kap82b, Kem86a, KT85, LP81, Lee87b, LWA89, Lin87b, Liu89, Mar88a, Mon87, mN83b, O’C87a, PD84, PS86, RD88, SL88, Som88, Spa88b, Sri87b, Ste86b, Sun88, Vos86, dlHRR88]. methodology [KK83, Mas86, Pag81, WL82]. Methods [CGO85, Kim85, O’C87b, SKSF83, Anb84, CB85, CM86b, Chu82, Czi89a, Czi89c, DL84, DW87b, FF86, Fra80, GA89, Gla88, KK84, Kan82, KM84b, LM84a, MZ88, SK88, SSKK88, Sim84, SW83a, ST83, Ano83a, Ano83z]. microcomputers [Ber85]. mid [MBA88]. midranges [JP88c]. migration [Des84]. Miller [Dor88]. minimal [WT83]. Minimax [Bol87, Gut82, Kot89a, LM86, MS86b, Ber89, CZ87, CL83, EKR86, HZ82, Kem86b, SL87, SL89b, Ter81]. minimaxity [Dah87]. minimize [Ali89a]. minimizing [Ang83a]. Minimum [BWD86, Boo82, EkBL88, HI89, KR88, KC80, Wil81, Cha82b, Dav85b, FG81, Fri81, KPS83, Kum80, LR82, PD81, PH85, Sam88, Sch86a, SI81]. minimum-distance [Sch86a]. minimum-type [FG81, Fri81, PH85]. mining [LM84a, Mar81b]. MINQUE [Gie83, Kap82a, Kle88, RS84a]. misclassification [AK86b, BT81a, BFR80, HT83a, eL82a]. misclassifications [GA89]. miser [Pre89]. Mises [AC82, Ang83b, BWMD83, PD81, YB82]. Mises-norm [PD81]. mismeasured [BA84]. Missing [Abr81, DR82, Fei82, HZ83, DK85, FS86, Gen88, Gup87b, HGB89, KD89, MK82, Mur89, Rad82, Rat89, Sri85, SC86]. misspecification [Dey83a, Lai89, LT82, OH84]. misspecified [KG89]. Mixed [Sea88, Sin83, AM87, BR82, GD88, Hen88, Hoc88, JH88, KH81, KS81a, Lin86, Lin87a, Lin87b, Lis84, MS86a, Mit81, MM89a, Muk87b, Nae86, OH84, RB83a, Sen81, TT88, Ter81, Ter88, TS88]. mixed-model [MS86a]. mixing [BP83, Fro89, RR87]. Mixture [Czi88, Bev84, CE85, Czi89b, EP85, Fro89, GM80a, GdT87, GWM89, GW84, Gup80, HB86, MM83, MM89b, PG89, RR87, RHB87, SA82, Sri83, SH84, Umb81, Wal80]. mixtures [AB87, Alb88, Amo84, BK85b, Har83, JL89, Kim88a, KK81, MC87b, SS88a]. ML [AH86a, AH86b]. MLANS [BSC80]. MLE [BG84, Gho89a, Hig82, Nic89, SL82a, SN82, Yan88]. MML [BK85a, FB89, KB84, TB84a, TB89]. mobile [Wan89a]. mobility [MC89]. mod [Sea80]. modal [BR82, Kem86a, RHB87]. mode [Bor85, SI85]. Model [AR87, GBS83, McD87, NT87, Pan83, WM87, AM82, AV88, AM87, AAY86, AAST88, Ali89a, AH86a, AH86b, AJG82, AR81, BKZ88, BE83, Bar89, BBH82, BG80, BR85, Ber86, BS89, Bon85, Bon88, Bor83, Bra84, BU84, Bro82, BSC80, Buo85, CC82, CF89a, CLMFC87, Cha80, CG87, Cha82b, CZ87, Che87a, Chr86, Chu82, CF86, CEA85, Con85, DFG87b, DK81, Ebr86, EN80a, Elp83, EB89, Far86, Far87, Far88, Fic84, FP86, FME89, Fri81, Fu80, Gal88, GD85, GWBG87, GR87b, Gho89a, Gho89b, Gho83b, Gho87a, GV89a, Gio88, GW84, Gun84, GM84, GMM84, GN87, HY86, Ham89, HH86b, Haw89b, Hel88a, Hel88b, Hen88, HS86, HL80b, IM88, JP88a, JP88b, Jos88, Kab86, Kap82b, KK86a, Ket82, KW83, Kha83a, KS81a, KY81, Kon81, LT84, LS81]. model [LP84, Mag89, ML88, MJ82, Mar85, Mar82, MC86, MN89, MA83, MS86a, Mit81, MB84b, MS80b, Mos87, Muk87b, MS85, MAT82, Nai89, Nan89, NDM84, NH87, Nig88, NS86, NM88, OW86, Pap89, PK81, PRR86, PNS89, PR88, Pep89, Phi80, PC82, Pra83, PH85, PRPR88, Pro86, RS84a, RP88, Rat89, Rei84, Sad82, SBC81, SS86b, SD84b, SM86a, Shi87, SOL81, Sin80a, SJ89b, SMS84, Spu82, Sri87d, SK87b, SA86, TSM86, Tom87, TP85, VS86, WG81, Wei81b, Wil88a, Wil86, Wu89, Yan89a, Yas81b, Yum85, ZS89, DK89]. model-based [CF89a]. Model-robust [GBS83]. modeling [ARP86, Tuc82, ZF88]. Modelling [JL89, Bor85]. Models [AK83, AG83, Liu85, SKSF83, Won89, AH88, AM84b, Arb85, Arc82, BA88, BT86, BL84, Bar84, BS83, BR82, BSC80, BG88b, CY86, CB85, Cha86a, Cha86b, CM86b, CE85, Chu83, CGO85, Cna85, Con88, Cru82, Dal81, DG81, Die84a, DS89, Elp89, Eub84, EN80b, Fei82, Fla89, GR83, Gel87, GD88, Gho87b, Goo80, GR82, Gun85, Hab85, Han80, Hen88, Hig82, Hoc88, HH82, HP82, Hut84, JI84, JH88, KG89, KH81, Kal81, Kay82, Kha83b, KM82, Khu80, KM81, Kle88, KT85, KD83, KD89, Kub89, Kul87b, LMP88, Lee87a, Lee87b, Lee88, LT80, LH82, Liu89, Llo87b, Loy87, MP89, ML83, Mar88a, MRS84, Mat85, MA84, MN83a, MSB85, MB85, MS81b, Mos89, MC89, Mur89, Nae86, NN82b, Nai82, Nar83, Nor82, Nyq83, OS86, PNA85, PM82]. models [Pei86a, Pei86b, PB85, RB83a, Ram84, Ray85, RHB87, RG88, Roc82, SL88, SL89a, SS87a, Sam89, Sea88, SH82, Shi85, SJM86, Sie83, SU86, SN89a, Smo83, ST86b, Ste89a, Ste89b, SJ88, SdB89, Sza85, TA83, TT88, TTB88, Ter81, TS88, Tra83a, Val86, VI89, gW82, WK82, Wei81a, Wei89, WT89, YN86, Yao86, YAH89]. models.with [IC80b]. moderate [GR81, SD87c]. modern [God85, Sch85a]. modes [Sag83]. modification [BB80, BS80c, DN80, Joh84, KJG88, Koz89, MSVS80, Ney80b, Ram84, RF84b, Smi80]. Modified [Ars87, BWMD83, CW82, HJK89, Per85, PG83, RR87, Tik88a, TG89, Blu85, Con81, FGP88, KC80, Ng89a, San87, Swa86a, TB88c, TGG86, TG87, Vad83]. module [Dav87]. modulus [GE88, Vos88]. modulus-ratio [Vos88]. Mohan [Joh86a]. Moment [HY86, JK84, And83, Bel88, CW82, Dav85b, JC84, SD86, TT85]. Moments [BMA88, Mat80, MW85, Tra84, Tub86, Bal87a, BA88, Bal88, Bha87, Cha84, Cha86a, CT83, DN80, Fir87, Gov89, GG83, Hin86, IV89, MBA88, MP82a, Shi83b, SC89, SM82b, TGG86]. money [BP84]. Monitoring [Can84, CA86]. monotone [Bil89, DS86b, Isa81, Koc85, Lin89, Mar82]. Monotonicity [Kub88, Rav89]. Monte [Ber86, DD85a, Fri81, Kul88, Mar85]. mood [SG86c, Wec85]. Moore [AP86b, Nor89]. mortality [RMW86]. most [CS83, Goo89, Jan88, KSR88, Kim88a, Mar88b, MS81a, PP87, RA80, WR87]. motor [Gap84]. motor-tax [Gap84]. Moving [NT87, DN80, Mit88, SL88, SL89a, SB84a]. MPSD [Rao81a]. MRPP [OM80, TT85, TT86]. MSE [Shi87]. muitivariate [VS81]. Multi [AK83, Bas89, BD87, Bri87, Gup88b, Haw89a, MI82, aS89a, Sin82, SKC83, YS82a]. multi-auxiliary [Sin82, SKC83]. multi-criteria [Bas89]. Multi-Dimensional [AK83]. multi-factor [BD87]. multi-parameter [Haw89a]. multi-response [Gup88b, MI82]. multi-sample [Bri87, aS89a]. multi-state [YS82a]. multicensored [Win82]. multicollinear [Gun83, HZ82]. multicollinearity [BF81, HZ83, Kal84, Wei89]. Multidimensional [Gen88, Mic83, AL80, AL82b, BWD86, Dav86]. multifactor [Che83b, MM89b]. multifactorial [KS84]. multimodal [Gho81d, Har85b]. Multinomial [CS87b, AM86, Alb81, Bar83, BK84, BB89b, Bon85, Cau83, CD87, CB82, Che86b, CC85, Dan88, EB89, FP86, GL81, Gun84, Hos87a, Koo85, kL82b, Pon83, RA80, RF84a, Sam89, Wan86, WJT89, Wil86]. multinormal [Wal84a]. multinormality [MF83]. multiparameter [CAR89, Dey83b, Min80]. Multiple [BLW87, GN83, JP88a, SB82b, Aic83, Ben89, Bor83, Bor85, CK83b, DW87a, DW89a, DB86, DR82, Dun86, Dun82b, FP84, Geo86, Gup87b, GH85, HL80b, Hsu85, KKB87, Lam89, LWA89, LH82, Mor87, SQ82, SS86b, Sri83, TCJ89, Tra83b, Von83, ZF87]. multiple-input [LH82]. Multiplicative [Chu83, Arb85, BJ88, Chu82, CGO85, MJ82, Mar85, Sch86b]. Multiplicative-interaction [Chu83, Chu82, CGO85]. multiply [Ger88, Mar84c]. multiprocess [Wil88a]. multisample [DM89, GN88, SW86a]. multisource [Gat89]. multistage [Abe80, OP86, Odh86]. multistate [ENP84, Kay82]. multitype [KJ83]. Multivariable [HA83, DA80]. Multivariate [AR87, BKZ88, Byr88, CHW87, Cor84, Die89, Gho81b, Har86, Jac85, Kha83c, Kha89a, NT87, PH80, aS89a, Sri85, VW83, Yat86, Alw86, tFAH84, AH86a, AH86b, BM82, Bar89, BF86, BJ83, Bha82, BT81b, Bou82, CG87, Cha83a, CP84, CK83a, CL83, CA81, DPS81, Dah87, Dey89, DLP85, Die84b, DW89b, DT81, DS80, DM89, EN81, FKN83, Fox89, GP82, GR83, GH85, Iso84, ISS82, KN85, Kan88a, Kem86b, KS81b, Koz86a, Koz86b, Kub88, KP84, Lee87c, LP87b, MYT87, MB84b, MSB85, MS81c, Nae86, Nai89, NDM84, NvdM86, Pei88, PF84, PR86, Pre89, Pro86, Pro88, RK81, Rad82, Raf84, RV82, Rei84, Rob89, SL88, SL89a, SS83, SL80a, SW86a, SYT89, SU86, Six81, SCT83, SZC85, SAH87, SA82, SA86, Suz83, Tan89, TS82, TB84a, TB88b, Tub80]. multivariate [Urz88, Wan83, WC87, Xek86, ZLMK82]. Multivaritate [CH88]. Multiyear [SG84a]. MVUE [Asr85].

Narrower [Ars87]. natural [EkBL88, Vog86]. nature [Kee80]. NBU [PS87]. near [NJ89]. nearest [Che87b, Gab87]. Nearness [Kha87b, LP87a, PK86, RRM86]. necessary [GC84, Roj86]. need [Sri87a]. negative [BS86, Bon85, Con89, Ebr82, EB89, Gho81b, Har83, HG83, Muk88, OL86, Ver88b, Wal84b, Zac81b]. neighbor [Che87b]. neighbours [Goo87]. nested [BG85, LL82, Lun82, MM89a, Ver88a, WG81]. Neumann [Bar84]. Neyman [MSVS80, Smi80, FG88, Gab80, KK85, Mie80]. Nile [IS86, JN87, JN89, Wit87b]. No [DG83, MSVS80, Ney80b, PO84b, Smi80]. Noemal [SB82b]. Noise [Gup87a, And88, DW87b, KG87]. Nominal [AL80, Gil81]. Nominal-scale [AL80]. Non [ES88, Gho87b, Joh85, SG85a, SJM86, SW83a, WR87, BS86, CS83, CV89b, CVDS85, Del81, Fra83, Gad83, GgT86, GK88b, Gup84, Hig82, Hor81, KR89a, KP89c, KN83a, KN83b, LDH84, Loy87, MS81a, Mik85, MN83a, NNQ84, NS85, OL86, PS86, Saw88, Sch85b, Sch86c, STSM89, SG84b, Swa87, TW81, TS88, Ver88b, Zac81b]. Non-adaptive [WR87]. Non-Bayesian [SW83a, GK88b]. non-central [NNQ84, OL86]. Non-conjugate [SJM86]. non-homogeneous [KR89a]. non-informative [CV89b]. Non-isomorphism [Joh85]. non-isotropic [Sch85b, Sch86c]. non-linear [Fra83, Loy87, TW81]. non-negative [BS86, Ver88b, Zac81b]. non-normal [Hig82, MS81a, Mik85, NS85, STSM89]. non-normality [GgT86, MN83a, TS88]. non-null [CS83, KN83a, KN83b, Saw88]. Non-orthogonal [Gho87b, Gup84]. Non-parametric [SG85a, Hor81]. Non-random [ES88]. non-regular [SG84b]. non-response [Gad83]. non-sequential [LDH84, PS86]. non-singularity [KP89c]. non-smooth [Swa87]. noncentral [Bet87, CARO84, Gel83, HH86a, Lam86, Mat80, MP82b, YM85]. nonhomogeneous [TL89a]. nonidentically [BB89a, Spa85]. Nonincreasing [Win87]. nonintegral [RS88]. Nonisomorphic [SFR84]. Nonlinear [And89a, GM85a, GM85b, Bon88, GdT87, Hug80, KM82, Ste89b, Val86, WT89]. nonmetric [Rav81]. nonnegative [And89a, LP89]. nonnegligible [Gho87a]. nonnormal [KK85, KKB87, PL82]. nonnormality [BK85b, JP81b, SdB89]. Nonnull [GN87, GN88]. Nonparametric [Che83a, DP83, Dav84, Ebr85, FH84, Gho89b, Gho85, GLT89, Hsi84, Iso89, LP88, LP85, PM84, PS87, RU89, SK82, SS83, SW88, Shi82b, SGP83, SU86, Smo83, TT89a, Win87, Yam87, Zeh85, Bar88b, Ber82, BW82, Che84b, DM89, Elp83, FR88, GHK85, Hol89b, Hsu80, JV87, LaR84, LP87d, MS86a, PW81, Pra89a, Pra89b, Rub86a, SQ82, Sch82, Sch83, Sch85c, Shi82a, SW83a, SD87c, SG85b, Swa86b, Tri88, Vad83, WW84]. Nonproper [PP88a]. nonrandom [FP89b]. nonregular [GS81b, KR88]. nonrespondents [Gun84]. nonresponse [Alb85]. Nonsampling [Iac83]. nonstationary [Fin85]. Norm [Shi82c, DS89, GM85a, GM85b, GdT87, Gra89, Nyq83, PD81]. norm-restricted [DS89]. Normal [Ang83a, CS87a, DS82, IM88, Joa83, LS83b, tFAH84, AH86a, AH86b, BK85b, BF86, BD87, BHL82, Ben89, Bis80, Bof88, BCK85, BW82, Cha81, CT83, CP84, CL89, CO84, CO86, CL83, DeP83, Dey83a, DLP85, DW89b, DS80, Ebr84, FKN83, FS80, Gho80, GN86a, Gle82, Gov89, GV89b, GM83b, Hay89, HD89a, Hig82, HM81, Hut84, IB89, Joh82, JS82, KS85, Kan88a, KL80, Kel86, Kem86b, Kim85, KK81, KKB87, Kor85, Kub88, KSC81, Kul85, KS89, Lee87c, LM86, Man80, MT83a, Mee88, MG87, MS81a, Mik85, Mil82b, MM88, NS85, Oht87a, OSH88, PS81, PP87, PG82, PKN89, PR86, Pro88, Rad82, RLN86, Ray85, RC85, Sak87, SC80, SL80a, SS88a, Sha80, Sin86, SW89, STSM89, SM82a, Six81, Sob89a, SAH87, SA82, Ste89b, Sun88, Tal80]. normal [Tik88a, TG89, Tub80, WSR84, Wie85, YI88, Yat86, dPV87, eL82a]. normality [AB87, BJ83, BM81b, Che86a, Chm81, GP82, GgT86, KL84, Kou81b, Koz86a, MY85, MN83a, MS81c, Mos83b, OM80, Tan82, TS88, TYM86, Won88]. Normalizing [Kun83, LO89]. normals [GWM89, Sri83, SH84]. Note [Cha82a, Dav88b, GS81a, KA87b, Lem81, Som84, Tre83, AM81, AT84, AG89a, Ang89b, Ang89c, Bal87a, BFR80, Beg81b, Bol89b, Bre82, Cer89, Cha88b, Che83b, CK83a, Cho83a, Chr86, CV89b, DL86, Dor88, Eub83, Fra88, FS80, GR87a, Gal80, Gil89, Gil81, Gil82, Gov89, Gun86, Has80, HL80a, Hig82, HG83, IS85, Koz86b, Kum80, Li82, LP89, Loh85, MYT87, Mur89, NO85, Par82, Per87, PH85, Pre89, RV82, RS89, SAN89a, SL80a, Sch86b, SW86a, Shi87, Sho82b, Sil86, Sin82, Sin86, SD87b, Swa86a, Tak88, TO85, TYM86, VS89, Whi89b, Wri88a, WH88, YI88]. notes [Woo88]. notion [Dav86]. nuisance [BM89, Spa89, Tar88, Wit89a]. null [CS83, CW87, Dav88b, Del81, KN83a, KN83b, MS81a, Pep88, Sam85, Saw88]. number [Bha89, JP81a, MS82, MCN81, Par81, Sob89b, YKY85, Yu85]. numbers [CS88b, Loh85, Loy89, Rao81a, Sad82]. Numerical [GdT87, SL82a, Vog89, Wah81, Rav81, Pet82].

O [Kai88, HP80, Rus80]. O-statistics [Kai88]. observation [Bha86, DK85, Gho82, KJS^{+}83, MYT87, Rau82, YAH89]. observations [Abr81, Adk86, Ali87, Bal87d, Bar81, Ebr85, Gho83a, GS85a, GS87, GHO86, Gup87b, JI84, KD82, KR82, KD89, LP85, Nar83, OM87, OP87, Pen85a, Pen85b, Rad82, Sha86, SS88b, Sri85, Wal89, Zeh85]. observed [PRR86, Wu88]. obtain [Duf83, WK82]. obtainable [Gup88c]. obtained [Ban80a]. obtaining [KP89c]. occupancy [MP85, PM85, Sic86]. occupational [MC89]. occurrence [Scl80]. occurring [McC86]. Odds [Hau89, Haf86, Hei88, HK83, Mar89a, Mic87, YN86]. OLS [SA89, Yum85]. Omnibus [MF83]. One [Adk86, Ars87, Ber84, Bha87, CO86, Cho88, KRS83, Ach84, AH86b, Ang83b, Bof88, Bre82, BMC86, CF89a, DW87a, DW89a, DK85, DK89, EA89, FH81, GC84, Gup87b, Hor85b, JK85, JP87, KC84a, Kam88, Kap82a, KP80b, KRB89, KJS^{+}83, LT85, LWA89, Mar82, Mar89b, MC85, Nel86, PRPR88, RS84a, Sak87, Sel89, aS82, SJ89b, Yeh88, Zac81b]. one- [DW89a]. one-parameter [Kam88, Mar89b, Nel86]. one-population [AH86b]. one-sample [Ang83b, Hor85b]. One-Sided [Ars87, CO86, Cho88, KRS83, Sak87, Zac81b]. one-stage [CF89a]. one-step [JP87]. one-truncation [KRB89, Sel89]. one-way [BMC86, DW87a, DK89, KC84a, Kap82a, Mar82, PRPR88, RS84a, aS82, SJ89b, Yeh88]. Op [SB82b]. operating [MCR88]. Operational [PCP87, Fir87]. Operator [Fin85]. operators [Cha84]. Optimal [BF81, BV89, Che88, DF89, EHA84, Fro89, GS87, HT83b, Kem86b, Maj88, MH86, PG89, RH88, Swa87, Ven80, VB80, BW82, CA86, Czi88, Dal88, DS87, ER80, EZ81, EKR86, Eub83, GV89b, Hsu82, Huf85, Kan88b, KM87, Lee87a, MH85, Neu88, Nyq83, PG82, PG83, Ray82, Roj86, Sin85, Swa88, Wit87b, Wri88a]. optimalities [Muk87b]. Optimality [BM89, Llo87a, O’C87b, BP81, CM84, Gal80, GS83, Hos87b, Li87, O’C87a, PP88a, Rus80, Sil86, Sin87, SS87b, Yeh88]. optimally [AUSH83]. Optimizing [Wha87]. Optimum [Kag88, PO84a, PO87, SHB85, Swa80, YS84, AS84a, LC85]. Order [Bal86, BM86a, BMA88, Har86, Tik88b, ZFP89, AUSH83, And86, BM85a, Bal87a, Bal87b, BA88, Bal88, BB89a, Bar88a, Bar88c, BD86a, CVV82, CW88, Cho88, Cor84, DMH81, Dav88a, DS85, EKR86, EK85, Gov89, Gro82, Has80, KA87a, KPS88, Koo85, KP86, KP87, Kos87, KY82, KM87, MT82, MBA88, Mar89b, MDH81, MH88, Muk87a, Nag80, NP88, PK85a, PTA89, RG84, RG88, Roh87, RS88, SS85a, Sat88, Sen81, Tak88, Wu88, ZF88]. Ordered [FH87, AJK89, Bar83, Ber82, BS83, BS84, Die89, Gil82, HL87, Hoc81, KI88, Ken81, KS84, Kul85, KS88, KS89, Man89, Mar80, MT83a, MA83, Miy87, SQ81, SW89]. ordering [BK86, Den89, Ebr82, EN81, Fer87, Nor89, Wri88a]. Ordinal [AK83, Ier85, SKSF83, AL82b, Nan89]. ordinal-scale [AL82b]. Ordinary [Krä85, BU84, TG87]. ordinates [Dey83a]. origin [Cha81, Iwa89]. original [Nar83]. Ornstein [Wit84]. orthant [Sun88]. orthogonal [Czi88, Gho87b, Gup84, Muk86, StR82, SFR84, Sri87a, yS89]. orthogonality [Sha88a]. other [AH83, CO84, Jai81, SD86]. outcomes [BA84]. Outlier [BT86, Bal87b, BA88, Bal88, CK87, FD85, Gat86, JP81b, Jos88, NM88, PCP87, SYT89, Sim87, SL83, SL85]. Outliers [Bar88a, Bal89, Ben89, Bha89, CK83b, Dix89, FKML89, GK88a, JP81a, Jai81, JPD82, KN87, Mil82a, MN83a, Nae86, Nai89, PB87, Sim84, Spa88b, Ste86a]. outlying [Rau82]. output [Ven80]. overall [Duf83, Sen87]. overflow [Mon87, Ram87]. overlap [IB89]. Overlapping [MSL86, IB89, Sin88]. overview [DL84, Spu84]. ozone [JBF^{+}82].

packing [SW86b]. Padé [Tuc82]. pair [BE82, Ray83]. pair-matching [Ray83]. paired [Che84b, CW87, Dav88b, EHA84, FP84, Fro87, GBS83, McC87, Sad82]. pairs [Eri83, LPW82, MB82, Per82, Per85, PR85, Wu82]. pairwise [AH84, BG88a, Ete84, FP84]. Panel [OEJ^{+}85, CM86a, KO82, MC86, MC87a]. paper [Joh86a, KJ90]. parabolic [CHW87]. paradigm [SSS^{+}85]. parallel [AB87, EG88, Gup88c, SD87b]. parallelism [LSS85a, SS86a]. Parameter [Dan88, EP85, FG81, Fri81, LaR84, LT80, SBC81, Sic86, Arb85, Asr85, Bed88, Beg80, Beg81a, Beg81b, Bel88, Ber89, Blu87, Blu88, BG84, Cha83b, CB87, CB88, CV89a, CW82, CHS86, Dah86, EH87, EA89, EL82b, EN80b, GM88, GS80, Gut82, HW88, Haw89a, Hel88a, HN84, How85, Isa81, Iwa86b, Jel87, Joh82, JC84, JN87, JN89, Kad81, Kam88, Kan82, KK80, Ket82, Khu80, KB80, Kim86, KRB89, LK83, LT85, Mag89, Mar89b, Mat85, Mil82b, MC85, NN82a, Nel86, PM89a, PD81, Pau85, PP88b, Rau87, RM89, Rus88, Sam84, Sel89, Sim84, SL82a, Sir84, SvW82, TS81b, Tom87, Wei81c, Wei82, Yan88, ZF88]. parameterizations [Sza85]. Parameters [NT87, Sin83, Wan09, AC82, Amo84, BM87a, BT81a, Ber81b, BC88, Bof84, BM88, BR82, BWMD83, CW88, CQ82, CC88, DB86, DD85b, Dix89, EZ81, Fro89, Gho83b, Gho87a, Gle82, HN86, HP87, HK87, Haw89b, HD89b, HD89a, Hel88b, HG83, Huf85, JKK86, KI88, Kha83b, KB82b, Kub89, Kul87a, KS88, Lam89, Llo87b, MD89, MB85, MM89a, MC87b, MD87, PR88, RK81, Rad85, Ray85, RHB87, Rei84, Rob81, SS85a, SHB85, SS85b, SR88a, SD83, SD84a, Sau82, SM86a, SJ87a, SJ89a, Spa89, SA86, Tar88, Tea84b, TG86, Tub80, Wei80, Wit89a, Yam86, Yam87, dWM85]. Parametric [LaR88, Che83b, Ebr86, Hor81, OA87, Pon83, SG85a, Tun85, Won89]. parent [Kha87a]. Pareto [AS84a, DMH81, HP87, Iwa86a, Kou81a, MT82, MT85, Mar84c, NH87, WA84, Win82]. parsimonious [Smi84]. part [GdT87, Kaf82, BM86a, GM85a, GM85b, Sam89, SW83a, ZS81]. Partial [KP89a, ZF87, ARS86, BB80, BHW81, Kha87a, KF89, Loy87, Pro88, Shi87, Slu86, SM87]. partial-autocorrelation [SM87]. Partially [GR89, KSR88, Loh84a, MP82d, BK85c, KK83, MN89, Miy87, OW80, PK85b, Rol88, SM82b, Wu82]. Partially-adaptive [Loh84a]. particle [JL89]. particles [PM82]. particular [GN86b, WA84]. partitioned [Kik83]. Pascal [Pan83]. passage [GN86b]. past [JC87]. patch [Pro84]. path [FV89]. Patricia [Dav88b]. pattern [Gup88d, JK84, PL86]. Patterned [BK85c, tFAH84, Sza85]. paucity [GR87c]. Pavlov [Ger88]. PBIB [KM84b]. peak [Hor81]. Pearson [BJ83, FG88, Hos86, Ros89]. PEB [KP85]. penalized [LP85]. Penrose [AP86b, Mik85, Nor89]. Percentage [FD85, PF84, Mee88]. percentile [Blu85, Blu88, Che83a, JP83b]. percentiles [CO84, Kea83, OSH88, SB82a]. Performance [Dre85, HF80, AB87, Kok85, NS85, Odh86, Odh87, SS87a]. period [KJS^{+}83, PH80, Pat83, RB83b, WF88]. Periodic [Spu83b, Che87a]. permitting [Bla84]. Permutation [MI82, Som85, Bra88, MB82, Sch84b]. Perspective [BS80c, BP87]. petroleum [Dav81]. pharmaceutical [Lai84].

Phase [HA83, Fre84, Anb84]. phases [Cna85]. phenomenon [BH86]. phenotype [DH89]. phi [Ros89]. phi-squared [Ros89]. piece [BS88, Joh82, Kim85]. piecewise [Bra81, Che88]. Pietra [Moo89]. Pillai [Mik85]. Pitman [Kha87b, KS89, LP87a, PK86, RRM86, SG84b]. Pivotal [Llo87b, SMS84]. placebo [Chr88]. placement [OW80]. plan [CB81]. planning [Gho83b]. Plans [Lun82, GC84, KB84, SM85, SV89, SM89]. plate [VW83, Wah81]. play [CLl86]. players [HH80]. plot [KP84]. plots [Cer89, MK82, WT89]. plotting [Har84]. plus [GC84, Sim87]. Point [DA80, LS80, AB89, BHW81, Boh83, Cha89a, DeM84, DC84a, Ebr86, Gho81c, HH89, Hsi84, Hsu81, IB89, JP87, Ken81, PL88, Pep88, Pra89a, Pra89b, SW88, Sen80, SM85, SM89, SD87c, Wol80, Yao86]. Point-of-expansion [DA80]. points [FD85, McC86, MZ88, Nel82, OS88, Ped85a, Ped85b, PF84, SLH83]. Poisson [AM88, Alb88, Bon88, BS80b, CS84b, CF88, Con88, CF89b, DS86a, Huf85, Job87, KK80, Kim88a, KR89a, OW86, PB87, PH89, PM85, Scl80, Sha81, Sha84, Sho82c, Sin80a, SJ87b, SJ88, Tal80, TG86, Umb81, WS82, Wri83]. Poisson-inverse [OW86, Sha81]. policies [CA86, Kot89a]. poly [LM84c, Won88]. poly- [LM84c, Won88]. Pólya [JR82, PTA89, Wal84b]. polychoric [LP87b]. polynomial [BBH82, BC88, Bra81, Che88, DA80, EKR86, EC83, Fin85, LT84, OS86, SM86a]. polynomials [GCH83, PM89b]. pooled [GR89]. Population [Hin86, WM87, AR83, Alb85, AH86a, AH86b, BT81a, BLW87, BE88, Bol89b, BW82, Cha81, Fra89, Gho81c, Hay89, HM81, Hsi81, HT83b, Hun89, KM80, MS82, NR83, Off86, PS81, Par81, PS87, PS83, Pra89c, Pra91, Ris85, Sha80, Sin86, SSK89, TS81a, TB84b, TL89a, Tub80, VS88, VSD88, VS89, Ven88, ZS81, Zac81a]. Populations [CS87a, SB82b, AR83, Ama88, Bjø85, BM86b, Bof88, Bol87, Bol89a, Che80, Cho80, Cho83b, GC80, GK81, GM83b, Hsu82, JP86, JP88c, Kaf82, Kan88a, KL80, Kim88b, KP80a, KP80b, KP81, KP82, KK81, KS82a, KN83a, KN83b, LS83c, Lia88, LM81, Muk88, O’C87a, PP87, RC85, Sak87, San89b, SM82a, SAH87, SA82, Ste89c, Str84, TB88b, TGB89, Tse88, Val86, Wal80, dPV87]. portmanteau [DR86]. positions [Har84]. Positive [Gro84, KS81b, AP86b, CT83, CK83a, Kik83, Sha84, Smi85, TG83]. positive-definite [Kik83]. positively [Ber81b, IS85, Sob87, dlHRR88]. possibly [Sch87]. Post [SF85, Dav89, Leu88a, Leu88b]. Post-hoc [SF85]. post-stratification [Leu88b]. post-stratified [Dav89, Leu88a]. posterior [FCP^{+}82, Gun86, Siv88, Spa88a, VW83]. potency [EH83a]. potential [dR88]. Power [ARP86, BE82, HK87, BH84, Ber86, Con81, EA89, Fri87, FD85, GE88, HH86b, Jai81, Kap89, KDK85, KR82, KC80, Kum80, MT82, OA87, RS81, Rig89, SW89, Sri87a, TT86, TGG86, TG87, VB80, Zei89]. powerful [CS83, Fei88, Goo89, Jan88, Khu82, Kim88a, Mar88b, MS81a]. powers [CS88a]. Practical [SG84c, Kle87a, SD86]. practice [KKC82, Mar81b]. practitioner [DeM84]. pre [Oht88]. pre-test [Oht88]. precision [Gil89, KM80, Pal88]. precisions [GM83b]. Predictability [Fla89]. predicted [Ang83a, KM84a]. predicting [Fu80, LM84a].

Prediction [AR87, Ami89, JH88, JP88b, Ng89b, Nig88, OS86, Pat89, Bal89, Bol87, Bol89a, Bol89b, CO86, Cho88, Dav84, EN80b, GG86, Gel87, GH85, Hen88, JI84, KH81, Kul87b, Kul87a, LP84, NH87, Nig89, PU86, Pei86a, Pei88, SB85]. Prediction-interval [JH88]. predictions [AR81, KO82]. Predictive [Men88, De 89, Kap83, NDM84, NM88, Tra83b]. predictor [Gun83, Hur87, KA80, Mas86, OH84]. predictors [JC87, Rus88, Sch80]. Preference [KM88, CVDS85, STJ88]. Preliminary [SK82, LSS85a, LSS85b, Lee87c, Oht87a, PU86, RS81, SS83, SS87a]. presence [AC85, BM89, Bal87b, Bha89, BF81, Dix89, FP89b, GK88a, GD85, JI84, Men88, RM88, SLH83, SL85]. present [BG88b, Wit89a]. preservation [CVDS85]. preserving [ZFP89]. pretreatment [GP86]. previous [Bal87d]. price [FV89, PG84, Vog87, Vog89]. prices [MS81c, Vog89]. primary [Spa89]. Principal [Oks88, Soo88, BP84, Cer89, Hsu81, NO85, Sch85b, Sch86c, Tan88]. principal-component [Cer89]. principle [CPW88, NS84]. principles [OM80, Sen81]. Prior [AG83, FS86, AM87, BF81, CD87, CP84, DGG89, DC84b, Gho81a, Lee88, PO84b, Pli87, Rao82, SJM86, Ste86b, Tre88, Wan86, Wri88a]. Prior-likelihood [FS86]. prior-valued [Wan86]. priori [Fer87, Jel87, OP84]. priority [SG86b]. priors [Alb88, CV89b, Dey83a, GK88b, Siv88, TSM86, Zac81a]. probabilistic [Fu80]. Probabilities [Cha84, AM86, BT81a, BK84, Cha86a, CB82, De 81, Den89, GN86b, Joe88, KM88, Mon87, OP84, PO84b, Pat85, Shi88, Tar86b, VW83, WCK85, YS82b, YS82a]. Probability [BE80, MT82, YSTM88, ARS86, BB80, BFR80, Bev84, Con81, FB82, Ger88, Hut84, IB89, Iwa86b, Jen83, KA87a, Li88, MP85, Nis89, PW81, PT86, Ram87, Rol88, Sam88, SL80a, Spa88a, Sun88, TZ85, TT89a, WH87, Wit89b, YN86, Yan89b, Zac81b, ZTJ86]. probable [RA80]. probing [Sri87a]. Probit [AA81, Ier85]. Problem [CS87b, FH87, AG89b, BCDE85, Bed89a, Bed89b, Beg81a, BR86, BD86c, BS87, CG81, CW87, CB81, Dah87, Dav88b, DT81, ES84, FR83, FB82, GK81, GL87, HG86, Hor85b, IS86, JN87, JN89, KG89, Kaf82, Kha87a, Kok85, MP85, ND81, NvdM86, Neu88, NS86, PG83, Per87, Pur86, Ram84, Rob89, RF84b, SS86a, SRP82, Sar85, SD86, SW88, Sch83, SG85a, Shi80, SW86b, SG86c, Tam86, VW83, Wit87b, eL82a]. problems [BYD85, CP82, Che86b, DR82, GdT87, GH85, KJS^{+}83, LC87, McC82, MH88, Muk88, Nag80, NR83, SG88b, Sri87b, Tan89, YO86]. procedure [AH83, Bor83, Bor85, Bra81, BG88a, CF89a, Dal81, Fri87, Gie83, Gra87, GL87, HP87, JP81a, Kab84, KW83, Kim88b, KP80a, KPS83, Mar80, Mos83a, MS88, PS87, PG83, PKG87, RA80, RS81, San87, Seh88, Spa85, Sri87d, TGB89, WS87]. Procedures [GSS87, BE83, BT82, BK84, Bjø85, Bro81, BF85, CK87, Dun82a, Dun86, EH83b, FP84, GM80b, HGO85, HM81, Hsu80, Hsu82, HH82, HT83b, HJK89, JP81b, Jen83, JH88, KH81, KI88, KL80, Kay89, KKB87, KSJ86, KJG88, LaR88, LY89, Lyb85, MS80a, Mor87, NR83, OW80, PST87, PL82, PL86, Per85, PS86, RS87, Sim87, STSM89, Smi82, ST86a, Swa83, TB88a, TB84a, TB89, TS88, Tse89, Wu89, Zac81b]. process [BD86b, BE88, BS87, CO89, CA86, Czi88, Czi89b, DD85b, De 81, Des84, Gho81a, GRY81, GN86b, Has80, JB86, Job87, KR89a, KP84, Mar84d, Mar89b, Mus89, Pur86, Rig89, Rus88, Scl80, SI87, Wil88b, Wit84, Wol80, Yam86, Yam87, ZF88]. processes [AM84a, AM88, And88, And89a, BJ88, BK82, DN80, HP80, Iso89, Kay82, LM84b, Mit88, PH88, Pei88, SB84a, SB85, SM87]. processing [Bha89]. Product [Wan87a, Box85, CDS84, CS84a, Kot89a, MT85, Rao87, SG87, Shi83b, SIU85, ST83]. production [MD87]. products [Khu82]. Professor [Mie80]. Profile [Sri87c, Nel87, Saw87]. program [MS87]. programming [DS87, DF89, Kab83]. progressive [TJZ88]. progressively [DS81]. projection [BE88]. Projectors [Kal81]. proof [AP86b]. proper [Ste86b]. Properties [AR87, MR87, Nel86, CD87, CW80, CA81, DM87, DS89, Gup82, HY86, HB86, JP83b, JP83a, JB84, Kag80, KK86a, Kha83c, Kim86, Kle87a, LW85, LS83a, LT82, Mac82, OH84, Oht86, Oht87b, OSH88, RG84, Shi85, Sho84, Sim87, SKG80, SDCB83, Swa88, Won82, Wri88b, Yum85]. property [Fin82, Gup88d, MH85, MH86, Rie82, Sha86, Xek83b, ZFP89]. proportional [AS84a, CB85, CEA85, Gab87, GR87b, Gho89a, Gho89b, GV89a, Kay82, RMW86, YS84, YAH89]. proportionally [Ebr85]. proportions [Alb81, AD80, AJG82, BP83, Bur81, Cau83, Cau84, Che84a, DGG89, GL81, Hab89, Hey84, KW86, OSG81, PH84, Tar86a, Wal80]. proposed [Ram84]. prospective [Cla82, YN86]. prototypes [HB86]. proving [Swa86a]. proxy [Oht87b]. Pruning [CT89]. Pseudo [Alb81, CC85, Fen83, Chi81, WAS88]. Pseudo-Bayes [Alb81]. pseudo-complete [WAS88]. Pseudo-Replicate [Fen83]. public [Ney80c]. purchasing [Cha86b]. purely [Che87a]. put [Sob89b].

Qdf [BT81a]. quadrant [Ebr82, dlHRR88]. quadratic [BCL80, Cha89a, Cha82a, CT83, El 81, Elp89, HGO85, Kab83, Kub89, Li88, tL84, LM86, MW85, NS85, RS87, SA82, SB89b, TC81, VJ88]. quadratic-discriminant [BCL80]. qualitative [PT86]. Quality [Har85a, Lyb85, Tag85, Alw86, BM85c, DS87, God85, HL85, Jac85, LC82, Nat85, Net85, Rod85, Sch85a, Sri87a, You85]. Quandt [Fre83, Hor81]. quantal [AD83, Tam86]. Quantile [GV89a, Kai83, AR83, Ang83a, KL82a, Kou81a, LaR88, LP87c, LP87d, Ste89c, Tam86, WT89]. quantiles [AUSH83, And83, BDF^{+}84, DMH81, Eva82, Fal86, Fro89, GS81b, Ima87, JV87, LWA80, MDH81, PS83, Sha88b, SD85]. quantitative [Kle87a, PT86]. quantitative-qualitative [PT86]. quantities [Llo87b]. quantity [SMS84, Vog87]. quartic [Li88]. quartiles [JKK86]. quasi [Mal80]. quasi-range [Mal80]. questions [Pie83]. queue [Gup82]. queueing [Kum81]. queues [RB83b]. Quota [JKW89]. quoted [Duf83]. quotient [MT82, MT85]. quotients [Loh84b].

R [Din87a, Din87b]. radiation [Ney80c]. rainbow [Utt82]. rainfall [Gab80].

Random [AH88, Ete87, Loh84b, AM82, AM87, BM81b, BG88b, BA84, CY86, CS84a, Den89, DJD89, ES88, Ete84, GR83, GV89a, Gil81, GQ84, Han80, HH86a, HH86b, Hug80, JP88b, KK84, Kel86, KK83, Kim88b, Kle87b, LMP88, Li88, LP87d, MT83b, MB84a, Mil82b, Nis88, PS85a, Ped85a, Ped85b, Ped86, PR88, PC82, PRPR88, RS84a, RP88, Roh87, Sat88, Sch80, SR88c, Shi85, SJ89b, Sir86, Spr89, Sri87d, TTB88, Tar86b, TT89b, WEA80, Wit89b, Wri88b, YS84, YSTM88, ZS81, Zac81b, ZF88]. random-parameter [ZF88]. Randomization [Lor84, Dav89, SW83b, Vad83]. Randomized [Bou82, CLl86, BM85b, Ber82, Bra88, BW88, FY86, Fer87, Fra89, Gab80, Gun85, HT83b, Kor87, Mat81, MP85, MI82, Per81, PM85, SF85, SC85, VB80, Von83]. randomly [Che87b, EL82b, Kim86, Mus89, Neu88]. randomness [Bar83, DR86]. Range [Siv88, Gro82, Mal80]. ranges [UV84]. Rank [AR87, FH87, NR83, SKSF83, zS88, Shi80, Wu82, AC85, AJK89, BCDE85, Bar84, BH84, BS83, BS84, BW88, CH83, Che83b, CI80, DW87a, DW89a, Die84b, Fra88, Fro87, Gov86, GR82, Jan88, Khu80, LW85, LSB89, MP82a, Man89, Mar84b, Mar88b, MZ88, MS81a, MS80b, ND81, Neu88, OA87, Rie82, SW86a, Sen81, aS82, aS89a, SSS87, TT86, gW82]. ranked [Lam89]. ranking [Coh82, DT81, FP84, FF86, McC82, RS87, Sri87b]. rankings [FY86, Fer87, SQ81, SQ80]. ranks [AP86a, CH88, FP89a, Lem80]. Rao [CM84, Sin87, Sta86, Tal86, WE80]. rare [RG89]. Rate [MR87, AOA89, BM86a, BS89, Che87b, DS86b, Duf83, Ebr86, Koc85, LWA89, Lin89, MS81d, PW81, SW83a, Yao86]. rates [BK86, BDK89, Che88, Duf83, GL89, Job87, Mad85, Muk81, Tak88]. ratio [BE80, Bar81, BB80, BD86a, BR86, Cha83a, CT83, CDS84, ĆM89, Dun82a, El 81, Fei84, Feu87, Gro84, GN87, GN88, Hau89, Hei88, Hos86, Hos87a, HLN88, Isa81, IS85, Kab86, KG87, KR89b, LMP88, LP87a, Mar89a, Mil82b, MM88, MW85, Off86, Ped83, Pra89c, Pra91, Pro88, Rol88, Sak87, SG87, SJ89a, SIU85, SW89, SS81b, SDCB83, ST83, SSK89, SB89b, Tub86, Ver88a, Vos88, WG81, YN86, ZG80]. rational [Mar88a, Som84]. Ratios [Buo85, Chi81, HH86a, HK83, Joh86b, Mic87, RM88, SB82a, Wef87, Yat86, ZLMK82]. ray [BV89, WS84]. Rayleigh [Sho82b]. RC [Chr86]. RCBD [Per81]. Re [MY85, Chm81]. re-appraisal [Chm81]. Re-examining [MY85]. real [Li88]. Recapture [WM87, GL89, SMS84, VS86]. receiver [MCR88]. recognition [Ano85k]. Recognizing [Hem80]. Reconsideration [Lec85]. reconsidered [Czi89a]. Record [Nag88]. recovered [WK86]. Recovering [MS81b]. recovery [PP81]. RECPAM [CT89]. recreationists [Gap84]. rectangular [KM88]. Recurrence [BMA88, MBA88, BB89a]. recursion [BR85]. Recursive [Hol87, Iso89, JP81b, Loy87, PG89, SU86]. Reduce [AR87]. reduced [MS81b, SM82b]. reduction [Phi80, Sam89, SA81, Sun88, Wit87a]. redundancy [CDS82, Kha89b]. reexamination [AB86]. Referee [Ano85k]. Reflecting [AG83]. reflection [Wal84b]. Reflections [Wat87]. regarding [Gab80, Hab89]. Region [Ars87, ID80, MM83, MM89b]. Regional [HW87b]. regions [Chi81, JKK86, KN85, KM88, Kul87a, Sch80, aS89b, Tub80, Wie85, ZLMK82].

Regression [Ber84, Gun83, HA83, Lis84, Spu83b, AB87, AAY86, AAST88, Ali89b, Atk82, BT86, Bar88b, Bar84, BP84, BC88, Bil89, BU84, Bru80, CC82, CY86, Cer89, Cha89a, CB85, CM86b, Che83a, CT89, CL83, Dal81, DD85a, Die84a, DR82, Dor88, Dor89, DA80, DS89, EZ81, EC83, ES84, Eub84, FBV84, Far83, Far84, FW86, Fir87, Fla89, Fra83, Fre84, FD85, Gad83, Gal82, Gal80, Gap84, GD85, Gel87, GWBG87, GH84, Gue87, GH85, HY86, Hal84, Ham89, HP83, Hol89b, Hor85a, Hos87b, Hos88, HL80b, HH82, Hsu81, HG83, HP82, Iwa89, JV87, Joh84, JC87, JP87, Kad81, Kay82, Ket82, KM82, KM81, KKC82, KB82b, LSS85a, LC85, Lee87a, LS81, LB84, MM86, Mas86, MT84, MA80, MB84b, MSB85, MH88, Nai89, Nar83, NJ84, NDM84]. regression [NO85, Nor82, Nyq83, OH84, Oht87b, Oht88, Old87, Pag81, PRPR88, Pri82, Ram84, RD88, RP88, RM89, SBC81, SK82, SS83, SS86b, Sam89, SND84, SB89a, Sho82a, SG86a, Soo88, SZC85, Spa88b, ST86b, SC89, SL87, SL89b, Ste86a, Str84, SK87b, SA86, SdB89, Tah89, Ter88, TS88, Tra83b, Tre80, öT85, Utt82, Vin84, Wan87b, YAH89, Yu88, dR88]. regressions [CP82, KO82, SQ82]. Regressor [Ste86a, AAST88, HZ83]. regressors [KR86, Rus85]. regret [Yu86]. regular [AT81, KL84, RB88, StR82, SG84b]. regularity [Cha87]. rejection [AGR89, Sim87]. rejection-plus-mean [Sim87]. rejective [Bro81]. Rejoinder [IC80a, KC87b, KC87c, MSVS80, Smi80]. related [BA88, Cho80, Cho83b, Del81, Ebr85, KM81, KV80, MR86, Nag88, Ney80c, Nor82, SI81, Tan89]. relating [BDF^{+}84, Kap82a]. Relation [Con81, Ram84]. Relations [Bal88, BMA88, JSB89, BB89a, MBA88]. relationship [Abd89, BR85, HG80, Iso84, Ken81, MR86, Pat81, Pro86, Sza85, Tuc82, Yum85]. Relationships [BA88, McC85, MM88, NV84, ZF87]. Relative [SS87a, DK81, HJ83, SAN89a, SS85a, STJ88]. relatively [SN89a]. relevant [BE83, Sob87]. Reliability [MA84, Sin83, Cha83b, EN80a, God85, GR87c, Gup80, GG83, How85, HW89a, HW89b, JSB89, KPS88, KP81, Lai85, Obr82, Som89a, Som89b]. reliable [Tse88]. rely [Poc82]. Remainder [Lin86]. remark [Far87]. Remarks [DPS81, And89b, Kim88a, Tre88]. remote [Gla88]. remotely [Hai88, HGB89, Hey84, KP89b]. renewable [EG88]. Renewal [Kee80, BK82, Hug80, TW81, Xie89]. repairable [SD83, SD84a, SD88]. Reparameterisation [SlR87]. reparameterization [BSC80]. Repeated [HH89, SS81b, AP86a, AH84, HF80, Mag80, Maj88, PH80, PK81, Som88, Str88, Von83, Wan83]. repetitions [Sad82]. Replicate [Fen83, Cha87]. replicated [Yum85]. replicates [NJ89]. replication [Iso84, SND84]. representation [Fin84, GS81b, Ped86, Wan89a]. Representations [Six81, Yas81b, CARO84, DL88, YM85]. rerandomization [GR87a]. Resampling [SS88b]. Residents [Gap84]. Residual [AR87, GR87b, GdT87, Gup80, GG83, JP83b, JP83a, Kun84, LP88, LR87, Mus89, SK87a]. residuals [BBH82, Hor81, Loy87, NN82b]. resolution [Czi89c, Gup88a, Kle87a, Kuw83, KM84c, Kuw89, MCN81, Shi82c]. resolvable [yS89]. respect [De 89, DW87b, KP82, Mat85, PP87, Roi87, Tan82, TYM86]. Response [Bra84, KO82, AL80, AL82b, AD83, Bou82, Cru82, EKR86, Fra89, Gad83, Gun84, Gun85, Gup88b, Ham89, KP87, MK82, Mie80, MI82, Net85, RU89, Roc82, Smo83, Tam86]. response-limited [Roc82]. responses [BV89, GQ84]. rest [Bof88]. Restricted [MM83, MM89b, SW83b, Ber89, Dav89, DS89, Far84, KG89, Kha88b, KP80a, KY82, Mag87, Mat85, MC84, MH85, MD87, Shi80, SM82b, VS81]. restrictions [PD84]. result [De 89, FG88, Neu89]. results [Bal87d, BG83, Cha87, Del81, EHA84, GgT86, GM83a, GG83, HS86, Kok85, Krä85, KM87, Lem80, LP84, Mar82, MP82b, MT84, Rus80, SL82a, Som89b, SL87, Suz83, TT85, WC87, Xie89, YB83]. retrospective [Haw89b, YN86]. reuse [Kap83]. revealing [Sri87a]. revenues [Gap84]. Review [CS88b, Kle87b, SW86b, BE83, HK89, HL84, Iac83, Nag88, NJ84, Pat86, Pat89, RR88a]. reviews [PB83d, PB84c, PB86]. revisited [TP86]. Ridge [AM84b, Hsu81, LB84, MD87, PG84, BP84, Cha88b, CL83, FBV84, Far83, Far84, Fir87, Fla89, Gal80, Gap84, Gib84, GH84, Gue87, Hal84, HZ82, HG83, Kad81, LM84a, LT82, LC85, MT84, NO85, NM89, Obe84, Oht86, Pag81, PNS89, Pli87, SRW84, Sil86, Ste86a, SM82b, Tre88, Vin84]. ridge-like [SM82b]. ridge-type [NM89, Pli87]. ridge-type-estimator [Tre88]. right [Cha88a, KP89a, LP87d, LP85, MP82c, Nic89, Sir84, Wu82]. right-censored [LP85, MP82c, Sir84, Wu82]. right-censoring [Cha88a]. right-censorship [LP87d]. risk [FBV84, Hal84, HI89, HK83, IC80b, Kem86b, Kub88, Mag87, RMW86, RM88, Rob89, SS86c, Woo80]. risks [BDK89, BG80, Ebr86, GW84, Lee87c, RMW86, RM88]. road [Har82]. robots [Wan89a]. Robson [Sin87]. Robust [Atk82, BK85a, Dun82b, Gat86, Hsu80, LM84a, LC82, MM86, PB85, PRPR88, SM86b, TS81a, TS82, Tik82, TB84a, TB88c, TB89, BM81a, Dor89, Eub83, FB89, Gho82, GBS83, KW83, Loh84a, Loh84b, OSH88, PG89, Rie80, RM89, Sho82a, Sim84, Swa83, TB88a, TB82, TB84b, TB85, TB88b, TGB89, TL89b, Tlk83, Wan88]. Robustness [BK85b, PYO82, Sho84, Spr82, BE82, CC82, Hei88, JP81b, KM81, KM84c, LC82, Rie82, SYT89, SA82, SH84, SL85, Tan82, TYM86, Yas81b]. roc [MCR88]. Roeloffs [KJG88]. role [Bar88c, Kle88, Sch85a]. Rolle [RV82]. room [GS83]. root [De 81, Mas86, PF84]. roots [Li88, RV82]. rotability [NP88]. rotatability [KP89c, PK85a]. rotatable [KP86, KP87, Kos87, Muk87a]. rotation [Kat80]. row [ER80, Joe88, Sha88a, WK86]. row-column [ER80, Sha88a]. RSMR [RMW86]. Rubin [Sta86, Tal86]. rule [CZ87, Che88, ESP84, SS87a, Sch89, Sen86, TS88]. rules [BHL82, Dav89, DeM84, GS80, Hal84, HZ82, Lia88, SW83b]. Rulf [McC80]. runs [Smi84, Wha87]. Russian [Gla88].

S [Dav88b, Ars87]. Sample [Che84a, DN80, Fen83, SAH87, Ali87, Ang83b, Bal88, BCDE85, BP81, Beg81a, Ben89, Bri87, BG88a, CY86, CF89a, Cha88a, CG87, CB85, CDS82, Cha83a, CO86, CF88, CM82, Dav89, DS89, DS82, Eis81, ER83, FCP^{+}82, Fal86, FH81, FB82, GR81, Ger88, GM85b, Gov89, GS80, GMM84, HL87, Han80, Hin86, Hor85b, IC80b, Iso84, JP88a, JP88c, JK85, JKK86, Kaf82, Kap83, KH86, KT84, KS82a, KS82b, Kou81a, Kul87b, Kul87a, LW85, Lam86, LY89, Leu88a, LWA80, LM81, Mic85, MSW89, Nag85, ND81, Neu88, Off86, OH84, Oht86, PS87, PS85a, PYO82, Pra89c, Pra91, Rao86, Ris85, Roh87, RS88, Rol88, Rub86a, Rub86b, RC85, Sha84, SG85a, SS89, Sha86, SG88a, Shi80, aS89a, SG84a, SKC83, Sin86, Som85, Spa88a, Spr89, SDCB83]. sample [ST83, SG85b, SG86c, TB82, Tlk83, Tra84, Tse89, WB80, WW84, WT89, WAS88, WS82, Wit89b, Wri83, YSD85]. sampled [AL82a]. samples [Ang89c, BA88, Bar86, Bed87, Bur81, CK83b, CK87, Cho88, DS87, DF89, Dun82a, Ebr84, Fed87, GP82, Gil89, Hau89, Hor85b, KRS83, KK81, KS82b, KW86, KRB89, Loy89, Mar84c, Meh82, Mus89, PS80, Par81, PB87, SR88a, Sel89, SB89a, SJ87a, SJ89a, Sin85, STSM89, SM86b, SD85, SD87c, TB88a, TG89, TL89b, WAS88, Win82]. Sampling [CHLL84, Fre83, O’C87b, Abe80, Alb85, BS80a, BP87, CF89a, CG88, Cau84, CAR89, Cna85, CB81, CA86, DC84b, EN80a, FP89b, Fra89, Gab87, Gho81c, Gun85, GL89, HP87, HS82, HH89, HT83a, Kar88, Kat80, KB84, KPS83, LS80, LS83a, Lac84, O’C87a, Oht87b, PX86, RM88, SNR84, San87, Sch85a, Sel89, SM85, SV89, SM89, SA82, Sri83, SH84, WS84, Wei81a, Wri88b, XP89, YS84, ZS81]. Sargan [KS85]. SAS [PVB86]. satellite [DeP83, Hey84, HH84]. Satterthwaite [MSW89]. Scale [DG83, AL80, AL82b, Asr85, BM87a, Bed88, Blu87, Blu88, Bof84, BG84, BWMD83, CW88, CH87, CV89b, DB86, Dey89, DM89, EL82b, Gat89, Gil81, HP87, HI89, Hsi84, JKK86, JN89, KI88, KR84, Nai82, OAH89, PS80, Pra89a, Pra89b, SS85a, SHB85, SR88a, Sau82, aS89a, Sho84, Sir84, SSS87, SG86c]. scale-parameters [DB86]. scales [Bar81]. scenario [Kle86]. scheme [FG81, KP89a, YS84]. schemes [BS86, HT83a, JKR89, LC82, Luc85, PX86, YAH89]. Scholz [RD88]. Schur [Som89a, Yeh88]. sciences [Hut84]. scientific [Lai84, Sri87a]. scientists [Nar81]. score [Bro82, Tar88]. scores [BH84]. Screening [HGO85, Fed87, Gho87b, HJK89, JKR89, Kle87b, KJ82, KJG88, MB84a, Mor87, OP86, Odh86, Odh87, OM87, OP87, OP84, PO84a, PO84b, PO87, Smi84, Wat87]. screens [PH84]. Search [Cha89c, EZ81, Gho87a, Gho87b, Gup88a, PST87]. season [SG84a]. seasonal [Liu89]. Second [Koo85, DS85, Gov89, HF88, SKK85]. seemingly [ST86b]. segmentation [ZS89]. segmented [GD85]. segments [SG84a]. seismology [VJS81]. selected [AUSH83, CW88, Hsi81, SS85a, VS88, VSD88, VS89, Ven88]. Selecting [Ber81a, BB89b, Cha81, CP84, CS87a, Hay89, KP80b, LC85, LM81, Ris85, Tse88, AR83, AAST88, BA87, BK84, Bjø85, Che80, Dal81, GK81, HM81, HT83b, Kap82b, Lia88, PS87, RA80, RG88, SKC83, Tam86, Yan89b]. Selection [CS87b, JP88c, KKB87, AD80, BE80, BE83, BT82, BD87, Bjø81, BM86b, Bro81, BG88a, CF89a, Che86b, CZ87, Che88, Cho80, DA80, DT81, Eub83, Gia86, GS80, GL87, Hsu80, Hsu82, HH82, Hur87, Jen83, JP86, KL80, Kha88a, Kha88b, Kim88b, KP80a, McC82, McD87, Men81b, MS80a, NS86, Nor82, PKG87, RF84a, San87, San89b, SF85, Seh88, SZC85, SA81, Swa83, Tra83b, Wu89, YO86]. selector [Hus88]. Semi [Pil85, AP86b, AD83, BD86b, Kha83a, Lec85, MC89, Smi85, Sob87]. Semi- [Pil85]. semi-Bayesian [Lec85]. semi-definite [AP86b]. semi-definiteness [Smi85]. semi-Markov [BD86b, MC89]. semi-quantal [AD83]. semi-symmetric [Kha83a]. sensed [Hai88, HGB89, Hey84, KP89b]. sensing [Gla88]. sensitive [Hal84]. sensitivities [PG84]. Sensitivity [Tan88, TO89, IC80b, MN83a]. separation [PKN89]. sequence [GQ84, MR88, Sob89b, Tar86b]. sequences [KL80, Ng89b]. Sequential [AM84a, CQ82, GM80b, Gho81c, GN86a, Hor80, JN89, KL80, Muk88, Sha80, AM88, BJ88, BB80, BK84, CW80, DL84, EH83b, EZ81, GL81, Gho87b, HT81, HES84, HH82, JB86, Kar89a, Kee80, Kle86, Lai84, LDH84, Man80, Muk81, OW80, PG83, PS86, RA80, Rol88, RP89, SC80, Sen80, SW83b, Tah89, Wij81, Wu88]. sequentially [Bro81]. Serial [Aic83, Goo87, MC86, Tub80]. Series [Liu85, Pan83, ARP86, Abr81, AL82a, AM82, AV88, Byr88, Con81, EA89, Fin85, HS86, JP88a, JBF^{+}82, KN85, KN87, KC80, Kum80, LP81, Mic85, NS86, Ped86, Pou83, RR88b, RJ84, Rav81, Roc82, Som88, Som89b, TG85, TGG86, TG87, WNK80, WA81, Wit87a, Wol80, YS81, ZF87, ZF88]. set [KK83, Mat85]. sets [Bar89, CH86, Coh82, JK85, KN84, LaR88, Men81a, Sch86a, SFR84, Wij81, WT83]. settings [GG86]. several [Bla84, CP82, CHS86, DF89, GP82, JI84, KP80b, KW86, Rao86, RMW86, SG85a, Sin80a, Sri87c, Tan82, VI89, Wal87]. severity [SA89]. shape [PM89a, Sau82, Wei82]. shapes [Woo80]. Shapiro [Ang89c]. sharper [SS86c]. Shewhart [CAR89, GRY81]. shift [Far88, HN84, Hsi84, Pra89a, Pra89b, RP89, Sch83]. shifting [Haw89b]. short [Pap86]. show [Pur86]. Shrinkage [PS80, PM89a, AWS87, AG89a, Ali89b, EH87, Geo86, GN86a, Gok89, JC87, Kub88, Lem81, NM89, Rob89, SS85b, SS86b, SS86a, Str84, UV84, WSR84]. shuffling [ES88]. sib [BE82]. sib-pair [BE82]. sibships [BE82]. Sidak [RMW86]. Sidak-type [RMW86]. Sided [Ars87, AGR89, BK85a, Bha87, CO86, Cho88, FB89, KRS83, KB84, Sak87, SC80, Zac81b]. Sievers [RD88]. sign [LW85, RP89, YS81]. signal [Bha89, KG87]. Signals [Gup87a, Bha89]. signed [Die84b, Fro87, LW85, Mar84b]. significance [BF81, DK89, FGV89, KT85, Lec85, Pau88, RS81, SQ80, SS81b, SS88c]. significance-testing [DK89]. significant [Bof88, Coh80]. similarity [GS85b]. Simple [GHK85, eL82a, AG89b, BT86, BH84, dSBPR84, BS83, EN80a, EC83, GQ84, HT81, Ket82, Kon81, KF89, KM84c, LT82, Lin87b, Man80, MYT87, Mar88a, OSH88, Sha88b, Sin85, SK87b, WK82, Wri88b, YS84, ZS81]. simpler [KP89c, SN89a]. simplex [MM83]. simplified [AP86b, MP82a]. simply [SW89]. Simpson [CG88]. Simulataneous [Moo88]. Simulation [GM83a, Far83, Gad83, Pag81, Wha87]. Simultaneous [CV89a, Chi81, EK88, Fou81, KB83, Kay89, KM82, Khu80, KS88, Pep89, Wan87b, BF85, Cho88, Dey83a, GR87a, Gra87, Kab83, Krä85, KS84, Mar82, Phi80, Rao86, RMW86, TP85, Tsu85, ZLMK82]. simultaneously [GV89b]. Single [CL89, SM85, Bal88, Cha87, FGV89, Jos88, KKB87, Rad85, Sch85b, SYT89]. Single-stage [CL89]. singly [BT83, DeP83, DC84a, Sir84]. singular [MRS84, Mat85, Pei86b, Rao81b]. singularity [KP89c]. situation [FD85]. situations [Gat86, KK85]. sixteen [Joh85]. Size [WM87, BG88a, Che84a, Eis81, Gab87, JL89, Par81, Rao81a, Rao86, Roh87, RS88, Sol83, SM86b, Tse89, WB80, WS82, Wri88b]. sized [SD87c]. sizes [CO86, CM82, GR81, OP87, PO87, Ris85, SAH87]. skew [Sch87]. Skewness [Mac82, SS88a, Gro86, MF83, Tar86a]. slippage [Sri87b]. slope [Pat81]. slopes [RD88]. Small [IC80b, Che84a, DS89, GR81, GG81, GMM84, KRS83, LW85, Loy89, OH84, Oht86, SD85]. smallest [Cha81, GHO86, Sat88, Tam86]. Smirnov [Bri87, Cal87, FGP88, JK85, KH86]. Smooth [BT83, Bar86, RB88, Gho89a, Swa87]. smoothed [Wan89b]. smoothing [Gu87, RM89]. smoothness [TSM86]. software [Ber85]. Sojourn [PM83]. solution [BD86c, CG81, DT81, EC83, ES84, McC82, NvdM86, Nic89]. solutions [AG89b, BYD85, SD86]. solving [GdT87]. Some [Abe80, And89b, AG89b, BM85a, Bar80, Ber82, BT81b, BYD85, BH86, BF85, CD87, Cha87, Den89, DM87, EA89, Fra80, Gib84, Gor87, Har86, Hos87b, HP80, HH82, JB84, KK84, Kar89b, Kim88a, Kle83, Kot89b, KJ83, KJ90, Kum81, KM87, Lem80, Mar82, Nag80, Nai82, Nor89, O’C87b, Off86, Oht87b, OSH88, PM88, PD84, Pra89c, Pra91, RF84a, SKSF83, Sen81, SW83b, Som89b, Str88, SG86c, yS89, Suz83, Tre88, Vog87, WC87, Wit88, Woo88, Xie89, AH84, BA88, BS83, Cha80, CI80, Con88, Del81, Gal82, Gal80, GG83, HH86a, HL84, HS86, Hsu82, Isa86, Jai81, KT84, KS82a, KS82b, KK85, KR88, Kuw83, LP87a, ML88, MP82a, Mar88b, McC85, McD87, MK82, MG87, NJ89, NM89, PX86, Ped86, Pro86, Rad82, Ram84, Rob89]. some [Rub86b, SS85a, Sch85b, Sil86, Sta86, Tlk83, TG86, WW84, Wri88b, XP89]. something [Kal84]. source [Bre82]. Sow [CC82]. space [Ber89, KD83, NN82b, Ste86a]. space-time [NN82b]. Spacing [EC83, Bof88, Eub83, Fro89]. spacings [Aly87, FKML89, Meh82, SHB85]. spanned [Tan88]. sparse [Mic85]. Spatial [Mar84a, Hai88, HGB89, HH89, Mar89b, Wan89a]. spatially [GS85a]. Spearman [Fra88]. special [Fin82, KP85, MSVS80, Ney80b, Six81, Smi80]. Specialized [Ano83x, Ano83y]. species [EN80a, OW86, TT89a]. Specific [BMA88]. specification [Han80, HRR88, Mit81, Moo88, SU86]. specified [Mee88, Nel86, Obr82, Sri87d]. spectral [Coo85, Pou83, WK82]. sphere [Woo88]. sphericity [CS83, CS88a, FKN83, GN88, KN83c]. spline [Che87a, Gu87, RM89, Shi87, VW83]. splines [Eub84]. spread [Hor85b, KK86a]. spring [Swa80]. spurious [JI84, MYT87]. SQC [Ber85]. Square [AR87, Ala87, BB83, Gim87, BWD86, CARO84, De 81, ES87, Fin84, Gel87, Gru85, HH86a, HG80, Hos86, Joh85, Kha89a, KW86, Lis88, McC85, Mit81, MS81c, Pri82, öT85, WH88]. squared [BA87, Cha83b, Coo87, Dav85b, Kem86b, Lee87c, LP87d, MS86b, Ros89, Swa88, Tre80, Val86]. squares [BC88, BHW81, BU84, Bru80, CM86b, CH83, Dre85, EK82, ES84, ES87, Far84, Fri81, GgT86, GS83, Gie83, KW84, KKOV87, Krä85, KD89, LSS85a, MC84, MS81b, PM89b, Phi80, SS86a, SS87a, SFR84, SB84b, SJ89b, Ver88b]. squares-fix [BHW81]. stability [FBV84]. stable [EkBL88, Kou81b, KB82b]. Stage [O’C87b, AWS87, AG89a, BF86, CF89a, CL89, CH88, CO86, CHS86, Gok89, HS83, HM81, KH81, Mar80, MB84a, MS80a, MM89a, Mor87, O’C87a, Odh87, OP87, Oht87b, OP84, PO84a, PO84b, PO87, Phi80, Seh88, Spu82, Ver88a, WSR84]. stand [Fu80]. Standard [FP86, Kan82, EH83b, Far86, Far87, GHK85, Spr89, ST86b]. standardized [And83, RM88]. standards [Nat85, You85]. standby [Lai85]. state [Ang86, SW84b, YS82b, YS82a]. states [Kee80]. stationarity [Fin85]. stationary [Byr88, Iso89, MR88, SI87]. Statistic [Gim87, AV88, And86, BW82, CM84, DW87a, DW89a, DL88, Fre83, GN88, Hoc81, ID80, IV89, Kha89a, Koz89, Lee85, Leu88a, LR82, Mik85, Muk81, NNQ84, RF84b, SR88c, Sha83, She86, Sin87, SH89, TB88b, Wha87]. Statistical [BM85c, Cla81, Dav81, Gho83b, KK83, Nat85, Net85, PB83d, PB84c, PB86, SD88, CS88b, Elp89, FME89, Gla88, Hog82, HL84, HP80, Kot89b, tL84, MH85, MH86, OA87, Rod85, Sri89, Vog87, WT89, Yan89a, YO86]. statisticians [Llo81, Nar81]. Statistics [Ano83a, Ano83z, BMA88, FH87, Har86, VJS81, AAY86, AUSH83, AC85, Ang83b, BM85a, Bal86, BM86a, Bal87a, Bal87b, BA88, Bal88, BB89a, BL84, Bar88a, Bar88c, Bro82, CVV82, CW88, Chh84, CK83b, Cho88, Cor84, DMH81, Dav85a, Dav88a, DS81, DK83, DL88, DR86, Gho81c, Gro82, HR81, Hus88, Kab84, Kai88, KPS88, Kum83, Llo81, MT82, MBA88, MP82a, Mar81a, Mar84b, McC85, MS81a, MDH81, MH88, Nag85, Nag80, Nag88, OM80, Par82, Pre80, Pro86, RG84, Roh87, RS88, SS85a, Sat88, Sch87, Sch89, Sen81, Sen84, Sha89, zS88, SS81b, SS88c, SC86, Tak88, TS82, Tik82, Tik88b, TB88c, WW84, WYF87, Wit88, YM86, Yam87, YM89, ZG80, BM85c, Lyb85, MSVS80, Smi80, Ney80b]. status [BG88b]. steady [Ang86]. Stein [AB86, BH86, CH86, HZ82, Hos88, KT86, Lam86, LlHK82, Li82, SS87a, SS86c, Sen84, Sen86, TS88]. Stein-like [HZ82]. Stein-rule [SS87a, Sen86, TS88]. Stein-type [KT86]. Step [OM87, BS89, JP87, Kab84, Lai89, LP87b, MS88]. step-down [MS88]. step-stress [BS89]. Step-wise [OM87]. stepwise [BA87, Che87a, Sch86b]. stereology [Cla81]. Sterrett [HJK89, JKR89]. stimulation [Gab80]. Stirling [CS88b, HF88]. Stochastic [Anb84, EN81, FF86, KPS88, Llo81, PM82, ARP86, BHL82, BC88, BE88, Bon88, DS85, KR86, Kay82, Lee88, Obr82, PM83, PK81, Rus85, UV84]. stochastically [AJK89, BA84]. stochasticity [Pra83]. stock [MS81c]. stocks [FBV84]. Stopping [DeM84, Zac81b]. straight [Tan82, Vog86]. strata [YS84]. strategies [ES88]. Stratification [WA84, AS84a, Gil89, JKW89, Leu88b, YS84]. stratified [CG88, DF89, Dav89, Hau89, Leu88a, TL89b, WS84]. stratospheric [JBF^{+}82]. stratum [Wri83]. strength [Cho83a, YKY85]. strengthening [Gia86]. stress [BS89]. string [BS88, MH85, MH86]. strings [ZS89]. Stronger [HH80]. Structural [Gup82, MSB85, CF86, DG81, EKR86, Iso84, KW83, Lee87b, Lee88, MB85, SA85, Tsu85]. structure [AH86a, AH86b, Bas89, BLW87, Byr88, DA80, MC86, Pei88, SD87a, SM86a]. structured [KK83]. structures [Six81]. Student [BG84]. Studentization [GH85]. Studentized [Leu89]. studies [Bro87, Cla82, Eri83, Gup80, LPW82, Mar85, MC87a, Ney80c, Per82, RMW86, RM88, Ray83, YN86]. Study [Kha89b, BT86, Ber86, CM86a, Chr88, Fic84, FKML88, Fri81, KO84, LC82, MB84a, NM89, Pap89, Rao86, Wan87a]. subgroup [Dav89]. Subject [Ano83z, Cau83, Cau84, DB86, Gal82, LY89, SB89a, Spu83a]. Subjective [Bev84]. subpopulation [HB86]. subsample [SG85a]. subsamples [SIU85, SG85b]. subsampling [Bed87, BG85, VB80]. Subset [BD87, BM86b, Cho80, Bjø81, Bro81, Che80, Hsu80, JP86, Kim88b, KP80a, San87, YO86]. subsets [Bha86, Khu80]. subspace [Tan88]. subspaces [El 81]. successive [Kha89a]. Sufficiency [Sha83]. sufficient [IS86, Roj86, Sch89, Sha83]. suggested [LR87]. suggestions [Vog87]. sum [JRF80, KD89, Luc85, Pat85, SW86a]. Sums [Aly87, Ete87, MS81b, SJ89b]. superadditivity [KR89a]. Superiority [Ter82, Ter86, Ter88, Tre83, KG89, Mit81]. supersaturated [RS84b]. supplementary [Dan88]. support [Sch85c, YSTM88]. surface [Bra84, KP87, MK82]. surfaces [Min80]. sursets [Sob87]. survey [Ber85, ENP84, Gla88, HS83, Pra91, Sch84b, SW83a, Tam86, TB82, Tlk83]. surveys [CHLL84, Fei84, GG81, HH84, Iac83, Off86, Pra89c, ST83]. Survival [Cna85, SL84, Arb85, BE84, CEA85, CT89, Duf83, Ebr85, FW86, FH81, FH84, GHK85, GM83a, Ger88, Gho89a, Gho89b, GW84, LP88, ME86, MS81d, PW81, Sch84b, TJZ88, Wei81b, Wol81, Yan89b, Zeh85]. survivorship [Arb85]. swamping [Ben89]. Sweden [Lyb85]. switching [BU84, Ram84]. symbol [YKY85]. symbols [yS89]. symmetric [CS84a, IV89, Kaf82, Kap86, Kha83a, Kik83, KLV85, KL86, YM89]. symmetrical [Gup88c, SD87b]. Symmetrically [SH82]. symmetry [FG88, HR81, KHB^{+}88, KR84, Koz83, Ros89, Sil88, Tom87]. symposium [Vin84]. synthesis [TS88]. system [BG83, ENP84, EG88, Isa86, IS87, Kot89a, Lai85, PM83, SD83, SD84a, SD88, Som89b, Tsu85, Vog87, You85]. Systematic [Bla84, FP89b, Buo89, FBV84, Wei81a]. systematically [AL82a]. systems [BHW81, BYD85, BD86c, DH89, ENP84, Krä85, PR84, Som88, Som89a].

t [Kha89a]. t’- [Kha89a]. T.V. [Pur86]. table [CQ82, Gro84, Hab85, Hab89, Kha83a, NS83, Ram87, SG84c, SS86d, Tuc82, WNK80]. table/urn [Ram87]. Tables [AG83, DG83, Ima87, Mic83, Alb87, AL80, AL82b, Ber81a, Bjø82, BWD86, Bra84, Bre82, Cau83, Cau84, Chr86, Chu83, Chu84, CGO85, Coh80, DW87a, DW89a, Dav86, GR81, Gen88, Güi82, Hos86, Joe88, Mar89a, MP86, Mic85, PVB86, Shi88, TO85, Wal87]. Taguchi [Czi89a, Czi89c]. Tail [Ars87, BK86, OSH88, Sam85, SL80a, Som84, SS88c, Wei80]. tail-ordering [BK86]. tailed [SYT89]. tails [CS81]. tampered [BS89]. target [Elp83, UV84]. tau [AC84, DPS81, FR83, SR88b]. tax [Gap84]. Taylor [Wit87a]. technique [KP84, Smi84, WAS88]. techniques [Cla81, Dav81, GP82, Hal84, HL84, IC80b, Kot89b, MN89, MB82, MI82, MD87, Nai82, Sim84]. teeth [CVDS85]. telephone [PG84]. term [Arb85, Can84, Ket82, LDH84]. Terminating [BK82]. terms [SK87a]. Test [Ala87, BB83, KSC81, KN84, PL82, PL86, Sin83, Aly87, Ang83b, Ang86, Ang89c, AGR89, AJK89, BDK89, BA87, BCDE85, Bar81, Bar84, BB80, BF86, BH84, BD86a, BS89, BWD86, dSBPR84, BS88, Bri87, Bro82, BM87b, Bur81, Cal87, CS83, CDS82, CW80, CK83b, CH88, CM82, Dav87, Die89, FKN83, Fei88, Feu87, FGP88, Fro87, GL81, GQ84, GMM84, GV89b, HL87, HK87, Hei88, HL80a, Hor81, Hos87a, Joe88, Kab86, KH86, Ken81, Kle83, KK85, KT85, KR89b, KS84, KN83a, KN83b, KN83c, Kum83, Lam86, LSS85a, LSS85b, Lec85, Lee87c, Lis88, Maj88, MP86, MS86a, MS81a, MN83a, MCR88, Moo88, MSW89, NN82b, NS84, Oht87a, Oht87b, Oht88, PU86, PS85a, PG82, PYO82, Pro86, Pro88, RS81, Ray85, Rol88, Rub86a, Sak87, SQ81, SK82]. test [SS83, SS87a, SR88b, SLH83, Sch82, Sch84a, SG82, SYT89, SR88c, Sha84, SG88a, SJ89a, zS88, Shi88, Shi85, Sho82b, Sin85, SK86, Smi82, SS88c, SA82, SC86, Sri87d, SH89, SG86c, SdB89, TS81a, TYM86, Utt82, Wec85, WJT89, Wil86, Wol80, YS81, Zei89]. testability [MS85]. Testable [Pei86b, dR89]. tested [GK88b, HSC80]. testimation [AG89a, Gok89]. testimator [AWS87, WSR84]. testimators [PM89a]. Testing [Ahm82, Ang86, Bar83, BE84, Ber84, BD86b, BM81a, DS85, EH83a, EZ89, Far86, Fla85, GP82, Gat89, GWM89, Koc85, KR89a, Kul85, Lin89, Loy89, MC87a, McC86, MB84b, Nay88, NJ84, Oht88, Per81, PKN89, Rig89, SND84, aS82, SK87a, TS81b, TB84b, TB85, TP85, WS89, Wol81, AH83, BD86a, BM87b, BW82, CEA85, DK83, DK89, Far87, Fri87, GN87, GN88, GV89b, Haw89b, HJK89, JKR89, Jos88, JN89, KI88, Kap85, Kar89b, KC84b, Kha84, Khu80, Kle88, Kum83, LaR88, LC87, Man80, Mar80, MT83a, MS87, Miy87, MS80b, NCLN82, NM88, PST87, Pau88, PH89, Pep88, Sha80, SJ89a, STSM89, SG88b, SL82a, SC86, TS82, Tik82, TGB89, WR87]. Tests [BM82, BJ83, FB82, Goo80, HR81, JP83b, JP83a, Kou81b, Koz83, LSS85b, McC87, Mos89, Pat88, PL83, PR85, Ros89, Sch86c, SCG85, Tar86b, Wij81, AB87, Aic83, AH86a, BT83, Bar86, Ben89, Ber82, Ber86, BE82, BS83, BS84, BF81, BW88, BWMD83, Cha88a, Cha80, CS88a, Che84b, Che83b, Chm81, Dah88, Del81, DS86b, Die84b, DR86, DM89, EA89, Eis81, El 81, EP85, Eri83, EK85, FY86, Fer87, FH81, Fou81, FD85, FGV89, GR87a, GHK85, Goo89, Gov86, Gro84, HH86b, HT81, HS83, HN84, Hor85a, Hos86, HL80b, Hsi84, HF80, Jan88, KR86, KO84, KT84, Ken81, KR84, KM81, Kim88a, Kle83, KB81, KB82a, KK85, KK86b, KW86, LW85, LMP88, Man89, Mar85, MF83, Mar88b, Mar84c, MY85, Meh82, MS81c, Muk81]. tests [NV84, NJ89, Neu88, O.’89, OSG81, Per82, PD84, PH84, Pra89a, Pra89b, PS86, PRPR88, Ram84, RS81, RLN86, Ray82, RB88, RP89, Rub86b, RS89, SK82, SS83, SS85a, SHB85, SC80, SD84b, Sch84b, Sch85b, Sch86b, SW86a, Sen87, SG85a, Shi80, aS89a, SOL81, SW89, SSS87, SS81b, Som85, Spu82, Spu83a, Spu84, SL85, SG85b, Tar86a, Tar88, Tik88b, TO85, TT86, Vad83, VI89, Ver88a, VJ88, Vos88, WCK85, WL82, Wu82]. th [Sat88]. their [ARP86, Anb84, Bha86, CS88b, DL84, Fin85, GS85a, Gro82, ISS82, Kai88, KM88, Kha83c, LR87, Moo89, OSH88, PRR86, PK85b, RG84, Sau82, SJM86, TC81, Yas81a]. themes [Dav85a]. theorem [AP86b, Ang89b, Bev84, CVV82, Dun81, El 82, Hug80, Kap82a, LP89, Poc82, Wri88b, Lin86, RV82]. theorems [RF84a]. theoretic [Bjø81, DC84b, HG86, Men81b, Pal88, Sam89, Ven80]. theoretical [Mit88, SL87]. Theory [Ano83a, Ano83z, Liu85, Ang89a, BLW87, Cha83a, Ebr86, ENP84, Gal88, HS83, Iso84, Kee80, KKC82, Kum81, Loy87, Mik85, Nis88, RF84a, SS87a, Sam89, SK88, SSKK88, Sic86, Som89a, SW84b, Sri89, Suz83, TW81, Vad83]. there [Fla85, KB81, Mil82a]. thin [VW83, Wah81]. third [CO86]. third-stage [CO86]. those [MB85]. Three [HM81, KM84b, OP87, Chu83, CW82, Güi82, HH86b, Joh82, JC84, Kan82, KP86, LK83, NS85, Odh87, Pau85, Rau87, SM86b, Ver88a, WCK85]. three-parameter [CW82, Joh82, JC84, Kan82, LK83, Pau85]. Three-stage [OP87, Odh87, Ver88a]. three-way [Chu83, Güi82]. threshold [Wei81c, Wei82]. Thurstone [Sad82]. Thurstone-type [Sad82]. ties [AC85]. Tighter [MP82a]. Tiku [O.’89]. Time [JBF^{+}82, Liu85, Sin83, Abr81, AL82a, AM82, AV88, AM84a, And89b, BD86a, Blu87, Byr88, CEA85, DF89, Ebr85, Ebr86, Fin85, GM84, HS86, Hsi84, JP88a, KN85, KN87, Kle83, NN82b, NS86, PM83, PS85a, Pei86a, Pei88, Pou83, Pra89a, Pra89b, Pra83, Rav81, Roc82, Sen80, WNK80, WA81, Wol80, Xek83a, YS81, ZF87, ZF88]. Time-Censored [Sin83]. time-lagged [ZF87]. time-sequential [Sen80]. time-series [Roc82]. times [Kay82, ME86]. titration [Chr88]. toe [Isa86]. Tolerance [Pat86, Sch80, BK85a, FB89, Wha87]. Tong [McC82]. tool [Khu82]. tools [Rod85]. top [Ima87]. top-down [Ima87]. topologies [Whi89a]. total [AJG82, BG85, Fu80, Kag88, Kle83, Swa80, WK86]. totals [Hab85, Hab89, YS84]. tournaments [HH80]. tow [KN84, Mar80]. tow-stage [Mar80]. toxicological [Ng89a]. trace [Bet87, Gib84, Mat80, MP82b]. tracking [FME89]. traditional [HF80]. Training [God85, Leu88a]. transect [LS83a]. transfer [KT85, LH82]. Transform [MR87, Kle83, Llo87b, ND81, OA87]. Transformation [Kun83, BFR80, CI80, LK81, LO89, Mos83b, Sar85, Wan87c]. transformations [Fin85, GE88, Kap89, SL84, SL83, Wan87b]. transformed [Hun89, Shi83a, Shi83b]. transforms [KS82a, KS82b]. transition [Kay82, YS82b, YS82a]. treated [SR88c]. Treatment [WT83, Ach84, Alb85, Can84, FY86, SB89a, Sin80a, Wal87, WK86, Yan89b]. treatments [Ber81a, CP84, DW87a, DW89a, EH83b, HT81, Maj88, Mat81, WT83]. trees [CT89]. Trend [YB83, BBH82, PH84, aS82, SW89, Stu88, Tar86a, Whi89b].

Trend-free [YB83, Whi89b]. trended [Krä85]. trends [Bla84, FP89b, Har85a, LP81, mN83b]. Trenkler [Neu89]. trial [FGV89, LDH84]. trials [Ach84, Anb84, AH83, Can84, CLl86, DL84, Gra87, GQ84, KB80, Lai84, PH80, PK81, TA83]. trigonometric [KS87]. Trimmed [DD85b, CVV82, DD85a]. trimming [BCL80, KRS83]. Triple [HP87, HT83a]. triplets [Wec85]. triserial [Ken81]. true [Ban80a]. Truncated [RP89, BL84, BS80a, Blu85, Blu87, Blu88, CO84, CF89b, DS81, DeP83, DC84a, HD89b, HD89a, Mar84c, MD89, Tit83, Tra85, Wal84a, WB80]. truncation [Beg80, Beg81a, Beg81b, Dey83b, DC84a, Iwa86b, KRB89, Sel89]. Tukey [FKML88, FKML89, Gro86, JB84, Mac82]. turning [Cha89a]. Two [Ahs87, AD80, Bal87b, Chr88, CHS86, Dav89, DS86b, Fre84, Gho80, HA83, KB84, LP87b, LS83b, Mor87, Ney80c, O’C87b, OP84, PO84b, Spu82, SG85b, AWS87, AG89a, AG83, AP86b, AGR89, AJG82, BDK89, BM89, BK85a, BA88, BF86, Bed89a, Bed89b, Beg81a, Bjø82, BS80b, Bra84, BH80, BG85, BW82, Cha83b, CT83, CQ82, Che84a, CV89a, CH88, CH87, Chr86, Chu82, Chu84, CGO85, Cna85, Coh80, Con89, DMH81, Dal81, Dav85a, DK83, DR82, ESP84, EN80b, FV89, FH81, FV87, Fon89, FB82, Fro89, FB89, GM88, GS85a, Gil81, GC80, Gok89, Gro84, GL87, HL87, HT81, HS83, Hos86, How85, Huf85, IB89, Joh82, Joh85, KH81, Kaf82, KSR88, KH86, KT84, Kee80, KK83, Kha87a, Kha83b, Kim85, Kor85, Kot89a, Lai85, Lai89, Lam86, LMP88]. two [LC87, Lem80, LS83c, Lis84, LP89, MY85, MB84a, MS80a, MDH81, MM88, MM89a, MW85, MSW89, MAT82, NN82a, ND81, Nel87, Neu88, O’C87a, Oht87a, Oht87b, PH80, Pat83, PO84a, PO87, PKN89, Phi80, PYO82, RS80, Rao81a, RD88, Rav89, Rol88, RS84b, Rub86a, Rub86b, Sak87, SQ82, SC80, SD88, Sch86c, Scl80, Seh88, Sel89, SG88a, SJ89a, aS89b, SA85, SM82a, Sir84, Sir86, SL84, SAH87, Spu83a, SA82, Sri83, SH84, SC86, SSK89, SG86c, yS89, SB89b, Sza85, TS81b, TGB89, TT86, Ver88a, WSR84, Wal80, WF88, WG81, WR87, WW84, WT89, WK81, Yan88, YS81, YG82, YS82b, ZG80, dWM85]. two- [Rol88]. two-and [Ver88a]. two-dimensional [Scl80]. two-factor [Bed89a, Bed89b, aS89b, WG81]. two-fold [BG85, MM89a]. two-group [LC87]. two-level [RS84b]. two-location [Kot89a]. two-location/two-product [Kot89a]. two-parameter [Cha83b, CV89a, EN80b, NN82a, Yan88]. two-parent [Kha87a]. two-period [PH80, Pat83, WF88]. Two-Phase [HA83, Fre84]. two-piece [Joh82, Kim85]. two-product [Kot89a]. Two-sample [Dav89, FB82, Kaf82, KH86, MSW89, ND81, Neu88, PYO82, Rub86a, Rub86b, SG88a, SG86c, WW84]. two-sided [SC80]. Two-Stage [O’C87b, Mor87, PO84b, Spu82, BF86, KH81, MB84a, MS80a, O’C87a, Oht87b, Seh88]. two-state [YS82b]. Two-step [LP87b]. two-unit [SD88]. two-way [AG83, AJG82, Bjø82, Bra84, BH80, CQ82, Chr86, Chu82, Chu84, CGO85, Coh80, Gro84, GL87, KK83, Lem80, MAT82, RS80, WK81]. type [CVV82, CB82, Dun86, Elp89, ER83, FG88, FG81, Fri81, FGP88, HY86, HM81, KT86, KSJ86, KB82b, KRB89, LaR88, Lam86, Lee87c, NM89, Pli87, PC82, PH85, Pra89c, Pra91, RS84a, RMW86, Rus80, Sad82, Sen84, Sin85, Spu83a, SSK89, TT83, TB88c, Tre88, GM88, Gup88b, KM87, Meh82, PM89a, Ped85a, PH84, SR88a, Sel89, Sir84, SH89, TG89, Wei80]. type- [ER83]. Type-1 [Ped85a]. types [NS85, PVB86, Rao81a].

U [Haw89b, HP80]. U.S. [LM84a, Nat85]. Uhlenbeck [Wit84]. umbrella [zS88, Shi88]. Umvu [Iwa86a, Iwa86b, SI85, Shi83b, ISS82, Kun83, LO89, Shi83a]. unaffected [Bla84]. unbalanced [BKZ88, BH80, BG85, BMC86, BG88b, Fra80, Hem80, HSC80, MM89a, PRPR88, Sea80, Sea88, SJ89b, TT88, TTB88, Ver88a]. Unbiased [Far85, KRB89, PL83, San89b, Sel89, SI87, Ste89b, Wit87b, BM87a, CQ82, EkBL88, Hos87b, IS86, KS81a, Kim88a, KPS83, KR88, KC80, Kum80, Moo89, Off86, Oht86, Rao87, Sam88, SH82, SI81, SIU85]. Unbiasedness [KH81, DM80]. unblocked [Goo87]. uncertainty [MB85, OS86]. uncommon [Kul85]. unconstrained [Hsu85]. Undiminished [LR87]. unequal [Bed87, BG88a, Cha88a, CM82, Gho83b, Kan88a, Mar80, OP87, OP84, PO87, Ris85, Sad82, TS81a, TGB89, Tse89, YO86]. uneven [Fu80]. uneven-aged [Fu80]. unified [JPD82, Sam88, TG86]. Uniform [Spu83b, Wef87, Bar80, Bil89, CP84, Cho80, Cho83b, DB86, Gor87, JP88c, VSD88]. uniformity [MM89a]. Uniformly [SI81, WK81, Ped85b, Sha88a]. unifying [CPW88]. unimodal [MP82c, YSTM88]. Unimodality [Win82, AM81]. Union [BS83, BS84, NS84]. Union-intersection [BS83, BS84, NS84]. uniqueness [RV82]. unit [Off86, Riv88, SD88]. units [Bed87, Gro82]. Univariate [Bha82, Bru80, LM84b, Mil82a, Ray82, SM82a, TYM86]. Universal [KT86]. unknown [BF86, Bed88, Beg81a, BWMD83, DC84a, EZ81, GN86a, HM81, Hsi81, Hsi84, Iwa86b, Koz83, Kul85, LM86, Man80, McC86, Pra89a, Pra89b, TZ85, Wei82]. unobserved [Par81, Yu88]. unordered [MP86]. unrelated [ST86b]. unreplicated [LY89]. unrestricted [Mag87]. unweighted [FR88, Kap82a]. updated [Fra87]. upon [LaR84]. Upper [DJD89, CV89a, Cho83a, Sam85, SS88c]. urn [Cha86a, EN80a, IM88, Ram87]. urns [Sob89b]. usages [Rav89]. Use [Ach84, HH84, KP84, McC80, Pat83, Rao82, SKC83, Swa82, Anb84, Ang89c, BU84, CD87, Hal84, Hen88, Kap83, Kap89, Mas86, PRR86, PD84, PKG87, RD88, SJM86, Sin82, WS84]. Used [LS83b, BE84, dSBPR84, DK83, EZ89, EN81, KC84b, KY82, Kum83, Sch86b, SL82a, ST86a]. Useful [SKG80, Ber82]. Usefulness [TB82]. User [Dou86, Ber85]. user-friendly [Ber85]. Ushakov [Ger88]. Using [DeP83, Hey84, Hut84, Kaf82, Kha87b, Kun83, Lin86, RG88, SQ81, Sin86, Alb88, AUSH83, BHL82, Ber81a, Bri87, CS88a, Cha84, DMH81, Dan88, Din87a, Din87b, EA89, FW86, GP82, GR83, GK88b, Gho87b, Gue87, HGO85, Ham89, Hor85a, Jel87, JP88a, Kay82, KG87, KC87a, KC87b, KC87c, Kot89b, KR82, Lee81, LP87a, LK81, LO89, Liu89, LM81, MN89, MZ88, Mor87, MC89, NM88, PK81, Pep89, Phi80, Pro84, Rad84, Sen87, SG84a, SQ80, SSK89, Swa86b, Tit83, Wal80, Wha87, Wil88b, Woo80]. utilisation [Mad85]. utility [Ban80a, Hun89].

validation [Hur87, Tut89]. validatory [Loy87, RM89]. validity [Gio88]. value [Ang89a, Cha81, Die84a, DR82, Gal88, HY86, Jel87, KR89a, Koz83, Zac81b]. valued [Isa81, Wan86, Zac81b]. values [Bri87, CM82, HZ83, Miy87, Nag88, Nis88, Rat89, Sch86b]. variability [ML88, NS83, Sen87]. Variable [CAR89, BB89a, CO84, Kan88b, KD83, MP85, Men81b, Nor82, Oht87b, Ray83, SKC83, SZC85, SA81, SA86, TT89b, Wri88b, Zac81b]. Variables [AK83, AAST88, BHW81, BM81b, Buo89, BA84, CC82, CT83, CS84a, CB81, Czi88, Czi89b, Den89, DJD89, DR82, Ete84, Gal82, Gil81, Gio88, Gov89, Gun83, Gup87b, HH86a, Hug80, Hun89, KW84, Kel86, Kem86a, Ket82, KW83, Kha89b, KKB87, Kon81, Mas86, Mat83, Mil82b, Mos89, PC82, RG88, Sat88, SD87a, Shi83a, Shi83b, SA85, SKC83, Tar86b, Tra83b, Yat86, YG82, YSTM88, Yu88, Zac81b]. Variance [AM87, Fen83, Joa83, Lac84, RP88, AK86a, And86, AJG82, BM82, BP81, BHL82, Bre82, BH80, BG85, BG88b, CF88, DN80, DF89, DK89, EkBL88, Fra80, GG86, GgT86, GD88, GRY81, Gie83, GS85b, GV89b, Han80, Hoc81, Hsi81, Kap82a, KPS83, Kor85, KR88, KSC81, Kul85, KC80, Kum80, Kuw89, LMP88, Lec85, Lee81, Lee85, Lem80, Man80, Mar80, Miy87, MR86, MAT82, Oht88, Pal88, PP88a, PS81, PK86, Pen85a, Pen85b, PVB86, PG82, PYO82, PRPR88, RB83a, Rao82, RM88, Ray85, SKK85, Sam88, SD87a, SS89, Sha86, SI81, Shi85, Sin86, Sob89a, ST86a, TB85, Ver88a, Ver88b, WEA80, WG81, WK82, Yu85, ZS81, Zac81a]. variance-covariance [SD87a, TB85]. Variances [Kun83, SB82b, BG88a, BW82, CM82, Del81, DJD89, GN86a, HL87, Hay89, HM81, ISS82, KB83, KT84, Kel86, KR89b, Mar80, McC87, MM88, Moo89, Oht87a, Ped83, Sak88, Sha88b, SJ89b, TT88, TS81a, Tik82, TB84b, TL89b, WG81]. variant [Sta86]. Variants [Gal88, Tal86]. Variate [MR87, JG85, Sun88]. variate- [JG85]. variates [AO82, Bal87a, Bar81, Con89, DS80, IS85, KN84, Mos87, SB82a, Tub86, Wu88]. variation [Ebr84, HI89, Iwa89, JS82, Lee81, Sam84, Sha80, Sin86, YI88]. Variations [Dav85a]. varieties [Joh85]. various [Kul88, vdMN87]. Vector [CD89, Ali89b, Isa81, Kha89b]. Vectors [Ete87, KW86, Lin87c, Riv88, Sch80, TS82, TGB89]. verdicts [Sol83]. version [RV82, SG82]. versions [DS81]. versus [BE80, Bof88, DW89a, FP84, KR84, MB85, MSW89, Off86, Von83]. very [Che84a]. via [tFAH84, BP83, Cha86a, Loh84a]. view [DeM84, Gho83a, Hsu81, Tam86]. VII [KM84c]. violations [PYO82]. vol [MSVS80, Ney80b, Smi80]. volume [Lac84, Ped86]. volumes [Ano83a, Ano83z]. voluntary [Gun85]. vote [CF86]. vs [BD86b, DeM84, HN84].

Wald [BS88, Wha87]. wall [Wan88]. Wallis [RF84b, Shi85]. Wang [Fir87].

Waring [Xek86]. Watson [HR81, Koz89, She86, Zei89]. way [Adk86, AG83, AJG82, Bjø82, Bra84, BH80, BMC86, CQ82, Chr86, Chu82, Chu83, Chu84, CGO85, Coh80, DW87a, DW89a, DK89, Gro84, Güi82, GL87, KC84a, Kap82a, KK83, KP89c, Lem80, Mar82, MAT82, PRPR88, RS80, RS84a, aS82, SJ89b, TB84b, TB85, WK82, WK81, Yeh88]. Weak [Mus89, Mit81]. weather [BS80c, MSVS80, Ney80b, Smi80]. Weibull [Wan09, Arc82, AR81, BE80, BE83, BT83, BWMD83, CC88, CW82, DC84a, EK82, EN80b, FKML89, Hsu82, KP80a, KP80b, KP81, KP82, KB81, KB82a, MT82, MD89, MR87, Nig89, PU86, PM89a, SG88b, Sir84, Sir86, Sri89, TS81b, Tse88]. weighing [Kag88, Swa80, Swa82]. weight [Kag88, Swa80]. Weighted [CH83, Fei84, Kep86, KA87b, SSS87, Aly87, CG81, Dre85, FY86, Fer87, GgT86, HN84, JSB89, KKOV87, MR83, MS86b, PRR86, SQ81, SQ82, Sie83, SQ80, Tom87, WCK85]. weighted- [HN84]. weights [Dor88, Dor89, LM86]. Weiss [ML88]. well [Vog86]. wet [Llo81]. where [GM85a]. whereat [Sea88]. wherefrom [Sea88]. whereto [Sea88]. Whether [Ber84, BE84, Kum83]. which [BK84, Che80, McC82, Sob87, Tan89]. white [And88]. Whither [Sen86]. Whitney [DS81, Hoc81]. Whittemore [Dav86]. width [Bis80, CHS86, RH88, SKN84, SvW82]. Wilcoxon [DS81, GM84, Hoc81, Sch84a]. Wilk [Ang89c]. Wilks [NNQ84]. Williams [KM87]. win [HH80]. winner [CLl86]. wise [BM81b, OM87]. Wishart [HLN88, Mat80, MP82b, TG83]. within [Che83b, Koo85]. Without [SB82b, Ket82, MB85, SND84, SL83, SC89]. Wolfe [Dav88b]. Wolfowitz [Wha87]. word [Sic86]. workloads [DF89]. works [Swa86a]. world [Har85a]. worths [Bas89].

XB [Rus80].

Yntema [Moo89]. Yule [Xek83b].

Zero [HL85, KK80, Ken81, Vos88]. zone [Gia86].

References

Albert:1981:PLD

[AA81] A. Albert and J. A. Anderson. Probit and logistic discriminant functions. Communications in Statistics: Theory and Methods, 10 (7):641–657, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1988:MSL

[AAST88] M. A. Ali, M. S. Abu-Salih, and S. K. Titi. A method of selecting the levels of regressor variables in lognormal regression model.

Communications in Statistics: Theory and Methods, 17(6):1741– 1761, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Abood:1986:IDG

[AAY86] A. M. Al-Abood and D. H. Young. Improved deviance goodness of fit statistics for a gamma regression model. Communications in Statistics: Theory and Methods, 15(6):1865–1874, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angers:1986:SEB

[AB86] Jean-François Angers and James O. Berger. The Stein effect and Bayesian analysis: a reexamination. Communications in Statistics: Theory and Methods, 15(7):2005–2023, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:1987:CPT

[AB87] Lalit K. Aggarwal and Steve M. Bajgier. Comparative performance of tests of normality in detecting mixtures of parallel regression lines. Communications in Statistics: Theory and Methods, 16(9): 2541–2563, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Achcar:1989:CHA

[AB89] Jorge Alberto Achcar and Heleno Bolfarine. Constant hazard against a change-point alternative: a Bayesian approach with censored data. Communications in Statistics: Theory and Methods, 18 (10):3801–3819, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdullah:1989:ELF

[Abd89] Mokhtar B. Abdullah. On W -estimators of a linear functional relationship. Communications in Statistics: Theory and Methods, 18(1):287–314, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abel:1980:SEC

[Abe80] Volker Abel. Some essentially complete classes for Bayesian multistage sampling inspection. Communications in Statistics: Theory and Methods, 9(14):1483–1489, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abraham:1981:MOT

[Abr81] B. Abraham. Missing observations in time series. Communications in Statistics: Theory and Methods, 10(16):1643–1653, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abeyasekera:1982:EPM

[AC82] S. Abeyasekera and D. Collett. On the estimation of the parameters of the von Mises distribution. Communications in Statistics: Theory and Methods, 11(18):2083–2090, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvo:1984:CAD

[AC84] Mayer Alvo and Paul Cabilio. A comparison of approximations to the distribution of average Kendall tau. Communications in Statistics: Theory and Methods, 13(25):3191–3216, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvo:1985:ARC

[AC85] Mayer Alvo and Paul Cabilio. Average rank correlation statistics in the presence of ties. Communications in Statistics: Theory and Methods, 14(9):2095–2108, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Achcar:1984:UBA

[Ach84] Jorge Alberto Achcar. Use of Bayesian analysis to design of clinical trials with one treatment. Communications in Statistics: Theory and Methods, 13(14):1693–1707, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amer:1980:TNG

[AD80] Paul D. Amer and Edward J. Dudewicz. Two new goals for selection based on proportions. Communications in Statistics: Theory and Methods, 9(14):1461–1472, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ashford:1983:DRA

[AD83] J. R. Ashford and Hilary J. Doyle. Dosage-response analysis for a semi-quantal response derived from an aqueous biological assay. Communications in Statistics: Theory and Methods, 12(23):2713– 2722, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1986:OWA

[Adk86] S. R. Adke. One way ANOVA for dependent observations. Communications in Statistics: Theory and Methods, 15(5):1515–1528, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1983:MRP

[AG83] James H. Albert and Arjun K. Gupta. Models for reflecting prior beliefs of association in two-way contingency tables. Communications in Statistics: Theory and Methods, 12(11):1241–1259, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1989:NSF

[AG89a] S. R. Adke and D. V. Gokhale. A note on shrinkage factors in two stage testimation. Communications in Statistics: Theory and Methods, 18(2):633–637, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asiribo:1989:SSA

[AG89b] Osebekwin Asiribo and John Gurland. Some simple approximate solutions to the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 18(4):1201–1216, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arabatzis:1989:CIE

[AGR89] Alexandras A. Arabatzis, Timothy G. Gregoire, and Marion R. Reynolds, Jr. Conditional interval estimation of the mean following rejection of a two sided test. Communications in Statistics: Theory and Methods, 18(12):4359–4373, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1983:NPT

[AH83] Sharon Anderson and Walter W. Hauck. A new procedure for testing equivalence in comparative bioavailability and other clinical trials. Communications in Statistics: Theory and Methods, 12(23): 2663–2692, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alberton:1984:ADM

[AH84] Yael Alberton and Yosef Hochberg. Approximations for the distribution of a maximal pairwise t in some repeated measure designs. Communications in Statistics: Theory and Methods, 13(22): 2847–2854, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrade:1986:MEL

[AH86a] Dalton F. Andrade and Ronald W. Helms. ML estimation and LR tests for the multivariate normal distribution with general linear model mean and linear-structure covariance matrix: k-population complete-data case. Communications in Statistics: Theory and Methods, 15(1):89–107, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrade:1986:MEM

[AH86b] Dalton F. Andrade and Ronald W. Helms. ML estimation for the multivariate normal distribution with general linear model mean and linear-structure covariance matrix; one-population, completedata case. Communications in Statistics: Theory and Methods, 15 (6):1927–1955, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1988:REG

[AH88] Dorothy A. Anderson and John P. Hinde. Random effects in generalized linear models and the EM algorithm. Communications in Statistics: Theory and Methods, 17(11):3847–3856, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmad:1982:TIT

[Ahm82] Hanzoor Ahmad. Testing independence through conditional and marginal data. Communications in Statistics: Theory and Methods, 11(3):291–296, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:1987:TCE

[Ahs87] M. Ahsanullah. Two characterizations of the exponential distribution. Communications in Statistics: Theory and Methods, 16 (2):375–381, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aickin:1983:STM

[Aic83] Mikel Aickin. Serial tests of multiple hypotheses. Communications in Statistics: Theory and Methods, 12(13):1535–1551, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arteaga:1982:CIP

[AJG82] Carmen Arteaga, S. Jeyaratnam, and Franklin A. Graybill. Confidence intervals for proportions of total variance in the two-way cross component of variance model. Communications in Statistics: Theory and Methods, 11(15):1643–1658, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aras:1989:RTS

[AJK89] Girish Aras, S. Rao Jammalamadaka, and Mohammad Kafai. A rank test for stochastically ordered alternatives. Communications in Statistics: Theory and Methods, 18(6):2263–2277, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agresti:1983:AMM

[AK83] Alan Agresti and Abbas Kezouh. Association models for multidimensional cross-classifications of ordinal variables. Communications in Statistics: Theory and Methods, 12(11):1261–1276, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agarwal:1986:CVB

[AK86a] G. G. Agarwal and Sanjay Kumar. On a class of variance balanced incomplete block designs. Communications in Statistics: Theory and Methods, 15(5):1529–1533, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amoh:1986:EMA

[AK86b] R. K. Amoh and K. Kocherlakota. Errors of misclassification associated with the inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 15(2):589–612, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1980:CMC

[AL80] Robert J. Anderson and J. Richard Landis. CATANOVA for multidimensional contingency tables: Nominal-scale response. Communications in Statistics: Theory and Methods, 9(11):1191–1206, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abraham:1982:FES

[AL82a] Bovas Abraham and Johannes Ledolter. Forecast efficiency of systematically sampled time series. Communications in Statistics:

Theory and Methods, 11(24):2857–2868, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1982:CMC

[AL82b] Robert J. Anderson and J. Richard Landis. CATANOVA for multidimensional contingency tables: ordinal-scale response. Communications in Statistics: Theory and Methods, 11(3):257–270, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alalouf:1987:CST

[Ala87] I. S. Alalouf. The chi square test with both margins fixed. Communications in Statistics: Theory and Methods, 16(1):29–43, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1981:PBE

[Alb81] James Albert. Pseudo-Bayes estimation of multinomial proportions. Communications in Statistics: Theory and Methods, 10(16): 1587–1611, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1985:BTN

[Alb85] James H. Albert. A Bayesian treatment of nonresponse when sampling from a dichotomous population. Communications in Statistics: Theory and Methods, 14(4):947–961, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1987:EBE

[Alb87] James H. Albert. Empirical Bayes estimation in contingency tables. Communications in Statistics: Theory and Methods, 16(8):2459– 2485, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1988:EPM

[Alb88] James H. Albert. Estimation of a Poisson mean using discrete mixtures of conjugate priors. Communications in Statistics: Theory and Methods, 17(1):245–261, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1987:DSM

[Ali87] Mukhter M. Ali. Distribution of sample mean when observations are correlated. Communications in Statistics: Theory and Methods, 16(5):1321–1343, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1989:AAM

[Ali89a] M. A. Ali. An alternative approach to minimize model inadequacy. Communications in Statistics: Theory and Methods, 18(8):2801– 2811, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1989:CSE

[Ali89b] M. A. Ali. On a class of shrinkage estimators of vector of regression coefficients the. Communications in Statistics: Theory and Methods, 18(12):4491–4500, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alwan:1986:CQC

[Alw86] Layth C. Alwan. Cusum quality control-multivariate approach. Communications in Statistics: Theory and Methods, 15(12):3531– 3543, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aly:1987:SWS

[Aly87] Emad-Eldin A. A. Aly. Sums of weighted spacings and a test for the logistic distribution. Communications in Statistics: Theory and Methods, 16(4):1075–1086, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abouammoh:1981:NUD

[AM81] A. M. Abouammoh and A. F. Mashhour. A note on the unimodality of discrete distributions. Communications in Statistics: Theory and Methods, 10(13):1345–1354, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abraham:1982:TSM

[AM82] B. Abraham and Ch E. Minder. A time series model with random coefficients. Communications in Statistics: Theory and Methods, 11(12):1381–1391, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1984:SEC

[AM84a] S. R. Adke and S. M. Manjunath. Sequential estimation for continuous time finite Markov processes. Communications in Statistics:

Theory and Methods, 13(9):1089–1106, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anup:1984:RED

[AM84b] Anup K. Anup and G. S. Maddala. Ridge estimators for distributed lag models. Communications in Statistics: Theory and Methods, 13(2):217–225, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1986:EBE

[AM86] Khursheed Alam and Amitava Mitra. An empirical Bayes estimate of multinomial probabilities. Communications in Statistics: Theory and Methods, 15(10):3103–3127, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmad:1987:VCE

[AM87] R. Ahmad and S. M. Mostafa. Variance components estimation for the balanced random effects model under mixed prior distributions. Communications in Statistics: Theory and Methods, 16(9):2605– 2621, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1988:ESE

[AM88] S. R. Adke and E. S. Murphree. Efficient sequential estimation for compound Poisson processes. Communications in Statistics: Theory and Methods, 17(2):443–460, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amato:1988:GKM

[Ama88] David A. Amato. A generalized Kaplan–Meier estimator for heterogeneous populations. Communications in Statistics: Theory and Methods, 17(1):263–286, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aminzadeh:1989:PIE

[Ami89] Mostafa S. Aminzadeh. Prediction intervals in exponential families. Communications in Statistics: Theory and Methods, 18(1):343– 353, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amoh:1984:EPM

[Amo84] R. K. Amoh. Estimation of parameters in mixtures of inverse Gaussian distributions. Communications in Statistics: Theory and Methods, 13(8):1031–1043, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anbar:1984:SAM

[Anb84] Dan Anbar. Stochastic approximation methods and their use in bioassay and Phase I clinical trials. Communications in Statistics: Theory and Methods, 13(19):2451–2467, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1983:CMG

[And83] Keaven Anderson. Convergence of the moment generating function of standardized quantiles. Communications in Statistics: Theory and Methods, 12(21):2445–2454, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1986:IVF

[And86] W. J. Anderson. An inequality for the variance of the first order statistic. Communications in Statistics: Theory and Methods, 15 (4):1153–1162, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andel:1988:PEW

[And88] J. Andel. On AR(1) processes with exponential white noise. Communications in Statistics: Theory and Methods, 17(5):1481–1495, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andel:1989:NNP

[And89a] J. Andel. Nonlinear nonnegative AR(1) processes. Communications in Statistics: Theory and Methods, 18(11):4029–4037, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1989:SRD

[And89b] W. J. Anderson. Some remarks on the duality of continuous time Markov chains. Communications in Statistics: Theory and Methods, 18(12):4533–4538, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1983:NQE

[Ang83a] J. E. Angus. Normal quantile estimation based on minimizing the error in predicted distribution functions. Communications in Statistics: Theory and Methods, 12(12):1345–1358, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1983:ADC

[Ang83b] John E. Angus. On the asymptotic distribution of Cramér–von Mises one-sample test statistics under an alternative. Communications in Statistics: Theory and Methods, 12(21):2477–2482, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1986:TSS

[Ang86] John E. Angus. Testing the steady state availability based on data from a fixed length test. Communications in Statistics: Theory and Methods, 15(4):1221–1234, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1989:AEV

[Ang89a] John E. Angus. Application of extreme value theory to the BEC family of distributions. Communications in Statistics: Theory and Methods, 18(12):4413–4419, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1989:NCL

[Ang89b] John E. Angus. A note on the central limit theorem for the bootstrap mean. Communications in Statistics: Theory and Methods, 18(5):1979–1982, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1989:NUS

[Ang89c] John E. Angus. A note on the use of the Shapiro–Wilk test of exponentiality for complete samples. Communications in Statistics: Theory and Methods, 18(5):1819–1830, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBa

[Ano80a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(1):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBb

[Ano80b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(2):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBc

[Ano80c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(3):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBd

[Ano80d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(4):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBe

[Ano80e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(5):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBf

[Ano80f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(6):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBg

[Ano80g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(7):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBh

[Ano80h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(8):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBi

[Ano80i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(9):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBj

[Ano80j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(10):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBk

[Ano80k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(11):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBl

[Ano80l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(12):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBm

[Ano80m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(13):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBn

[Ano80n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(14):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBo

[Ano80o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(15):ebi–ebii, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBp

[Ano80p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(16):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBq

[Ano80q] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(17):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1980:EBr

[Ano80r] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 9(18):ebi, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBa

[Ano81a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(1):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBb

[Ano81b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(2):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBc

[Ano81c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(3):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBd

[Ano81d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(4):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBe

[Ano81e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(5):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBf

[Ano81f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(6):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBg

[Ano81g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(7):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBh

[Ano81h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(8):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBi

[Ano81i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(9):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBj

[Ano81j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(10):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBk

[Ano81k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(11):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBl

[Ano81l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(12):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBm

[Ano81m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(13):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBn

[Ano81n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(14):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBo

[Ano81o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(15):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBp

[Ano81p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(16):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBq

[Ano81q] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(17):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBr

[Ano81r] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(18):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBs

[Ano81s] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(19):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBt

[Ano81t] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(20):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBu

[Ano81u] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(21):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBv

[Ano81v] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(22):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1981:EBw

[Ano81w] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 10(23):ebi, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBa

[Ano82a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(1):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBb

[Ano82b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(2):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBc

[Ano82c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(3):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBd

[Ano82d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(4):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBe

[Ano82e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(5):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBf

[Ano82f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(6):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBg

[Ano82g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(7):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBh

[Ano82h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(8):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBi

[Ano82i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(9):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBj

[Ano82j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(10):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBk

[Ano82k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(11):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBl

[Ano82l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(12):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBm

[Ano82m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(13):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBn

[Ano82n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(14):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBo

[Ano82o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(15):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBp

[Ano82p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(16):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBq

[Ano82q] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(17):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBr

[Ano82r] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(18):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBs

[Ano82s] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(19):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBt

[Ano82t] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(20):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBu

[Ano82u] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(21):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBv

[Ano82v] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(22):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBw

[Ano82w] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(23):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBx

[Ano82x] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(24):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1982:EBy

[Ano82y] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 11(25):ebi, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:AIC

[Ano83a] Anonymous. Author index for Communications in Statistics: Theory and Methods, volumes 1–12. Communications in Statistics: Theory and Methods, 12(25):2969–2985, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBa

[Ano83b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(1):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBb

[Ano83c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(2):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBc

[Ano83d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(3):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBd

[Ano83e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(4):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBe

[Ano83f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(5):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBf

[Ano83g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(6):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBg

[Ano83h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(7):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBh

[Ano83i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(9):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBi

[Ano83j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(10):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBj

[Ano83k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(11):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBk

[Ano83l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(12):ebi–ebii, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBl

[Ano83m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(13):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBm

[Ano83n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(15):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBn

[Ano83o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(16):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBo

[Ano83p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(17):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBp

[Ano83q] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(18):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBq

[Ano83r] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(20):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBr

[Ano83s] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(21):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBs

[Ano83t] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(22):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBt

[Ano83u] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(23):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBu

[Ano83v] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(24):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBv

[Ano83w] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 12(25):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBMa

[Ano83x] Anonymous. Editorial Board members specialized area. Communications in Statistics: Theory and Methods, 12(8):–, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:EBMb

[Ano83y] Anonymous. Editorial Board members specialized areas. Communications in Statistics: Theory and Methods, 12(14):1–11, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1983:SIC

[Ano83z] Anonymous. Subject index for Communications in Statistics: Theory and Methods, volumes 1–12. Communications in Statistics: Theory and Methods, 12(25):2987–3033, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBa

[Ano84a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(1):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBb

[Ano84b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(2):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBc

[Ano84c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(3):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBd

[Ano84d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(4):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBe

[Ano84e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(5):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBf

[Ano84f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(6):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBg

[Ano84g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(7):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBh

[Ano84h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(8):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBi

[Ano84i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(9):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBj

[Ano84j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(10):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBk

[Ano84k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(11):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBl

[Ano84l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(12):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBm

[Ano84m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(13):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBn

[Ano84n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(14):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBo

[Ano84o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(15):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBp

[Ano84p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(16):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBq

[Ano84q] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(17):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBr

[Ano84r] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(18):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBs

[Ano84s] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(19):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBt

[Ano84t] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(20):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBu

[Ano84u] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(21):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBv

[Ano84v] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(22):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBw

[Ano84w] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(23):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBx

[Ano84x] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(24):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1984:EBy

[Ano84y] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 13(25):ebi, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBa

[Ano85a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(1):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBb

[Ano85b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(2):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBc

[Ano85c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(3):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBd

[Ano85d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(4):i, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBe

[Ano85e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(7):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBf

[Ano85f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(8):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBg

[Ano85g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(9):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBh

[Ano85h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(10):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBi

[Ano85i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(11):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:EBj

[Ano85j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 14(12):ebi, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1985:RR

[Ano85k] Anonymous. Referee recognition. Communications in Statistics: Theory and Methods, 14(6):1507–1509, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:E

[Ano86a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 15(9):7–8, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBa

[Ano86b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(1):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBb

[Ano86c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(2):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBc

[Ano86d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(3):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBd

[Ano86e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(4):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBe

[Ano86f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(5):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBf

[Ano86g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(6):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBg

[Ano86h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(7):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBh

[Ano86i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(8):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBi

[Ano86j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(9):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBj

[Ano86k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(10):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBk

[Ano86l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(11):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1986:EBl

[Ano86m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 15(12):ebi, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(1):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(2):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(3):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBd

[Ano87d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(4):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBe

[Ano87e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(5):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBf

[Ano87f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(6):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBg

[Ano87g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(7):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBh

[Ano87h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(8):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBi

[Ano87i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(9):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBj

[Ano87j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(10):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBk

[Ano87k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(11):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:EBl

[Ano87l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 16(12):ebi, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1987:E

[Ano87m] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 16(11):3427–3428, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:E

[Ano88a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 17(8):vii, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBa

[Ano88b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(3):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBb

[Ano88c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(4):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBc

[Ano88d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(5):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBd

[Ano88e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(6):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBe

[Ano88f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(7):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBf

[Ano88g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(8):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBg

[Ano88h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(9):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBh

[Ano88i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(11):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1988:EBi

[Ano88j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 17(12):ebi, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:E

[Ano89a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 18(7):vii, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBa

[Ano89b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(1):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBb

[Ano89c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(2):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBc

[Ano89d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(3):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBd

[Ano89e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(4):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBe

[Ano89f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(5):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBf

[Ano89g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(6):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBg

[Ano89h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(7):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBh

[Ano89i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(8):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBi

[Ano89j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(9):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBj

[Ano89k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(10):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBk

[Ano89l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(11):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1989:EBl

[Ano89m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 18(12):ebi, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1982:DLC

[AO82] Mir M. Ali and M. Obaidullah. Distribution of linear combination of exponential variates. Communications in Statistics: Theory and Methods, 11(13):1453–1463, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Osh:1989:IFR

[AOA89] M. A. Al-Osh and A. A. Alzaid. On the increasing failure rate criteria. Communications in Statistics: Theory and Methods, 18 (12):4421–4436, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agresti:1986:CMR

[AP86a] Alan Agresti and Jane Pendergast. Comparing mean ranks for repeated measures data. Communications in Statistics: Theory and Methods, 15(5):1417–1433, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrews:1986:SPT

[AP86b] Donald W. K. Andrews and Peter C. B. Phillips. A simplified proof of a theorem on the difference of the Moore–Penrose inverses of two positive semi-definite matrices. Communications in Statistics: Theory and Methods, 15(10):2973–2975, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ashour:1981:BPC

[AR81] S. K. Ashour and D. R. Rashwan. Bayesian predictions for compound Weibull model. Communications in Statistics: Theory and Methods, 10(16):1613–1624, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1983:SFP

[AR83] Khursheed Alam and M. Haseeb Rizvi. On selecting from k finite populations the population with the largest α-quantile. Communications in Statistics: Theory and Methods, 12(3):355–362, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:1987:DRA

[AR87] Sung K. Ahn and Gregory C. Reinsel. Distribution of residual autocovariances and prediction mean square error properties the multivariate reduce rank autoregressive model. Communications in Statistics: Theory and Methods, 16(1):61–78, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arbutiski:1985:FMS

[Arb85] Thomas Arbutiski. A family of multiplicative survival models incorporating a long-term survivorship parameter c as a function of covariates. Communications in Statistics: Theory and Methods, 14 (7):1627–1642, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Archer:1982:MLE

[Arc82] N. P. Archer. Maximam likelihood estimation with Weibull models when the data are grouped. Communications in Statistics: Theory and Methods, 11(2):199–207, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:1988:BEM

[Arn88] Barry C. Arnold. Bounds on the expected maximum. Communications in Statistics: Theory and Methods, 17(7):2135–2150, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdul-Razak:1986:PSD

[ARP86] Rafid S. Abdul-Razak and Ganapati P. Patil. Power series distributions and their conjugates in stochastic modeling and Bayesian inference. Communications in Statistics: Theory and Methods, 15 (3):623–641, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alzaid:1986:CDP

[ARS86] Abdulhamid A. Alzaid, C. Radhakrishna Rao, and D. N. Shanbhag. Characterization of discrete probability distributions by partial independence. Communications in Statistics: Theory and Methods, 15(3):643–656, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arsham:1987:MOS

[Ars87] Hossein Arsham. A modified one-sided K–S confidence region narrower in one tail. Communications in Statistics: Theory and Methods, 16(1):17–28, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiraishi:1982:THA

[aS82] Taka aki Shiraishi. Testing homogeneity against trend based on rank in one-way layout. Communications in Statistics: Theory and Methods, 11(11):1255–1268, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:1984:OSP

[AS84a] Vijay Aggarwal and Ravindra Singh. On optimum stratification with proportional allocation for a class of Pareto distributions. Communications in Statistics: Theory and Methods, 13(24):3107– 3116, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1984:BLC

[AS84b] Khursheed Alam and K. M. Lal Saxena. Bounded length confidence interval for a common mean. Communications in Statistics: Theory and Methods, 13(17):2133–2142, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiraishi:1989:MMS

[aS89a] Taka aki Shiraishi. Multivariate multi-sample rank tests for location-scale alternatives. Communications in Statistics: Theory and Methods, 18(1):85–105, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiraishi:1989:ECR

[aS89b] Taka aki Shiraishi. R-estimators and confidence regions for main effects in a two-factor MANOVA. Communications in Statistics: Theory and Methods, 18(1):261–276, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asrabadi:1985:ECI

[Asr85] Badiollah R. Asrabadi. The exact confidence interval for the scale parameter and the MVUE of the Laplace distribution. Communications in Statistics: Theory and Methods, 14(3):713–733, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1981:CCR

[AT81] Mir M. Ali and J. G. C. Templeton. Common content of a regular hypersimplex and a concentric hypersphere. Communications in Statistics: Theory and Methods, 10(13):1297–1304, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abrahams:1984:NCB

[AT84] Julia Abrahams and John B. Thomas. A note on the characterization of bivariate densities by conditional densities. Communications in Statistics: Theory and Methods, 13(3):395–400, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atkinson:1982:RDR

[Atk82] Anthony C. Atkinson. Robust and diagnostic regression analyses. Communications in Statistics: Theory and Methods, 11(22):2559– 2571, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atwood:1980:C

[Atw80] Corwin I. Atwood. Comments. Communications in Statistics: Theory and Methods, 9(17):1847–1851, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1983:EQU

[AUSH83] M. Masoom Ali, Dale Umbach, A. K. Md Ehsanes Saleh, and Khatab M. Hassanein. Estimating quantiles using optimally selected order statistics. Communications in Statistics: Theory and Methods, 12(19):2261–2271, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abraham:1988:SCM

[AV88] B. Abraham and K. Vijayan. A statistic to check model adequacy in time series. Communications in Statistics: Theory and Methods, 17(12):4271–4278, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1987:TSS

[AWS87] S. R. Adke, V. B. Waikar, and F. J. Schurmann. A two stage shrinkage testimator for the mean of an exponential distribution. Communications in Statistics: Theory and Methods, 16(6):1821– 1834, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Butler:1984:COS

[BA84] J. S. Butler and Kathryn H. Anderson. Continuous outcomes stochastically mismeasured as discrete random variables: a maximum likelihood approach. Communications in Statistics: Theory and Methods, 13(16):2005–2012, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bajgier:1987:SAS

[BA87] Steve M. Bajgier and Lalit K. Aggarwal. A stepwise algorithm for selecting category boundaries for the chi-squared goodness-of-fit test. Communications in Statistics: Theory and Methods, 16 (7):2061–2081, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1988:RAM

[BA88] N. Balakrishnan and R. S. Ambagaspitiya. Relationships among moments of order statistics in samples from two related outlier models and some applications. Communications in Statistics: Theory and Methods, 17(7):2327–2341, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1986:OSD

[Bal86] N. Balakrishnan. Order statistics from discrete distributions. Communications in Statistics: Theory and Methods, 15(3):657–675, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1987:NMO

[Bal87a] N. Balakrishnan. A note on moments of order statistics from exchangeable variates. Communications in Statistics: Theory and Methods, 16(3):855–861, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1987:TII

[Bal87b] N. Balakrishnan. Two identities involving order statistics in the presence of an outlier. Communications in Statistics: Theory and Methods, 16(8):2385–2389, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balanda:1987:KCC

[Bal87c] Kevin P. Balanda. Kurtosis comparisons of the Cauchy and double exponential distributions. Communications in Statistics: Theory and Methods, 16(2):579–592, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baldessari:1987:EDD

[Bal87d] Bruno A. Baldessari. Experimental designs with dependent observations: New results and comments on previous results. Communications in Statistics: Theory and Methods, 16(3):785–803, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1988:RIM

[Bal88] N. Balakrishnan. Relations and identities for the moments of order statistics from a sample containing a single outlier. Communications in Statistics: Theory and Methods, 17(7):2173–2190, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balasooriya:1989:DOE

[Bal89] Uditha Balasooriya. Detection of outliers in the exponential distribution based on prediction. Communications in Statistics: Theory and Methods, 18(2):711–720, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1980:CCU

[Ban80a] K. S. Banerjee. A comparison of the constant-utility true index and the true index obtained through the factorial approach. Communications in Statistics: Theory and Methods, 9(1):87–105, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1980:CA

[Ban80b] K. S. Banerjee. Corrections and amendments. Communications in Statistics: Theory and Methods, 9(1):115, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnett:1980:SBU

[Bar80] Vic Barnett. Some bivariate uniform distributions. Communications in Statistics: Theory and Methods, 9(4):453–461, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barker:1981:AOT

[Bar81] Lawrence Barker. Allocating observations in a test for the ratio of scales of independent gamma variates. Communications in Statistics: Theory and Methods, 10(23):2469–2473, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baras:1983:TRA

[Bar83] Mario Baras. Testing randomness against ordered alternatives in a multinomial experiment with grouped frequencies. Communications in Statistics: Theory and Methods, 12(22):2575–2580, 1983.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bartels:1984:RNT

[Bar84] Robert Bartels. The rank von Neumann test as a test for autocorrelation in regression models. Communications in Statistics: Theory and Methods, 13(20):2495–2502, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bargal:1986:STF

[Bar86] Adel I. Bargal. Smooth tests of fit for censored gamma samples. Communications in Statistics: Theory and Methods, 15(2):537– 549, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnett:1988:OOS

[Bar88a] Vic Barnett. Outliers and order statistics. Communications in Statistics: Theory and Methods, 17(7):2109–2118, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barry:1988:AIN

[Bar88b] Daniel Barry. Asymptotic IMSE for a nonparametric Bayesian regression estimate. Communications in Statistics: Theory and Methods, 17(10):3277–3293, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barter:1988:HRO

[Bar88c] H. Leon Barter. History and role of order statistics. Communications in Statistics: Theory and Methods, 17(7):2091–2107, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barr:1989:MMD

[Bar89] Aiala Barr. A multivariate model for discrete data sets. Communications in Statistics: Theory and Methods, 18(2):445–459, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [KJ90].

Basak:1989:EMC

[Bas89] Indrani Basak. Estimation of the multi-criteria worths of the alternatives in a hierarchical structure of comparisons. Communications in Statistics: Theory and Methods, 18(10):3719–3738, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baruah:1980:MPS

[BB80] H. K. Baruah and G. P. Bhattacharjee. A modification of the partial sequential probability ratio test. Communications in Statistics: Theory and Methods, 9(6):639–648, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bemis:1983:BEC

[BB83] Kerry G. Bemis and Vasant P. Bhapkar. Ban estimation for chisquare test criteria in categorical data. Communications in Statistics: Theory and Methods, 12(11):1211–1223, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bapat:1989:IRR

[BB89a] R. B. Bapat and M. I. Beg. Identities and recurrence relations for order statistics corresponding to nonidentically distributed variable. Communications in Statistics: Theory and Methods, 18(5): 1993–2004, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandari:1989:SBC

[BB89b] Subir Kumar Bhandari and Arup Bose. Selecting the t-best cells in a multinomial distribution. Communications in Statistics: Theory and Methods, 18(9):3313–3326, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bartels:1982:PTM

[BBH82] Robert Bartels, Graham Bornholt, and Kevin Hanslow. The polynomial trend model with autocorrelated residuals. Communications in Statistics: Theory and Methods, 11(12):1393–1402, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhat:1988:LSE

[BC88] B. R. Bhat and K. Suresh Chandra. Least squares estimation of the parameters of a stochastic difference equation with polynomial regression component. Communications in Statistics: Theory and Methods, 17(10):3427–3446, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barbour:1985:NRT

[BCDE85] A. D. Barbour, D. I. Cartwright, J. B. Donnelly, and G. K. Eagleson. A new rank test for the k-sample problem. Communications in Statistics: Theory and Methods, 14(6):1471–1484, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boullion:1985:CEF

[BCK85] Thomas L. Boullion, Grace C. Cascio, and Jerome P. Keating. Comparison of estimators of the fraction defective in the normal distribution. Communications in Statistics: Theory and Methods, 14(7):1511–1529, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broffitt:1980:EHT

[BCL80] B. Broffitt, W. R. Clarke, and P. A. Lachenbruch. The effect of Huberizing and trimming on the quadratic-discriminant function. Communications in Statistics: Theory and Methods, 9(1):13–25, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhat:1986:LRT

[BD86a] B. R. Bhat and Sunita K. Deshpande. Likelihood ratio test for testing order of continuous time finite Markov chains. Communications in Statistics: Theory and Methods, 15(6):1751–1771, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhat:1986:TMP

[BD86b] B. R. Bhat and Sunita K. Deshpande. Testing for Markov process vs semi-Markov process. Communications in Statistics: Theory and Methods, 15(8):2375–2382, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1986:ABS

[BD86c] Lyle Broemeling and Joaquin Diaz. An adaptive Bayesian solution to the control problem for linear control systems. Communications in Statistics: Theory and Methods, 15(6):1803–1817, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bechhofer:1987:SSN

[BD87] Robert E. Bechhofer and Charles W. Dunnett. Subset selection for normal means in multi-factor experiments. Communications in Statistics: Theory and Methods, 16(8):2277–2286, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burk:1984:CRQ

[BDF^{+}84] F. Burk, L. Dion, D. Fridshal, E. Langford, C. O’Cinneide, and T. Parsons. In a conjecture relating χ^{2} and F quantiles. Communications in Statistics: Theory and Methods, 13(5):661–670, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagai:1989:DFT

[BDK89] Isha Bagai, Jayant V. Deshpandé, and Subhash C. Kochar. A distribution-free test for the equality of failure rates due to two competing risks. Communications in Statistics: Theory and Methods, 18(1):107–120, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bain:1980:PCS

[BE80] Lee J. Bain and Max Engelhardt. Probability of correct selection of Weibull versus gamma based on livelihood ratio. Communications in Statistics: Theory and Methods, 9(4):375–381, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blackwelder:1982:PRS

[BE82] William C. Blackwelder and Robert C. Elston. Power and robustness of sib-pair linkage tests and extension to larger sibships. Communications in Statistics: Theory and Methods, 11(5):449– 484, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bain:1983:RMS

[BE83] Lee J. Bain and Max Engelhardt. A review of model selection procedures relevant to the Weibull distribution. Communications in Statistics: Theory and Methods, 12(5):589–609, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basu:1984:TWS

[BE84] A. P. Basu and Nader Ebrahimi. Testing whether survival function is bivariate new better than used. Communications in Statistics: Theory and Methods, 13(15):1839–1849, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Biswas:1988:ADA

[BE88] Suddhendu Biswas and Nather Abas Ebraheem. On an application of the diffusion approximation of the generalized birth and death process for stochastic population projection. Communications in Statistics: Theory and Methods, 17(3):779–789, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bedi:1987:ISU

[Bed87] Pradeep Kumer Bedi. Interpenetrating subsampling of unequal samples and the estimator with distinct units. Communications in Statistics: Theory and Methods, 16(5):1505–1512, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beder:1988:ECF

[Bed88] Jay H. Beder. Estimating a covariance function having an un unknown scale parameter. Communications in Statistics: Theory and Methods, 17(5):1641–1657, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beder:1989:PCTa

[Bed89a] Jay H. Beder. The problem of confounding in two-factor experiments. Communications in Statistics: Theory and Methods, 18 (2):591–612, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beder:1989:PCTb

[Bed89b] Jay H. Beder. The problem of confounding in two-factor experiments. Communications in Statistics: Theory and Methods, 18 (6):2165–2188, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beg:1980:ETP

[Beg80] M. A. Beg. Estimation of Pr(y, x) for truncation parameter distributions. Communications in Statistics: Theory and Methods, 9 (3):327–345, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beg:1981:EFD

[Beg81a] M. A. Beg. Exponential family distribution with unknown truncation parameter: two sample problem. Communications in Statistics: Theory and Methods, 10(2):177–187, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beg:1981:NET

[Beg81b] M. A. Beg. A note on the estimation of P r{y, x} for truncation parameter distributions. Communications in Statistics: Theory and Methods, 10(7):687–691, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bell:1988:GGP

[Bel88] Bradley M. Bell. Generalized gamma parameter estimation and moment evaluation. Communications in Statistics: Theory and Methods, 17(2):507–517, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bendre:1989:MSE

[Ben89] S. M. Bendre. Masking and swamping effects on tests for multiple outliers in normal sample. Communications in Statistics: Theory and Methods, 18(2):697–710, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:1981:SAT

[Ber81a] Roger L. Berger. Selecting all treatments better than a control using existing tables. Communications in Statistics: Theory and Methods, 10(20):2025–2037, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berman:1981:MLE

[Ber81b] M. Berman. The maximum likelihood estimators of the parameters of the gamma distribution are always positively biased. Communications in Statistics: Theory and Methods, 10(7):693–697, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berenson:1982:SUN

[Ber82] Mark L. Berenson. Some useful nonparametric tests for ordered alternatives in randomized block experiments. Communications in Statistics: Theory and Methods, 11(15):1681–1693, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:1984:TWO

[Ber84] Roger L. Berger. Testing whether one regression function is larger than another. Communications in Statistics: Theory and Methods, 13(15):1793–1810, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:1985:SLC

[Ber85] Roger W. Berger. A survey of low-cost, user-friendly SQC software for microcomputers. Communications in Statistics: Theory and Methods, 14(11):2575–2588, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:1986:MCS

[Ber86] Martijn P. F. Berger. A Monte Carlo study of the power of alternative tests under the generalized MANOVA model. Communications in Statistics: Theory and Methods, 15(4):1251–1283, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berry:1989:BME

[Ber89] Calvin J. Berry. Bayes minimax estimation of a Bernoulli p in a restricted parameter space. Communications in Statistics: Theory and Methods, 18(12):4607–4616, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Betz:1987:AHL

[Bet87] M. Austin Betz. An approximation for the Hotelling–Lawley trace in the noncentral case. Communications in Statistics: Theory and Methods, 16(11):3169–3183, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bev:1984:SPD

[Bev84] Littlewood Bev. Subjective probability and the DFR-mixture closure theorem. Communications in Statistics: Theory and Methods, 13(7):859–863, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brook:1981:OSL

[BF81] R. J. Brook and R. H. Fletcher. Optimal significance levels of prior tests in the presence of multicollinearity. Communications in Statistics: Theory and Methods, 10(14):1401–1413, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bryant:1985:SCC

[BF85] John L. Bryant and G. Edward Fox. Some comments on a class of simultaneous inference procedures in ANCOVA. Communications in Statistics: Theory and Methods, 14(10):2511–2530, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bass:1986:TST

[BF86] Thomas A. Bass and Karen Yuen Fung. A two-stage test for the mean of a multivariate normal distribution with unknown covariance matrix. Communications in Statistics: Theory and Methods, 15(1):109–121, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beauchamp:1980:NEL

[BFR80] J. J. Beauchamp, J. E. Folkert, and D. S. Robson. A note on the effect of logarithmic transformation on the probability of misclassification. Communications in Statistics: Theory and Methods, 9 (8):777–794, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basu:1980:IDU

[BG80] A. P. Basu and J. K. Ghosh. Identifiability of distributions under competing risks and complementary risks model. Communications in Statistics: Theory and Methods, 9(14):1515–1525, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basu:1983:IRS

[BG83] A. P. Basu and J. K. Ghosh. Identifiability results for a k-out-of-p system. Communications in Statistics: Theory and Methods, 12 (2):199–205, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borwein:1984:BMS

[BG84] P. Borwein and G. Gabor. On the behavior of MLE of the scale parameter of the Student family. Communications in Statistics: Theory and Methods, 13(24):3047–3057, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burdick:1985:CIT

[BG85] Richard K. Burdick and Franklin A. Graybill. Confidence intervals on the total variance in an unbalanced two-fold nested classification with equal subsampling. Communications in Statistics: Theory and Methods, 14(4):761–774, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bruvold:1988:SSS

[BG88a] Norman T. Bruvold and Shanker Ganesan. A sample size selection procedure for pairwise comparisons of means with unequal variances. Communications in Statistics: Theory and Methods, 17 (5):1497–1505, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burdick:1988:PSC

[BG88b] Richard K. Burdick and Franklin A. Graybill. The present status of confidence interval estimation on variance components in balanced and unbalanced random models. Communications in Statistics: Theory and Methods, 17(4):1165–1195, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burdick:1980:CUT

[BH80] D. S. Burdick and D. G. Herr. Counterexamples in unbalanced twoway analysis of variance. Communications in Statistics: Theory and Methods, 9(2):231–241, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Behnen:1984:SAA

[BH84] Konrad Behnen and Marie Huskové. A simple algorithm for the adaptation of scores and power behavior of the corresponding rank test. Communications in Statistics: Theory and Methods, 13(3): 305–325, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1986:SLS

[BH86] L. D. Brown and J. T. Hwang. Some limitation on Stein’s phenomenon. Communications in Statistics: Theory and Methods, 15 (7):2025–2042, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:1981:ECL

[Bha81] C. G. Bhattacharya. Estimation of a common location. Communications in Statistics: Theory and Methods, 10(10):955–961, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhargava:1982:UMC

[Bha82] R. P. Bhargava. Univariate and multivariate categorical data analysis for block designs. Communications in Statistics: Theory and Methods, 11(11):1209–1231, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhargava:1986:COL

[Bha86] R. P. Bhargava. Classification of an observation into linear manifolds, their closed subsets and closed convex subsets. Communications in Statistics: Theory and Methods, 15(8):2363–2374, 1986.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharyya:1987:OSC

[Bha87] B. B. Bhattacharyya. One sided Chebyshev inequality when the first four moments are known. Communications in Statistics: Theory and Methods, 16(9):2789–2791, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandary:1989:DOS

[Bha89] Madhusudan Bhandary. Detection of outliers in signal processing and detection of number of signals in presence of outliers. Communications in Statistics: Theory and Methods, 18(6):2233–2261, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ben-Horim:1982:EEU

[BHL82] Moshe Ben-Horim and Haim Levy. Evaluating estimators using stochastic dominance rules: the variance of a normal distribution. Communications in Statistics: Theory and Methods, 11(9):1071– 1076, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boardman:1981:PLS

[BHW81] Anthony E. Boardman, Baldwin S. Hui, and Herman Wold. The partial least squares-fix point method of estimating interdependent systems with latent variables. Communications in Statistics: Theory and Methods, 10(7):613–639, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bilodeau:1989:MRD

[Bil89] Martin Bilodeau. On the monotone regression dependence for Archimedian bivariate uniform. Communications in Statistics: Theory and Methods, 18(3):981–988, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bishop:1980:IFW

[Bis80] Thomas A. Bishop. Improved fixed-width confidence intervals for the largest normal mean. Communications in Statistics: Theory and Methods, 9(3):243–253, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bera:1983:TMN

[BJ83] A. Bera and S. John. Tests for multivariate normality with Pearson alternatives. Communications in Statistics: Theory and Methods, 12(1):103–117, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:1988:ESE

[BJ88] D. S. Bai and J. S. Jang. Efficient sequential estimation in multiplicative counting processes. Communications in Statistics: Theory and Methods, 17(12):4371–4388, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bjornstad:1981:DTA

[Bjø81] Jan F. Bjørnstad. A decision-theoretic approach to subset selection. Communications in Statistics: Theory and Methods, 10(23): 2411–2433, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bjornstad:1982:CDT

[Bjø82] Jan F. Bjørnstad. Comparison of dependent two-way contingency tables. Communications in Statistics: Theory and Methods, 11 (6):673–686, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bjornstad:1985:ACP

[Bjø85] Jan F. Bjørnstad. Asymptotic consistency of procedures for selecting good populations. Communications in Statistics: Theory and Methods, 14(7):1659–1668, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Becker:1982:TMR

[BK82] Mark Becker and Anant M. Kshirsagar. Terminating Markov renewal processes. Communications in Statistics: Theory and Methods, 11(3):325–342, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bechhofer:1984:CSP

[BK84] Robert E. Bechhofer and Radhika V. Kulkarni. Closed sequential procedures for selecting the multinomial events which have the largest probabilities. Communications in Statistics: Theory and Methods, 13(24):2997–3031, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1985:RTS

[BK85a] N. Balakrishnan and S. Kocherlakota. Robust two sided tolerance limits based on MML estimators. Communications in Statistics: Theory and Methods, 14(1):175–184, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1985:RNL

[BK85b] N. Balakrishnan and Subrahmaniam Kocherlakota. Robustness to nonnormality of the linear discriminant function: mixtures of normal distributions. Communications in Statistics: Theory and Methods, 14(2):465–478, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1985:PCP

[BK85c] Bhagwandas Shakti Banerjee and Sanpei Kageyama. Patterned constructions of partially balanced incomplete block design. Communications in Statistics: Theory and Methods, 14(6):1259–1267, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagai:1986:TOC

[BK86] Isha Bagai and Subhash C. Kochar. On tail-ordering and comparison of failure rates. Communications in Statistics: Theory and Methods, 15(4):1377–1388, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:1988:MCC

[BKZ88] Z. D. Bai, P. R. Krishnaiah, and L. C. Zhao. Multivariate components of covariance model in unbalanced case. Communications in Statistics: Theory and Methods, 17(4):1311–1324, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:1984:IST

[BL84] Shaul K. Bar-Lev. On independence of statistics in truncated exponential models. Communications in Statistics: Theory and Methods, 13(17):2161–2172, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blasidell:1984:SDP

[Bla84] Edwin B. Blasidell. Systematic designs permitting determination of trends in several attributes each unaffected by the others. Communications in Statistics: Theory and Methods, 13(5):581–598, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumenthal:1985:AEM

[Blu85] Saul Blumenthal. Asymptotic expansions for modified maximum likelihood estimators with percentile truncated data. Communications in Statistics: Theory and Methods, 14(4):905–925, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumenthal:1987:ETG

[Blu87] Saul Blumenthal. Estimating n with time grouped and truncated data from a scale parameter distribution. Communications in Statistics: Theory and Methods, 16(4):1161–1179, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumentthal:1988:EPG

[Blu88] Saul Blumentthal. Estimating with percentile grouped and truncated date from a scale parameter distribution. Communications in Statistics: Theory and Methods, 17(11):3607–3628, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bentler:1987:MPC

[BLW87] P. M. Bentler, S. Y. Lee, and Li-Jen Weng. Multiple population covariance structure analysis under arbitrary distribution theory. Communications in Statistics: Theory and Methods, 16(7):1951– 1964, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braun:1981:TRA

[BM81a] Henry I. Braun and Donald R. McNeil. Testing in robust ANOVA. Communications in Statistics: Theory and Methods, 10(2):149– 165, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buhler:1981:WJI

[BM81b] Wolfgang J. Bühler and Klaus J. Miescke. On (n−l)-wise and joint independence and normality of n random variables: an example. Communications in Statistics: Theory and Methods, 10(9):927– 930, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Backhouse:1982:TDM

[BM82] A. R. Backhouse and R. J. McKay. Tests of dimensionality in multivariate analysis of variance. Communications in Statistics: Theory and Methods, 11(9):1003–1027, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1985:SGI

[BM85a] N. Balakrishnan and H. J. Malik. Some general identities involving order statistics. Communications in Statistics: Theory and Methods, 14(2):333–339, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basford:1985:CAR

[BM85b] K. E. Basford and G. J. McLachlan. Cluster analysis in a randomized complete block design. Communications in Statistics: Theory and Methods, 14(2):451–463, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Booth:1985:SQC

[BM85c] J. Booth and W. Mudryk. Statistical quality control at Statistics Canada. Communications in Statistics: Theory and Methods, 14 (11):2589–2603, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1986:OSL

[BM86a] N. Balakrishnan and H. J. Malik. Order statistics from the linearexponential distribution, Part I: increasing hazard rate case. Communications in Statistics: Theory and Methods, 15(1):179–203, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1986:SSB

[BM86b] Evc Bofinger and Kerrie Mangersen. Subset selection of the t best populations. Communications in Statistics: Theory and Methods, 15(10):3145–3161, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1987:BLU

[BM87a] N. Balakrishnan and H. J. Malik. Best linear unbiased estimation of location and scale parameters of the log-logistic distribution.

Communications in Statistics: Theory and Methods, 16(12):3477– 3495, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burns:1987:GTT

[BM87b] Kevin C. Burns and Carl A. Mauro. Group testing with test error as a function of concentration. Communications in Statistics: Theory and Methods, 16(10):2821–2837, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bravo:1988:IEP

[BM88] Gina Bravo and Brenda Macgibbon. Improved estimation for the parameters of an inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 17(12):4285–4299, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagchi:1989:OPT

[BM89] Sunanda Bagchi and A. C. Mukhopadhyay. Optimality in the presence of two factor interactions among the nuisance factors. Communications in Statistics: Theory and Methods, 18(3):1139–1152, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1988:RRI

[BMA88] N. Balakrishnan, Henrick J. Malick, and S. E. Ahmed. Recurrence relations and identities for moments of order statistics, II; specific continuous distributions. Communications in Statistics: Theory and Methods, 17(8):2657–2694, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burdick:1986:CII

[BMC86] Richard K. Burdick, Farooq Maqsood, and Franklin A. Craybill. Confidence intervals on the intraclass correlation in the unbalanced one-way classification. Communications in Statistics: Theory and Methods, 15(11):3353–3378, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1984:CDL

[Bof84] Eve Bofinger. Δ-correct decision for location and scale parameters. Communications in Statistics: Theory and Methods, 13(24):3117– 3121, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1985:ECI

[Bof85] Eve Bofinger. Expanded confidence intervals. Communications in Statistics: Theory and Methods, 14(8):1849–1864, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1988:LSS

[Bof88] Eve Bofinger. Least significant spacing for ‘one versus the rest’ normal populations. Communications in Statistics: Theory and Methods, 17(5):1697–1716, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bohning:1983:CFP

[Boh83] Dankmar Bohning. Convergence of a fixed point algorithm. Communications in Statistics: Theory and Methods, 12(22):2645–2653, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolfarine:1987:MPF

[Bol87] Heleno Bolfarine. Minimax prediction in finite populations. Communications in Statistics: Theory and Methods, 16(12):3683–3700, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolfarine:1989:EPF

[Bol89a] Heleno Bolfarine. Equivariant prediction in finite populations. Communications in Statistics: Theory and Methods, 18(3):927– 941, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolfarine:1989:NFP

[Bol89b] Heleno Bolfarine. A note on finite population prediction under asymmetric loss functions. Communications in Statistics: Theory and Methods, 18(5):1863–1869, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bonett:1985:NML

[Bon85] Douglas G. Bonett. The negative multinomial logit model. Communications in Statistics: Theory and Methods, 14(7):1713–1717, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bonett:1988:ESN

[Bon88] Douglas G. Bonett. Exact and stochastic nonlinear constrained estimation in the conditional Poisson log-linear model. Communications in Statistics: Theory and Methods, 17(10):3301–3311, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boos:1982:MAD

[Boo82] Dennis D. Boos. Minimum Anderson–Darling estimation. Communications in Statistics: Theory and Methods, 11(24):2747–2774, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borowiak:1983:MMD

[Bor83] Dale Borowiak. A multiple model discrimination procedure. Communications in Statistics: Theory and Methods, 12(24):2911–2921, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borowiak:1985:DMF

[Bor85] Dale Borowiak. A dependent multiple failure mode modelling procedure. Communications in Statistics: Theory and Methods, 14 (8):1837–1847, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bourke:1982:RRM

[Bou82] Patrick D. Bourke. Randomized response multivariate designs for categorical data. Communications in Statistics: Theory and Methods, 11(25):2889–2901, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Box:1985:EDP

[Box85] George E. P. Box. Experimental design for product improvement. Communications in Statistics: Theory and Methods, 14(11):2605, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baringhaus:1981:OSV

[BP81] Ludwig Baringhaus and Detlef Plachky. On the optimality of the sample variance. Communications in Statistics: Theory and Methods, 10(13):1277–1284, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bryant:1983:EMP

[BP83] J. L. Bryant and A. S. Paulson. Estimation of mixing proportions via distance between characteristic functions. Communications in Statistics: Theory and Methods, 12(9):1009–1029, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baye:1984:CRP

[BP84] Michael R. Baye and Darrell F. Parker. Combining ridge and principal component regression: a money demand illustration. Communications in Statistics: Theory and Methods, 13(2):197–205, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boswell:1987:PCS

[BP87] M. T. Boswell and G. P. Patil. A perspective of composite sampling. Communications in Statistics: Theory and Methods, 16(10): 3069–3093, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1982:EPM

[BR82] Lyle Broemeling and Muthiya Rajagopalan. Estimating the parameters of mixed linear models with modal estimators. Communications in Statistics: Theory and Methods, 11(25):2921–2935, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bednar:1985:RBL

[BR85] J. Bee Bednar and Brenda Roberts. On the relationship between Levinson recursion and the r and s arrays for ARMA model identification. Communications in Statistics: Theory and Methods, 14 (5):1217–1248, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bozdogan:1986:ALR

[BR86] Hamparsum Bozdogan and Donald E. Ramirez. An adjusted likelihood-ratio approach to the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 15(8):2405–2433, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brannigan:1981:APP

[Bra81] M. Brannigan. An adaptive piecewise polynomial curve fitting procedure for data analysis. Communications in Statistics: Theory and Methods, 10(18):1823–1848, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bradu:1984:RSM

[Bra84] Dan Bradu. Response surface model diagnosis in two-way tables. Communications in Statistics: Theory and Methods, 13(24):3059– 3106, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bradbury:1988:APD

[Bra88] I. S. Bradbury. Approximations to permutation distributions in the completely randomized design. Communications in Statistics: Theory and Methods, 17(2):543–555, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Breuer:1982:NGV

[Bre82] Joseph Breuer. A note on the greenwood variance estimator in cohort life tables: Isolating the effect of one source of bias. Communications in Statistics: Theory and Methods, 11(25):2947–2951, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brittain:1987:VMS

[Bri87] Erica H. Brittain. P -values for the multi-sample Kolmogorov– Smirnov test using the expanded Bonferroni approximation. Communications in Statistics: Theory and Methods, 16(3):821–835, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bronstrom:1981:SRS

[Bro81] Göran Bronström. On sequentially rejective subset selection procedures. Communications in Statistics: Theory and Methods, 10 (3):203–221, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1982:GFT

[Bro82] Charles C. Brown. On a goodness of fit test for the logistic model based on score statistics. Communications in Statistics: Theory and Methods, 11(10):1087–1105, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brooks:1987:DCL

[Bro87] R. J. Brooks. On the design of comparative lifetime studies. Communications in Statistics: Theory and Methods, 16(5):1221–1240, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunk:1980:BLS

[Bru80] H. D. Brunk. Bayesian least squares estimates of univariate regression functions. Communications in Statistics: Theory and Methods, 9(11):1101–1136, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumenthal:1980:ETI

[BS80a] Saul Blumenthal and Lalitha Sanathanan. Estimation with truncated inverse binomial sampling. Communications in Statistics: Theory and Methods, 9(10):997–1017, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bovett:1980:CTP

[BS80b] James M. Bovett and John G. Saw. On comparing two Poisson intensity functions. Communications in Statistics: Theory and Methods, 9(9):943–948, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bradley:1980:PWM

[BS80c] Ralph A. Bradley and Elton Scott. Perspective from a weather modification experiment. Communications in Statistics: Theory and Methods, 9(18):1941–1961, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boyd:1983:UIR

[BS83] Michael N. Boyd and Pranab K. Sen. Union-intersection rank tests for ordered alternatives in some simple linear models. Communications in Statistics: Theory and Methods, 12(15):1737–1753, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boyd:1984:UIR

[BS84] Michael N. Boyd and Pranab K. Sen. Union-intersection rank tests for ordered alternatives in a complete block design. Communications in Statistics: Theory and Methods, 13(3):285–303, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bell:1986:INN

[BS86] C. B. Bell and E. P. Smith. Inference for non-negative autoregressive schemes. Communications in Statistics: Theory and Methods, 15(8):2267–2293, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1987:CPA

[BS87] Lyle D. Broemeling and Mun S. Son. The classification problem with autoregressive process. Communications in Statistics: Theory and Methods, 16(4):927–936, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Breusch:1988:AFW

[BS88] Trevor S. Breusch and Peter Schmidt. Alternative forms of the Wald test: how long is a piece of string? Communications in Statistics: Theory and Methods, 17(8):2789–2795, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharyya:1989:TFR

[BS89] G. K. Bhattacharyya and Zanzawi Soejoeti. A tampered failure rate model for step-stress accelerated life test. Communications in Statistics: Theory and Methods, 18(5):1627–1643, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bryce:1980:EHL

[BSC80] G. Rex Bryce, Del T. Scott, and Melvin W. Carter. Estimation and hypothesis in linear models — a reparameterization approach to the cell MLANS model. Communications in Statistics: Theory and Methods, 9(2):131–150, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bayne:1981:QMP

[BT81a] C. K. Bayne and W. Y. Tan. Qdf misclassification probabilities for known population parameters. Communications in Statistics: Theory and Methods, 10(22):2315–2326, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Block:1981:SCM

[BT81b] Henry W. Block and Mei-Ling Ting. Some concepts of multivariate dependence. Communications in Statistics: Theory and Methods, 10(8):749–762, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bean:1982:BSP

[BT82] Steven Bean and Chris P. Tsokos. Bandwidth selection procedures for kernel density estimates. Communications in Statistics: Theory and Methods, 11(9):1045–1069, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bargal:1983:SGF

[BT83] Adel I. Bargal and David R. Thomas. Smooth goodness of fit tests for the Weibull distribution with singly censored data. Communications in Statistics: Theory and Methods, 12(12):1431–1447, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balasooriya:1986:ODL

[BT86] Uditha Balasooriya and Y. K. Tse. Outlier detection in linear models: a comparative study in simple linear regression. Communications in Statistics: Theory and Methods, 15(12):3589–3597, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brannas:1984:IUO

[BU84] Kurt Brännäs and Stig Uhlin. Improper use of the ordinary least squares estimator in the switching regression model. Communications in Statistics: Theory and Methods, 13(14):1781–1791, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buonaccorsi:1985:RLC

[Buo85] John P. Buonaccorsi. Ratios of linear combinations in the general linear model. Communications in Statistics: Theory and Methods, 14(3):635–650, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buonaccorsi:1989:EVS

[Buo89] John P. Buonaccorsi. Errors-in-variables with systematic biases. Communications in Statistics: Theory and Methods, 18(3):1001– 1021, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burstein:1981:BTI

[Bur81] Herman Burstein. Binomial 2×2 test for independent samples with independent proportions. Communications in Statistics: Theory and Methods, 10(1):11–29, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunden:1989:OFR

[BV89] Marshall N. Brunden and Thomas J. Vidmar. Optimal 3 × 3 factorial and 3-ray designs for drug interaction experiments with dichotomous responses. Communications in Statistics: Theory and Methods, 18(8):2835–2859, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bush:1982:AON

[BW82] A. Bush and H. S. Wieand. An asymptotically optimal nonparametric statistic for testing equality of two normal population means and variances. Communications in Statistics: Theory and Methods, 11(1):1–12, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burnett:1988:LRT

[BW88] Richard T. Burnett and R. Andrew Willan. Linear rank tests for randomized block design. Communications in Statistics: Theory and Methods, 17(8):2455–2470, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bonett:1986:MCS

[BWD86] Douglas G. Bonett, Arthur J. Woodward, and John L. Davidson. Minimum chi-square estimator and test for incomplete multidimensional contingency tables. Communications in Statistics: Theory and Methods, 15(4):1201–1207, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bush:1983:MCM

[BWMD83] John G. Bush, Brian W. Woodruff, Albert H. Moore, and Edward J. Dunne. Modified Cramér–von Mises and Anderson–Darling tests for Weibull distributions with unknown location and scale parameters. Communications in Statistics: Theory and Methods, 12 (21):2465–2476, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1985:SBS

[BYD85] Lyle Broemeling, Abdullah Mat Yusoff, and Joaquin Diaz. Some Bayesian solutions for problems of adaptive estimation in linear dynamic systems. Communications in Statistics: Theory and Methods, 14(2):401–418, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byrne:1988:MAS

[Byr88] Philip Byrne. Multivariate analysis with a stationary time series covariance structure. Communications in Statistics: Theory and Methods, 17(6):1935–1943, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cuadras:1981:CGM

[CA81] Carles M. Cuadras and Joan Augé. A continuous general multivariate distribution and its properties. Communications in Statistics: Theory and Methods, 10(4):339–353, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crigler:1986:EOM

[CA86] John R. Crigler and Jesse C. Arnold. Economically optimal Markovian sampling policies for process monitoring. Communications in Statistics: Theory and Methods, 15(6):1773–1803, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calitz:1987:AKS

[Cal87] Fred Calitz. An alternative to the Kolmogorov–Smirnov test for goodness of fit. Communications in Statistics: Theory and Methods, 16(12):3519–3534, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Canner:1984:MLT

[Can84] Paul L. Canner. Monitoring long-term clinical trials for beneficial and adverse treatment effects. Communications in Statistics: Theory and Methods, 13(19):2369–2394, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chengalur:1989:VSI

[CAR89] Indushobha N. Chengalur, Jesse C. Arnold, and Marion R. Reynolds, Jr. Variable sampling intervals for multiparameter Shewhart charts. Communications in Statistics: Theory and Methods, 18(5):1769–1792, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1984:NRN

[CARO84] Youn-Min Chou, Kathleen H. Arthur, Rebecca B. Rosenstein, and D. B. Owen. New representations of the noncentral chi-square density and cumulative. Communications in Statistics: Theory and Methods, 13(21):2673–2678, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Causey:1983:EPM

[Cau83] Beverley D. Causey. Estimation of proportions for multinomial contingency tables subject to marginal constraints. Communications in Statistics: Theory and Methods, 12(22):2581–2587, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Causey:1984:EUG

[Cau84] Beverley D. Causey. Estimation under generalized sampling of cell proportions for contingency tables subject to marginal constraints. Communications in Statistics: Theory and Methods, 13(20):2487– 2494, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Craig:1981:EBA

[CB81] J. A. Craig, Jr. and R. P. Bland. An empirical Bayes approach to a variables acceptance sampling plan problem. Communications in Statistics: Theory and Methods, 10(23):2399–2410, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1982:BTE

[CB82] Cheng Chen and Thomas C. Baker, Jr. On Bayes-type estimators of linearly or log-linearly constrained multinomial cell probabilities. Communications in Statistics: Theory and Methods, 11(18):2061– 2073, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chambless:1985:MLM

[CB85] Lloyd E. Chambless and Kerrie E. Boyle. Maximum likelihood methods for complex sample data: logistic regression and discrete proportional hazards models. Communications in Statistics: Theory and Methods, 14(6):1377–1392, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1987:ECB

[CB87] Shu-Mei Chen and Gouri K. Bhattacharyya. Exact confidence bounds for an exponential parameter under hybrid censoring. Communications in Statistics: Theory and Methods, 16(8):2429–2442, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1988:ECB

[CB88] Shu-Mei Chen and Gouri K. Bhattacharyya. Exact confidence bounds for an exponential parameter under hybrid censoring. Communications in Statistics: Theory and Methods, 17(6):1857–1870, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carroll:1982:SAR

[CC82] R. J. Carroll and Paul P. Callo. Sow aspects of robustness in the functional errors-in-variables regression model. Communications in Statistics: Theory and Methods, 11(22):2573–2585, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1985:PML

[CC85] Christy Chuang and Christopher Cox. Pseudo maximum likelihood estimation for the Dirichlet-multinomial distribution. Communications in Statistics: Theory and Methods, 14(10):2293–2311, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1988:EWP

[CC88] K. F. Cheng and C. H. Chen. Estimation of the Weibull parameters with grouped data. Communications in Statistics: Theory and Methods, 17(2):325–341, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [Wan09].

Chaloner:1987:SPD

[CD87] Kathryn Chaloner and George T. Duncan. Some properties of the Dirichlet-multinomial distribution and its use in prior elicitation. Communications in Statistics: Theory and Methods, 16(2):511– 523, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cleroux:1989:VCE

[CD89] Robert Cléroux and Gilles R. Ducharme. Vector correlation for elliptical distributions. Communications in Statistics: Theory and Methods, 18(4):1441–1454, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandra:1982:SRT

[CDS82] Mahesh Chandra, Tertius De Wet, and Nozer D. Singpurwalla. On the sample redundancy and a test for exponentiality. Communications in Statistics: Theory and Methods, 11(4):429–438, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:1984:ECP

[CDS84] Y. P. Chaubey, T. D. Dwivedi, and M. Singh. An efficiency comparison of product and ratio estimator. Communications in Statistics: Theory and Methods, 13(6):699–709, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chinchilli:1985:MMG

[CE85] Vernon M. Chinchilli and Ronald K. Elswick. A mixture of the MANOVA and GMANOVA models. Communications in Statistics: Theory and Methods, 14(12):3075–3089, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ciampi:1985:GMT

[CEA85] A. Ciampi and J. Etezadi-Amoli. A general model for testing the proportional hazards and the accelerated failure time hypotheses in the analysis of censored survival data with covariates. Communications in Statistics: Theory and Methods, 14(3):651–667, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cerdan:1989:NBA

[Cer89] Santiago Velilia Cerdán. A note on the behaviour of augmented principal-component plots in regression. Communications in Statistics: Theory and Methods, 18(1):331–342, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1986:SMM

[CF86] Christy Chuang and Steven F. Francom. A structural model for the Markov chain of change in vote intention. Communications in Statistics: Theory and Methods, 15(12):3475–3487, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1988:MLE

[CF88] P. C. Consul and Felix Famoye. Maximum likelihood estimation for the generalized Poisson distribution when sample mean is larger than sample variance. Communications in Statistics: Theory and Methods, 17(1):299–309, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carter:1989:MBS

[CF89a] Nancy J. Carter and G. David Faulkenberry. A model-based sample selection procedure for one-stage sampling. Communications in Statistics: Theory and Methods, 18(8):3135–3147, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1989:TGP

[CF89b] P. C. Consul and Felix Famoye. The truncated generalized Poisson distribution and its estimation. Communications in Statistics: Theory and Methods, 18(10):3635–3648, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:1981:ALF

[CG81] Y. P. Chaubey and G. Gabor. Another look at Fisher’s solution to the problem of the weighted mean. Communications in Statistics: Theory and Methods, 10(12):1225–1237, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakravorti:1987:LSA

[CG87] S. R. Chakravorti and Irwin Guttman. A large sample analysis of the magnitudinal model in multivariate analysis. Communications in Statistics: Theory and Methods, 16(7):1995–2002, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Caso:1988:GSI

[CG88] Covadonga Caso and Maria Angeles Gil. The Gini–Simpson index of diversity: estimation in the stratified sampling. Communications in Statistics: Theory and Methods, 17(9):2981–2995, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1985:MDM

[CGO85] Christy Chuang, David Gheva, and Charles Odoroff. Methods for diagnosing multiplicative-interaction models for two-way contingency tables. Communications in Statistics: Theory and Methods, 14(9):2057–2080, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1983:WLS

[CH83] Kuo-Sheuan Cheng and Thomas P. Hettmansperger. Weighted least-squares rank estimates. Communications in Statistics: Theory and Methods, 12(9):1069–1086, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Casella:1986:CSS

[CH86] George Casella and Jiunn Tzon Hwang. Confidence sets and the Stein effect. Communications in Statistics: Theory and Methods, 15(7):2043–2063, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christen:1987:DTC

[CH87] C. Christen and F. K. Hwang. Detecting the two complementary defectives with a balance scale. Communications in Statistics: Theory and Methods, 16(10):3025–3032, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chin:1988:MTS

[CH88] Eugenia Bair Chin and John E. Hewett. Multivaritate two stage test based on ranks. Communications in Statistics: Theory and Methods, 17(12):4099–4119, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakravorti:1980:STM

[Cha80] S. R. Chakravorti. On some tests in MANOCOVA model with different dispersion matrices. Communications in Statistics: Theory and Methods, 9(3):291–308, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chattopadhyay:1981:SNP

[Cha81] A. K. Chattopadhyay. Selecting the normal population with largest (smallest) value of Mahalanobis distance from the origin. Communications in Statistics: Theory and Methods, 10(1):31–37, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chattopadhyay:1982:NGD

[Cha82a] A. K. Chattopadhyay. Note on gammaization of dependent quadratic forms. Communications in Statistics: Theory and Methods, 11(2):217–221, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:1982:BMB

[Cha82b] Yogendra P. Chaubey. Best minimum bias linear estimators in Gauss–Markoff model. Communications in Statistics: Theory and Methods, 11(17):1959–1963, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1983:LST

[Cha83a] D. S. Chang. On the large sample theory of the multivariate ratio estimator. Communications in Statistics: Theory and Methods, 12 (12):1381–1390, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chao:1983:AMS

[Cha83b] Anne Chao. Approximate mean squared errors of estimations of reliability in the two-parameter exponential case. Communications in Statistics: Theory and Methods, 12(6):633–644, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:1984:PMG

[Cha84] Ch A. Charalambides. Probabilities and moments of generalized discrete distributions using finite difference operators. Communications in Statistics: Theory and Methods, 13(25):3225–3241, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:1986:DPM

[Cha86a] Charalambos A. Charalambides. Derivation of probabilities and moments of certain generalized discrete distributions via urn models. Communications in Statistics: Theory and Methods, 15(3): 677–696, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatfield:1986:DDP

[Cha86b] C. Chatfield. Discrete distributions and purchasing models. Communications in Statistics: Theory and Methods, 15(3):697–708, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chauhan:1987:SRR

[Cha87] Chand K. Chauhan. Some results on the regularity in a single replicate design. Communications in Statistics: Theory and Methods, 16(8):2263–2268, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborti:1988:LST

[Cha88a] S. Chakraborti. Large sample tests for equality of medians under unequal right-censoring. Communications in Statistics: Theory and Methods, 17(12):4075–4084, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chawia:1988:NGR

[Cha88b] J. S. Chawia. A note on general ridge estimator. Communications in Statistics: Theory and Methods, 17(3):739–744, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaloner:1989:BDE

[Cha89a] Kathryn Chaloner. Bayesian design for estimating the turning point of a quadratic regression. Communications in Statistics: Theory and Methods, 18(4):1385–1400, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1989:CHD

[Cha89b] Cheng-Tao Chang. On the construction of hypercubic design. Communications in Statistics: Theory and Methods, 18(9):3365–3371, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1989:SDG

[Cha89c] Kashinath Chatterjee. Search designs for general asymmetric factorials. Communications in Statistics: Theory and Methods, 18 (6):2189–2200, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1980:SSW

[Che80] Hubert J. Chen. On selecting a subset which contains all populations better than a control. Communications in Statistics: Theory and Methods, 9(8):851–864, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1983:NEP

[Che83a] K. F. Cheng. Nonparametric estimators for percentile regression functions. Communications in Statistics: Theory and Methods, 12 (6):681–692, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chester:1983:NPE

[Che83b] S. Theodore Chester, Jr. A note on a parametric extension of rank tests to the analysis of multifactor experiments within blocks. Communications in Statistics: Theory and Methods, 12(21):2493– 2502, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1984:SSD

[Che84a] Hubert J. Chen. Sample size determinations when two binomial proportions are very small. Communications in Statistics: Theory and Methods, 13(21):2707–2712, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1984:ANT

[Che84b] K. F. Cheng. Asymptotically nonparametric tests with censored paired data. Communications in Statistics: Theory and Methods, 13(12):1453–1470, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1986:ANL

[Che86a] Chan-Fu Chen. Asymptotic normality for limiting discrete distributions. Communications in Statistics: Theory and Methods, 15 (5):1699–1707, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1986:LFC

[Che86b] Pinyuen Chen. On the least favorable configuration in multinomial selection problems. Communications in Statistics: Theory and Methods, 15(2):367–385, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1987:SAP

[Che87a] Zehua Chen. A stepwise approach for the purely periodic interaction spline model. Communications in Statistics: Theory and Methods, 16(3):877–895, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1987:NNH

[Che87b] Philip E. Cheng. A nearest neighbor hazard rate estimator for randomly censored data. Communications in Statistics: Theory and Methods, 16(3):613–625, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1988:ORC

[Che88] Ken-Wei Chen. Optimal rates of convergence for the piecewise polynomial estimator with the index chosen by the FPE selection rule. Communications in Statistics: Theory and Methods, 17(9): 2887–2906, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chhikara:1984:CS

[Chh84] Raj S. Chhikara. Communications in statistics. Communications in Statistics: Theory and Methods, 13(23):7–8, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chikuse:1981:SPC

[Chi81] Yasuko Chikuse. Simultaneous pseudo confidence regions for ratios of the discriminant coefficients. Communications in Statistics: Theory and Methods, 10(6):531–547, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chhikara:1984:SDL

[CHLL84] Raj S. Chhikara, Cecil R. Hallum, and T. G. Lycthuan-Lee. Sampling designs for LANDSAT crop surveys. Communications in Statistics: Theory and Methods, 13(23):2905–2935, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chmielewski:1981:RAT

[Chm81] Margaret A. Chmielewski. A re-appraisal of tests for normality. Communications in Statistics: Theory and Methods, 10(20):2005– 2014, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chotai:1980:SSB

[Cho80] Jayanti Chotai. Subset selection based on likelihood from uniform and related populations. Communications in Statistics: Theory and Methods, 9(11):1147–1164, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chopra:1983:NUB

[Cho83a] D. V. Chopra. A note on an upper bound for the constraints of balanced arrays with strength t. Communications in Statistics: Theory and Methods, 12(15):1755–1759, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chotai:1983:IIP

[Cho83b] Jayanti Chotai. Isotonic inference for populations related to the uniform distribution. Communications in Statistics: Theory and Methods, 12(18):2109–2118, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1988:OSS

[Cho88] Youn-Min Chou. One-sided simultaneous prediction intervals for the order statistics of a future samples from an exponential distribution. Communications in Statistics: Theory and Methods, 17 (11):3995–4003, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christy:1986:NAR

[Chr86] Chuang Christy. A note on the applications of the RC model to two-way contingency tables. Communications in Statistics: Theory and Methods, 15(1):1–6, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christy:1988:TAA

[Chr88] Chuang-Stein Christy. Two approaches for the analysis of a titration study with a placebo control. Communications in Statistics: Theory and Methods, 17(3):821–832, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Costanza:1986:TSF

[CHS86] M. C. Costanza, H. I. Hamdy, and M. S. Son. Two stage fixed width confidence intervals for the common location parameter of several exponential distributions. Communications in Statistics: Theory and Methods, 15(8):2305–2322, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1982:EBM

[Chu82] Christy Chuang. Empirical Bayes methods for a two-way multiplicative-interaction model. Communications in Statistics:

Theory and Methods, 11(25):2977–2989, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1983:MIL

[Chu83] Christy Chuang. Multiplicative-interaction logit models for i × j × 2 three-way tables. Communications in Statistics: Theory and Methods, 12(24):2871–2885, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chuang:1984:DTW

[Chu84] Christy Chuang. On the decomposition of X^{2} for two-way contingency tables. Communications in Statistics: Theory and Methods, 13(9):1121–1126, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carter:1987:MPB

[CHW87] E. M. Carter, J. J. Hubert, and M. N. Walsh. Multivariate parabolic bioassay. Communications in Statistics: Theory and Methods, 16(9):2779–2788, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Conover:1980:RTM

[CI80] W. J. Conover and Ronald L. Iman. The rank transformation as a method of discrimination with some examples. Communications in Statistics: Theory and Methods, 9(5):465–487, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chhetry:1983:NPD

[CK83a] Devendra Chhetry and George Kimeldorf. A note on positive dependence in multivariate distributions. Communications in Statistics: Theory and Methods, 12(22):2601–2603, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chikkagoudar:1983:DTS

[CK83b] M. S. Chikkagoudar and S. H. Kunchur. Distributions of test statistics for multiple outliers in exponential samples. Communications in Statistics: Theory and Methods, 12(18):2127–2142, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chikkagoudar:1987:CMO

[CK87] M. S. Chikkagoudar and S. H. Kunchur. Comparison of many outlier procedures for exponential samples. Communications in Statistics: Theory and Methods, 16(3):627–645, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cooley:1983:BME

[CL83] Edward A. Cooley and Pi-Erh Lin. Bayes minimax estimators of a multivariate normal mean, with application to generalized ridge regression. Communications in Statistics: Theory and Methods, 12 (24):2861–2869, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1989:SSI

[CL89] Hubert J. Chen and Kin Lam. Single-stage interval estimation of the largest normal mean under heteroscedasticity. Communications in Statistics: Theory and Methods, 18(10):3703–3718, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clarke:1981:STS

[Cla81] K. R. Clarke. Statistical techniques in stereology. Communications in Statistics: Theory and Methods, 10(15):1459–1478, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clayton:1982:APS

[Cla82] D. G. Clayton. The analysis of prospective studies of disease aetiology. Communications in Statistics: Theory and Methods, 11(19): 2129–2155, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1986:RPW

[CLl86] Richard G. Cornell, Bryce D. Landenberger, and Robert H. Bart left. Randomized play the winner clinical trials. Communications in Statistics: Theory and Methods, 15(1):159–178, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castilio:1987:DFL

[CLMFC87] Enrique Castilio, Alberto Luceno, Alejandro Montalban, and Alfonso Fernandez-Canteli. A dependent fatigue lifetime model. Communications in Statistics: Theory and Methods, 16(4):1181– 1193, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cyr:1982:ACV

[CM82] J. L. Cyr and E. B. Manoukian. Approximate critical values with error bounds for Bartlett’s test of homogeneity of variances for unequal sample sizes. Communications in Statistics: Theory and Methods, 11(15):1671–1680, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandra:1984:ORS

[CM84] Tapas K. Chandra and Rahul Mukerjee. On the optimality of Rao’s statistic. Communications in Statistics: Theory and Methods, 13 (12):1507–1515, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carroll:1986:ACE

[CM86a] Steven S. Carroll and Lawrence S. Mayer. Analysis of the crosseffects in a cross-lagged panel study. Communications in Statistics: Theory and Methods, 15(11):3379–3399, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1986:FLR

[CM86b] S. Chatterjee and D. L. McLeish. Fitting linear regression models to censored data by least squares and maximum likelihood methods. Communications in Statistics: Theory and Methods, 15(11): 3227–3243, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cwik:1989:EDR

[ĆM89] J. Ćwik and J. Mielniczuk. Estimating density ratio with application to discriminant analysis. Communications in Statistics: Theory and Methods, 18(8):3057–3069, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cnaan:1985:SMT

[Cna85] Avital Cnaan. Survival models with two phases and length biased sampling. Communications in Statistics: Theory and Methods, 14 (4):861–886, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1984:APV

[CO84] Youn-Min Chou and D. B. Owen. An approximation to percentiles of a variable of the bivariate normal distribution when the other variable is truncated, with applications. Communications in Statistics: Theory and Methods, 13(20):2535–2547, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1986:OSN

[CO86] Youn-Min Chou and D. B. Owen. One-sided normal prediction intervals and sample sizes for a third-stage sample. Communications in Statistics: Theory and Methods, 15(8):2435–2448, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1989:DEP

[CO89] Youn-Min Chou and D. B. Owen. On the distributions of the estimated process capability indices. Communications in Statistics: Theory and Methods, 18(12):4549–4560, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cohen:1980:GDS

[Coh80] Ayala Cohen. On the graphical display of the significant components in two-way contingency tables. Communications in Statistics: Theory and Methods, 9(10):1025–1041, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cohen:1982:ALS

[Coh82] Ayala Cohen. Analysis of large sets of ranking data. Communications in Statistics: Theory and Methods, 11(3):235–256, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1981:RMP

[Con81] P. C. Consul. Relation of modified power series distributions to Lagrangian probability distributions. Communications in Statistics: Theory and Methods, 10(20):2039–2046, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Connell:1985:BLM

[Con85] Fred J. Connell. The bivariate linear model: a geometric approach. Communications in Statistics: Theory and Methods, 14(7):1605– 1614, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1988:SML

[Con88] P. C. Consul. On some models leading to the generalized Poisson distribution. Communications in Statistics: Theory and Methods, 17(2):423–442, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1989:DTG

[Con89] P. C. Consul. On the differences of two generalized negative binomial variates. Communications in Statistics: Theory and Methods, 18(2):673–690, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:1985:BAS

[Coo85] Peyton Cook. Bayesian autoregressive spectral analysis. Communications in Statistics: Theory and Methods, 14(5):1001–1018, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:1987:BAF

[Coo87] Peyton Cook. Bayesian autoregressive filter and squared gain estimation. Communications in Statistics: Theory and Methods, 16 (9):2697–2715, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1982:BE

[Cor82] Richard G. Cornell. Biostatistics and epidemiology. Communications in Statistics: Theory and Methods, 11(5):445–448, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Corley:1984:MOS

[Cor84] H. W. Corley. Multivariate order statistics. Communications in Statistics: Theory and Methods, 13(10):1299–1304, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1982:CSL

[CP82] Hubert J. Chen and Yuan S. P. On concurrence of several linear regressions and applications to problems of enzyme kinetics. Communications in Statistics: Theory and Methods, 11(4):395– 409, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1984:SAT

[CP84] Hubert J. Chen and John R. Pickett. Selecting all treatments better than a control under a multivariate normal distribution and a uniform prior distribution. Communications in Statistics: Theory and Methods, 13(1):59–80, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:1988:LDU

[CPW88] R. Dennis Cook, Daniel Peña, and Sanford Weisberg. The likelihood displacement: a unifying principle for influence measures. Communications in Statistics: Theory and Methods, 17(3):623– 640, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1982:SUE

[CQ82] C. C. Chen and A. J. Quinzi. Sequential unbiased estimation of parameters in a two-way contingency table. Communications in Statistics: Theory and Methods, 11(18):2091–2103, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1989:BEB

[CR89] Myron N. Chang and P. V. Rao. Berry–Esséen bound for the Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 18(12):4647–4664, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crump:1982:DDB

[Cru82] Kenny S. Crump. Designs for discriminating between binary dose response models with applications to animal carcinocenicity experiments. Communications in Statistics: Theory and Methods, 11 (4):375–393, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charles:1981:DNM

[CS81] Locke Charles and John D. Spurrier. On the distribution of a new measure of heaviness of tails. Communications in Statistics: Theory and Methods, 10(19):1967–1980, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carter:1983:ANN

[CS83] E. M. Carter and M. S. Srivastava. Asymptotic non-null distribution for the locally most powerful invariant test for sphericity. Communications in Statistics: Theory and Methods, 12(15):1719– 1725, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1984:PSR

[CS84a] Robert W. Chen and Eric V. Slud. On the product of symmetric random variables. Communications in Statistics: Theory and Methods, 13(5):611–615, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1984:MLE

[CS84b] P. C. Consul and M. M. Shoukri. Maximum likelihood estimation for the generalized Poisson distribution. Communications in Statistics: Theory and Methods, 13(12):1533–1547, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1987:IFS

[CS87a] Pinyuen Chen and Milton Sabel. An integrated formulation for selecting the t best of k normal populations. Communications in Statistics: Theory and Methods, 16(1):121–146, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1987:NFM

[CS87b] Pinyuen Chen and Milton Sabel. A new formulation for the multinomial selection problem. Communications in Statistics: Theory and Methods, 16(1):147–180, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1988:CPS

[CS88a] Y. M. Chan and M. S. Srivastava. Comparison of powers for the sphericity tests using both the asymptotic distribution and the bootstrap. Communications in Statistics: Theory and Methods, 17 (3):671–690, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:1988:RSN

[CS88b] Ch. A. Charalambides and Jagbir Singh. Review of the Stirling numbers, their generalizations and statistical applications. Communications in Statistics: Theory and Methods, 17(8):2533–2595, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:1983:EMR

[CT83] Yogendra P. Chaubey and A. B. M. Nur Enayet Talukder. Exact moments of a ratio of two positive quadratic forms in normal variables. Communications in Statistics: Theory and Methods, 12 (6):675–679, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ciampi:1989:PRT

[CT89] Antonio Ciampi and Johanne Thiffault. Pruning regression trees for censored survival data: the RECPAM approach. Communications in Statistics: Theory and Methods, 18(9):3373–3388, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1989:SUC

[CV89a] Hubert J. Chen and Kanlaya Vanichbuncha. Simultaneous upper confidence bounds for distances from the best two-parameter exponential distribution. Communications in Statistics: Theory and Methods, 18(8):3019–3031, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consonni:1989:NCI

[CV89b] Guido Consonni and Piero Veronese. A note on coherent invariant distributions as non-informative priors for exponential and location-scale families. Communications in Statistics: Theory and Methods, 18(8):2883–2907, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coolen:1985:NLA

[CVDS85] H. C. C. H. Coolen, J. L. A. Van Rijckevorsel, H. J. Duivenvoorden, and A. H. B. Schuurs. A non linear analysis of the preference for preservation of the teeth. Communications in Statistics: Theory and Methods, 14(2):437–450, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Callaert:1982:CTL

[CVV82] H. Callaert, M. Vandemaele, and N. Veraverbeke. A Cramér type large deviation theorem for trimmed linear combinations of order statistics. Communications in Statistics: Theory and Methods, 11 (23):2689–2698, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1980:BSE

[CW80] Evan E. Chen and Robert L. Wardrop. Bayes sequential estimation in a life test and asymptotic properties. Communications in Statistics: Theory and Methods, 9(6):659–672, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cohen:1982:MML

[CW82] Clifford A. Cohen and Betty Whitten. Modified maximum likelihood and modified moment estimators for the three-parameter Weibull distribution. Communications in Statistics: Theory and Methods, 11(23):2631–2656, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Costello:1987:END

[CW87] Patricia S. Costello and Douglas A. Wolfe. An extreme null distribution approach to the problem of paired comparisons. Communications in Statistics: Theory and Methods, 16(8):2327–2341, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See note [Dav88b].

Chan:1988:LEL

[CW88] Lai K. Chan and Cheng W. Linear estimation of the location and scale parameters based on selected order statistics. Communications in Statistics: Theory and Methods, 17(7):2259–2278, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carter:1986:LSI

[CY86] Randy L. Carter and Mark C. K. Yang. Large sample inference in random coefficient regression models. Communications in Statistics: Theory and Methods, 15(8):2507–2525, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1987:AMM

[CZ87] Keh-Wei Chen and Z. Zheng. Admissible minimax model selection rule. Communications in Statistics: Theory and Methods, 16(11): 3393–3401, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Czitrom:1988:MEP

[Czi88] Veronica Czitrom. Mixture experiments with process variables: d-optimal orthogonal experimental designs. Communications in Statistics: Theory and Methods, 17(1):105–121, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Czitrom:1989:ATM

[Czi89a] Veronica Czitrom. An application of Taguchi’s methods reconsidered. Communications in Statistics: Theory and Methods, 18(11):

4105–4119, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Czitrom:1989:EDF

[Czi89b] Veronica Czitrom. Experimental design for four mixture components with process variables. Communications in Statistics: Theory and Methods, 18(12):4561–4581, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Czitrom:1989:TML

[Czi89c] Veronica Czitrom. Taguchi methods: linear graphs of high resolution. Communications in Statistics: Theory and Methods, 18(12): 4583–4606, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Driscoll:1980:PES

[DA80] Michael F. Driscoll and Dennis J. Anderson. Point-of-expansion, structure, and selection in multivariable polynomial regression. Communications in Statistics: Theory and Methods, 9(8):821–836, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahiya:1986:IPE

[Dah86] Ram C. Dahiya. Integer-parameter estimation in discrete distributions. Communications in Statistics: Theory and Methods, 15 (3):709–725, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahel:1987:MML

[Dah87] S. Dahel. On the minimaxity of the maximum likelihood estimator in a multivariate problem. Communications in Statistics: Theory and Methods, 16(5):1289–1296, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahel:1988:ATM

[Dah88] S. Dahel. Admissible tests for the mean with additional information. Communications in Statistics: Theory and Methods, 17(6): 1965–1972, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dale:1981:PSB

[Dal81] Borowiak Dale. A procedure for selecting between two regression models. Communications in Statistics: Theory and Methods, 10 (12):1197–1203, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Daly:1988:COH

[Dal88] James E. Daly. The construction of optimal histograms. Communications in Statistics: Theory and Methods, 17(9):2921–2931, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Danaher:1988:PED

[Dan88] Peter J. Danaher. Parameter estimation for the Dirichletmultinomial distribution using supplementary beta-binomial data. Communications in Statistics: Theory and Methods, 17(6):1777– 1788, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1981:STP

[Dav81] John C. Davis. Statistical techniques in petroleum exploration. Communications in Statistics: Theory and Methods, 10(15):1479– 1503, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1984:NPA

[Dav84] Barry R. Davis. Nonparametric prediction analysis for binary data. Communications in Statistics: Theory and Methods, 13(9):1161– 1175, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1985:VTF

[Dav85a] H. T. David. Variations on two familiar industrial statistics themes. Communications in Statistics: Theory and Methods, 14(11):2607– 2620, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1985:AMA

[Dav85b] Linda June Davis. Alternative moment approximations for Berkson’s minimum logit chi-squared estimator. Communications in Statistics: Theory and Methods, 14(7):1693–1708, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1986:WNC

[Dav86] Linda June Davis. Whittemore’s notion of collapsibility in multidimensional contingency tables. Communications in Statistics:

Theory and Methods, 15(8):2541–2554, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1987:BEC

[Dav87] H. T. David. Bayes extraction of component failure information from Boolean module test data. Communications in Statistics: Theory and Methods, 16(10):2873–2883, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1988:GBI

[Dav88a] H. A. David. General bounds and inequalities in order statistics. Communications in Statistics: Theory and Methods, 17(7):2119– 2134, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1988:NEN

[Dav88b] H. A. David. Note on “An extreme null distribution approach to the problem of paired comparisons” by Patricia S. Costello and Douglas A. Wolfe. Communications in Statistics: Theory and Methods, 17(11):4005–4009, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [CW87].

Davis:1989:TSP

[Dav89] Charles S. Davis. Two-sample post-stratified or subgroup analysis with restricted randomization rules. Communications in Statistics: Theory and Methods, 18(1):367–378, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dasgupta:1986:EMG

[DB86] Anirban Dasgupta and James Berger. Estimation of multiple gamma scale-parameters: Bayes estimation subject to uniform domination. Communications in Statistics: Theory and Methods, 15(7):2065–2086, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duran:1983:CEE

[DBO83] B. S. Duran, T. L. Boullion, and P. L. Odell. On a criterion for evaluating an estimator. Communications in Statistics: Theory and Methods, 12(10):1191–1208, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dusit:1984:EST

[DC84a] Charernkavanich Dusit and A. Clifford Cohen. Estimation in the singly truncated Weibull distribution with an unknown truncation point. Communications in Statistics: Theory and Methods, 13(7): 843–857, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1984:ITA

[DC84b] Danny Dyer and Paul Chiou. An information-theoretic approach to incorporating prior information in binomial sampling. Communications in Statistics: Theory and Methods, 13(17):2051–2083, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeJongh:1985:MCC

[DD85a] P. J. De Jongh and T. De Wet. A Monte Carlo comparison of regression trimmed means. Communications in Statistics: Theory and Methods, 14(10):2457–2472, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeJongh:1985:TMB

[DD85b] P. J. De Jongh and T. De Wet. Trimmed mean and bounded influence estimators for the parameters of the AR(1) process. Communications in Statistics: Theory and Methods, 14(6):1361–1375, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLong:1981:CPS

[De 81] David M. De Long. Crossing probabilities for a square root boundary by a Bessel process. Communications in Statistics: Theory and Methods, 10(21):2197–2213, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DelaHorra:1989:ARR

[De 89] J. De la Horra. An asymptotic result with respect to predictive distributions. Communications in Statistics: Theory and Methods, 18(2):515–521, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DelPino:1981:ANN

[Del81] Guido E. Del Pino. On the asymptotic non null distribution of some tests for homogeneity of variances and related results. Communications in Statistics: Theory and Methods, 10(3):255–267, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeMets:1984:SGV

[DeM84] David L. DeMets, Ph.D. Stopping guidelines vs stopping rules: a practitioner’s point of view. Communications in Statistics: Theory and Methods, 13(19):2395–2417, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dennis:1989:SPA

[Den89] Samuel Y. Dennis. Some probabilities associated with the ordering of random variables having a Dirichlet distribution. Communications in Statistics: Theory and Methods, 18(1):157–170, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DePriest:1983:UST

[DeP83] Douglas J. DePriest. Using the singly truncated normal distribution to analyze satellite data. Communications in Statistics: Theory and Methods, 12(3):263–272, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshmukh:1984:MLE

[Des84] Shailaja R. Deshmukh. Maximum likelihood estimation for branching migration process. Communications in Statistics: Theory and Methods, 13(14):1769–1780, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1983:CCO

[Dey83a] Dipak K. Dey. On the choice of co-ordinates in simultaneous estimation of normal means under misspecification of normal priors. Communications in Statistics: Theory and Methods, 12(6): 661–673, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1983:TME

[Dey83b] Dipak K. Dey. On truncation of multiparameter estimator in discrete exponential families. Communications in Statistics: Theory and Methods, 12(1):89–101, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1989:ESM

[Dey89] Dipak K. Dey. On estimation of the scale matrix of the multivariate F distribution. Communications in Statistics: Theory and Methods, 18(4):1373–1383, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1989:OAS

[DF89] Miles Davis and Robert H. Finch, Jr. Optimal allocation of stratified samples with several variance constraints and equal workloads over time by geometric programming. Communications in Statistics: Theory and Methods, 18(4):1507–1520, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dobrigai:1987:LCC

[DFG87a] A. Dobrigai, D. A. S. Fraser, and R. Gebotys. Linear calibration and conditional inference. Communications in Statistics: Theory and Methods, 16(4):1037–1048, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dobriyal:1987:ACI

[DFG87b] A. Dobriyal, D. A. S. Fraser, and R. Gebotys. Approximate conditional inference and the linear functional model. Communications in Statistics: Theory and Methods, 16(12):3729–3737, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dawid:1981:CBI

[DG81] A. P. Dawid and I. Guttman. Conjugate Bayesian inference for structural models. Communications in Statistics: Theory and Methods, 10(8):739–748, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dana:1983:HNI

[DG83] Jacqueline Lan Dana and D. V. Gokhale. On the hypothesis of no interaction on a linear scale in contingency tables. Communications in Statistics: Theory and Methods, 12(11):1225–1240, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Denzau:1989:BEP

[DGG89] Arthur T. Denzau, Patrick C. Gibbons, and Edward Greenberg. Bayesian estimation of proportions with a cross-entropy prior. Communications in Statistics: Theory and Methods, 18(5):1843– 1861, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1989:EGH

[DH89] Danny Dyer and Larry F. Heath. Estimation of gene/haplotype frequencies in genetic marker systems based on phenotype data. Communications in Statistics: Theory and Methods, 18(11):3927– 3947, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dielman:1984:LAV

[Die84a] Terry E. Dielman. Least absolute value estimation in regression models: an annotated bibliography. Communications in Statistics: Theory and Methods, 13(4):513–541, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dietz:1984:LSR

[Die84b] E. Jacquelin Dietz. Linear signed rank tests for multivariate location. Communications in Statistics: Theory and Methods, 13(12): 1435–1451, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dietz:1989:MGJ

[Die89] E. Jacquelin Dietz. Multivariate generalizations of Jonckheere’s test for ordered alternatives. Communications in Statistics: Theory and Methods, 18(10):3763–3783, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dinse:1987:DHC

[Din87a] Gregg E. Dinse. Discussion of “Handling cause of death in equivocal cases using the EM algorithm” by R. L. Kodel and J. J. Chen. Communications in Statistics: Theory and Methods, 16(9):2587– 2592, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dinse:1987:FDH

[Din87b] Gregg E. Dinse. Further discussion of “Handling cause of death in equivocal cases using the EM algorithm” by R. L. Kodell and J. J. Chen. Communications in Statistics: Theory and Methods, 16 (9):2597–2601, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dixit:1989:EPG

[Dix89] Ulhas J. Dixit. Estimation of parameters of the gamma distribution in the presence of outliers. Communications in Statistics: Theory and Methods, 18(8):3071–3085, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dharmadhikari:1989:UBV

[DJD89] S. W. Dharmadhikari and Kumar Joag-Dev. Upper bounds for the variances of certain random variables. Communications in Statistics: Theory and Methods, 18(9):3235–3247, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1981:RBE

[DK81] Danny Dyer and P. Keating. On the relative behavior of estimators of the characteristic life in the exponential failure model. Communications in Statistics: Theory and Methods, 10(5):489– 501, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshpande:1983:LCT

[DK83] Jayant V. Deshpande and Subhash C. Kochar. A linear combination of two U-statistics for testing new better than used. Communications in Statistics: Theory and Methods, 12(2):153–159, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deo:1985:COM

[DK85] S. S. Deo and A. V. Kharshikar. On consequences of one missing observation in a balanced design. Communications in Statistics: Theory and Methods, 14(12):3091–3100, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donner:1989:EIS

[DK89] Allan Donner and John J. Koval. The effect of imbalance on significance-testing in one-way Model II analysis of variance. Communications in Statistics: Theory and Methods, 18(4):1239–1250, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demets:1984:OSM

[DL84] David L. Demets, Ph.D. and Gordon K. K. Lan, Ph.D. An overview of sequential methods and their application in clinical trials. Communications in Statistics: Theory and Methods, 13(19):2315–2338, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dean:1986:NCG

[DL86] A. M. Dean and S. M. Lewis. A note on the connectivity of generalized cyclic designs. Communications in Statistics: Theory and Methods, 15(11):3429–3433, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Driscoll:1988:CSR

[DL88] Michale F. Driscoll and W. D. Laigh. Characterizations and O-statistic representations of LU-statistics. Communications in Statistics: Theory and Methods, 17(8):2763–2777, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaHorra:1988:BMI

[dlHRR88] J. de la Horra and C. Ruiz-Rivas. A Bayesian method for inferring the degree of dependence for a positively quadrant dependent distribution. Communications in Statistics: Theory and Methods, 17 (12):4357–4370, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dickey:1985:BED

[DLP85] J. M. Dickey, D. V. Lindley, and S. J. Press. Bayesian estimation of the dispersion matrix of a multivariate normal distribution. Communications in Statistics: Theory and Methods, 14(5):1019–1034, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Don:1980:UIG

[DM80] F. J. Henk Don and Jan R. Magnus. On the unbiasedness of iterated GLS estimators. Communications in Statistics: Theory and Methods, 9(5):519–527, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ducharme:1987:SAP

[DM87] Gilles R. Ducharme and Philip Milasevic. Some asymptotic properties of the circular median. Communications in Statistics: Theory and Methods, 16(3):659–664, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duran:1989:CMM

[DM89] Benjamin S. Duran and Chi-Chung Mitchell. A class of multisample multivariate nonparametric tests for location-scale. Communications in Statistics: Theory and Methods, 18(1):67–84, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dale:1981:EPQ

[DMH81] Umbach Dale, Masoom Ali Mir, and Khatab M. Hassanein. Estimating Pareto quantiles using two order statistics. Communications in Statistics: Theory and Methods, 10(19):1933–1941, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davies:1980:SMA

[DN80] Neville Davies and Paul Newbold. Sample moments of the autocorrelations of moving average processes and a modification to Bartlett’s asymptotic variance formula. Communications in Statistics: Theory and Methods, 9(14):1473–1481, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dorfman:1988:NME

[Dor88] Alan Dorfman. A note on Miller’s empirical weights for heteroscedastic linear regression. Communications in Statistics: Theory and Methods, 17(10):3521–3535, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dorsett:1989:BLW

[Dor89] Dovalee Dorsett. Bounded-leverage weights for robust regression estimators. Communications in Statistics: Theory and Methods, 18 (8):2785–2800, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Douglas:1986:UFC

[Dou86] J. B. Douglas. User friendly computerized maximum likelihood estimation. Communications in Statistics: Theory and Methods, 15(3):727–745, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dalai:1983:NBI

[DP83] S. R. Dalai and E. G. Phadia. Nonparametric Bayes inference for concordance in bivariate distributions. Communications in Statistics: Theory and Methods, 12(8):947–963, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dabrowska:1981:RMA

[DPS81] D. Dabrowska, E. Pleszczyńska, and W. Szczesny. Remarks on multivariate analogues of Kendall’s tau. Communications in Statistics: Theory and Methods, 10(23):2435–2445, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

duPreez:1987:EMN

[dPV87] J. P. du Preez and J. H. Venter. An estimator for the means of normal populations. Communications in Statistics: Theory and Methods, 16(2):593–612, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donner:1982:MVP

[DR82] Allan Donner and Bernard Rosner. Missing value problems in multiple linear regression with two independent variables. Communications in Statistics: Theory and Methods, 11(2):127–140, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dufour:1986:GPS

[DR86] Jean-Marie Dufour and Roch Roy. Generalized portmanteau statistics and tests of randomness. Communications in Statistics: Theory and Methods, 15(10):2953–2972, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delRio:1988:PEL

[dR88] Manuel del Rio. On the potential in the estimation of linear functions in regression. Communications in Statistics: Theory and Methods, 17(3):729–738, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delRio:1989:CFA

[dR89] Manuel del Rio. A coordinate-free approach to testable hypotheses. Communications in Statistics: Theory and Methods, 18(9):3493– 3503, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drew:1985:PWL

[Dre85] James H. Drew. Performance of the weighted least squares approach to categorical data analysis. Communications in Statistics: Theory and Methods, 14(8):1963–1979, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunn:1980:CCM

[DS80] C. L. Dunn and W. B. Smith. Combinatoric classification of multivariate normal variates. Communications in Statistics: Theory and Methods, 9(13):1317–1340, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLong:1981:EBP

[DS81] Elizabeth Ray DeLong and Pranab Kumar Sen. Estimation of P {X > Y } based on progressively truncated versions of the Wilcoxon–Mann–Whitney statistics. Communications in Statistics: Theory and Methods, 10(10):963–981, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunn:1982:NCC

[DS82] C. L. Dunn and W. B. Smith. Normal combinatoric classification: The sample case. Communications in Statistics: Theory and Methods, 11(3):271–289, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshpande:1985:TSO

[DS85] Jayant V. Deshpande and Harshinder Singh. Testing for second order stochastic dominance. Communications in Statistics: Theory and Methods, 14(4):887–893, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1986:EPL

[DS86a] Dipak K. Dey and C. Srinivasan. On estimation of Poisson logits. Communications in Statistics: Theory and Methods, 15(7):2087– 2097, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diana:1986:TTM

[DS86b] Giancarlo Diana and Alessandra Salvan. Two tests for monotone failure rate with incomplete data. Communications in Statistics: Theory and Methods, 15(12):3545–3558, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1987:GPO

[DS87] Miles Davis and Rudolph E. Schwartz. Geometric programming for optimal allocation of integrated samples in quality control. Communications in Statistics: Theory and Methods, 16(11):3235–3254, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duffy:1989:SSP

[DS89] Diane E. Duffy and Thomas J. Santner. On the small sample properties of norm-restricted maximum likelihood estimators for logistic regression models. Communications in Statistics: Theory and Methods, 18(3):959–980, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borges:1984:STN

[dSBPR84] Wagner de Souza Borges, Frank Proschan, and Josemar Rodrigues. A simple test for new better than used in expectation. Communications in Statistics: Theory and Methods, 13(25):3217–3223, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1981:EBB

[DSS81] A. Dey, M. Singh, and G. M. Saha. Efficiency balanced block designs. Communications in Statistics: Theory and Methods, 10 (3):237–247, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dudewicz:1981:MSM

[DT81] Edward J. Dudewicz and Vidya S. Taneja. A multivariate solution of the multivariate ranking and selection problem. Communications in Statistics: Theory and Methods, 10(18):1849–1868, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duffy:1983:MQS

[Duf83] Stephen W. Duffy. On the merging of quoted survival rates to obtain an overall rate. Communications in Statistics: Theory and Methods, 12(23):2753–2760, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunkl:1981:ADB

[Dun81] Charles F. Dunkl. The absorption distribution and the q-binomial theorem. Communications in Statistics: Theory and Methods, 10 (19):1915–1920, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunn:1982:CCL

[Dun82a] Charles L. Dunn. Comparison of combinatoric and likelihood ratio procedures for classifying samples. Communications in Statistics: Theory and Methods, 11(21):2361–2377, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunnett:1982:RMC

[Dun82b] Charles W. Dunnett. Robust multiple comparisons. Communications in Statistics: Theory and Methods, 11(22):2611–2629, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunn:1986:AMC

[Dun86] C. L. Dunn. Application of multiple comparison type procedures to the eigenvalues of Σ_{1}^{−1}Σ_{2}. Communications in Statistics: Theory and Methods, 15(2):451–471, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Damico:1987:ETE

[DW87a] Joseph A. Damico and Douglas A. Wolfe. Extended tables of the exact distribution of a rank statistic for all treatments multiple comparisons in one-way layout designs. Communications in Statistics: Theory and Methods, 16(8):2343–2360, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DuBien:1987:CAC

[DW87b] Janice L. DuBien and William D. Warde. A comparison of agglomerative clustering methods with respect to noise. Communications in Statistics: Theory and Methods, 16(5):1433–1460, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Damico:1989:ETE

[DW89a] Joseph A. Damico and Douglas A. Wolfe. Extended tables of the exact distribution of a rank statistic for treatments versus control multiple comparisons in one- way layout designs. Communications in Statistics: Theory and Methods, 18(9):3327–3353, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drezner:1989:AMB

[DW89b] Zvi Drezner and George O. Wesolowsky. An approximation method for bivariate and multivariate normal equiprobability contours.

Communications in Statistics: Theory and Methods, 18(6):2331– 2344, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deWaal:1985:EBA

[dWM85] D. J. de Waal and J. S. Maritz. An empirical Bayes approach in estimating the parameters of two or more geometric distributions. Communications in Statistics: Theory and Methods, 14(12):2997– 3005, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eideh:1989:STP

[EA89] Abdul-Hakeem Abdul-Hay Eideh and Mohamed Salahuddin Ahmed. Some tests for the power series distributions in one parameter using the Kullback–Leibler information measure. Communications in Statistics: Theory and Methods, 18(10):3649–3663, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Easterling:1980:C

[Eas80] Robert G. Easterling. Comments. Communications in Statistics: Theory and Methods, 9(17):1853–1857, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1989:MLE

[EB89] Marc A. Evans and Douglas G. Bonett. Maximum likelihood estimation for the negative multinomial log-linear model. Communications in Statistics: Theory and Methods, 18(11):4059–4065, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1982:ONQ

[Ebr82] Nader Ebrahimi. The ordering of negative quadrant dependence. Communications in Statistics: Theory and Methods, 11(21):2389– 2399, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1984:MLE

[Ebr84] Nader Ebrahimi. Maximum likelihood estimation, from doubly censored samples, of the mean of a normal distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 13(5):651–659, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1985:NES

[Ebr85] Nader Ebrahimi. Nonparametric estimation of survival functions for incomplete observations when the life time distribution is proportionally related to the censoring time distribution. Communications in Statistics: Theory and Methods, 14(12):2887–2898, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1986:PAT

[Ebr86] Nader Ebrahimi. A parametric approach to theory of competing risks when the time of death from each cause follows a changepoint hazard rate model. Communications in Statistics: Theory and Methods, 15(1):215–229, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escobar:1983:SIP

[EC83] Luis A. Escobar and James L. Cornette. Spacing of information in polynomial regression: a simple solution. Communications in Statistics: Theory and Methods, 12(17):2035–2042, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elperin:1988:LCL

[EG88] T. Elperin and I. Gertsbakh. Lower confidence limit for the availability of a parallel system with renewable components. Communications in Statistics: Theory and Methods, 17(2):311–323, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1983:TLC

[EH83a] Nader Ebrahimi and Mohamed Habibullah. Testing for linearity and convexity of potency curve. Communications in Statistics: Theory and Methods, 12(6):693–700, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Edwards:1983:ASP

[EH83b] Don Edwards and Jason C. Hsu. Adaptive sequential procedures for comparing new treatments with a standard. Communications in Statistics: Theory and Methods, 12(10):1135–1145, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1987:SEE

[EH87] Nader Ebrahimi and Balakrishna Hosmane. On shrinkage estimation of the exponential location parameter. Communications in Statistics: Theory and Methods, 16(9):2623–2637, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Helbawy:1984:ODR

[EHA84] Abdalla T. El-Helbawy and Essam A. Ahmed. Optimal design results for 2^{n} factorial paired comparison experiments. Communications in Statistics: Theory and Methods, 13(22):2827–2845, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eisenberg:1981:SSC

[Eis81] Bennett Eisenberg. On the sample size of curtailed tests. Communications in Statistics: Theory and Methods, 10(21):2177–2196, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Engeman:1982:GLS

[EK82] Richard M. Engeman and Thomas J. Keefe. On generalized least squares estimation of the Weibull distribution. Communications in Statistics: Theory and Methods, 11(19):2181–2193, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1985:HOG

[EK85] Merran A. Evans and Maxwell L. King. Higher order generalisation of first order autoregressive tests. Communications in Statistics: Theory and Methods, 14(12):2907–2918, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eve:1988:SCC

[EK88] Bofinger Eve and Mengersen Kerrie. Simultaneous comparisons with a control and with the “best”. Communications in Statistics: Theory and Methods, 17(5):1507–1540, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Euhyoon:1988:MVU

[EkBL88] Park Euhyoon and Shaul k Bar-Lev. Minimum variance unbiased estimation in natural exponential families generated by stable distributions. Communications in Statistics: Theory and Methods, 17 (12):4301–4313, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ellerton:1986:COO

[EKR86] Roger R. W. Ellerton, Christos P. Kitsos, and Stefan Rinco. Choosing the optimal order of a response polynomial — structural approach with minimax criterion. Communications in Statistics: Theory and Methods, 15(1):129–136, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElBassiouni:1981:LRT

[El 81] Mohamed Y. El Bassiouni. Likelihood ratio tests for covariance hypotheses generating commutative quadratic subspaces. Communications in Statistics: Theory and Methods, 10(23):2461–2468, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElBassiouni:1982:TG

[El 82] Mohamed Y. El Bassiouni. On a theorem of Graybill. Communications in Statistics: Theory and Methods, 11(13):1519–1522, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

elLozy:1982:SCB

[eL82a] M. D. Mohamed el Lozy. Simple computation of a bivariate normal integral arising from a problem of misclassification with applications to the diagnosis of hypertension. Communications in Statistics: Theory and Methods, 11(19):2195–2205, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eubank:1982:LSP

[EL82b] R. L. Eubank and V. N. LaRiccia. Location and scale parameter estimation from randomly censored data. Communications in Statistics: Theory and Methods, 11(25):2869–2888, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elphinstone:1983:TDM

[Elp83] C. D. Elphinstone. A target distribution model for nonparametric density estimation. Communications in Statistics: Theory and Methods, 12(2):161–198, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elpeit:1989:LSM

[Elp89] Barbel Elpeit. On linear statistical models of commutative quadratic type. Communications in Statistics: Theory and Methods, 18(9):3407–3450, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Neweihi:1980:ESM

[EN80a] Emad El-Neweihi. Extensions of a simple model: with applications in reliability, extinction of species, inventory depletion and urn sampling. Communications in Statistics: Theory and Methods, 9 (4):399–414, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1980:BPT

[EN80b] I. G. Evans and A. M. Nigm. Bayesian prediction for two-parameter Weibull lifetime models. Communications in Statistics: Theory and Methods, 9(6):649–658, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Neweihi:1981:SOC

[EN81] Emad El-Neweihi. Stochastic ordering and a class of multivariate new better than used distributions. Communications in Statistics: Theory and Methods, 10(16):1655–1672, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Neweihi:1984:DSS

[ENP84] Emad El-Neweihi and Frank Proschan. Degradable systems: a survey of multistate system theory. Communications in Statistics: Theory and Methods, 13(4):405–432, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Epps:1985:PET

[EP85] T. W. Epps and L. R. Pulley. Parameter estimates and tests of fit for infinite mixture distributions. Communications in Statistics: Theory and Methods, 14(12):3125–3145, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eccleston:1980:ORC

[ER80] J. A. Eccleston and K. G. Russell. (M, S)-optimal row-column designs. Communications in Statistics: Theory and Methods, 9 (4):449–452, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1983:BIG

[ER83] I. G. Evans and A. S. Ragab. Bayesian inferences given a type-2 censored sample from a Burr distribution. Communications in Statistics: Theory and Methods, 12(13):1569–1580, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eric:1983:ETM

[Eri83] Peritz Eric. Exact tests for matched pairs: studies with covariates. Communications in Statistics: Theory and Methods, 12(10):1209– 1210, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escobar:1984:CFS

[ES84] Luis A. Escobar and Bradley Skarpness. A closed form solution for the least squares regression problem with linear inequality constraints. Communications in Statistics: Theory and Methods, 13 (9):1127–1134, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escobar:1987:MSE

[ES87] Luis A. Escobar and Bradley Skarpness. Mean square error and efficiency of the least squares estimator over interval constraints. Communications in Statistics: Theory and Methods, 16(2):397– 406, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Emanuel:1988:NRS

[ES88] David C. Emanuel and Kenneth H. Sutrick. Non-random shuffling strategies in blackjack. Communications in Statistics: Theory and Methods, 17(9):2953–2958, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eakin:1984:CBD

[ESP84] Mark Eakin, Craig Slinkman, and Elinor S. Pape. A cost based decision rule for deciding between two estimators of μ^{y} −μ^{x}. Communications in Statistics: Theory and Methods, 13(1):45–58, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Etemadi:1984:MIA

[Ete84] Nasrollah Etemadi. On the maximal inequalities for the average of pairwise i.i.d. random variables. Communications in Statistics:

Theory and Methods, 13(22):2749–2756, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Etermsdi:1987:SIR

[Ete87] N. Etermsdi. On sums of independent random vectors. Communications in Statistics: Theory and Methods, 16(1):241–252, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eubank:1983:NOR

[Eub83] R. L. Eubank. A note on optimal and robust spacing selection. Communications in Statistics: Theory and Methods, 12(21):2483– 2491, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eubank:1984:ARM

[Eub84] R. L. Eubank. Approximate regression models and splines. Communications in Statistics: Theory and Methods, 13(4):433–484, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1982:CBB

[Eva82] Michael Evans. Confidence bands for bivariate quantiles. Communications in Statistics: Theory and Methods, 11(13):1465–1474, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eichhorn:1981:BSS

[EZ81] B. H. Eichhorn and S. Zacks. Bayes sequential search of an optimal dosage: linear regression with both parameters unknown. Communications in Statistics: Theory and Methods, 10(10):931– 953, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimi:1989:TBG

[EZ89] Nader Ebrahimi and Hassan Zahedi. Testing for bivariate Gumbel against bivariate new better than used in expectation. Communications in Statistics: Theory and Methods, 18(4):1357–1371, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Falk:1986:ABA

[Fal86] Micheal Falk. On the accuracy of the bootstrap approximation of the joint distribution of sample quantiles. Communications in Statistics: Theory and Methods, 15(10):2867–2876, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1983:ERC

[Far83] R. W. Farebrother. An examination of recent criticisms of ridge regression simulation designs. Communications in Statistics: Theory and Methods, 12(22):2549–2555, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1984:RLS

[Far84] R. W. Farebrother. The restricted least squares estimator and ridge regression. Communications in Statistics: Theory and Methods, 13 (2):191–196, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1985:UE

[Far85] R. W. Farebrother. Unbiased L_{1} and L_{∞} estimation. Communications in Statistics: Theory and Methods, 14(8):1941–1962, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1986:TLI

[Far86] R. W. Farebrother. Testing linear inequality constraints in the standard linear model. Communications in Statistics: Theory and Methods, 15(1):07–31, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1987:FRT

[Far87] R. W. Farebrother. A further remark on testing linear inequality constraints in the standard linear model. Communications in Statistics: Theory and Methods, 16(7):2003–2005, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1988:ELS

[Far88] R. W. Farebrother. Elemental location shift estimators in the constrained linear model. Communications in Statistics: Theory and Methods, 17(1):79–85, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Foutz:1982:TTS

[FB82] Robert V. Foutz and Jeffrey B. Birch. Tests for the two-sample problem based on empirical probability measures. Communications in Statistics: Theory and Methods, 11(16):1839–1853, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1989:RTS

[FB89] K. Y. Fung and N. Balakrishnan. On the robust two sided tolerance limits based on MML estimators. Communications in Statistics: Theory and Methods, 18(3):1135–1138, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fabozzi:1984:SSR

[FBV84] Frank J. Fabozzi, Raj Baldev, and Hrishikesh D. Vinod. The stability of the systematic risk of individual stocks: an application of ridge regression. Communications in Statistics: Theory and Methods, 13(2):151–161, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fahmy:1982:ISP

[FCP^{+}82] Salwa Fahmy, A. Carlos, B. Pereira, Frank Proschan, and Moshe Shaked. The influence of the sample on the posterior distribution. Communications in Statistics: Theory and Methods, 11(16):1757– 1768, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fuchs:1985:PPP

[FD85] Camil Fuchs and Norman R. Draper. Percentage points and power calculations for outlier tests in a regression situation. Communications in Statistics: Theory and Methods, 14(10):2473–2487, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Federer:1987:SSL

[Fed87] Walter T. Federer. On screening samples in the laboratory and factors in factorial investigations. Communications in Statistics: Theory and Methods, 16(10):3033–3049, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feingold:1982:MDL

[Fei82] Marcia Feingold. Missing data in linear models with correlated errors. Communications in Statistics: Theory and Methods, 11 (24):2831–2843, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feiveson:1984:WRE

[Fei84] Alan H. Feiveson. Weighted ratio estimation in agricultural surveys. Communications in Statistics: Theory and Methods, 13(23): 2937–2959, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feingold:1988:MPT

[Fei88] Marcia Feingold. A more powerful test for incomplete block designs. Communications in Statistics: Theory and Methods, 17(9): 3107–3119, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fenech:1983:PRE

[Fen83] Alan Paul Fenech. A pseudo-replicate estimator of the variance of a function of sample means. Communications in Statistics: Theory and Methods, 12(7):777–798, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferretti:1987:ETB

[Fer87] Nélida E. Ferretti. Efficiency of tests based on weighted rankings for randomized blocks when the alternatives have an a priori ordering. Communications in Statistics: Theory and Methods, 16 (6):1629–1653, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feuerverger:1987:NDM

[Feu87] Andrey Feuerverger. New derivations of the maximum likelihood estimator and the likelihood ratio test. Communications in Statistics: Theory and Methods, 16(5):1241–1251, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fligner:1986:SER

[FF86] Michael A. Fligner and David Fairley. Stochastic equivalence of ranking methods. Communications in Statistics: Theory and Methods, 15(6):1855–1863, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Friedman:1981:PEMa

[FG81] L. Friedman and I. Gertsbakh. Parameter estimations in a minimum-type scheme. Communications in Statistics: Theory and Methods, 10(5):439–462, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feuerverger:1988:ANP

[FG88] Andrey Feuerverger and Sucharita Ghosh. An asymptotic Neyman–Pearson type result under symmetry constraints. Communications in Statistics: Theory and Methods, 17(5):1557–1564, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Friedman:1988:MKS

[FGP88] Yaakov Friedman, Alexander Gelman, and Eswar Phadia. A modified Kolmogorov–Smirnov type goodness of fit test. Communications in Statistics: Theory and Methods, 17(12):4147–4162, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Furnival:1989:CIS

[FGV89] George M. Furnival, Timothy G. Gregoire, and Harry T. Valentine. Confidence intervals and significance tests for a single trial. Communications in Statistics: Theory and Methods, 18(10):3749–3761, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fleming:1981:CHT

[FH81] Thomas R. Fleming and David P. Harrington. A class of hypothesis tests for one and two sample censored survival data. Communications in Statistics: Theory and Methods, 10(8):763–794, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fleming:1984:NES

[FH84] Thomas R. Fleming and David P. Harrington. Nonparametric estimation of the survival distribution in censored data. Communications in Statistics: Theory and Methods, 13(20):2469–2486, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fairley:1987:LRS

[FH87] David Fairley and Michael Higner. Linear rank statistics for the ordered alternatives problem. Communications in Statistics: Theory and Methods, 16(1):1–16, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fick:1984:SLM

[Fic84] Gordon H. Fick. A study of linear model analysis by an error contour emulation. Communications in Statistics: Theory and Methods, 13(16):2029–2047, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Findley:1982:SPE

[Fin82] David F. Findley. A special property of the expected log likelihood. Communications in Statistics: Theory and Methods, 11(21):2379– 2387, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Findley:1984:MSC

[Fin84] David F. Findley. On the mean square convergence of the convolution representation of linear filters. Communications in Statistics: Theory and Methods, 13(9):1073–1087, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Findley:1985:BOP

[Fin85] David F. Findley. On backshift-operator polynomial transformations to stationarity for nonstationary time series and their aggregates. Communications in Statistics: Theory and Methods, 14(1): 49–61, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Firinguetti:1987:EML

[Fir87] Lusi Firinguetti. Exact moments of Lawless and Wang’s operational ridge regression estimator. Communications in Statistics: Theory and Methods, 16(3):731–745, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freimer:1988:SGT

[FKML88] Marshall Freimer, Georgia Kollia, Govind S. Mudholkar, and C. Thomas Lin. A study of the generalized Tukey lambda family. Communications in Statistics: Theory and Methods, 17(10): 3547–3567, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freimer:1989:EES

[FKML89] Marshall Freimer, Georgia Kollia, Govind S. Mudholkar, and C. Thomas Lin. Extremes, extreme spacings and outliers in the Tukey and Weibull families. Communications in Statistics: Theory and Methods, 18(11):4261–4274, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:1983:ADS

[FKN83] C. Fang, P. R. Krishnaiah, and B. N. Nagarsenker. Asymptotic distributions of the sphericity test in a complex multivariate normal distribution. Communications in Statistics: Theory and Methods, 12(3):288, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flack:1985:THW

[Fla85] Virginia Foard Flack. Testing hypotheses when there could be an initial effect. Communications in Statistics: Theory and Methods, 14(10):2345–2363, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flack:1987:CII

[Fla87] Virginia Foard Flack. Confidence intervals for the interrater agreement measure kappa. Communications in Statistics: Theory and Methods, 16(4):953–968, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flack:1989:PMR

[Fla89] Virginia F. Flack. Predictability measures for ridge regression models. Communications in Statistics: Theory and Methods, 18(2): 755–766, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Foster:1989:SMT

[FME89] Theda A. Foster, Donna L. Mohr, and Robert C. Elston. A statistical model of tracking. Communications in Statistics: Theory and Methods, 18(8):2861–2881, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fong:1989:BAC

[Fon89] Duncan K. H. Fong. A Bayesian approach to the comparison of two means. Communications in Statistics: Theory and Methods, 18 (10):3785–3800, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Foutz:1981:STF

[Fou81] Robert V. Foutz. Simultaneous tests for finite families of hypotheses. Communications in Statistics: Theory and Methods, 10(13):

1263–1275, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fox:1989:CMC

[Fox89] David R. Fox. Conditional multivariate calibration. Communications in Statistics: Theory and Methods, 18(6):2311–2330, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fairley:1984:BEP

[FP84] David Fairley and Dennis K. Pearl. The Bahadur efficiency of paired versus joint ranking procedures for pairwise multiple comparisons. Communications in Statistics: Theory and Methods, 13 (12):1471–1481, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fomby:1986:SEM

[FP86] Thomas B. Fomby and James E. Pearce. Standard errors in the multinomial logit model. Communications in Statistics: Theory and Methods, 15(8):2555–2568, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feigin:1989:DCE

[FP89a] Paul D. Feigin and A. N. Pettitt. Detection of concordance for extreme ranks. Communications in Statistics: Theory and Methods, 18(5):1805–1818, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:1989:SNS

[FP89b] Robert L. Fountain and Pramod K. Pathak. Systematic and nonrandom sampling in the presence of linear trends. Communications in Statistics: Theory and Methods, 18(7):2511–2526, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fligner:1983:IPK

[FR83] Michael A. Fligner and Steven W. Rust. On the independence problem and Kendall’s tau. Communications in Statistics: Theory and Methods, 12(14):1597–1607, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franklin:1988:CNU

[FR88] L. A. Franklin and G. D. Richardson. A comparison of nonparametric unweighted and length-biased density estimation of fibres.

Communications in Statistics: Theory and Methods, 17(11):3629– 3644, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frane:1980:SCM

[Fra80] James W. Frane. Some computing methods for unbalanced analysis of variance and covariance. Communications in Statistics: Theory and Methods, 9(2):151–166, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraiman:1983:GEA

[Fra83] Ricardo Fraiman. General M-estimators and applications to bounded influence estimation for non-linear regression. Communications in Statistics: Theory and Methods, 12(22):2617–2631, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frangos:1987:UBJ

[Fra87] Chris C. Frangos. An updated bibliography on the jackknife method. Communications in Statistics: Theory and Methods, 16 (6):1543–1584, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franklin:1988:NAC

[Fra88] L. A. Franklin. A note on approximations and convergence in distribution for Spearman’s rank correlation coefficient. Communications in Statistics: Theory and Methods, 17(1):55–59, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franklin:1989:CER

[Fra89] Le Roy A. Franklin. A comparison of estimators for randomized response sampling with continuous distributions from a dichotomous population. Communications in Statistics: Theory and Methods, 18(2):489–505, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freeman:1983:SEQ

[Fre83] James M. Freeman. Sampling experiments and the Quandt statistic. Communications in Statistics: Theory and Methods, 12(16): 1879–1887, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freeman:1984:TPR

[Fre84] James M. Freeman. Two-phase regression and goodness of fit. Communications in Statistics: Theory and Methods, 13(11):1321– 1334, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Friedman:1981:PEMb

[Fri81] Lea Friedman. Parameter estimation in a minimum-type model by the least-squares method — a Monte Carlo study. Communications in Statistics: Theory and Methods, 10(5):463–487, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1987:LPA

[Fri87] Hans Frick. On level and power of Anderson and Hauck’s procedure for testing equivalence in comparative bioavailability. Communications in Statistics: Theory and Methods, 16(9):2771–2778, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Froda:1987:SRT

[Fro87] Sorana Froda. A signed rank test for censored paired lifetimes. Communications in Statistics: Theory and Methods, 16(12):3497– 3517, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:1989:OSQ

[Fro89] Steven G. From. Optimal spacing of quantiles for the estimation of the mixing parameters in a mixture of two exponential distributions. Communications in Statistics: Theory and Methods, 18 (6):2201–2223, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraser:1980:FNB

[FS80] D. A. S. Fraser and F. Streit. A further note on the bivariate normal distribution. Communications in Statistics: Theory and Methods, 9(10):1097–1099, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraser:1986:PLF

[FS86] D. A. S. Fraser and S. E. Sackey. Prior-likelihood factorization and missing data. Communications in Statistics: Theory and Methods, 15(11):3321–3331, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:1980:PMP

[Fu80] James C. Fu. A probabilistic model for predicting total basal area growth of closed uneven-aged stand. Communications in Statistics: Theory and Methods, 9(8):865–874, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fligner:1987:ATG

[FV87] Michael A. Fligner and Joseph S. Verducci. Aspects of two group concordance. Communications in Statistics: Theory and Methods, 16(5):1479–1503, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fassler:1989:AID

[FV89] Albert Fässler and Arthur Vogt. Analytic integration of the Divisia price index on the exponential path for two commodities. Communications in Statistics: Theory and Methods, 18(9):3473–3476, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finkelstein:1986:IRI

[FW86] Dianne M. Finkelstein and Robert A. Wolfe. Isotonic regression for interval censored survival data using an E–M algorithm. Communications in Statistics: Theory and Methods, 15(8):2493–2505, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferretti:1986:ETB

[FY86] Nélida E. Ferretti and Víctor J. Yohai. Efficiency of tests based on weighted rankings for equality of treatment effects in a complete randomized blocks layout. Communications in Statistics: Theory and Methods, 15(4):1179–1200, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geng:1989:BEM

[GA89] Zhi Geng and Chooichiro Asano. Bayesian estimation methods for categorical data with misclassifications. Communications in Statistics: Theory and Methods, 18(8):2935–2954, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabriel:1980:NCR

[Gab80] K. Ruben Gabriel. On Neyman’s “comments” regarding the first Israeli randomized rainfall stimulation experiment. Communications in Statistics: Theory and Methods, 9(18):1963–1973, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabler:1987:NPS

[Gab87] Siegfried Gabler. The nearest proportional to size sampling design. Communications in Statistics: Theory and Methods, 16(4):1117– 1131, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gad:1983:SCE

[Gad83] Nathan Gad. A simulation comparison of estimators for a regression coefficient under differential non-response. Communications in Statistics: Theory and Methods, 12(6):645–659, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galpin:1980:NSC

[Gal80] J. S. Galpin. A note on some conditions for optimality of certain ridge regression estimates. Communications in Statistics: Theory and Methods, 9(10):1019–1024, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gallo:1982:CRE

[Gal82] Paul P. Gallo. Consistency of regression estimates when some variables are subject to error. Communications in Statistics: Theory and Methods, 11(9):973–983, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galambos:1988:VGD

[Gal88] Janes Galambos. Variants of the graph dependent model in extreme value theory. Communications in Statistics: Theory and Methods, 17(7):2211–2221, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gapinski:1984:RRR

[Gap84] James H. Gapinski. Residents, recreationists, ridge regression, and revenues: the five r’s of Florida’s motor-tax. Communications in Statistics: Theory and Methods, 13(2):163–171, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gather:1986:REM

[Gat86] Ursula Gather. Robust estimation of the mean of the exponential distribution in outlier situations. Communications in Statistics: Theory and Methods, 15(8):2323–2345, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gather:1989:TMC

[Gat89] Ursula Gather. Testing for multisource contamination in location/ scale families. Communications in Statistics: Theory and Methods, 18(1):1–34, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gokhale:1983:MRA

[GBS83] D. V. Gokhale, R. J. Beaver, and B. W. Sirotnik. Model-robust analysis of paired comparison experiments. Communications in Statistics: Theory and Methods, 12(1):25–36, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glazebrook:1980:DEE

[GC80] K. D. Glazebrook and T. F. Cox. On the design of efficient experiments for choosing between two Bernoulli populations. Communications in Statistics: Theory and Methods, 9(3):255–264, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1984:NCE

[GC84] B. C. Gupta and Ramirez S. S. Carvajal. A necessary condition for the existence of main effect plus one plans for 2^{m} factorials. Communications in Statistics: Theory and Methods, 13(5):567– 580, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gurland:1983:NDD

[GCH83] John Gurland, Evan E. Chen, and Fabian M. Hernandez. A new discrete distribution involving Laguerre polynomials. Communications in Statistics: Theory and Methods, 12(17):1987–2004, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gbur:1985:ELL

[GD85] E. E. Gbur and P. F. Dahm. Estimation of the linear-linear segmented regression model in the presence of measurement error.

Communications in Statistics: Theory and Methods, 14(4):809– 826, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gelfand:1988:IEV

[GD88] Alan E. Gelfand and Dipak K. Dey. Improved estimation of variance components in mixed models. Communications in Statistics: Theory and Methods, 17(10):3313–3331, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonin:1987:NAS

[GdT87] R. Gonin and S. H. C. du Toit. Numerical algorithms for solving nonlinear L_{p}-norm estimation problems: part II — a mixture method for large residual and ill-conditioned problems. Communications in Statistics: Theory and Methods, 16(4):969–986, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

George:1988:GMP

[GE88] Varghese T. George and Robert C. Elston. Generalized modulus power transformations. Communications in Statistics: Theory and Methods, 17(9):2933–2952, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gelfand:1983:END

[Gel83] Alan E. Gelfand. Estimation in noncentral distributions. Communications in Statistics: Theory and Methods, 12(4):463–475, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gelfand:1987:MSE

[Gel87] Alan E. Gelfand. Mean square error behavior for prediction in linear regression models. Communications in Statistics: Theory and Methods, 16(6):1717–1724, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geng:1988:MCT

[Gen88] Zhi Geng. Multidimensional contingency tables with missing data. Communications in Statistics: Theory and Methods, 17(12):4137– 4146, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

George:1986:FBM

[Geo86] Edward I. George. A formal Bayes multiple shrinkage estimator. Communications in Statistics: Theory and Methods, 15(7):2099– 2114, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gertsbakh:1988:EPU

[Ger88] I. Gertsbakh. On the equivalence of Pavlov–Ushakov and Kaplan– Meier estimators of survival probability in multiply censored sample. Communications in Statistics: Theory and Methods, 17(9): 2881–2886, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghangurde:1981:ESA

[GG81] P. D. Ghangurde and G. B. Gray. Estimation for small areas in household surveys. Communications in Statistics: Theory and Methods, 10(22):2327–2338, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1983:MRL

[GG83] Pushpa Lata Gupta and Ramesh C. Gupta. On the moments of residual life in reliability and some characterization results. Communications in Statistics: Theory and Methods, 12(4):449–461, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gambino:1986:BAP

[GG86] G. Jack Gambino and I. Guttman. A Bayesian approach to prediction in analysis of variance settings. Communications in Statistics: Theory and Methods, 15(2):493–504, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:EEC

[GG88] R. D. Gupta and Ramesh C. Gupta. Estimation of p{y_{p} > max(y_{1}, y_{2}, …, y_{p−1})} in the exponential case. Communications in Statistics: Theory and Methods, 17(3):911–924, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gart:1986:FRE

[GgT86] J. J. Gart, H. M. Petti grew, and D. G. Thomas. Further results on the effect of bias, variance estimation, and non-normality of the empirical logit on weighted least squares analyses. Communications in Statistics: Theory and Methods, 15(3):755–782, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerard:1984:MEC

[GH84] John B. Guerard, Jr. and Raymond L. Horton. The management of executive compensation in large, dynamic firms: a ridge regression estimation. Communications in Statistics: Theory and Methods, 13(2):183–190, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guttman:1985:SPP

[GH85] Irwin Guttman and P. Hougaard. Studentization and prediction problems in multivariate multiple regression. Communications in Statistics: Theory and Methods, 14(6):1251–1258, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gatsonis:1985:SNT

[GHK85] Constantine Gatsonis, H. K. Hsieh, and Ramesh Korwar. Simple nonparametric tests for a known standard survival based on censored data. Communications in Statistics: Theory and Methods, 14 (9):2137–2162, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1980:TNA

[Gho80] B. K. Ghosh. Two normal approximations to the binomial distribution. Communications in Statistics: Theory and Methods, 9 (4):427–438, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1981:EBE

[Gho81a] J. K. Ghorai. Empirical Bayes estimation of a distribution function with a gamma process prior. Communications in Statistics: Theory and Methods, 10(12):1239–1248, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:MND

[Gho81b] Malay Ghosh. Multivariate negative dependence. Communications in Statistics: Theory and Methods, 10(4):307–337, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:SPE

[Gho81c] Malay Ghosh. Sequential point estimation of the means of Ustatistics in finite population sampling. Communications in Statistics: Theory and Methods, 10(21):2215–2229, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:MD

[Gho81d] Pankaj Ghosh. On multimodal distributions. Communications in Statistics: Theory and Methods, 10(17):1789–1793, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1982:IOR

[Gho82] Subir Ghosh. Information in an observation in robust designs. Communications in Statistics: Theory and Methods, 11(10):1173– 1184, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1983:IOV

[Gho83a] Subir Ghosh. Influential observations in view of design and inference. Communications in Statistics: Theory and Methods, 12(14): 1675–1683, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1983:SPW

[Gho83b] Subir Ghosh. Statistical planning when the parameters in the model are of unequal importance. Communications in Statistics: Theory and Methods, 12(16):1841–1853, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1985:NEB

[Gho85] Malay Ghosh. Nonparametric empirical Bayes estimation of certain funtionals. Communications in Statistics: Theory and Methods, 14 (9):2081–2094, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1986:CCB

[GHO86] Alan J. Gross, Hurshell H. Hunt, and Robert E. Odeh. The correlation coefficient between the smallest and largest observations when (N − 1) of the n observations are iid exponentially distributed. Communications in Statistics: Theory and Methods, 15(4):1113– 1123, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1987:INP

[Gho87a] Subir Ghosh. Influential nonnegligible parameters under the search linear model. Communications in Statistics: Theory and Methods, 16(4):1013–1025, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1987:NOD

[Gho87b] Subir Ghosh. Non-orthogonal designs for measuring dispersion effects in sequential factor screening experiments using search linear models. Communications in Statistics: Theory and Methods, 16 (10):2839–2850, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1989:DMS

[Gho89a] J. K. Ghorai. Deficiency of the MLE of a smooth survival function under the proportional hazard model. Communications in Statistics: Theory and Methods, 18(8):3047–3056, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1989:NBE

[Gho89b] J. K. Ghorai. Nonparametric Bayesian estimation of a survival function under the proportional hazard model. Communications in Statistics: Theory and Methods, 18(5):1831–1842, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giani:1986:SIZ

[Gia86] Guido Giani. On a strengthening of the indifference zone approach to a generalized selection goal. Communications in Statistics: Theory and Methods, 15(10):3163–3171, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gibbons:1984:SCR

[Gib84] Diane I. Gibbons. Some characterizations of the ridge trace. Communications in Statistics: Theory and Methods, 13(2):173–182, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giesbrecht:1983:EPC

[Gie83] F. G. Giesbrecht. An efficient procedure for computing MINQUE of variance components and generalized least squares estimates of fixed effects. Communications in Statistics: Theory and Methods, 12(18):2169–2177, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gilula:1981:NMD

[Gil81] Zvi Gilula. A note on measuring the degree of dependence between two discrete random variables measured on a nominal scale. Communications in Statistics: Theory and Methods, 10(20):2047–2055, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gilula:1982:NAA

[Gil82] Zvi Gilula. A note on the analysis of association in crossclassifications having ordered categories. Communications in Statistics: Theory and Methods, 11(11):1233–1240, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gil:1989:NSG

[Gil89] Maria Angeles Gil. A note on stratification and gain in precision in estimating diversity from large samples. Communications in Statistics: Theory and Methods, 18(4):1521–1526, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gimotty:1987:ADG

[Gim87] Phyllis A. Gimotty. The asymptotic distribution of the goodnessof-fit chi-square statistic computed from imputed data. Communications in Statistics: Theory and Methods, 16(1):45–60, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giorgio:1988:VIL

[Gio88] Vittadini Giorgio. On the validity of the indeterminate latent variables in the LISREL model. Communications in Statistics: Theory and Methods, 17(3):861–874, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1981:PSG

[GK81] Shanti S. Gupta and Woo-Chul Kim. On the problem of selecting good populations. Communications in Statistics: Theory and Methods, 10(11):1043–1077, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gather:1988:MLE

[GK88a] U. Gather and B. K. Kale. Maximum likelihood estimation in the presence of outliers. Communications in Statistics: Theory and Methods, 17(11):3767–3784, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

George:1988:NBA

[GK88b] Varghese T. George and S. K. Katti. A non-Bayesian approach to estimation using tested priors. Communications in Statistics: Theory and Methods, 17(5):1629–1639, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:STH

[GL81] B. K. Ghosh and S. Keith Lee. A sequential test for homogeneity of multinomial proportions. Communications in Statistics: Theory and Methods, 10(9):883–895, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1987:ADF

[GL87] Shanti S. Gupta and Lii-Yuh Leu. An asymptotic distribution-free selection procedure for a two-way layout problem. Communications in Statistics: Theory and Methods, 16(8):2313–2325, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1989:EFH

[GL89] Rajan Gupta and Larry Lee. Estimating fault hitting rates by recapture sampling. Communications in Statistics: Theory and Methods, 18(6):2139–2153, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glaz:1988:SSM

[Gla88] Joseph Glaz. A survey of statistical methods on remote sensing in Russian literature from 1976 to 1985. Communications in Statistics: Theory and Methods, 17(8):2419–2454, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gleit:1982:EFP

[Gle82] Alan Gleit. Estimation of functions of the parameters of a normal distribution. Communications in Statistics: Theory and Methods, 11(24):2845–2855, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1989:NEB

[GLT89] Malay Ghosh, Parthasarthi Lahiri, and Ram C. Tiwari. Nonparametric empirical Bayes estimation of the distribution function and the mean. Communications in Statistics: Theory and Methods, 18 (1):121–146, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ganesalingam:1980:CMC

[GM80a] S. Ganesalingam and G. J. McLachlan. A comparison of the mixture and classification approaches to cluster analysis. Communications in Statistics: Theory and Methods, 9(9):923–933, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1980:SPI

[GM80b] J. K. Ghosh and N. Mukhopadhyay. Sequential procedures in identification. Communications in Statistics: Theory and Methods, 9 (15):1625–1639, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaver:1983:JKM

[GM83a] Donald P. Gaver and Rupert G. Miller, Jr. Jackknifing the Kaplan– Meier survival estimator for censored data: Simulation results and asymptotic analysis. Communications in Statistics: Theory and Methods, 12(15):1701–1718, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guttman:1983:ECM

[GM83b] Irwin Guttman and Ulrich Menzefricke. Estimating the common mean of r normal populations with different precisions. Communications in Statistics: Theory and Methods, 12(10):1147–1159, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1984:GWA

[GM84] Ramesh C. Gupta and Joel E. Michalek. On Gehan’s Wilcoxon and the accelerated failure time model. Communications in Statistics: Theory and Methods, 13(5):563–566, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonin:1985:NNEa

[GM85a] R. Gonin and A. H. Money. Nonlinear L_{P} -norm estimation: Part I — on the choice of the exponent, p, where the errors are additive. Communications in Statistics: Theory and Methods, 14(4):827– 840, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonin:1985:NNEb

[GM85b] R. Gonin and A. H. Money. Nonlinear L_{p}-norm estimation: Part II: The asymptotic distribution of the exponent, p, as a function of the sample kurtosis. Communications in Statistics: Theory and Methods, 14(4):841–849, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1988:MLE

[GM88] Malay Ghosh and Lily Liorens Mantelle. Maximum likelihood estimation for two parameter exponentials under Type I censoring. Communications in Statistics: Theory and Methods, 17(9):2859– 2879, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1984:SST

[GMM84] Ramesh C. Gupta, Kishan G. Mehrotra, and Joel E. Michalek. A small sample test for an absolutely continuous bivariate exponential model. Communications in Statistics: Theory and Methods, 13 (14):1735–1740, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glaz:1983:MCL

[GN83] Joseph Glaz and Joseph Naus. Multiple clusters on the line. Communications in Statistics: Theory and Methods, 12(17):1961–1986, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1986:SSE

[GN86a] Malay Ghosh and David M. Nickerson. Sequential shrinkage estimation of independent normal means with unknown variances. Communications in Statistics: Theory and Methods, 15(7):2115– 2130, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glaz:1986:APF

[GN86b] Joseph Glaz and Joseph Naus. Approximating probabilities of first passage in a particular Gaussian process. Communications in Statistics: Theory and Methods, 15(6):1709–1722, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1987:NDL

[GN87] A. K. Gupta and D. K. Nagar. Nonnull distribution of the likelihood ratio criterion for testing equality of covariance matrices under intraclass correlation model. Communications in Statistics: Theory and Methods, 16(11):3323–3341, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:AEN

[GN88] A. K. Gupta and D. K. Nagar. Asymptotic expansion of the nonnull distribution of likelihood ratio statistic for testing multisample sphericity. Communications in Statistics: Theory and Methods, 17 (9):3145–3155, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Godfrey:1985:TEQ

[God85] A. Blanton Godfrey. Training & education in quality & reliability — a modern approach. Communications in Statistics: Theory and Methods, 14(11):2621–2638, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gokhale:1989:TSS

[Gok89] D. V. Gokhale. On two stage shrinkage testimation for exponential. Communications in Statistics: Theory and Methods, 18(9):3203– 3213, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goodnight:1980:THF

[Goo80] J. H. Goodnight. Tests of hypotheses in fixed effects linear models. Communications in Statistics: Theory and Methods, 9(2):167–180, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1987:SUD

[Goo87] I. J. Good. Serial unblocked designs balanced for correlation between neighbours. Communications in Statistics: Theory and Methods, 16(4):1153–1159, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1989:AMP

[Goo89] Phillip Good. Almost most powerful tests against composite alternatives. Communications in Statistics: Theory and Methods, 18 (5):1913–1925, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goria:1987:SCU

[Gor87] M. N. Goria. Some characterizations of the uniform distribution. Communications in Statistics: Theory and Methods, 16(3):813– 819, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:1986:ALA

[Gov86] Z. Govindarajulu. Augmenting Lehmann alternatives for improving asymptotic efficiency of rank tests. Communications in Statistics: Theory and Methods, 15(4):1209–1220, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govaerts:1989:NMC

[Gov89] Bernadette Govaerts. A note on a method to compute the asymptotic distribution of the sample second order moments of dynamic linear normal variables. Communications in Statistics: Theory and Methods, 18(2):479–488, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gallardo:1982:TMN

[GP82] S. Rincon Gallardo and Quesenberry C. P. Testing multivariate normality using several samples: applications techniques. Communications in Statistics: Theory and Methods, 11(4):343–358, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geng:1986:HDB

[GP86] Shu Geng and S. K. Perng. Homogeneous designs based on pretreatment information. Communications in Statistics: Theory and Methods, 15(1):59–72, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grimson:1984:STE

[GQ84] Roger C. Grimson and Dana Quade. A simple test for equally likely random responses in a sequence of Bernoulli trials. Communications in Statistics: Theory and Methods, 13(9):1063–1071, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gans:1981:BEM

[GR81] Lydla P. Gans and C. A. Robertson. The behavior of estimated measures of association in small and moderate sample sizes for 2×3 tables. Communications in Statistics: Theory and Methods, 10(16): 1673–1686, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gruber:1982:BEL

[GR82] Marvin H. J. Gruber and Poduri S. R. S. Rao. Bayes estimators for linear models with less than full rank. Communications in Statistics: Theory and Methods, 11(1):59–69, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaver:1983:ERC

[GR83] Kenneth M. Gaver and Ramon Rabinovitch. Evaluating random choice models using multivariate correlation analysis. Communications in Statistics: Theory and Methods, 12(16):1901–1939, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabriel:1987:NSI

[GR87a] K. R. Gabriel and J. Robinson. A note on simultaneous inference with rerandomization tests. Communications in Statistics: Theory and Methods, 16(7):2007–2015, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1987:EMR

[GR87b] J. K. Ghorai and L. Rejtö. Estimation of mean residual life with censored data under the proportional hazard model. Communications in Statistics: Theory and Methods, 16(7):2097–2114, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1987:IRB

[GR87c] Alan J. Gross and Philip F. Rust. Item reliability based on a paucity of item failures. Communications in Statistics: Theory and Methods, 16(10):2981–2990, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Greene:1989:PPC

[GR89] Tom Greene and William S. Rayens. Partially pooled covariance matrix estimation in discriminant analysis. Communications in Statistics: Theory and Methods, 18(10):3679–3702, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gray:1987:SIP

[Gra87] Robert J. Gray, Ph.D. A simultaneous inference procedure for clinical trials. Communications in Statistics: Theory and Methods, 16(2):499–510, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gray:1989:INA

[Gra89] J. Brian Gray. The internal norm approach to influence diagnostics. Communications in Statistics: Theory and Methods, 18(3):943– 958, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Groeneveld:1982:BBO

[Gro82] Richard A. Groeneveld. Best bounds for order statistics and their expectations in range and mean units with applications. Communications in Statistics: Theory and Methods, 11(16):1809–1815, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grove:1984:PAT

[Gro84] Daniel M. Grove. Positive association in a two-way contingency table: likelihood ratio tests. Communications in Statistics: Theory and Methods, 13(8):931–945, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Groeneveld:1986:MSK

[Gro86] Richard A. Groeneveld. Measurement of skewness and kurtosis for the generalized Tukey lambda distributions. Communications in Statistics: Theory and Methods, 15(2):329–343, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gruber:1985:CBE

[Gru85] Marvin H. J. Gruber. A comparison of Bayes estimators and constrained least square estimators. Communications in Statistics: Theory and Methods, 14(2):479–489, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:SCE

[GRY81] B. K. Ghosh, Marion R. Reynolds, Jr., and Van Hui Yer. Shewhart x-charts with estimated process variance. Communications in Statistics: Theory and Methods, 10(18):1797–1822, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1980:RBS

[GS80] Shanti S. Gupta and Ashok K. Singh. On rules based on sample medians for selection of the largest location parameter. Communications in Statistics: Theory and Methods, 9(12):1277–1298, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:EN

[GS81a] B. K. Ghosh and P. K. Sen. Editorial note. Communications in Statistics: Theory and Methods, 10(21):v, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1981:BRQ

[GS81b] Malay Ghosh and Shashikala Sukhatme. On Bahadur’s representation of quantiles in nonregular cases. Communications in Statistics: Theory and Methods, 10(3):269–282, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1983:OGR

[GS83] S. Ghosh and K. P. Shah. On the optimality of the generalized room squares. Communications in Statistics: Theory and Methods, 12 (18):2119–2125, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:1985:EDT

[GS85a] P. S. Gill and G. K. Shukla. Experimental designs and their efficiencies for spatially correlated observations in two dimensions. Communications in Statistics: Theory and Methods, 14(9):2181– 2197, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1985:VCG

[GS85b] I. J. Good and E. P. Smith. The variance and covariance of a generalized index of similarity especially for a generalization on an index of Hellinger and Bhattacharyya. Communications in Statistics: Theory and Methods, 14(12):3053–3061, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:1987:OCD

[GS87] P. S. Gill and G. K. Shukla. Optimal change-over designs for correlated observations. Communications in Statistics: Theory and Methods, 16(8):2243–2261, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Groggel:1987:PAB

[GSS87] D. J. Groggel, R. L. Schaefer, and J. H. Skillings. Procedures for analyzing block designs with censored data. Communications in Statistics: Theory and Methods, 16(2):431–444, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gu:1987:WHW

[Gu87] Chong Gu. What happens when bootstrapping the smoothing spline. Communications in Statistics: Theory and Methods, 16 (11):3275–3284, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerard:1987:CFU

[Gue87] John B. Guerard, Jr. Composite forecasting using ridge regression. Communications in Statistics: Theory and Methods, 16(4):937– 952, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guinel:1982:BFT

[Güi82] E. Güinel. Bayes factors in three-way contingency tables. Communications in Statistics: Theory and Methods, 11(8):910–933, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunst:1983:RAM

[Gun83] Richard F. Gunst. Regression analysis with multicollinear predictor variables: definition, direction, and effects. Communications in Statistics: Theory and Methods, 12(19):2217–2260, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunel:1984:BAM

[Gun84] Erdogan Gunel. A Bayesian analysis of the multinomial model for a dichotomous response with nonrespondents. Communications in Statistics: Theory and Methods, 13(6):737–751, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunel:1985:BCR

[Gun85] Erdogan Gunel. A Bayesian comparison of randomized and voluntary response sampling models. Communications in Statistics: Theory and Methods, 14(10):2411–2435, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunel:1986:NPE

[Gun86] Erdogan Gunel. A note on the posterior expectation. Communications in Statistics: Theory and Methods, 15(4):1337–1345, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1980:MRL

[Gup80] Ramesh C. Gupta. On the mean residual life function of a mixture in reliability studies. Communications in Statistics: Theory and Methods, 9(13):1365–1369, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1982:SPE

[Gup82] Pushpa Lata Gupta. Structural properties and estimation in m—E_{K} —1 queue. Communications in Statistics: Theory and Methods, 11(6):711–719, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1984:ANO

[Gup84] S. C. Gupta. The analysis of non-orthogonal designs with many classifications. Communications in Statistics: Theory and Methods, 13(12):1483–1492, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1986:ECD

[Gup86] Sudhir C. Gupta. Efficiency-consistency in designs. Communications in Statistics: Theory and Methods, 15(4):1315–1318, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1987:CGS

[Gup87a] Pushpa Lata Gupta. A characterization for GPSD signals in additive noise. Communications in Statistics: Theory and Methods, 16(1):233–240, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1987:EMC

[Gup87b] R. P. Gupta. Estimation of multiple correlation coefficient when observations are missing on one of the variables. Communications in Statistics: Theory and Methods, 16(2):353–360, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1987:GGC

[Gup87c] Sudhir C. Gupta. Generating generalized cyclic designs with factorial balance. Communications in Statistics: Theory and Methods, 16(7):1885–1900, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:BCF

[Gup88a] B. C. Gupta. A bound connected with 2^{6} factorial search designs of resolution 3.1. Communications in Statistics: Theory and Methods, 17(9):3137–3144, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:TIM

[Gup88b] S. C. Gupta. Type A incomplete multi-response designs. Communications in Statistics: Theory and Methods, 17(9):3089–3105, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:DSP

[Gup88c] Sudhir Gupta. Designs for symmetrical parallel line assays obtainable through group divisible designs. Communications in Statistics: Theory and Methods, 17(11):3865–3868, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1988:PBP

[Gup88d] Sudhir Gupta. On the property of balance and the pattern of Cmatrix in block designs. Communications in Statistics: Theory and Methods, 17(9):3125–3135, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gutmann:1982:MLE

[Gut82] Samuel Gutmann. Minimax linear empirical Bayes estimation of the binomial parameter. Communications in Statistics: Theory and Methods, 11(18):2075–2082, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gijbels:1989:QEP

[GV89a] I. Gijbels and N. Veraverbeke. Quantile estimation in the proportional hazards model of random censorship. Communications in Statistics: Theory and Methods, 18(5):1645–1663, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1989:LOT

[GV89b] Ashis Sen Gupta and Lea Vermeire. Locally optimal test for simultaneously testing the mean and variance of a normal distribution. Communications in Statistics: Theory and Methods, 18(12): 4519–4532, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1982:BLE

[gW82] Song gui Wang. On biased linear estimators in models with arbitrary rank. Communications in Statistics: Theory and Methods, 11 (14):1571–1581, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Greenhouse:1984:CRD

[GW84] Joel B. Greenhouse and Robert A. Wolfe. A competing risks derivation of a mixture model for the analysis of survival data. Communications in Statistics: Theory and Methods, 13(25):3133–3154, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

George:1987:GML

[GWBG87] Yinsheng Qu George, George W. Williams, Gerald J. Beck, and Marlene Goormastic. A generalized model of logistic regression for clustered data. Communications in Statistics: Theory and Methods, 16(12):3447–3476, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gray:1989:TMM

[GWM89] H. L. Gray, Wayne A. Woodward, and Gary McCartor. Testing for the maximum mean in a mixture of normals. Communications in Statistics: Theory and Methods, 18(11):4011–4027, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heath:1983:EMT

[HA83] A. B. Heath and J. A. Anderson. Estimation in a multivariable two-phase regression. Communications in Statistics: Theory and Methods, 12(7):809–828, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haber:1985:LLM

[Hab85] Michael Haber. Log-linear models for correlated marginal totals of a contingency table. Communications in Statistics: Theory and Methods, 14(12):2845–2856, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haber:1989:DMT

[Hab89] Michael Haber. Do the marginal totals of a 2×2 contingency table contain information regarding the table proportions? Communications in Statistics: Theory and Methods, 18(1):147–156, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haff:1986:LLO

[Haf86] L. R. Haff. On linear log-odds and estimation of discriminant coefficients. Communications in Statistics: Theory and Methods, 15 (7):2131–2144, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haining:1988:ESM

[Hai88] Rober Haining. Estimating spatial means with an application to remotely sensed data. Communications in Statistics: Theory and Methods, 17(2):573–597, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hallett:1984:URR

[Hal84] A. J. Hughes Hallett. The use of ridge regression techniques to generate risk sensitive decision rules. Communications in Statistics: Theory and Methods, 13(2):127–138, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamada:1989:CUI

[Ham89] Michael Hamada. The costs of using incomplete response data for the exponential regression model. Communications in Statistics: Theory and Methods, 18(5):1691–1714, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:1980:EVS

[Han80] Chien-Pai Han. On estimating the variance of the sample mean in random models based on conditional specification. Communications in Statistics: Theory and Methods, 9(14):1417–1425, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harris:1982:BDR

[Har82] R. R. Harris. Biostatistics: a difficult road ahead. Communications in Statistics: Theory and Methods, 11(19):2125–2128, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harris:1983:FMG

[Har83] Carl M. Harris. On finite mixtures of geometric and negative binomial distributions. Communications in Statistics: Theory and Methods, 12(9):987–1007, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harter:1984:ALP

[Har84] H. Leon Harter. Another look at plotting positions. Communications in Statistics: Theory and Methods, 13(13):1613–1633, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harrington:1985:QTA

[Har85a] H. James Harrington. Quality trends around the world. Communications in Statistics: Theory and Methods, 14(11):2639, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hart:1985:CCC

[Har85b] Jeffrey D. Hart. A counterexample to a claim concerning the convolution of multimodal distributions. Communications in Statistics: Theory and Methods, 14(12):2943–2945, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harter:1986:SAO

[Har86] H. Leon Harter. Some applications of order statistics to multivariate analysis II. Communications in Statistics: Theory and Methods, 15(9):2609–2649, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hasza:1980:NML

[Has80] David P. Hasza. A note on maximum likelihood estimation for the first-order autoregressive process. Communications in Statistics: Theory and Methods, 9(13):1411–1415, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hauck:1989:ORI

[Hau89] Walter W. Hauck. Odds ratio inference from stratified samples. Communications in Statistics: Theory and Methods, 18(2):767– 800, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1989:ECM

[Haw89a] D. L. Hawkins. Estimating changes in a multi-parameter exponential family. Communications in Statistics: Theory and Methods, 18 (10):3595–3623, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1989:UAR

[Haw89b] D. L. Hawkins. A U–I approach to retrospective testing for shifting parameters in a linear model. Communications in Statistics: Theory and Methods, 18(8):3117–3134, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayter:1989:SNP

[Hay89] A. J. Hayter. Selecting the normal population with the largest mean when the variances are bounded. Communications in Statistics: Theory and Methods, 18(4):1455–1467, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hathaway:1986:APF

[HB86] Richard J. Hathaway and James C. Bezdek. On the asymptotic properties of fuzzy c-means cluster prototypes as estimators of mixture subpopulation centers. Communications in Statistics: Theory and Methods, 15(2):505–513, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hegde:1989:EPTb

[HD89a] Laxman M. Hegde and Ram C. Dahiya. Estimation of the parameters in a truncated normal distribution. Communications in Statistics: Theory and Methods, 18(11):4177–4195, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hegde:1989:EPTa

[HD89b] Laxman M. Hegde and Ram C. Dahiya. Estimation of the parameters of a truncated gamma distribution. Communications in Statistics: Theory and Methods, 18(2):561–577, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heilbron:1988:ETC

[Hei88] David C. Heilbron. Exact test of a common odds ratio: robustness and comparison with an alternative test. Communications in Statistics: Theory and Methods, 17(6):1889–1902, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Helms:1988:CPH

[Hel88a] Ronald W. Helms. Comparisons of parameter and hypothesis definitions in a general linear model. Communications in Statistics: Theory and Methods, 17(8):2725–2753, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Helms:1988:DLM

[Hel88b] Ronald W. Helms. Definitions of linear model parameters and hypotheses as functions of e[y]. Communications in Statistics: Theory and Methods, 17(8):2715–2723, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hemmerle:1980:RBU

[Hem80] William J. Hemmerle. Recognizing balance with unbalanced data. Communications in Statistics: Theory and Methods, 9(2):201–211, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henderson:1988:UEL

[Hen88] C. R. Henderson. The use of equivalent linear models in mixed model estimation and prediction. Communications in Statistics: Theory and Methods, 17(4):969–982, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hearron:1984:BAG

[HES84] A. E. Hearron, G. L. Elfring, and J. R. Schultz. Biopharmaceutical applications of group sequential designs. Communications in Statistics: Theory and Methods, 13(19):2419–2450, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heydorn:1984:USR

[Hey84] Richard P. Heydorn. Using satellite remotely sensed data to estimate crop proportions. Communications in Statistics: Theory and Methods, 13(23):2881–2903, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huynh:1980:PTT

[HF80] Huynh Huynh and Leonard S. Feldt. Performance of traditional F tests in repeated measures designs under covariance heterogeneity. Communications in Statistics: Theory and Methods, 9(1):61–74, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1988:MGS

[HF88] Mei-Ling Huang and Karen Yuen Fung. A more generalized Stirling distribution of the second kind. Communications in Statistics: Theory and Methods, 17(12):4337–4356, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hess:1980:RBG

[HG80] J. L. Hess and R. F. Gunst. A relationship between generalized and integrated mean square errors. Communications in Statistics: Theory and Methods, 9(3):321–326, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hua:1983:GRR

[HG83] Tsushung A. Hua and Richard F. Gunst. Generalized ridge regression: a note on negative ridge parameters. Communications in Statistics: Theory and Methods, 12(1):37–45, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haack:1986:DTA

[HG86] D. Haack and Z. Govindarajulu. A decision theoretic approach to a distribution-free compound classification problem. Communications in Statistics: Theory and Methods, 15(10):2877–2897, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haining:1989:MLE

[HGB89] Robert Haining, Daniel Griffith, and Robert Bennett. Maximum likelihood estimation with missing spatial data and with an application to remotely sensed data. Communications in Statistics: Theory and Methods, 18(5):1875–1894, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haas:1985:SPU

[HGO85] R. W. Haas, R. F. Gunst, and D. B. Owen. Screening procedures using quadratic forms. Communications in Statistics: Theory and Methods, 14(6):1393–1404, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1980:SPW

[HH80] R. K. Hwang and F. Hsuan. Stronger players win more knockout tournaments in average. Communications in Statistics: Theory and Methods, 9(1):107–113, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1982:SSS

[HH82] Tong An Hsu and Deng Yuan Huang. Some sequential selection procedures for good regression models. Communications in Statistics: Theory and Methods, 11(4):411–421, 1982. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Houston:1984:USD

[HH84] A. G. Houston and F. G. Hall. Use of satellite data in agricultural surveys. Communications in Statistics: Theory and Methods, 13 (23):2857–2880, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1986:BDS

[HH86a] D. L. Hawkins and Chien-Pai Han. Bivariate distributions of some ratios of independent noncentral chi-square random variables. Communications in Statistics: Theory and Methods, 15(1):261– 277, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1986:PCT

[HH86b] D. L. Hawkins and Chien-Pai Han. A power comparison of three tests for design effects in a random effects covariance model. Communications in Statistics: Theory and Methods, 15(11):3401–3418, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hines:1989:RSS

[HH89] R. J. O’ Hara Hines and W. G. S. Hines. Repeated sampling of spatial point distributions. Communications in Statistics: Theory and Methods, 18(7):2599–2614, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hirano:1989:MRS

[HI89] Katuomi Hirano and Kosei Iwase. Minimum risk scale equivariant estimator: estimating the mean of an inverse Gaussian distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 18(1):189–197, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hietikko:1985:ACD

[Hie85] Harri Hietikko. On the approximation to the ARMA(P, Q) covariance determinant. Communications in Statistics: Theory and Methods, 14(2):341–353, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Higgins:1982:NML

[Hig82] James J. Higgins. A note on MLE’s in linear models with nonnormal or dependent errors. Communications in Statistics: Theory and Methods, 11(2):141–149, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1985:KED

[Hil85] Peter D. Hill. Kernel estimation of a distribution function. Communications in Statistics: Theory and Methods, 14(3):605–620, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hinde:1986:PMM

[Hin86] Ray Hinde. Population moments and moments of the sample mean. Communications in Statistics: Theory and Methods, 15(10):3129– 3143, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hall:1983:ARE

[HJ83] David L. Hall and Brian L. Joiner. Asymptotic relative efficiencies of R-estimators of location. Communications in Statistics: Theory and Methods, 12(7):739–763, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1989:MDS

[HJK89] Qishen Huang, Norman L. Johnson, and Samuel Kotz. Modified Dorfman–Sterrett screening (group testing) procedures and the effects of faulty inspection. Communications in Statistics: Theory and Methods, 18(4):1485–1495, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hora:1983:BIO

[HK83] Stephen C. Hora and Gary D. Kelley. Bayesian inference on the odds and risk ratios. Communications in Statistics: Theory and Methods, 12(6):725–738, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Handa:1987:PTE

[HK87] B. R. Handa and N. S. Kambo. Power of a test for equality of k exponential location parameters. Communications in Statistics: Theory and Methods, 16(9):2755–2769, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hardin:1989:CFA

[HK89] J. Michael Hardin and Badrig M. Kurkjian. The calculus for factorial arrangements: a review and bibliography. Communications in Statistics: Theory and Methods, 18(4):1251–1277, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hickman:1980:NGT

[HL80a] Edgar P. Hickman and Clayton S. Long. A note on Glejser’s test for heteroscedasticity. Communications in Statistics: Theory and Methods, 9(12):1209–1220, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hosmer:1980:GFT

[HL80b] David W. Hosmer and Stanley Lemesbow. Goodness of fit tests for the multiple logistic regression model. Communications in Statistics: Theory and Methods, 9(10):1043–1069, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1984:RSA

[HL84] Robert V. Hogg and Russell V. Lenth. A review of some adaptive statistical techniques. Communications in Statistics: Theory and Methods, 13(13):1551–1579, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsiang:1985:ZDQ

[HL85] Thomas C. Hsiang and Hau L. Lee. Zero defects: a quality costs approach. Communications in Statistics: Theory and Methods, 14 (11):2641–2655, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Halperin:1987:TSO

[HL87] Max Halperin and K. K. Gordon Lan. A two sample ordered alternative test for means and variances. Communications in Statistics: Theory and Methods, 16(5):1297–1313, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1988:DRI

[HLN88] Chyi Hung Hsu, Martin S. Levy, and James W. Neill. The distribution of the ratio of independent central Wishart determinants. Communications in Statistics: Theory and Methods, 17(9):3177– 3190, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hochberg:1981:TSE

[HM81] Yosef Hochberg and Ruth Marcus. Three stage elimination type procedures for selecting the best normal population when variances are unknown. Communications in Statistics: Theory and Methods, 10(6):597–612, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hiroki:1984:BTP

[HN84] Tsurumi Hiroki and Sheflin Neil. Bayesian tests of a parameter shift under heteroscedasticity: weighted-t vs. double-t approaches. Communications in Statistics: Theory and Methods, 13(8):1003– 1013, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamdan:1986:MLE

[HN86] M. A. Hamdan and M. O. Nasro. Maximum likelihood estimation of the parameters of the bivariate binomial distribution. Communications in Statistics: Theory and Methods, 15(3):747–754, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hochberg:1981:VEW

[Hoc81] Yosef Hochberg. On the variance estimate of a Wilcoxon–Mann– Whitney statistic for group ordered data. Communications in Statistics: Theory and Methods, 10(17):1719–1732, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hocking:1988:CMA

[Hoc88] R. R. Hocking. A cell means analysis of mixed linear models. Communications in Statistics: Theory and Methods, 17(4):983– 1010, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1982:ASI

[Hog82] Robert V. Hogg. On adaptive statistical inferences. Communications in Statistics: Theory and Methods, 11(22):2531–2542, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holst:1987:REL

[Hol87] U. Holst. Recursive M-estimators of location. Communications in Statistics: Theory and Methods, 16(8):2201–2226, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holgate:1989:LCF

[Hol89a] P. Holgate. The lognormal characteristic function. Communications in Statistics: Theory and Methods, 18(12):4539–4548, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holiday:1989:EDN

[Hol89b] D. B. Holiday. The estimation of derivatives of a nonparametric regression function when the data are correlated. Communications in Statistics: Theory and Methods, 18(6):2107–2124, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hora:1980:SD

[Hor80] Stephen C. Hora. Sequential discrimination. Communications in Statistics: Theory and Methods, 9(9):905–916, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horn:1981:HRN

[Hor81] Paul Horn. Heteroscedasticity of residuals: a non-parametric alternative to the Goldfeld–Quandt peak test. Communications in Statistics: Theory and Methods, 10(8):795–808, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hora:1985:GFT

[Hor85a] Stephen C. Hora. Goodness of fit tests using regression. Communications in Statistics: Theory and Methods, 14(2):307–332, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horn:1985:BSA

[Hor85b] Paul S. Horn. On borrowing spread from auxiliary samples in the one-sample problem. Communications in Statistics: Theory and Methods, 14(12):3107–3124, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hosmane:1986:ILR

[Hos86] B. S. Hosmane. Improved likelihood ratio tests and Pearson chisquare tests for independence in two dimensional contingency tables. Communications in Statistics: Theory and Methods, 15(6): 1875–1888, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hosmane:1987:ILR

[Hos87a] Balakrishna Hosmane. Improved likelihood ratio test for multinomial goodness of fit. Communications in Statistics: Theory and Methods, 16(11):3185–3198, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hosmane:1987:SCO

[Hos87b] Balakrishna Hosmane. Some conditions for optimality of the almost unbiased estimators of regression coefficients. Communications in Statistics: Theory and Methods, 16(6):1725–1731, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hosmane:1988:GSE

[Hos88] Balakrishna S. Hosmane. On a generalized Stein estimator of regression coefficients. Communications in Statistics: Theory and Methods, 17(6):1735–1740, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Howlader:1985:ABE

[How85] H. A. Howlader. Approximate Bayes estimation of reliability of two parameter inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 14(4):937–946, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1980:SSI

[HP80] Y. S. Hsu and W. J. Park. Some statistical inferences on O–U processes. Communications in Statistics: Theory and Methods, 9 (5):529–540, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1982:EIR

[HP82] D. Y. Huang and S. Panchapakesan. On eliminating inferior regression models. Communications in Statistics: Theory and Methods, 11(7):751–759, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hocking:1983:RD

[HP83] R. R. Hocking and O. J. Pendleton. The regression dilemma. Communications in Statistics: Theory and Methods, 12(5):497– 527, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamdy:1987:TSP

[HP87] H. I. Hamdy and W. J. Pallotta. Triple sampling procedure for estimating the scale parameters of Pareto distributions. Communications in Statistics: Theory and Methods, 16(7):2155–2164, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1981:TSB

[HR81] D. L. Hill and P. V. Rao. Tests of symmetry based on Watson statistics. Communications in Statistics: Theory and Methods, 10 (11):1111–1125, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:1988:IBC

[HRR88] Chien-Pai Han, C. V. Rao, and J. Ravichandran. Inference based on conditional specification. Communications in Statistics: Theory and Methods, 17(6):1945–1964, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harishchandra:1982:LSA

[HS82] K. Harishchandra and P. Srivenkataramana. Link sampling for attributes. Communications in Statistics: Theory and Methods, 11 (16):1855–1868, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hewett:1983:STS

[HS83] John E. Hewett and John D. Spurrier. A survey of two stage tests of hypotheses: theory and application. Communications in Statistics: Theory and Methods, 12(20):2307–2425, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holbert:1986:BTS

[HS86] Don Holbert and Mun-Shig Son. Bootstrapping a time series model: some empirical results. Communications in Statistics: Theory and Methods, 15(12):3669–3691, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hocking:1980:HTU

[HSC80] R. R. Hocking, F. M. Speed, and A. T. Coleman. Hypotheses to be tested with unbalanced data. Communications in Statistics: Theory and Methods, 9(2):117–129, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsieh:1981:EMS

[Hsi81] Hui-Kuang Hsieh. On estimating the mean of the selected population with unknown variance. Communications in Statistics: Theory and Methods, 10(18):1869–1878, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsieh:1984:NTS

[Hsi84] H. K. Hsieh. Nonparametric tests for scale shift at an unknown time point. Communications in Statistics: Theory and Methods, 13 (11):1335–1355, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1980:RNS

[Hsu80] Jason C.. Hsu. Robust and nonparametric subset selection procedures. Communications in Statistics: Theory and Methods, 9(14): 1439–1459, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsuan:1981:RRP

[Hsu81] Francis C. Hsuan. Ridge regression from principal component point of view. Communications in Statistics: Theory and Methods, 10 (19):1981–1995, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1982:SOS

[Hsu82] Tong An Hsu. On some optimal selection procedures for Weibull populations. Communications in Statistics: Theory and Methods, 11(23):2657–2668, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsu:1985:MUM

[Hsu85] Jason C. Hsu. A method of unconstrained multiple comparisons with the best. Communications in Statistics: Theory and Methods, 14(9):2009–2028, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayre:1981:CSA

[HT81] Lakhbir S. Hayre and Bruce W. Turnbull. A class of simple approximate sequential tests for adaptive comparison of two treatments. Communications in Statistics: Theory and Methods, 10(22):2339– 2360, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hochbeg:1983:TSS

[HT83a] Yosef Hochbeg and Aaron Tenenbein. On triple sampling schemes for estimating from binomial data with misclassification errors. Communications in Statistics: Theory and Methods, 12(13):1523– 1533, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1983:ODP

[HT83b] Deng-Yuan Huang and Sheng-Tsaing Tseng. Γ-optimal decision procedures for selecting the best population in randomized complete block design. Communications in Statistics: Theory and Methods, 12(3):341–354, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huffman:1985:CTP

[Huf85] Michael D. Huffman. Comparing two Poisson parameters: what to do when the optimal isn’t done. Communications in Statistics: Theory and Methods, 14(12):3063–3074, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hugwood:1980:NRT

[Hug80] Robert Charles Hugwood. A nonlinear renewal theorem for discrete random variables. Communications in Statistics: Theory and Methods, 9(16):1677–1698, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hung:1989:UTA

[Hun89] Hsien-Ming Hung. On the utility of transformed auxiliary variables in finite population inference. Communications in Statistics: Theory and Methods, 18(1):171–188, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hurvich:1987:ASL

[Hur87] Clifford M. Hurvich. Automatic selection of a linear predictor through frequency domain cross-validation. Communications in Statistics: Theory and Methods, 16(11):3199–3234, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Husler:1988:ABS

[Hus88] Jürg Husler. On the asymptotic behaviour of selector statistics. Communications in Statistics: Theory and Methods, 17(10):3569– 3590, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hutchinson:1984:UBN

[Hut84] T. P. Hutchinson. Using the bivariate normal distribution to construct probability models in the health sciences. Communications in Statistics: Theory and Methods, 13(14):1723–1733, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holland:1987:DFC

[HW87a] Paul W. Holland and Yuchung J. Wang. Dependence function for continuous bivariate densities. Communications in Statistics: Theory and Methods, 16(3):863–876, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holland:1987:RDC

[HW87b] Paul W. Holland and Yuchung J. Wang. Regional dependence for continuous bivariate densities. Communications in Statistics: Theory and Methods, 16(1):193–206, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:1988:BEB

[HW88] G. G. Hamedani and G. G. Walter. Bayes estimation of the binomial parameter n. Communications in Statistics: Theory and Methods, 17(6):1829–1843, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Howlader:1989:BERa

[HW89a] H. A. Howlader and G. Weiss. Bayes estimators of the reliability of the logistic distribution. Communications in Statistics: Theory and Methods, 18(1):245–259, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Howlader:1989:BERb

[HW89b] H. A. Howlader and G. Weiss. Bayes estimators of the reliability of the logistic distribution. Communications in Statistics: Theory and Methods, 18(4):1339–1355, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haddow:1986:MPE

[HY86] A. A. Haddow and D. H. Young. Moment properties of estimators for a type 1 extreme-value regression model. Communications in Statistics: Theory and Methods, 15(8):2527–2539, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1982:AGR

[HZ82] R. Carter Hill and Rod F. Ziemer. The application of generalized ridge and Stein-like general minimax rules to multicollinear data. Communications in Statistics: Theory and Methods, 11(6):623– 638, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1983:MRV

[HZ83] R. Carter Hill and Rod F. Ziemer. Missing regressor values under conditions of multicollinearity. Communications in Statistics: Theory and Methods, 12(22):2557–2573, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iachan:1983:NES

[Iac83] Ronaldo Iachan. Nonsampling errors in surveys — a review. Communications in Statistics: Theory and Methods, 12(19):2273–2287, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Inman:1989:OCM

[IB89] Henry F. Inman and Edwin L. Bradley, Jr. The overlapping coefficient as a measure of agreement between probability distributions and point estimation of the overlap of two normal densities. Communications in Statistics: Theory and Methods, 18(10):3851–3874, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1980:RC

[IC80a] Ronald L. Iman and W. J. Conover. Rejoinder to comments. Communications in Statistics: Theory and Methods, 9(17):1863–1874, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1980:SSS

[IC80b] Ronald L. Iman and W. J. Conover. Small sample sensitivity analysis techniques for computer models.with an application to risk assessment. Communications in Statistics: Theory and Methods, 9 (17):1749–1842, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1980:ACR

[ID80] Ronald L. Iman and James M. Davenport. Approximations of the critical region of the Friedman statistic. Communications in Statistics: Theory and Methods, 9(6):571–595, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ierza:1985:OPG

[Ier85] Joseph V. Ierza. Ordinal probit: A generalization. Communications in Statistics: Theory and Methods, 14(1):1–11, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Indira:1988:NAU

[IM88] N. K. Indira and V. V. Menon. Normal approximation in an urn model with indistinguishable balls. Communications in Statistics: Theory and Methods, 17(10):3389–3402, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1987:TEQ

[Ima87] Ronald L. Iman. Tables of the exact quantiles of the top-down correlation coefficient for n = 3(1)14. Communications in Statistics: Theory and Methods, 16(5):1513–1540, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iubbs:1985:NRP

[IS85] J. D. Iubbs and O. E. Smith. A note on the ratio of positively correlated gamma variates. Communications in Statistics: Theory and Methods, 14(1):13–23, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwase:1986:ISU

[IS86] Kosei Iwase and Noriaki Set. Incomplete sufficient unbiased estimators in the problem of the Nile. Communications in Statistics: Theory and Methods, 15(1):279–289, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isaki:1987:ECM

[IS87] Cary T. Isaki and Linda K. Schultz. The effects of correlation and matching error on dual system estimation. Communications in Statistics: Theory and Methods, 16(8):2405–2427, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isao:1981:EDM

[Isa81] Hashimoto Isao. Estimation for distributions with monotone likelihood ratio: case of vector valued parameter. Communications in Statistics: Theory and Methods, 10(19):1901–1913, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isaki:1986:MTD

[Isa86] Cary T. Isaki. Mas of toe dual system estimator and some alternatives. Communications in Statistics: Theory and Methods, 15 (5):1435–1450, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogawa:1984:LST

[Iso84] Yoshiko Isogawa. Large sample theory for a multivariate structural relationship with replication. Communications in Statistics: Theory and Methods, 13(17):2111–2118, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogai:1989:NRE

[Iso89] Eiichi Isogai. Nonparametric recursive estimation in stationary Markov processes. Communications in Statistics: Theory and Methods, 18(4):1309–1323, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwase:1982:UEM

[ISS82] Kosei Iwase, Kunio Shimizu, and Manibu Suzuki. On UMVU estimators for the multivariate lognormal distribution and their variances. Communications in Statistics: Theory and Methods, 11(6): 687–697, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iyer:1989:EMS

[IV89] Hari K. Iyer and Dominic F. Vecchia. Exact moments of the symmetric cubic assignment statistic. Communications in Statistics: Theory and Methods, 18(12):4309–4320, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwase:1986:UEC

[Iwa86a] Kosei Iwase. Umvu estimation for the cumulative hazard function in a Pareto distribution of the first kind. Communications in Statistics: Theory and Methods, 15(12):3559–3566, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwase:1986:UEP

[Iwa86b] Kōsei Iwase. Umvu estimator of the probability in the geometric distribution with unknown truncation parameter. Communications in Statistics: Theory and Methods, 15(8):2449–2453, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwase:1989:LRT

[Iwa89] Kōsei Iwase. Linear regression through the origin with constant coefficient of variation for the inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 18(10):3587–3593, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jackson:1985:MQC

[Jac85] J. Edward Jackson. Multivariate quality control. Communications in Statistics: Theory and Methods, 14(11):2657–2688, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1981:DOP

[Jai81] Ram B. Jain. Detecting outliers: power and some other considerations. Communications in Statistics: Theory and Methods, 10 (22):2299–2314, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janssen:1988:LMP

[Jan88] Arnold Janssen. Locally most powerful rank tests. Communications in Statistics: Theory and Methods, 17(5):1325–1349, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jorge:1984:SPT

[JB84] Martinez Jorge and Iglewicz Boris. Some properties of the Tukey g and h family of distributions. Communications in Statistics: Theory and Methods, 13(3):353–369, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jang:1986:ESE

[JB86] J. S. Jang and D. S. Bai. Efficient sequential estimation in a Markov branching process with immigration. Communications in Statistics: Theory and Methods, 15(5):1681–1697, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1982:TSA

[JBF^{+}82] D. S. St John, S. P. Bailey, W. H. Fellner, J. M. Minor, R. D. Snee, and E. I. du Pont de. Time series analysis of stratospheric ozone. Communications in Statistics: Theory and Methods, 11(12): 1293–1333, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1984:AEM

[JC84] G. Jones and R. C. H. Cheng. On the asymptotic efficiency of moment and maximum likelihood estimators in the three-parameter inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 13(18):2307–2314, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1987:EAA

[JC87] M. C. Jones and J. B. Copas. Empirical assessment of the applicability of shrinkage predictors in regression when past and future data differ. Communications in Statistics: Theory and Methods, 16 (8):2165–2185, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jelihovschi:1987:EBP

[Jel87] Enio Jelihovschi. Estimation of the binomial parameter, using an a priori hypothesis about its value. Communications in Statistics:

Theory and Methods, 16(3):665–692, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jennison:1983:EPC

[Jen83] Christopher Jennison. Equal probability of correct selection for Bernoulli selection procedures. Communications in Statistics: Theory and Methods, 12(24):2887–2896, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Javier:1985:MVD

[JG85] W. R. Javier and A. K. Gupta. On matric variate-t distribution. Communications in Statistics: Theory and Methods, 14(6):1413– 1425, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeske:1988:PIP

[JH88] Daniel R. Jeske and David A. Harville. Prediction-interval procedures and (fixed-effects) confidence-interval procedures for mixed linear models. Communications in Statistics: Theory and Methods, 17(4):1053–1087, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jack:1984:BAP

[JI84] Gambino Jack and Guttman Irwin. A Bayesian approach to prediction in the presence of spurious observations for several models. Communications in Statistics: Theory and Methods, 13(7): 791–812, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1984:MED

[JK84] Norman L. Johnson and Samuel Kotz. Moment estimators of defect pattern frequencies, allowing for inspection error. Communications in Statistics: Theory and Methods, 13(10):1215–1226, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1985:KSC

[JK85] Bruce McK Johnson and Timothy J. Killeen. Kolmogorov–Smirnov confidence sets: the case of more than one sample. Communications in Statistics: Theory and Methods, 14(12):2873–2885, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1986:ECR

[JKK86] Bruce McK. Johnson, Timothy J. Killeen, and Timothy J. Killeen. Exact confidence regions for scale and location parameters based on sample quartiles. Communications in Statistics: Theory and Methods, 15(12):3515–3529, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1989:DSS

[JKR89] Norman L. Johnson, Samuel Kotz, and Robert N. Rodriguez. Dorfman–Sterrett screening (group testing) schemes and the effects of faulty inspection. Communications in Statistics: Theory and Methods, 18(4):1469–1484, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1980:FID

[JKS80] N. L. Johnson, Samuel Kotz, and H. L. Sorkin. “Faulty inspection” distributions. Communications in Statistics: Theory and Methods, 9(9):917–922, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1989:QFE

[JKW89] Norman L. Johnson, Samuel Kotz, and Xizhi Wu. Quota fulfilment and errors in stratification. Communications in Statistics: Theory and Methods, 18(4):1497–1505, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1989:MMS

[JL89] P. N. Jones and G. J. Mc Lachlan. Modelling mass-size particle data by finite mixtures. Communications in Statistics: Theory and Methods, 18(7):2629–2646, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1987:EPP

[JN87] S. Joshi and S. Nabar. Estimation of the parameter in the problem of the Nile. Communications in Statistics: Theory and Methods, 16 (11):3149–3156, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1989:SAA

[JN89] S. M. Joshi and S. P. Nabar. Sequential analysis applied to testing the scale parameter in the problem of the Nile. Communications in Statistics: Theory and Methods, 18(1):239–243, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joanes:1983:ECF

[Joa83] D. N. Joanes. Estimators of certain functions of the variance of a normal distribution. Communications in Statistics: Theory and Methods, 12(9):1031–1037, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jobe:1987:ERP

[Job87] J. Marcus Jobe. Error rates for Poisson process discrimination. Communications in Statistics: Theory and Methods, 16(3):647– 658, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joe:1988:EPC

[Joe88] Harry Joe. Extreme probabilities for contingency tables under row and column independence with application to Fisher’s exact test. Communications in Statistics: Theory and Methods, 17(11):3677– 3685, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1982:TPT

[Joh82] S. John. The three-parameter two-piece normal family of distributions and its fitting. Communications in Statistics: Theory and Methods, 11(8):879–885, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1984:BMK

[Joh84] Rice John. Boundary modification for kernel regression. Communications in Statistics: Theory and Methods, 13(7):893–900, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1985:NIT

[Joh85] Peter W. M. John. Non-isomorphism of two latin square designs for sixteen varieties. Communications in Statistics: Theory and Methods, 14(1):117–121, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1986:PKM

[Joh86a] Peter W. M. John. On a paper by Kageyama and Mohan. Communications in Statistics: Theory and Methods, 15(6):1995–2001, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1986:CBR

[Joh86b] Robert E. Johnson. On the correlation between F -ratios. Communications in Statistics: Theory and Methods, 15(8):2483–2492, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1988:ETU

[Jos88] P. C. Joshi. Estimation and testing under exchangeable exponential model with a single outlier. Communications in Statistics: Theory and Methods, 17(7):2315–2326, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joun:1988:BDE

[Jou88] M. Joun. Bootstrapping density estimates. Communications in Statistics: Theory and Methods, 17(1):61–78, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ja:1981:PEN

[JP81a] Ram B. Ja’ and Louís A. Pingel. A procedure for estimating the number of outliers. Communications in Statistics: Theory and Methods, 10(10):1029–1041, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1981:RRO

[JP81b] Ram B. Jain and Louis A. Pingel. On the robustness of recursive outlier detection procedures to nonnormality. Communications in Statistics: Theory and Methods, 10(13):1323–1334, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joe:1983:TPR

[JP83a] Harry Joe and Frank Proschan. Tests for properties of residual life. Communications in Statistics: Theory and Methods, 12(10): 1121–1134, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joe:1983:TPP

[JP83b] Harry Joe and Frank Proschan. Tests for properties of the percentile residual life function. Communications in Statistics: Theory and Methods, 12(10):1087–1119, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeyaratnam:1986:EAS

[JP86] S. Jeyaratnam and S. Panchapakesan. Estimation after subset selection from exponential populations. Communications in Statistics: Theory and Methods, 15(11):3459–3473, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jureckova:1987:AOS

[JP87] Jana Jurecková and Stephen Portnoy. Asymptotics for one-step M-estimators in regression with application to combining efficiency and high breakdown point. Communications in Statistics: Theory and Methods, 16(8):2187–2199, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeon:1988:MTS

[JP88a] T. J. Jeon and S. J. Park. Multiple time series model identification using concatenated sample cross-correlations. Communications in Statistics: Theory and Methods, 17(1):1–16, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeyaratnam:1988:PIR

[JP88b] S. Jeyaratnam and S. Panchapakesan. Prediction intervals for the random intercept linear model. Communications in Statistics: Theory and Methods, 17(9):3067–3073, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeyaratnam:1988:SUP

[JP88c] S. Jeyaratnam and S. Panchapakesan. Selection from uniform populations based on sample midranges and an associated estimation after selection. Communications in Statistics: Theory and Methods, 17(7):2303–2314, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1982:UAE

[JPD82] Ram B. Jain, Louis A. Pingel, and John L. Davidson. A unified approach for estimation and detection of outliers. Communications in Statistics: Theory and Methods, 11(25):2953–2976, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1982:CGM

[JR82] K. G. Janardan and B. Raja Rao. Characterization of generalized Markov–Pólya and generalized Pólya–Eggenberger distributions.

Communications in Statistics: Theory and Methods, 11(18):2113– 2124, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joiner:1980:SCF

[JRF80] J. Joiner, B. L. Raktoe, and W. T. Federer. On sum composition of fractional factorial designs. Communications in Statistics: Theory and Methods, 9(18):1875–1899, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1982:AEE

[JS82] Suneeta Joshi and Sathe T. S. On approximating the equivariant estimator of the mean of a normal distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 11(2):209–215, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1989:RRM

[JSB89] Kanchan Jain, Harshinder Singh, and Isha Bagai. Relations for reliability measures of weighted distributions. Communications in Statistics: Theory and Methods, 18(12):4393–4412, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janssen:1987:NRE

[JV87] P. Janssen and N. Veraverbeke. On nonparametric regression estimators based on regression quantiles. Communications in Statistics: Theory and Methods, 16(2):383–396, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kirmani:1980:CGD

[KA80] S. N. U. A. Kirmani and S. N. Alam. Characterization of the geometric distribution by the form of a predictor. Communications in Statistics: Theory and Methods, 9(5):541–547, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:1987:CPD

[KA87a] A. H. Khan and M. Masoom Ali. Characterization of probability distributions through higher order gap. Communications in Statistics: Theory and Methods, 16(5):1281–1287, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kirmani:1987:NWD

[KA87b] S. N. U. A. Kirmani and M. Ahsanullah. A note on weighted distributions. Communications in Statistics: Theory and Methods, 16(1):275–280, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1983:QPA

[Kab83] D. G. Kabe. A quadratic programming approach to the construction of simultaneous confidence intervals. Communications in Statistics: Theory and Methods, 12(17):2053–2058, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1984:MIM

[Kab84] D. G. Kabe. On the maximal invariance of MANOVA step down procedure statistics. Communications in Statistics: Theory and Methods, 13(20):2571–2581, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1986:GML

[Kab86] D. G. Kabe. On a GMANOVA model likelihood ratio test criterion. Communications in Statistics: Theory and Methods, 15(11):3419– 3427, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadiyala:1981:BBP

[Kad81] Krishna Kadiyala. Bounds on the biasing parameter in ridge regression. Communications in Statistics: Theory and Methods, 10 (22):2369–2372, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kafadar:1982:UBE

[Kaf82] Karen Kafadar. Using biweight M-estimates in the two-sample problem part 1: symmetric populations. Communications in Statistics: Theory and Methods, 11(17):1883–1901, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1980:PEB

[Kag80] Sanpei Kageyama. On properties of efficiency-balanced designs. Communications in Statistics: Theory and Methods, 9(6):597–616, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1981:CEB

[Kag81] Sanpei Kageyama. Constructions of efficiency-balanced designs. Communications in Statistics: Theory and Methods, 10(6):559– 580, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1988:OCB

[Kag88] Sanpei Kageyama. Optimum chemical balance weighing designs for estimating the total weight. Communications in Statistics: Theory and Methods, 17(8):2697–2704, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaigh:1983:QIE

[Kai83] W. D. Kaigh. Quantile interval estimation. Communications in Statistics: Theory and Methods, 12(21):2427–2443, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaigh:1988:STA

[Kai88] W. D. Kaigh. O-statistics and their applications. Communications in Statistics: Theory and Methods, 17(7):2191–2210, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kala:1981:PLE

[Kal81] Radosław Kala. Projectors and linear estimation in general linear models. Communications in Statistics: Theory and Methods, 10 (9):849–873, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalman:1984:WCD

[Kal84] R. E. Kalman. We can do something about multicollinearity. Communications in Statistics: Theory and Methods, 13(2):115–125, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamps:1988:DMO

[Kam88] U. Kamps. Distance measures in a one-parameter class of density functions. Communications in Statistics: Theory and Methods, 17 (6):2013–2019, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kane:1982:SGF

[Kan82] V. E. Kane. Standard and goodness-of-fit parameter estimation methods for the three-parameter lognormal distribution. Communications in Statistics: Theory and Methods, 11(17):1935–1957, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanto:1984:CIA

[Kan84] Antti J. Kanto. A characterization of the inverse autocorrelation function. Communications in Statistics: Theory and Methods, 13 (20):2503–2510, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanazawa:1988:DID

[Kan88a] Mitsuyo Kanazawa. On the distributions of the inner distance functions in the multivariate normal populations with unequal covariance matrices. Communications in Statistics: Theory and Methods, 17(12):4025–4038, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanazawa:1988:OVC

[Kan88b] Yuichiro Kanazawa. An optimal variable cell histogram. Communications in Statistics: Theory and Methods, 17(5):1401–1422, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaplan:1982:TRM

[Kap82a] Jack Kaplan. A theorem relating MINQUE and unweighted means estimators of variance components in the one-way design. Communications in Statistics: Theory and Methods, 11(4):423–428, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1982:MSD

[Kap82b] Russell F. Kappenman. On a method for selecting a distributional model. Communications in Statistics: Theory and Methods, 11 (6):663–672, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1983:IUP

[Kap83] Russell F. Kappenman. On the inferential use of predictive sample reuse. Communications in Statistics: Theory and Methods, 12(3):

363–380, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1985:TAE

[Kap85] Russell F. Kappenman. A testing approach to estimation. Communications in Statistics: Theory and Methods, 14(10):2365–2377, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1986:AES

[Kap86] Russell F. Kappenman. Adaptive m estimation of symmetric distribution location. Communications in Statistics: Theory and Methods, 15(10):2935–2951, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1989:UPT

[Kap89] Russell F. Kappenman. The use of power transformations for improved entropy estimation. Communications in Statistics: Theory and Methods, 18(9):3355–3364, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karakostas:1988:ECL

[Kar88] K. X. Karakostas. On the efficiency or a centrally located sampling design. Communications in Statistics: Theory and Methods, 17(8): 2755–2762, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karunamuni:1989:EBS

[Kar89a] R. J. Karunamuni. On empirical Bayes sequential estimation. Communications in Statistics: Theory and Methods, 18(7):2533– 2552, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karunamuni:1989:SIE

[Kar89b] R. J. Karunamuni. Some improvements on empirical Bayes testing in Lebesgue-exponential families. Communications in Statistics: Theory and Methods, 18(9):3157–3176, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katiyar:1980:RSD

[Kat80] Anand S. Katiyar. A rotation sampling design for inventory control. Communications in Statistics: Theory and Methods, 9(3):

309–320, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kay:1982:ATT

[Kay82] Richard Kay. The analysis of transition times in multistate stochastic processes using proportional hazard regression models. Communications in Statistics: Theory and Methods, 11(15):1743–1756, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kay:1989:SPC

[Kay89] R. J. Mc Kay. Simultaneous procedures for covariance matrices. Communications in Statistics: Theory and Methods, 18(2):429– 443, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1980:PEB

[KB80] Seong-In Kim and Do Sun Bai. On parameter estimation in Bernoulli trials with dependence. Communications in Statistics: Theory and Methods, 9(13):1401–1410, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:1981:WAL

[KB81] John P. Klein and Asit P. Basu. Weibull accelerated life tests when there are competing causes of failure. Communications in Statistics: Theory and Methods, 10(20):2073–2100, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:1982:ALT

[KB82a] John P. Klein and Asit P. Basu. Accelerated life tests under competing Weibull causes of failure. Communications in Statistics: Theory and Methods, 11(20):2271–2286, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koutrouvelis:1982:ADR

[KB82b] Ioannis A. Koutrouvelis and David F. Bauer. Asymptotic distribution of regression type estimators of parameters of stable laws. Communications in Statistics: Theory and Methods, 11(23):2715– 2730, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaiser:1983:SCI

[KB83] Lee D. Kaiser and David C. Bowden. Simultaneous confidence intervals for all linear contrasts of means with heterogeneous variances. Communications in Statistics: Theory and Methods, 12(1): 73–88, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1984:TSA

[KB84] S. Kocherlakota and N. Balakrishnan. Two sided acceptance sampling plans based on MML estimators. Communications in Statistics: Theory and Methods, 13(24):3123–3131, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1980:MVU

[KC80] Ashish Kumar and P. C. Consul. Minimum variance unbiased estimation for modified power series distribution. Communications in Statistics: Theory and Methods, 9(12):1261–1275, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

K:1984:ABD

[KC84a] Shah Arvind K. and P. L. Claypool. Analysis of binary data for the balanced one-way classification. Communications in Statistics: Theory and Methods, 13(11):1375–1390, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1984:AEF

[KC84b] Jerome P. Keating and Youn-Min Chou. An adaptive estimator frequently used in material fatigue testing. Communications in Statistics: Theory and Methods, 13(9):1047–1061, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kodell:1987:HCD

[KC87a] Ralph L. Kodell and James J. Chen. Handling cause of death in equivocal cases using the em algorithm. Communications in Statistics: Theory and Methods, 16(9):2565–2585, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kodell:1987:RDG

[KC87b] Ralph L. Kodell and James J. Chen. Rejoinder to discussion by G. E. Dinse of “Handling cause of death in equivocal causes using the EM algorithm”. Communications in Statistics: Theory and Methods, 16(9):2593–2595, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). Check: em or EM or something else??

Kodell:1987:RFD

[KC87c] Ralph L. Kodell and James J. Chen. Rejoinder to further discussion by G. E. Dinse of handling cause of death in equivocal cases using the EM algorithm. Communications in Statistics: Theory and Methods, 16(9):2603, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korwar:1982:EBD

[KD82] Ramesh M. Korwar and Ram C. Dahiya. Estimation of a bivariate distribution function from incomplete observations. Communications in Statistics: Theory and Methods, 11(8):887–897, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1983:DFD

[KD83] Anant M. Kshirsagar and Thomas M. Davis. Discriminant functions from the dummy variable space for growth curve models. Communications in Statistics: Theory and Methods, 12(4):427– 440, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1989:DBH

[KD89] Anant M. Kshirsagar and Sheela Deo. Distribution of the biased hypothesis sum of squares in linear models with missing observations. Communications in Statistics: Theory and Methods, 18(8): 2747–2754, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleijnen:1985:EEP

[KDK85] J. P. C. Kleijnen, H. M. M. T. Denis, and R. M. G. Kerckhoffs. Efficient estimation of power functions. Communications in Statistics: Theory and Methods, 14(8):1865–1878, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1983:EPB

[Kea83] Jerome P. Keating. Estimators of percentiles based on absolute loss. Communications in Statistics: Theory and Methods, 12(4): 441–447, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keener:1980:RTS

[Kee80] Robert W. Keener. Renewal theory and the sequential design of experiments with two states of nature. Communications in Statistics: Theory and Methods, 9(16):1699–1726, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kella:1986:DMB

[Kel86] Offer Kella. On the distribution of the maximum of bivariate normal random variables with general means and variances. Communications in Statistics: Theory and Methods, 15(11):3265–3276, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:1986:MMG

[Kem86a] C. D. Kemp. A modal method for generating binomial variables. Communications in Statistics: Theory and Methods, 15(3):805– 813, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kempthorne:1986:OMS

[Kem86b] Peter J. Kempthorne. Optimal minimax squared error risk estimation of the mean of a multivariate normal distribution. Communications in Statistics: Theory and Methods, 15(7):2145–2157, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kenneth:1981:TZP

[Ken81] Schechtman Kenneth. A test for zero point triserial correlation coefficient and its relationship with tests for homogeneity of means with ordered alternatives. Communications in Statistics: Theory and Methods, 10(12):1167–1182, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kepm:1986:WDM

[Kep86] A. W. Kepm. Weighted discrepancies and maximum likelihood estimation for discrete distributions. Communications in Statistics: Theory and Methods, 15(3):783–803, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ketellapper:1982:ERP

[Ket82] Ronald H. Ketellapper. On estimating the regression parameter in the simple errors-in-variables model without constant term. Communications in Statistics: Theory and Methods, 11(10):1147–1153, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koronacki:1989:DEU

[KF89] J. Koronacki and W. Feluch. Density estimation under simple partial censoring. Communications in Statistics: Theory and Methods, 18(5):1927–1942, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1987:ESN

[KG87] Ravindra Khattree and D. S. Gill. Estimation of signal to noise ratio using Mahalanobis distance. Communications in Statistics: Theory and Methods, 16(3):897–907, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1989:SPM

[KG89] D. G. Kabe and A. K. Gupta. On a superiority problem in misspecified restricted linear models. Communications in Statistics: Theory and Methods, 18(5):1753–1757, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kackar:1981:UTS

[KH81] Raghu N. Kackar and David A. Harville. Unbiasedness of twostage estimation and prediction procedures for mixed linear models. Communications in Statistics: Theory and Methods, 10(13): 1249–1261, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katzenbeisser:1986:AKS

[KH86] Walter Katzenbeisser and Peter Hackl. An alternative to the Kolmogorov–Smirnov two-sample test. Communications in Statistics: Theory and Methods, 15(4):1163–1177, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khamiss:1983:LLM

[Kha83a] H. J. Khamiss. Log-linear model analysis of the semi-symmetric intraclass contingency table. Communications in Statistics: Theory and Methods, 12(23):2723–2752, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1983:ECL

[Kha83b] C. G. Khatri. Estimation of the common location parameters for the two linear models. Communications in Statistics: Theory and Methods, 12(8):929–946, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1983:MDE

[Kha83c] C. G. Khatri. Multivariate discrete exponential family of distributions and their properties. Communications in Statistics: Theory and Methods, 12(8):877–893, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:1984:DAD

[Kha84] M. K. Khan. Discrete adaptive design in attribute life testing. Communications in Statistics: Theory and Methods, 13(12):1423– 1433, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khamis:1987:MPA

[Kha87a] H. J. Khamis. A measure of partial association between disease and genotype for the two-parent problem. Communications in Statistics: Theory and Methods, 16(7):2029–2060, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1987:CED

[Kha87b] Ravindra Khattree. On comparison of estimates of dispersion using generalized Pitman nearness criterion. Communications in Statistics: Theory and Methods, 16(1):263–274, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1988:ASI

[Kha88a] Ravindra Khattree. Asymmetric selection indices. Communications in Statistics: Theory and Methods, 17(11):3981–3993, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1988:IRL

[Kha88b] Ravindra Khattree. Inequality restricted linear selection indices. Communications in Statistics: Theory and Methods, 17(9):2959– 2980, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1989:MGS

[Kha89a] C. G. Khatri. Multivariate generalization of t’- statistic based on the mean square successive difference. Communications in Statistics: Theory and Methods, 18(5):1983–1992, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1989:SRV

[Kha89b] C. G. Khatri. Study of redundancy of vector variables in canonical correlations. Communications in Statistics: Theory and Methods, 18(4):1425–1440, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1988:DSL

[KHB^{+}88] Russell F. Kappenman, R. V. Hogg, G. K. Bril, S. M. Han, and L. Yuh. Detection of symmetry or lack of it and applications. Communications in Statistics: Theory and Methods, 17(12):4163– 4177, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khuri:1980:STP

[Khu80] A. I. Khuri. Simultaneous testing of parameter subsets in less than full rank models. Communications in Statistics: Theory and Methods, 9(6):617–627, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khuri:1982:DPP

[Khu82] A. I. Khuri. Direct products: a powerful tool for the analysis of balanced data. Communications in Statistics: Theory and Methods, 11(25):2903–2920, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalpana:1988:NCD

[KI88] Kusum Kalpana and Bagai Isha. A new class of distribution free procedures for testing homogeneity of scale parameters against ordered alternatives. Communications in Statistics: Theory and Methods, 17(5):1365–1376, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kikuchi:1983:IPP

[Kik83] David A. Kikuchi. The inverse of a partitioned positive-definite symmetric matrix. Communications in Statistics: Theory and Methods, 12(16):1889–1900, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kimber:1985:MTP

[Kim85] A. C. Kimber. Methods for the two-piece normal distribution. Communications in Statistics: Theory and Methods, 14(1):235– 245, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1986:APM

[Kim86] Jee Soo Kim. Asymptotic properties of the maximum likelihood estimator of a randomly censored exponential parameter. Communications in Statistics: Theory and Methods, 15(12):3637–3646, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1988:SRL

[Kim88a] Byung Soo Kim. Some remarks on locally most powerful unbiased tests for the detection of mixtures of Poisson or binomial distributions. Communications in Statistics: Theory and Methods, 17 (11):3733–3741, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1988:SSP

[Kim88b] Jee Soo Kim. A subset selection procedure for exponential populations under random censoring. Communications in Statistics: Theory and Methods, 17(1):183–206, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kotz:1982:EIG

[KJ82] Samuel Kotz and Norman L. Johnson. Errors in inspection and grading: distributional aspects of screening and hierarchal screening. Communications in Statistics: Theory and Methods, 11(18): 1997–2016, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kotz:1983:SDA

[KJ83] Samuel Kotz and Norman L. Johnson. Some distributions arising from faulty inspection with multitype defectives, and an application to grading. Communications in Statistics: Theory and Methods, 12(24):2809–2821, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kotz:1990:SCP

[KJ90] Samuel Kotz and Norman L. Johnson. Some comments on a paper by Aiala Barr. Communications in Statistics: Theory and Methods, 19(3):815–819, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Bar89].

Kotz:1988:IEG

[KJG88] Samuel Kotz, Norman L. Johnson, and Nina Gunther. Inspection errors in Graff–Roeloffs modification of hierarchical Dorfman screening procedures. Communications in Statistics: Theory and Methods, 17(1):139–168, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kupper:1983:APC

[KJS^{+}83] Lawrence L. Kupper, Joseph M. Janis, Ibrahim A. Salama, Carl N. Yoshizawa, Bernard G. Greenberg, and H. H. Winsborough. Ageperiod-cohort analysis: an illustration of the problems in assessing interaction in one observation per cell data. Communications in Statistics: Theory and Methods, 12(23):2779–2807, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kupper:1984:A

[KJS^{+}84] Lawrence L. Kupper, Joseph M. Janis, Ibrahim A. Salama, Carl N. Yoshizawa, and Bernard G. Greenberg. Acknowledgement. Communications in Statistics: Theory and Methods, 13(9):1177, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katti:1980:EAE

[KK80] S. K. Katti and Magdy Khedr. Efficiency of an adaptive estimator of a log-zero-Poisson parameter. Communications in Statistics: Theory and Methods, 9(12):1221–1246, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1981:DSM

[KK81] K. Kocherlakota and S. Kocherlakota. On the distribution of_{r} in samples from the mixtures of bivariate normal populations. Communications in Statistics: Theory and Methods, 10(19):1943–1966, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kempthorne:1983:SMB

[KK83] W. Joan Kempthorne and Gary G. Koch. Statistical methodology for binary data from a set of partially structured two-way random cross-classifications. Communications in Statistics: Theory and Methods, 12(23):2761–2777, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalton:1984:SER

[KK84] Graham Kalton and Leslie Kish. Some efficient random imputation methods. Communications in Statistics: Theory and Methods, 13 (16):1919–1939, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1985:STI

[KK85] K. Kocherlakota and S. Kocherlakota. On some tests for independence in nonnormal situations: Neyman’s C(A) test. Communications in Statistics: Theory and Methods, 14(6):1453–1470, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:1986:DPM

[KK86a] A. W. Kemp and C. D. Kemp. Distributional properties of a model for the spread or drug abuse. Communications in Statistics: Theory and Methods, 15(11):3287–3298, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1986:GFT

[KK86b] S. Kocherlakota and K. Kocherlakota. Goodness of fit tests for discrete distributions. Communications in Statistics: Theory and Methods, 15(3):815–829, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1987:SPB

[KKB87] S. Kocherlakota, K. Kocherlakota, and N. Balakrishnan. Selection procedures based on normal and nonnormal variables with single and multiple characters. Communications in Statistics: Theory and Methods, 16(10):3007–3023, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleinbaum:1982:LRA

[KKC82] David G. Kleinbaum, Lawrence L. Kupper, and Lloyd E. Chambless. Logistic regression analysis of epidemiologic data: theory and practice. Communications in Statistics: Theory and Methods, 11 (5):485–547, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleijnen:1987:JEW

[KKOV87] Jack P. C. Kleijnen, Peter C. A. Karremans, Wim K. Oortwijn, and Willem J. H. Van Groenendaal. Jackknifing estimated weighted least squares: Jewls. Communications in Statistics: Theory and Methods, 16(3):747–764, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kao:1980:SSP

[KL80] Samuel C. Kao and Tze Leung Lai. Sequential selection procedures based on confidence sequences for normal populations. Communications in Statistics: Theory and Methods, 9(16):1657–1676, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalgh:1982:GQE

[KL82a] W. D. Kalgh and Peter A. Lachenbruch. A generalized quantile estimator. Communications in Statistics: Theory and Methods, 11 (19):2217–2238, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lam:1982:BMD

[kL82b] Hing kam Lam. Bounded multinomial distribution. Communications in Statistics: Theory and Methods, 11(16):1869–1880, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kindermann:1984:CBA

[KL84] R. P. Kindermann and M. S. Levy. A condition for best asymptotic normality in a regular family of distributions. Communications in Statistics: Theory and Methods, 13(17):2173–2179, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnaiah:1986:CES

[KL86] P. R. Krishnaiah and Jugan Lin. Complex elliptically symmetric distributions. Communications in Statistics: Theory and Methods, 15(12):3693–3718, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klefsjo:1983:STA

[Kle83] Bengt Klefsjö. Some tests against aging based on the total time on test transform. Communications in Statistics: Theory and Methods, 12(8):907–927, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleijnen:1986:SSE

[Kle86] Jack P. C. Kleijnen. A scenario for sequential experimentation. Communications in Statistics: Theory and Methods, 15(10):3067– 3078, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleijnen:1987:PCR

[Kle87a] Jack P. C. Kleijnen. A practical comment on resolution IV designs: Bad properties of quantitative factors. Communications in Statistics: Theory and Methods, 16(3):805–812, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleijnen:1987:RRG

[Kle87b] Jack P. C. Kleijnen. Review of random and group-screening designs. Communications in Statistics: Theory and Methods, 16(10): 2885–2900, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kleffe:1988:RMT

[Kle88] J. Kleffe. On the role of MINQUE in testing of hypotheses under hiked linear models. Communications in Statistics: Theory and Methods, 17(4):1287–1309, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kraft:1985:ESD

[KLV85] C. H. Kraft, Y. Lepage, and C. Van Eeden. Estimation of a symmetric density function. Communications in Statistics: Theory and Methods, 14(2):273–288, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubat:1980:EPB

[KM80] Peter Kubat and Philip McDunnough. Estimating the precision of a bivariate population. Communications in Statistics: Theory and Methods, 9(5):489–500, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khuri:1981:DRR

[KM81] A. I. Khuri and R. H. Myers. Design related robustness of tests in regression models. Communications in Statistics: Theory and Methods, 10(3):223–235, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khorasani:1982:SCB

[KM82] Fariborz Khorasani and George A. Milliken. Simultaneous confidence bands for nonlinear regression models. Communications in Statistics: Theory and Methods, 11(11):1241–1253, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1984:CEB

[KM84a] Jerome P. Keating and Rhonda C. Magel. Comments on estimators based on error in the predicted distribution function. Communications in Statistics: Theory and Methods, 13(15):1889–1900, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kegeyama:1984:TMC

[KM84b] Sanpei Kegeyama and R. N. Mohan. Three methods of constructing PBIB designs. Communications in Statistics: Theory and Methods, 13(25):3185–3189, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuwada:1984:RBF

[KM84c] Masahide Kuwada and Masaaki Matsuura. On robustness of balanced fractional 2^{m} factorial designs of resolution VII derived from simple arrays. Communications in Statistics: Theory and Methods, 13(14):1675–1691, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunert:1987:SRO

[KM87] J. Kunert and R. J. Martin. Some results on optimal design under a first-order autoregression and on finite Williams Type II designs. Communications in Statistics: Theory and Methods, 16(7):1901– 1922, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1988:PRT

[KM88] Jerome P. Keating and Robert L. Mason. Preference regions and their probabilities based on a rectangular grid. Communications in Statistics: Theory and Methods, 17(6):1973–1983, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulp:1983:ANNa

[KN83a] R. W. Kulp and B. N. Nagarsenker. Asymptotic non-null distribution of a test connected with exponential populations. Communications in Statistics: Theory and Methods, 12(24):2923–2930, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulp:1983:ANNb

[KN83b] R. W. Kulp and B. N. Nagarsenker. Asymptotic non-null distribution of a test of equality of exponential populations. Communications in Statistics: Theory and Methods, 12(24):2931–2936, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulp:1983:FAE

[KN83c] R. W. Kulp and B. N. Nagarsenker. Further asymptotic expansions of the distribution of the sphericity test. Communications in Statistics: Theory and Methods, 12(1):119–129, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulp:1984:TIB

[KN84] R. W. Kulp and B. N. Nagarsenker. Test of independence between tow sets of variates. Communications in Statistics: Theory and Methods, 13(6):685–698, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:1985:CRM

[KN85] P. Kabaila and G. Nelson. On confidence regions for the mean of a multivariate time series. Communications in Statistics: Theory and Methods, 14(3):735–753, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1987:DOB

[KN87] Ravindra Khattree and Dayanand N. Naik. Detection of outliers in bivariate time series data. Communications in Statistics: Theory and Methods, 16(12):3701–3714, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadiyala:1982:RPR

[KO82] Rao K. Kadiyala and Dennis Oberhelman. Response predictions in regressions on panel data. Communications in Statistics: Theory and Methods, 11(23):2699–2714, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadiyala:1984:ATH

[KO84] K. Rao Kadiyala and H. Dennis Oberhelman. Alternative tests for heteroscedasticity of disturbances: a comparative study. Communications in Statistics: Theory and Methods, 13(8):987–1001, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kochar:1985:TEA

[Koc85] Subhash C. Kochar. Testing exponentiality against monotone failure rate average. Communications in Statistics: Theory and Methods, 14(2):381–392, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1986:BIG

[Koc86] Subrahmaniam Kocherlakota. The bivariate inverse Gaussian distribution: an introduction. Communications in Statistics: Theory and Methods, 15(4):1081–1112, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kokolakis:1985:MAP

[Kok85] George E. Kokolakis. A martingale approach to the problem of dimensionality and performance of Bayesian classifiers: asymptotic results. Communications in Statistics: Theory and Methods, 14(4): 927–935, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Konijn:1981:MLE

[Kon81] H. S. Konijn. Maximum likelihood estimator and confidence intervals for a simple errors in variables model. Communications in Statistics: Theory and Methods, 10(10):983–996, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koorts:1985:SOE

[Koo85] Albertus S. Koorts. Second order efficiency for a class of estimators within the multinomial family. Communications in Statistics:

Theory and Methods, 14(10):2331–2343, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korwar:1985:BVG

[Kor85] R. M. Korwar. Bounds for the variance of the Graybill–Deal estimator of the common mean of two normal distributions. Communications in Statistics: Theory and Methods, 14(2):355–361, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korn:1987:DRC

[Kor87] E. L. Korn. Data-randomized confidence intervals for discrete distributions. Communications in Statistics: Theory and Methods, 16 (3):705–715, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koske:1987:FOR

[Kos87] J. K. Arap Koske. A fourth order rotatable design in four dimensions. Communications in Statistics: Theory and Methods, 16(9): 2747–2753, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koti:1989:MPT

[Kot89a] Kallappa M. Koti. Minimax policies for a two-location/two-product inventory system. Communications in Statistics: Theory and Methods, 18(12):4617–4630, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koti:1989:SII

[Kot89b] Kallappa M. Koti. Some inequalities in incomplete beta functions by using statistical techniques. Communications in Statistics: Theory and Methods, 18(2):691–696, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koutrouvelis:1981:LSQ

[Kou81a] Ioannis A. Koutrouvelis. Large-sample quantile estimation in Pareto laws. Communications in Statistics: Theory and Methods, 10(2):189–201, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koutrouvelis:1981:TNS

[Kou81b] Ioannis A. Koutrouvelis. Tests for normality in stable laws. Communications in Statistics: Theory and Methods, 10(6):581–595, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1983:TSA

[Koz83] James A. Koziol. Tests for symmetry about an unknown value based on the empirical distribution function. Communications in Statistics: Theory and Methods, 12(24):2823–2846, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1986:AMN

[Koz86a] James A. Koziol. Assessing multivariate normality: a compendium. Communications in Statistics: Theory and Methods, 15(9):2763– 2783, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1986:NAD

[Koz86b] James A. Koziol. A note on the asymptotic distribution of Mardia’s measure of multivariate kurtosis. Communications in Statistics: Theory and Methods, 15(5):1507–1513, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1989:MWS

[Koz89] James A. Koziol. A modification of Watson’s statistic for goodnessof-fit. Communications in Statistics: Theory and Methods, 18(10): 3739–3747, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kingston:1980:RSS

[KP80a] Joseph V. Kingston and Jagdish K. Patel. A restricted subset selection procedure for Weibull populations. Communications in Statistics: Theory and Methods, 9(13):1371–1383, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kingston:1980:SBO

[KP80b] Joseph V. Kingston and Jagdish K. Patel. Selecting the best one of several Weibull populations. Communications in Statistics: Theory and Methods, 9(4):383–398, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kingston:1981:IEL

[KP81] Joseph V. Kingston and Jagdish K. Patel. Interval estimation of the largest reliability of k Weibull populations. Communications in Statistics: Theory and Methods, 10(22):2279–2298, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kingston:1982:CWP

[KP82] Joseph V. Kingston and Jagdish K. Patel. Classifying Weibull populations with respect to a control. Communications in Statistics: Theory and Methods, 11(8):899–909, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:1984:UAF

[KP84] S. R. Kulkarni and S. R. Paranjape. Use of Andrews’ function plot technique to construct control curves for multivariate process. Communications in Statistics: Theory and Methods, 13(20):2511– 2533, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1985:NSC

[KP85] Sanpei Kageyama and P. D. Puri. A new special class of PEB designs. Communications in Statistics: Theory and Methods, 14 (7):1731–1744, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koske:1986:FOR

[KP86] J. K. Arap Koske and M. S. Patel. A fourth order rotatable design in three dimensions. Communications in Statistics: Theory and Methods, 15(11):3435–3444, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koske:1987:CFO

[KP87] J. K. Arap Koske and M. S. Patel. Construction of fourth order rotatable designs with estimation of corresponding response surface. Communications in Statistics: Theory and Methods, 16(5): 1361–1376, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaushik:1989:PDC

[KP89a] L. S. Kaushik and P. D. Puri. Partial diallel crosses based on generalized right angular association scheme. Communications in Statistics: Theory and Methods, 18(7):2501–2510, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:1989:CBC

[KP89b] Ruben Klein and S. James Press. Contextual Bayesian classification of remotely sensed data. Communications in Statistics: Theory and Methods, 18(9):3177–3202, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koske:1989:SWO

[KP89c] J. K. Arap Koske and M. S. Patel. A simpler way of obtaining non-singularity conditions of rotatability. Communications in Statistics: Theory and Methods, 18(7):2489–2500, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korwa:1983:GIP

[KPS83] Ramesh M. Korwa, Govind Prasad, and Ashok Sahai. A generalized improvement procedure for variance bounds for minimum variance unbiased estimator in inverse sampling. Communications in Statistics: Theory and Methods, 12(15):1807–1812, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1988:SCO

[KPS88] Jee Soo Kim, Frank Proschan, and Jayaram Sethuraman. Stochastic comparisons of order statistics, with applications in reliability. Communications in Statistics: Theory and Methods, 17(7):2151– 2172, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krafft:1982:GPW

[KR82] Olaf Krafft and Ludger Rüschendorf. On the gain in power when using additional observations. Communications in Statistics: Theory and Methods, 11(7):787–800, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kepner:1984:CTB

[KR84] James L. Kepner and Ronald H. Randles. Comparison of tests for bivariate symmetry versus location and/or scale alternatives. Communications in Statistics: Theory and Methods, 13(8):915– 930, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1986:TIB

[KR86] D. G. Kabe and U. L. Gouranga Rao. On tests of independence between stochastic regressors and disturbances. Communications in Statistics: Theory and Methods, 15(12):3617–3623, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnamoorthy:1988:MVU

[KR88] K. Krishnamoorthy and Vijay K. Rohatgi. Minimum variance unbiased estimation in some nonregular families. Communications in Statistics: Theory and Methods, 17(11):3757–3765, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kochar:1989:TSM

[KR89a] Subhash C. Kochar and T. Ramallingam. Testing for superadditivity of the mean value function of a non-homogeneous Poisson process. Communications in Statistics: Theory and Methods, 18 (4):1549–1562, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1989:LRT

[KR89b] Anant M. Kshirsagar and R. Radhakrishnan. A likelihood ratio test for interaction among variances. Communications in Statistics: Theory and Methods, 18(12):4511–4517, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kraer:1985:OLS

[Krä85] Walter Kräer. Ordinary least squares estimation of simultaneous equation systems with trended data: further results. Communications in Statistics: Theory and Methods, 14(8):1997–2005, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnamoorthy:1989:UET

[KRB89] K. Krishnamoorthy, Vijay K. Rohatgi, and Josef Blass. Unbiased estimation in type II censored samples from a one-truncation parameter density. Communications in Statistics: Theory and Methods, 18(3):1023–1030, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kafadar:1983:OST

[KRS83] Karen Kafadar, John Rice, and Clifford Spiegelman. One-sided trimming in small samples with asymmetric contamination. Communications in Statistics: Theory and Methods, 12(4):477–496, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatrl:1981:UEF

[KS81a] C. G. Khatrl and K. R. Shah. On the unbiased estimation of fixed effects in a mixed model for growth curves. Communications in Statistics: Theory and Methods, 10(4):401–406, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khursheed:1981:PDM

[KS81b] Alam Khursheed and K. M. Lai Saxena. Positive dependence in multivariate distributions. Communications in Statistics: Theory and Methods, 10(12):1183–1196, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1982:BSTa

[KS82a] Subrahmaniam Kocherlakota and M. Singh. On the behaviour of some transforms of the sample correlation coefficient: discrete bivariate populations. Communications in Statistics: Theory and Methods, 11(18):2017–2043, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1982:BSTb

[KS82b] Subrahmaniam Kocherlakota and M. Singh. On the behaviour of some transforms of the sample correlation coefficient in samples from the bivariate t and the bivariate X^{2} distribution. Communications in Statistics: Theory and Methods, 11(18):2045–2060, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulatunga:1984:TSH

[KS84] D. D. Sarath Kulatunga and Syoichi Sasabuchi. A test of simultaneous homogeneity against ordered alternatives in multifactorial design. Communications in Statistics: Theory and Methods, 13 (25):3173–3183, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kafaei:1985:ASD

[KS85] Mohammad-Ali Kafaei and Peter Schmidt. On the adequacy of the “Sargan distribution” as an approximation to the normal. Communications in Statistics: Theory and Methods, 14(3):509–526, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1987:EBT

[KS87] Anant M. Kshirsagar and Sharon A. Naberhuis Stehouwer. Efficiency of Bailey’s trigonometric levels in factorial experiments. Communications in Statistics: Theory and Methods, 16(6):1655– 1673, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1988:SEO

[KS88] Somesh Kumar and Divakar Sharma. Simultaneous estimation of ordered parameters. Communications in Statistics: Theory and Methods, 17(12):4315–4336, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1989:PEO

[KS89] Somesh Kumar and Divakar Sharma. On the Pitman estimator of ordered normal means. Communications in Statistics: Theory and Methods, 18(11):4163–4175, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kudo:1981:TEN

[KSC81] A. Kudo, S. Sasabuchi, and J. R. Choi. Test of equality of normal means in the absence of independent estimator of variance. Communications in Statistics: Theory and Methods, 10(7):659– 668, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kotz:1986:EIE

[KSJ86] Samuel Kotz, Mao Shisong, and Norman L. Johnson. Effects of inspection errors on curtailed Dorman-type procedures. Communications in Statistics: Theory and Methods, 15(3):831–838, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1988:PEB

[KSR88] Sanpei Kageyama, G. M. Saha, and Mukerjee Rahul. D^{−1}-partially efficiency-balanced designs with at most two efficiency classes. Communications in Statistics: Theory and Methods, 17(5):1669– 1683, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1984:CST

[KT84] Jerome P. Keating and Ram C. Tripathi. Comparison of some two sample tests for equality of variances. Communications in Statistics: Theory and Methods, 13(21):2691–2706, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koreisha:1985:ITF

[KT85] Sergio G. Koreisha and Stanley A. Taylor. Identification of transfer function models: an asymptotic test of significance for the corner method. Communications in Statistics: Theory and Methods, 14 (1):159–173, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ki:1986:UDS

[KT86] Fanny Ki and Kam-Wah Tsuf. Universal domination of Stein-type estimators. Communications in Statistics: Theory and Methods, 15 (7):2159–2173, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:1988:MRS

[Kub88] Tatsuya Kubokawa. Monotonicity of risk for a shrinkage estimator of a multivariate normal mean. Communications in Statistics: Theory and Methods, 17(2):499–506, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:1989:ECP

[Kub89] Tatsuia Kubokawa. Estimating common parameters of growth curve models under a quadratic loss. Communications in Statistics: Theory and Methods, 18(9):3149–3155, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulatunga:1985:THN

[Kul85] D. D. Sarath Kulatunga. Testing homogeneity of t normal means against ordered alternatives in 2 groups having uncommon and unknown variance. Communications in Statistics: Theory and Methods, 14(3):701–711, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:1987:LSP

[Kul87a] P. M. Kulkarni. Large sample prediction regions with estimated parameters. Communications in Statistics: Theory and Methods, 16 (6):1591–1605, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:1987:LSE

[Kul87b] Pandurang M. Kulkarni. Large sample estimation and prediction for explosive growth curve models. Communications in Statistics: Theory and Methods, 16(9):2677–2696, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulperger:1988:MCA

[Kul88] R. J. Kulperger. A Monte Carlo approximation of the distributions of the maximum of various Brownian bridges. Communications in Statistics: Theory and Methods, 17(12):4389–4397, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1980:NMV

[Kum80] Ashish Kumar. A note on minimum variance unbiased estimators of power series distributions. Communications in Statistics: Theory and Methods, 9(5):549–556, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1981:SAL

[Kum81] Ashish Kumar. Some applications of Lagrangian distributions in queueing theory and epidemiology. Communications in Statistics: Theory and Methods, 10(14):1429–1436, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumazawa:1983:CTS

[Kum83] Yoshiki Kumazawa. A class of test statistics for testing whether new is better than used. Communications in Statistics: Theory and Methods, 12(3):311–321, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunio:1983:VUE

[Kun83] Shimizu Kunio. Variances of UMVU estimators for means and variances after using a normalizing transformation. Communications in Statistics: Theory and Methods, 12(9):975–985, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunert:1984:DBC

[Kun84] Joachim Kunert. Designs balanced for circular residual effects. Communications in Statistics: Theory and Methods, 13(21):2665– 2671, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuwada:1983:BAD

[Kuw83] Masahide Kuwada. Best alias designs in some class of balanced fractional 3^{m} factorial designs of resolution V. Communications in Statistics: Theory and Methods, 12(9):1039–1051, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuwada:1989:AVB

[Kuw89] Masahide Kuwada. Analysis of variance of balanced fractional 2^{n} factorial designs of resolution 2l+1. Communications in Statistics: Theory and Methods, 18(10):3883–3905, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1980:EDD

[KV80] B. K. Kim and J. Van Ryzin. Estimation of the derivative of a density and related functions. Communications in Statistics: Theory and Methods, 9(7):697–709, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ketellapper:1983:FEA

[KW83] Ronald H. Ketellapper and Joop A. Weisbeek. Further evidence on the appropriateness of a robust estimation procedure for the structural errors-in-variables model. Communications in Statistics: Theory and Methods, 12(13):1511–1522, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kapteyn:1984:EVC

[KW84] Arie Kapteyn and Tom Wansbeek. Errors in variables: consistent adjusted least squares (CALS) estimation. Communications in Statistics: Theory and Methods, 13(15):1811–1837, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koehler:1986:CST

[KW86] Kenneth J. Koehler and Jeffrey R. Wilson. Chi-square tests for comparing vectors of proportions for several cluster samples. Communications in Statistics: Theory and Methods, 15(10):2977–2990, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:1981:ABI

[KY81] Ruben Klein and Victor J. Yohai. Asymptotic behavior of iterative M-estimators for the linear model. Communications in Statistics:

Theory and Methods, 10(23):2373–2388, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kudo:1982:AFU

[KY82] Akio Kudo and Jing Shing Yao. An approximate formula for p(l, k) used in order restricted inferences, and the analysis of freezing dates of Lake Mendota. Communications in Statistics: Theory and Methods, 11(6):607–614, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lackritz:1984:VAV

[Lac84] James R. Lackritz. Variance approximation for volume fraction estimators from areal sampling. Communications in Statistics: Theory and Methods, 13(1):81–92, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lai:1984:ISE

[Lai84] T. L. Lai. Incorporating scientific, ethical and economic considerations into the design of clinical trials in the pharmaceutical industry: a sequential approach. Communications in Statistics: Theory and Methods, 13(19):2355–2368, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lai:1985:RSS

[Lai85] C. D. Lai. On the reliability of a standby system composed of two dependent exponential components. Communications in Statistics: Theory and Methods, 14(4):851–860, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laitila:1989:AMB

[Lai89] Thomas Laitila. Asymptotic misspecification biases for Heckman’s two step estimator. Communications in Statistics: Theory and Methods, 18(2):743–753, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lam:1986:STS

[Lam86] Yuk-Miu Lam. Stein’s two sample test of a general linear hypothesis and a noncentral F -type distribution. Communications in Statistics: Theory and Methods, 15(10):3019–3033, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lam:1989:MCR

[Lam89] K. Lam. The multiple comparison of ranked parameters. Communications in Statistics: Theory and Methods, 18(4):1217–1237, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LaRiccia:1984:PEB

[LaR84] Vincent N. LaRiccia. Parameter estimation based upon nonparametric function estimators. Communications in Statistics: Theory and Methods, 13(22):2771–2792, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LaRiccia:1988:PTP

[LaR88] Vincent N. LaRiccia. Parametric testing procedures for type II censored data sets based on the quantile function. Communications in Statistics: Theory and Methods, 17(2):519–542, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawrence:1980:C

[Law80] Willard E. Lawrence. Comments. Communications in Statistics: Theory and Methods, 9(17):1843–1845, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1984:RRA

[LB84] Yih-Wu Liu and Richard H. Bee. Ridge regression: an application in criminology. Communications in Statistics: Theory and Methods, 13(2):263–271, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lucas:1982:RCR

[LC82] James M. Lucas and Ronald B. Crosier. Robust cusum: a robustness study for cusum quality control schemes. Communications in Statistics: Theory and Methods, 11(23):2669–2687, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1985:SOR

[LC85] Tze-San Lee and Don B. Campbell. Selecting the optimum k in ridge regression. Communications in Statistics: Theory and Methods, 14(7):1589–1604, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leeper:1987:HTT

[LC87] James D. Leeper and Hang Ling Chang. Hypothesis testing for twogroup incomplete-data growth curve problems. Communications in Statistics: Theory and Methods, 16(12):3715–3728, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lan:1984:MFS

[LDH84] K. K. Gordon Lan, David L. Detlets, and Max Halperin. More flexible sequential and non-sequential designs in long-term clinical trial. Communications in Statistics: Theory and Methods, 13(19): 2339–2353, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lecoutre:1985:RTA

[Lec85] B. Lecoutre. Reconsideration of the F -test of the analysis of variance: the semi-Bayesian significance test. Communications in Statistics: Theory and Methods, 14(10):2437–2446, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1981:EVM

[Lee81] Kok Huat Lee. Estimation of variance of mean using known coefficient of variation. Communications in Statistics: Theory and Methods, 10(5):503–514, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1985:EVS

[Lee85] A. J. Lee. On estimating the variance of a U-statistic. Communications in Statistics: Theory and Methods, 14(2):289–301, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1987:COD

[Lee87a] Carl M.-S. Lee. Constrained optimal designs for regression models. Communications in Statistics: Theory and Methods, 16(3):765– 783, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1987:DFM

[Lee87b] Sik-Yum Lee. A distribution-free method for structural equation models with incomplete data. Communications in Statistics: Theory and Methods, 16(4):1133–1151, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1987:EMS

[Lee87c] Youngjo Lee. The exact mean squared error risks of preliminary test type estimators for the multivariate normal mean. Communications in Statistics: Theory and Methods, 16(7):1965–1975, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1988:ESE

[Lee88] Sik-Yum Lee. Estimation of structural equation models with exact and stochastic prior information. Communications in Statistics: Theory and Methods, 17(5):1565–1589, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemmer:1980:SER

[Lem80] H. H. Lemmer. Some empirical results on the two-way analysis of variance by ranks. Communications in Statistics: Theory and Methods, 9(14):1427–1438, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemmer:1981:NSE

[Lem81] H. H. Lemmer. Note on shrinkage estimators for the binomial distribution. Communications in Statistics: Theory and Methods, 10 (10):1017–1027, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1988:ACS

[Leu88a] C. Y. Leung. Anderson’s classification statistic based on a poststratified training sample. Communications in Statistics: Theory and Methods, 17(5):1659–1667, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1988:CCP

[Leu88b] C. Y. Leung. Covariate classification by post-stratification. Communications in Statistics: Theory and Methods, 17(11):3869–3880, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1989:SLL

[Leu89] C. Y. Leung. The Studentized location linear discriminant function. Communications in Statistics: Theory and Methods, 18(11): 3977–3990, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levin:1983:CIM

[Lev83] Bruce Levin. On calculations involving the maximum cell frequency. Communications in Statistics: Theory and Methods, 12 (11):1299–1327, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1982:IMI

[LH82] Lon-Mu Liu and Dominique M. Hanssens. Identification of multiple-input transfer function models. Communications in Statistics: Theory and Methods, 11(3):297–314, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1982:NJS

[Li82] Tze Fen Li. A note on James–Stein and Bayes empirical Bayes estimators. Communications in Statistics: Theory and Methods, 11 (9):1029–1043, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1987:AOB

[Li87] Tze Fen Li. On asymptotic optimality of Bayes empirical Bayes estimators. Communications in Statistics: Theory and Methods, 16 (4):1087–1099, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1988:EPR

[Li88] Hung C. Li. The exact probability that the roots of quadratic, cubic, and quartic equations are all real if the equation coefficients are random. Communications in Statistics: Theory and Methods, 17(2):395–409, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1988:IRS

[Lia88] TaChen Liang. Isotonic rules for selecting good exponential populations. Communications in Statistics: Theory and Methods, 17 (6):1871–1887, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1986:UCR

[Lin86] Paul K. H. Lin. Using the Chinese Remainder Theorem in constructing confounded designs for mixed factorial experiments.

Communications in Statistics: Theory and Methods, 15(4):1389– 1398, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1987:CMF

[Lin87a] Paul K. H. Lin. Confounding in mixed factorial experiments through isomorphisms. Communications in Statistics: Theory and Methods, 16(2):421–429, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1987:SMC

[Lin87b] Paul K. H. Lin. A simple method for constructing confounded designs for mixed factorial experiments. Communications in Statistics: Theory and Methods, 16(2):407–419, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1987:MAB

[Lin87c] Pi-Erh Lin. Measures of association between vectors. Communications in Statistics: Theory and Methods, 16(2):321–338, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Link:1989:TEA

[Lin89] William A. Link. Testing for exponentiality against monotone failure rate average alternatives. Communications in Statistics: Theory and Methods, 18(8):3009–3017, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liski:1984:RAT

[Lis84] Erkki P. Liski. Regression analysis with two mixed forms of heteroscedasticity. Communications in Statistics: Theory and Methods, 13(8):1015–1030, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liski:1988:TMS

[Lis88] Erkki P. Liski. A test of the mean square error criterion for linear admissible estimators. Communications in Statistics: Theory and Methods, 17(11):3743–3756, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1985:TBT

[Liu85] Shu-Ing Liu. Theory of bilinear time series models. Communications in Statistics: Theory and Methods, 14(10):2549–2561, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1989:ISA

[Liu89] Lon-Mu Liu. Identification of seasonal ARIMA models using a filtering method. Communications in Statistics: Theory and Methods, 18(6):2279–2288, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1981:IEU

[LK81] Hoi M. Leung and Lawrence L. Kupper. Interval estimation using the folded logarithmic transformation. Communications in Statistics: Theory and Methods, 10(10):997–1015, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LaRiccia:1983:AEC

[LK83] Vincent N. La Riccia and Ross P. Kindermann. An asymptotically efficient closed form estimator for the three-parameter lognormal distribution. Communications in Statistics: Theory and Methods, 12(2):243–261, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lund:1982:DEN

[LL82] Richard E. Lund and M. Gail Linnell. Description and evaluation of a nested cube experimental design. Communications in Statistics: Theory and Methods, 11(20):2297–2313, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1982:GJS

[LlHK82] Tze Fen Li and lea Hon Kuo. Generalized James–Stein estimators. Communications in Statistics: Theory and Methods, 11(20):2249– 2257, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lloyd:1981:SHI

[Llo81] Emlyn Lloyd. Stochastic hydrology: An introduction to wet statistics for dry statisticians. Communications in Statistics: Theory and Methods, 10(15):1505–1522, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lloyd:1987:OML

[Llo87a] C. J. Lloyd. Optimality of marginal likelihood estimating equations. Communications in Statistics: Theory and Methods, 16(6):

1733–1741, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lloyd:1987:PQB

[Llo87b] C. J. Lloyd. Pivotal quantities based on the conditional distribution function transform in models with accessory parameters. Communications in Statistics: Theory and Methods, 16(2):361– 374, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lorenzen:1981:SLP

[LM81] Thomas J. Lorenzen and Gary C. McDonald. Selecting logistic populations using the sample medians. Communications in Statistics: Theory and Methods, 10(2):101–124, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawrence:1984:RRE

[LM84a] Kenneth D. Lawrence and Lawrence C. Marsh. Robust ridge estimation methods for predicting U.S. coal mining fatalities. Communications in Statistics: Theory and Methods, 13(2):139–149, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyle:1984:FUA

[LM84b] Broemeling Lyle and Land Margaret. On forecasting with univariate autoregressive processes: a Bayesian approach. Communications in Statistics: Theory and Methods, 13(11):1305–1320, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyle:1984:APD

[LM84c] Broemeling Lyle and Yusoff Abdullah Mat. An approximation to the poly-t distribution. Communications in Statistics: Theory and Methods, 13(11):1407–1422, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1986:MEI

[LM86] Pi-Erh Lin and Amany M. Mousa. Minimax estimation of independent normal means under a quadratic loss function with unknown weights. Communications in Statistics: Theory and Methods, 15 (7):2175–2189, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lamotte:1988:CIT

[LMP88] Lynn Roy Lamotte, Archer McWhorter, Jr., and Ram A. Prasad. Confidence intervals and tests on the variance ratio in random models with two variance components. Communications in Statistics: Theory and Methods, 17(4):1135–1164, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1989:UEA

[LO89] Huaixiang Li and D. B. Owen. On the UMVU estimators after using a normalizing transformation. Communications in Statistics: Theory and Methods, 18(3):801–816, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loh:1984:PAR

[Loh84a] Wei-Yin Loh. Partially-adaptive robust estimators of location via exponential embedding. Communications in Statistics: Theory and Methods, 13(20):2549–2570, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loh:1984:RQR

[Loh84b] Wei-Yin Loh. Random quotients and robust estimation. Communications in Statistics: Theory and Methods, 13(22):2757–2769, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loh:1985:NGK

[Loh85] Wei-Yin Loh. A note on the geometry of Kullback–Leibler information numbers. Communications in Statistics: Theory and Methods, 14(4):895–904, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lorenzen:1984:RBD

[Lor84] Thomas J. Lorenzen. Randomization and blocking in the design of experiments. Communications in Statistics: Theory and Methods, 13(21):2601–2623, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loynes:1987:CVR

[Loy87] R. M. Loynes. Cross-validatory, recursive and partial residuals for non-linear models: theory. Communications in Statistics: Theory and Methods, 16(4):997–1011, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loynes:1989:TDL

[Loy89] R. M. Loynes. Testing distributions in large numbers of small samples. Communications in Statistics: Theory and Methods, 18 (5):1759–1768, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1981:GMC

[LP81] Larry Lee and R. Pirie. A graphical method for comparing trends in series of events. Communications in Statistics: Theory and Methods, 10(9):827–848, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levy:1984:MLP

[LP84] Martin S. Levy and S. K. Perng. A maximum likelihood prediction function for the linear model with consistency results. Communications in Statistics: Theory and Methods, 13(10):1257–1273, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lubecke:1985:NMP

[LP85] André M. Lubecke and W. J. Padgett. Nonparametric maximum penalized likelihood estimation of a density from arbitrarily rightcensored observations. Communications in Statistics: Theory and Methods, 14(2):257–271, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1987:CSC

[LP87a] Carl M. S. Lee and Shyamal Das Peddada. Comparison of some common ratio estimators using Pitman measure of nearness. Communications in Statistics: Theory and Methods, 16(7):2017–2027, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1987:TSE

[LP87b] Sik-Yum Lee and Wai-Yin Poon. Two-step estimation of multivariate polychoric correlation. Communications in Statistics: Theory and Methods, 16(2):307–320, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lio:1987:GQE

[LP87c] Y. L. Lio and W. J. Padgett. A generalized quantile estimator under censoring. Communications in Statistics: Theory and Methods, 16(11):3301–3321, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lio:1987:MSE

[LP87d] Y. L. Lio and W. J. Padgett. On the mean squared error of nonparametric quantile estimators under random right-censorship. Communications in Statistics: Theory and Methods, 16(6):1617– 1628, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lahiri:1988:NBE

[LP88] Parthasarathi Lahiri and Dong Ho Park. Nonparametric Bayes and empirical Bayes estimation of the residual survival function at age t. Communications in Statistics: Theory and Methods, 17(12): 4085–4098, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liski:1989:FNT

[LP89] Erkki P. Liski and Simo Puntan. A further note on a theorem on the difference of the generalized inverses of two nonnegative definite matrices. Communications in Statistics: Theory and Methods, 18 (5):1747–1751, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lachenbruch:1982:AMP

[LPW82] Peter A. Lachenbruch, Mari Palta, and Robert F. Woolson. Analysis of matched pairs studies with censored data. Communications in Statistics: Theory and Methods, 11(5):549–565, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Littell:1982:MDE

[LR82] Ramon C. Littell and P. V. Rao. On minimum distance estimation of location based on the Kolmogorov statistic. Communications in Statistics: Theory and Methods, 11(16):1793–1807, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lakatos:1987:URE

[LR87] E. Lakatos and D. Raghavarao. Undiminished residual effects designs and their suggested applications. Communications in Statistics: Theory and Methods, 16(5):1345–1359, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lackritz:1980:PSA

[LS80] James R. Lackritz and Richard L. Scheaffer. Point sampling in areal-fraction analysis. Communications in Statistics: Theory and Methods, 9(4):355–374, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levin:1981:EBM

[LS81] Bruce Levin and Patrick E. Shrout. On extending Bock’s model of logistic regression in the analysis of categorical data. Communications in Statistics: Theory and Methods, 10(2):125–147, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lackritz:1983:APE

[LS83a] James R. Lackritz and Richard L. Scheaffer. Asymptotic properties of estimators derived from line transect sampling. Communications in Statistics: Theory and Methods, 12(8):849–864, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1983:ACT

[LS83b] C. Y. Leung and M. S. Srivastava. Asymptotic comparison of two discriminants used in normal covariate classification. Communications in Statistics: Theory and Methods, 12(14):1637–1646, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1983:CCT

[LS83c] C. Y. Leung and M. S. Srivastava. Covariate classification for two correlated populations. Communications in Statistics: Theory and Methods, 12(2):223–241, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Landenna:1989:ADG

[LSB89] G. Landenna, A. Scagni, and M. Boldrini. An approximated distribution of the Gini’s rank association coefficient. Communications in Statistics: Theory and Methods, 18(6):2017–2026, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lambert:1985:LSE

[LSS85a] Annick Lambert, A. K. Md. Ehsanes Saleh, and Pranab Kumar Sen. On least squares estimation of intercepts after a preliminary test on parallelism of regression lines. Communications in Statistics: Theory and Methods, 14(4):793–807, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lambert:1985:THI

[LSS85b] Annick Lambert, A. K. Md. Ehsanes Saleh, and Pranab Kumar Sen. Tests of homogeneity of intercepts after a preliminary test. Communications in Statistics: Theory and Methods, 14(9):2243– 2259, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1980:PED

[LT80] Lon-Mu Liu and George C. Tiao. Parameter estimation in dynamic models. Communications in Statistics: Theory and Methods, 9(5): 501–517, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1982:EMP

[LT82] B. M. S. Lee and P. K. Trivedi. Error misspecification and properties of the simple ridge estimator. Communications in Statistics: Theory and Methods, 11(14):1615–1624, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1984:DPG

[LT84] Jack W. Lee and W. Y. Tan. On the degree of polynomial in a general linear model. Communications in Statistics: Theory and Methods, 13(6):781–790, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lefante:1985:ALF

[LT85] John J. Lefante, Jr. and Malcolm E. Turner, Jr. The average likelihood and a fiducial approximation: one parameter members of the generalized gamma distributions. Communications in Statistics: Theory and Methods, 14(2):419–436, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lucas:1985:CSC

[Luc85] James M. Lucas. Cumulative sum (cusum) control schemes. Communications in Statistics: Theory and Methods, 14(11):2689–2704, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lund:1982:PBF

[Lun82] Richard E. Lund. Plans for blocking and fractions of nested cube designs. Communications in Statistics: Theory and Methods, 11 (20):2287–2296, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lachenbruch:1985:SSP

[LW85] Peter A. Lachenbruch and Robert F. Woolson. On small sample properties of the generalized signed rank and generalized sign tests. Communications in Statistics: Theory and Methods, 14(9):2109– 2127, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1980:AJD

[LWA80] Pi-Erh Lin, Ke-Tsai Wu, and Ibrahim A. Ahmad. Asymptotic joint distribution of sample quantiles and sample mean with applications. Communications in Statistics: Theory and Methods, 9(1):51–60, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lesaffre:1989:EER

[LWA89] E. Lesaffre, J. L. Willems, and A. Albert. Estimation of error rate in multiple group logistic discrimination. The approximate leaving-one-out method. Communications in Statistics: Theory and Methods, 18(8):2989–3007, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1989:LSC

[LY89] Seung-Hoon Lee and Bong-Jin Yum. Large-sample comparisons of calibration procedures when both measurements are subject to error: the unreplicated case. Communications in Statistics: Theory and Methods, 18(10):3821–3840, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyberg:1985:QCP

[Lyb85] Lars Lyberg. Quality control procedures at Statistics Sweden. Communications in Statistics: Theory and Methods, 14(11):2705– 2751, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mitra:1980:MER

[MA80] Amitava Mitra and Khursheed Alam. Measurement error in regression analysis. Communications in Statistics: Theory and Methods, 9(7):717–723, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeks:1983:MCO

[MA83] Stephen L. Meeks and Ralph B. D’ Agostino. A model for comparisons with ordered categorical data. Communications in Statistics: Theory and Methods, 12(8):895–906, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mickey:1984:RMC

[MA84] Max R. Mickey and Claude O. Archer. Reliability models for categorical data. Communications in Statistics: Theory and Methods, 13(15):1851–1869, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MacGillivray:1982:SPA

[Mac82] H. L. MacGillivray. Skewness properties of asymmetric forms of Tukey lambda distributions. Communications in Statistics: Theory and Methods, 11(20):2239–2248, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madan:1985:MCU

[Mad85] Dilip B. Madan. The measurement of capital utilisation rates. Communications in Statistics: Theory and Methods, 14(6):1301– 1314, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Magda:1980:CBR

[Mag80] Constantine Gregory Magda. Circular balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 9(18):1901–1918, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Magee:1987:ARC

[Mag87] Lonnie Magee. Asymptotic risk comparisions of restricted and unrestricted maximum likelihood estimators. Communications in Statistics: Theory and Methods, 16(2):545–558, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Magee:1989:BAC

[Mag89] Lonnie Magee. Bias approximations for covariance parameter estimators in the linear model with AR(1) errors. Communications in Statistics: Theory and Methods, 18(2):395–422, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Majumdar:1988:ORM

[Maj88] Dibyen Majumdar. Optimal repeated measurements designs for comparing test treatments with a control. Communications in Statistics: Theory and Methods, 17(11):3687–3703, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malik:1980:EFC

[Mal80] Henrick J. Malik. Exact formula for the cumulative distribution function of the quasi-range from the logistic distribution. Communications in Statistics: Theory and Methods, 9(14):1527–1534, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mann:1980:SSC

[Man80] J. E. Mann. A simple sequential criterion for testing a normal mean with unknown variance. Communications in Statistics: Theory and Methods, 9(15):1619–1624, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mansouri:1989:LRT

[Man89] Hossein Mansouri. Linear rank tests for homogeneity against ordered alternatives in ANOVA and ANOCOVA. Communications in Statistics: Theory and Methods, 18(12):4321–4334, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1980:TSP

[Mar80] Ruth Marcus. A tow-stage procedure for testing homogeneity of means against ordered alternatives in analysis of variance with unequal variances. Communications in Statistics: Theory and Methods, 9(9):949–963, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardia:1981:DSG

[Mar81a] K. V. Mardia. Directional statistics in geosciences. Communications in Statistics: Theory and Methods, 10(15):1523–1543, 1981.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marechal:1981:PMG

[Mar81b] Alain Maréchal. The practice of mining geostatistics in 1980. Communications in Statistics: Theory and Methods, 10(15):1545–1558, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1982:SRS

[Mar82] Ruth Marcus. Some results on simultaneous confidence intervals for monotone contrasts in one-way ANOVA model. Communications in Statistics: Theory and Methods, 11(6):615–622, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardia:1984:SDC

[Mar84a] K. V. Mardia. Spatial discrimination and classification maps. Communications in Statistics: Theory and Methods, 13(18):2181–2197, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Markowski:1984:MLS

[Mar84b] Edward P. Markowski. The mean of linear signed rank statistics under alternatives. Communications in Statistics: Theory and Methods, 13(15):1915–1918, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marlin:1984:GFT

[Mar84c] Paul G. Marlin. Goodness-of-fit tests for the Pareto and lognormal distributions based on multiply truncated samples. Communications in Statistics: Theory and Methods, 13(16):1965–1979, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1984:EML

[Mar84d] R. J. Martin. Exact maximum likelihood for incomplete data from a correlated Gaussian process. Communications in Statistics: Theory and Methods, 13(10):1275–1288, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marasinghe:1985:ATM

[Mar85] Mervyn G. Marasinghe. Asymptotic tests and Monte Carlo studies associated with the multiplicative interaction model. Communications in Statistics: Theory and Methods, 14(9):2219–2231, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Markku:1988:SME

[Mar88a] Rahiala Markku. A simple method for the estimation of rational distributed lag models. Communications in Statistics: Theory and Methods, 17(6):1903–1928, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Markowski:1988:CSL

[Mar88b] Edward P. Markowski. A characterization of some locally most powerful rank tests. Communications in Statistics: Theory and Methods, 17(3):887–891, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maritz:1989:EBE

[Mar89a] J. S. Maritz. Empirical Bayes estimation of the log odds ratio in 2 × 2 contingency tables. Communications in Statistics: Theory and Methods, 18(9):3215–3233, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1989:ILD

[Mar89b] R. J. Martin. Information loss due to incomplete data from a spatial Gaussian one-parameter first-order conditional process. Communications in Statistics: Theory and Methods, 18(12):4631–4645, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:1986:LRR

[Mas86] Robert L. Mason. Latent root regression: a biased regression methodology for use with collinear predictor variables. Communications in Statistics: Theory and Methods, 15(9):2651–2678, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1980:MTN

[Mat80] A. M. Mathai. Moments of the trace of a noncentral Wishart matrix. Communications in Statistics: Theory and Methods, 9(8): 795–801, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matloff:1981:AAC

[Mat81] Norman S. Matloff. An alternative to the analysis of covariance for comparing randomized treatments. Communications in Statistics: Theory and Methods, 10(20):2015–2024, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muse:1982:ACD

[MAT82] H. D. Muse, R. L. Anderson, and B. Thitakamol. Additional comparisons of designs to estimate variance components in a two-way classification model. Communications in Statistics: Theory and Methods, 11(13):1403–1425, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1983:LCI

[Mat83] A. M. Mathai. On linear combinations of independent exponential variables. Communications in Statistics: Theory and Methods, 12 (6):625–632, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathew:1985:ALE

[Mat85] Thomas Mathew. Admissible linear estimation in singular linear models with respect to a restricted parameter set. Communications in Statistics: Theory and Methods, 14(2):491–498, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McKay:1980:C

[MB80] M. D. McKay and R. J. Beckman. Comments. Communications in Statistics: Theory and Methods, 9(17):1859–1861, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1982:ECP

[MB82] Paul W. Mielke, Jr. and Kenneth J. Berry. An extended class of permutation techniques for matched pairs. Communications in Statistics: Theory and Methods, 11(11):1197–1207, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mauro:1984:CRB

[MB84a] Carl A. Mauro and Kevin C. Burns. A comparison of random balance and two-stage group screening designs: A case study. Communications in Statistics: Theory and Methods, 13(21):2625–2648, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moen:1984:TCL

[MB84b] David H. Moen and Lyle D. Broemeling. Testing for a change lit the regression matrix of a multivariate linear model. Communications in Statistics: Theory and Methods, 13(12):1521–1531, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moen:1985:UFM

[MB85] David H. Moen and Lyle D. Broemeling. The uncertainty of forecasting: models with structural change versus those without changing parameters. Communications in Statistics: Theory and Methods, 14(9):2029–2040, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1989:AFD

[MB89] Rahul Mukerjee and Mausumi Bose. Admissibility in factorial designs. Communications in Statistics: Theory and Methods, 18(3): 1153–1158, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malik:1988:RRM

[MBA88] J. Henrick Malik, N. Balakrishnan, and S. E. Ahmed. Recurrence relations mid identities for moments of order statistics, I: arbitrary continuous distribution. Communications in Statistics: Theory and Methods, 17(8):2623–2655, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittelhammer:1984:ARL

[MC84] Ron C. Mittelhammer and Roger K. Conway. On the admissibility of restricted least squares estimators. Communications in Statistics: Theory and Methods, 13(9):1135–1146, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittelhammer:1985:ECM

[MC85] Ron C. Mittelhammer and Roger K. Conway. Extending Chipman’s MARL estimator to cases of ignorance in one or more parameter dimensions. Communications in Statistics: Theory and Methods, 14(8):1879–1888, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mayer:1986:DSC

[MC86] Lawrence S. Mayer and Steven S. Carroll. Detecting serial correlation in the error structure of a cross-lagged panel model. Communications in Statistics: Theory and Methods, 15(2):345–366, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mayer:1987:THC

[MC87a] Lawrence S. Mayer and Steven S. Carroll. Testing for homogeneity in cross-lagged panel studies. Communications in Statistics: Theory and Methods, 16(9):2487–2516, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1987:EPF

[MC87b] S. P. Mukherjee and A. Chatterjee. Estimation of parameters in finite mixtures of exponential families from censored data. Communications in Statistics: Theory and Methods, 16(11):3133–3147, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1989:MOM

[MC89] S. P. Mukherjee and A. K. Chattopadhyay. Measurement of occupational mobility using semi-Markov models. Communications in Statistics: Theory and Methods, 18(5):1961–1978, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCabe:1980:URE

[McC80] George P. McCabe, Jr. Use of the 21% Rulf in experimental design. Communications in Statistics: Theory and Methods, 9(7):765–776, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCulloch:1982:CUW

[McC82] Charles E. McCulloch. Conditions under which F {N_{1}} = ∞ for Tong’s adaptive solution to ranking and selection problems. Communications in Statistics: Theory and Methods, 11(7):815–819, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCulloch:1985:RAS

[McC85] Charles E. McCulloch. Relationships among some chi-square goodness of fit statistics. Communications in Statistics: Theory and Methods, 14(3):593–603, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCabe:1986:THO

[McC86] B. P. M. McCabe. Testing for heteroscedasticity occurring at unknown points. Communications in Statistics: Theory and Methods, 15(5):1597–1613, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCulloch:1987:TEV

[McC87] Charles E. McCulloch. Tests for equality of variances with paired data. Communications in Statistics: Theory and Methods, 16(5): 1377–1391, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McDonald:1987:MSS

[McD87] James B. McDonald. Model selection: some generalized distributions. Communications in Statistics: Theory and Methods, 16(4): 1049–1074, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mount-Campbell:1981:NFF

[MCN81] Clark A. Mount-Campbell and John B. Neuhardt. On the number of 2^{n−p} fractional factorials of resolution III. Communications in Statistics: Theory and Methods, 10(20):2101–2111, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moise:1988:TCR

[MCR88] Alain Moise, Bernard Clement, and Marios Raissis. A test for crossing receiver operating characteristic (roc) curves. Communications in Statistics: Theory and Methods, 17(6):1985–2003, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1987:RER

[MD87] S. P. Mukherjee and U. Dasgupta. Ridge estimation of restricted parameters in production techniques. Communications in Statistics: Theory and Methods, 16(12):3439–3446, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittal:1989:EPT

[MD89] Mukul M. Mittal and Ram C. Dahiya. Estimating the parameters of a truncated Weibull distribution. Communications in Statistics: Theory and Methods, 18(6):2027–2042, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mir:1981:EQE

[MDH81] Masoom Ali Mir, Umbach Dale, and Khatab M. Hassanein. Estimation of quantiles of exponential and double exponential distributions based on two order statistics. Communications in Statistics: Theory and Methods, 10(19):1921–1932, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morgan:1986:ECA

[ME86] Timothy M. Morgan and Robert M. Elasoff. Effect of censoring on adjusting for covariates in comparison of survival times. Communications in Statistics: Theory and Methods, 15(6):1837–1854, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mee:1988:EPN

[Mee88] Robert W. Mee. Estimation of the percentage of a normal distribution lying outside a specified interval. Communications in Statistics: Theory and Methods, 17(5):1465–1479, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehrotra:1982:GFT

[Meh82] Kishan G. Mehrotra. On goodness of fit tests based on spacings for Type II censored samples. Communications in Statistics: Theory and Methods, 11(8):869–878, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1981:BCD

[Men81a] Ulrich Menzefricke. Bayesian clustering of data sets. Communications in Statistics: Theory and Methods, 10(1):65–77, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1981:DTA

[Men81b] Ulrich Menzefricke. A decision-theoretic approach to variable selection in discriminant analysis. Communications in Statistics: Theory and Methods, 10(7):669–686, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1988:PDP

[Men88] Ulrich Menzefricke. Predictive distributions in the presence of measurement error. Communications in Statistics: Theory and Methods, 17(1):229–243, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardia:1983:OTM

[MF83] K. V. Mardia and K. Foster. Omnibus tests of multinormality based on skewness and kurtosis. Communications in Statistics: Theory and Methods, 12(2):207–221, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeden:1987:CSB

[MG87] Glen Meeden and James E. Gentle. Computing some Bayes estimates for the mean of a normal distribution. Communications in Statistics: Theory and Methods, 16(8):2443–2457, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1985:OSD

[MH85] Rahul Mukerjee and S. Huda. D-optimal statistical designs with restricted string property. Communications in Statistics: Theory and Methods, 14(3):669–677, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1986:OSD

[MH86] Rahul Mukerjee and S. Huda. Optimal statistical designs with circular string property. Communications in Statistics: Theory and Methods, 15(5):1615–1626, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moussa-Hamouda:1988:IRP

[MH88] Effat Moussa-Hamouda. Inference in regression problems based on order statistics. Communications in Statistics: Theory and Methods, 17(7):2343–2367, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mjelke:1982:PTA

[MI82] P. W. Mjelke and H. K. Iyer. Permutation techniques for analyzing multi-response data from randomized block experiments. Communications in Statistics: Theory and Methods, 11(13):1427–1437, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Michael:1983:MIC

[Mic83] Haber Michael. Multidimensional intraclass contingency tables. Communications in Statistics: Theory and Methods, 12(7):765– 776, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mickey:1985:LSB

[Mic85] Ruth M. Mickey. Large sample bias of maximum likelihood estimates from a series of sparse i × j tables. Communications in Statistics: Theory and Methods, 14(3):755–759, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mickey:1987:MDC

[Mic87] Ruth M. Mickey. The magnitude of the difference of crude and adjusted log odds ratios. Communications in Statistics: Theory and Methods, 16(11):3403–3415, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1980:ARP

[Mie80] Paul W. Mielke, Jr. Another response to Professor Neyman. Communications in Statistics: Theory and Methods, 9(18):1985–1987, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mikhail:1985:PCP

[Mik85] N. N. Mikhail. Penrose’s criterion and Pillai’s statistic under nonnormal theory. Communications in Statistics: Theory and Methods, 14(12):3101–3106, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mill:1982:WTO

[Mil82a] Richard W. Mill. When there are outliers in the carriers: the univariate case. Communications in Statistics: Theory and Methods, 11(8):849–868, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Milliken:1982:CIA

[Mil82b] George A. Milliken. On a confidence interval about a parameter estimated by a ratio of normal random variables. Communications in Statistics: Theory and Methods, 11(17):1985–1995, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minder:1980:IML

[Min80] Ch. E. Minder. On inspecting multiparameter likelihood surfaces. Communications in Statistics: Theory and Methods, 9(18):1931– 1939, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittelhammer:1981:SEG

[Mit81] Ron C. Mittelhammer. On specification error in the general linear model and weak mean square error superiority of the mixed estimator. Communications in Statistics: Theory and Methods, 10 (2):167–176, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittnik:1988:DTA

[Mit88] Stefan Mittnik. Derivation of the theoretical autocovariance and autocorrelation function of autoregessive moving average processes. Communications in Statistics: Theory and Methods, 17(11):3825– 3831, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miyazaki:1987:CVT

[Miy87] Haruo Miyazaki. Critical values for testing partially ordered alternatives in the analysis of variance. Communications in Statistics: Theory and Methods, 16(5):1393–1431, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marasinghe:1982:EMI

[MJ82] Mervyn G. Marasinghe and Dallas E. Johnson. Estimation of σ^{2} in the multiplicative interaction model. Communications in Statistics: Theory and Methods, 11(3):315–324, 1982. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McKee:1982:EMP

[MK82] Bonnie McKee and A. M. Kshirsagar. Effect of missing plots in some response surface designs. Communications in Statistics: Theory and Methods, 11(14):1525–1549, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Margaret:1983:BFC

[ML83] Land Margaret and Broemeling Lyle. Bayesian forecasting with changing linear models. Communications in Statistics: Theory and Methods, 12(12):1421–1430, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malice:1988:SEV

[ML88] Marie-Pierre Malice and Claude Lefèvre. On some effects of variability in the Weiss epidemic model. Communications in Statistics:

Theory and Methods, 17(11):3723–3731, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murthy:1983:RRS

[MM83] M. S. Ramachandra Murthy and J. S. Murty. Restricted region simplex design for mixture experiments. Communications in Statistics: Theory and Methods, 12(22):2605–2615, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maronna:1986:RRT

[MM86] Ricardo Maronna and Stephan Morgenthaler. Robust regression through robust covariances. Communications in Statistics: Theory and Methods, 15(4):1347–1365, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Molinska:1988:RBC

[MM88] Anna Molińska and Krzysztof Moliński. Relationships between confidence intervals for the ratio of two variances of independent normal distributions. Communications in Statistics: Theory and Methods, 17(3):925–933, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Molinska:1989:IEP

[MM89a] Anna Molińska and Krzysztof Moliński. Interval estimation of parameters in an unbalanced mixed two-fold nested classification — the last-stage uniformity case. Communications in Statistics: Theory and Methods, 18(11):4121–4135, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murthy:1989:RRD

[MM89b] M. S. Ramachandra Murthy and J. S. Murty. Restricted region designs for multifactor mixture experiments. Communications in Statistics: Theory and Methods, 18(4):1279–1295, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miyashita:1983:SNN

[MN83a] H. Miyashita and P. Newbold. On the sensitivity to non-normality of a test for outliers in linear models. Communications in Statistics: Theory and Methods, 12(12):1413–1419, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nam:1983:EMI

[mN83b] Jun mo Nam. Efficient method for identification of cyclic trends in incidence. Communications in Statistics: Theory and Methods, 12 (9):1053–1068, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McDonald:1989:ABE

[MN89] James B. McDonald and Ray D. Nelson. Alternative beta estimation for the market model using partially adaptive techniques. Communications in Statistics: Theory and Methods, 18(11):4039– 4058, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monahan:1987:AMC

[Mon87] John F. Monahan. An alternative method for computing overflow probabilities. Communications in Statistics: Theory and Methods, 16(11):3355–3357, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moon:1988:SST

[Moo88] C. G. Moon. Simulataneous specification test in a binary logit. Communications in Statistics: Theory and Methods, 17(10):3361– 3387, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moothathu:1989:UEG

[Moo89] T. S. K. Moothathu. On unbiased estimation of GINI index and Yntema–Pietra index of lognormal distribution and their variances. Communications in Statistics: Theory and Methods, 18(2):661– 672, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morris:1987:TSF

[Mor87] Max D. Morris. Two-stage factor screening procedures using multiple grouping assignments. Communications in Statistics: Theory and Methods, 16(10):3051–3067, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moschopoulos:1983:GPD

[Mos83a] Panagis G. Moschopoulos. A general procedure for deriving distributions. Communications in Statistics: Theory and Methods, 12 (17):2005–2015, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moschopoulos:1983:NTN

[Mos83b] Panagis G. Moschopoulos. On a new transformation to normality. Communications in Statistics: Theory and Methods, 12(16):1873– 1878, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moser:1987:GVG

[Mos87] Barry Kurt Moser. Generalized F variates in the general linear model. Communications in Statistics: Theory and Methods, 16 (7):1867–1884, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moschopoulos:1989:THC

[Mos89] Panagis G. Moschopoulos. Tests of hypotheses on concomitant variables in linear models. Communications in Statistics: Theory and Methods, 18(5):1735–1746, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mann:1982:TBS

[MP82a] Barbara L. Mann and Walter R. Pirie. Tighter bounds and simplified estimation for moments of some rank statistics. Communications in Statistics: Theory and Methods, 11(10):1107–1117, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1982:FRT

[MP82b] A. M. Mathai and K. C. S. Pillai. Further results on the trace of a noncentral Wishart matrix. Communications in Statistics: Theory and Methods, 11(10):1077–1086, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McNichols:1982:MLE

[MP82c] D. T. McNichols and W. J. Padgett. Maximum likelihood estimation of unimodal and decreasing densities based on arbitrarily right-censored data. Communications in Statistics: Theory and Methods, 11(20):2259–2270, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minan:1982:PEB

[MP82d] A. K. Minan and N. N. Puri. Partially efficiency-balanced designs — II. Communications in Statistics: Theory and Methods, 11(24): 2817–2830, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menon:1985:PGF

[MP85] V. V. Menon and B. Prasad. The probability generating function of empty cell variable in a randomized occupancy problem. Communications in Statistics: Theory and Methods, 14(10):2287–2292, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehta:1986:HAF

[MP86] Cyrus R. Mehta and Nitin R. Patel. A hybrid algorithm for Fisher’s exact test in unordered r × c contingency tables. Communications in Statistics: Theory and Methods, 15(2):387–403, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mackinnon:1989:CGL

[MP89] Murray J. Mackinnon and Martin L. Puterman. Collinearity in generalized linear models. Communications in Statistics: Theory and Methods, 18(9):3463–3472, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menjoge:1983:IEW

[MR83] Shailendra Menjoge and Poduri S. R. S. Rao. Improved estimators with the weighted and compounded loss functions. Communications in Statistics: Theory and Methods, 12(2):131–139, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1986:FRB

[MR86] Debabrata Mukherjee and Makarand V. Ratnaparkhi. On the functional relationship between entropy and variance with related applications. Communications in Statistics: Theory and Methods, 15 (1):291–311, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1987:FRT

[MR87] S. P. Mukherjee and Dillip Roy. Failure rate transform of the Weibull variate and its properties. Communications in Statistics:

Theory and Methods, 16(1):281–291, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCormick:1988:ADM

[MR88] William P. McCormick and Jack H. Reeves. Approximating the distribution of the maximum of a dependent stationary sequence based on estimates from a generating function. Communications in Statistics: Theory and Methods, 17(11):3801–3823, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathew:1984:ALE

[MRS84] Thomas Mathew, C. R. Rao, and B. K. Sinha. Admissible linear estimation in singular linear models. Communications in Statistics: Theory and Methods, 13(24):3033–3045, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miescke:1980:CCL

[MS80a] Klaus-J. Miescke and Joachim Sehr. On a conjecture concerning the least favorable configuration of certain two-stage selection procedures. Communications in Statistics: Theory and Methods, 9 (15):1609–1617, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monlezun:1980:GEH

[MS80b] Charles J. Monlezun and F. M. Speed. The geometry of estimation and hypothesis testing in the constrained linear model — the full rank case. Communications in Statistics: Theory and Methods, 9 (2):213–230, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1981:ANN

[MS81a] Paul W. Mielke, Jr. and Pranab Kumar Sen. On asymptotic nonnormal null distributions for locally most powerful rank test statistics. Communications in Statistics: Theory and Methods, 10(11): 1079–1094, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monlezun:1981:RSS

[MS81b] Charles J. Monlezun and F. M. Speed. Recovering sums of squares between full and reduced linear models. Communications in Statistics: Theory and Methods, 10(9):875–882, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moore:1981:CST

[MS81c] David S. Moore and John B. Stubblebine. Chi-square tests for multivariate normality with application to common stock prices. Communications in Statistics: Theory and Methods, 10(8):713– 738, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mykytyn:1981:MLE

[MS81d] Stephen W. Mykytyn and Thomas J. Santner. Maximum likelihood estimation of the survival function based on censored data under hazard rate assumptions. Communications in Statistics: Theory and Methods, 10(14):1369–1387, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marchand:1982:ENE

[MS82] J. P. Marchand and F. E. Schroeck, Jr. On the estimation of the number of equally likely classes in a population. Communications in Statistics: Theory and Methods, 11(10):1139–1146, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murray:1985:ETC

[MS85] Leigh W. Murray and D. W. Smith. Estimability, testability, and connectedness in the cell means model. Communications in Statistics: Theory and Methods, 14(8):1889–1917, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meydrech:1986:NTM

[MS86a] Edward F. Meydrech and P. K. Sen. On a nonparametric test for the mixed-model. Communications in Statistics: Theory and Methods, 15(5):1587–1596, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miceli:1986:MEC

[MS86b] Robert J. Miceli and William E. Strawderman. Minimax estimation for certain independent component distributions under weighted squared error loss. Communications in Statistics: Theory and Methods, 15(7):2191–2200, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mitchell:1987:CPD

[MS87] Toby J. Mitchell and David S. Scott. A computer program for the design of group testing experiments. Communications in Statistics:

Theory and Methods, 16(10):2943–2955, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1988:FDS

[MS88] Govind S. Mudholkar and Perla Subbaiah. On a Fisherian detour of the step-down procedure for MANOVA. Communications in Statistics: Theory and Methods, 17(2):599–611, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moen:1985:SCM

[MSB85] David H. Moen, Diego Salazar, and Lyle D. Broemeling. Structural changes in multivariate regression models. Communications in Statistics: Theory and Methods, 14(8):1757–1768, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mishra:1986:OCG

[MSL86] Satya N. Mishra, Arvind K. Shah, and John J. Lefante. Overlapping coefficient: the generalized t approach. Communications in Statistics: Theory and Methods, 15(1):123–128, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1980:RCS

[MSVS80] A. J. Miller, D. E. Shaw, L. G. Veitch, and E. J. Smith. Rejoinder to “ ‘Comments on the special issue of Communications in Statistics (vol. 8, no. 10) concerned with weather modification experiments’ by Jerzy Neyman”. Communications in Statistics: Theory and Methods, 9(18):1975–1978, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Ney80b].

Moser:1989:TST

[MSW89] Barry K. Moser, Gary R. Stevens, and Christian L. Watts. The two-sample t test versus Satterthwaite’s approximate F test. Communications in Statistics: Theory and Methods, 18(11):3963–3975, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malik:1982:PDF

[MT82] Henrick J. Malik and Roger Trudel. Probability density function of quotient of order statistics from the Pareto, power and Weibull distributions. Communications in Statistics: Theory and Methods, 11(7):801–814, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1983:THN

[MT83a] Ruth Marcus and Hovav Talpaz. On testing homogeneity of t normal means against ordered alternatives in r groups. Communications in Statistics: Theory and Methods, 12(24):2897–2902, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1983:RCH

[MT83b] A. M. Mathai and D. S. Tracy. On a random convex hull in an nball. Communications in Statistics: Theory and Methods, 12(15): 1727–1736, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1984:FRC

[MT84] Timothy I. Miller and Ronald L. Tracy. Further results concerning ridge regression. Communications in Statistics: Theory and Methods, 13(2):251–262, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malik:1985:DPQ

[MT85] Henrick J. Malik and Roger Trudel. Distributions of the product and the quotient from bivariates t, f and Pareto distribution. Communications in Statistics: Theory and Methods, 14(12):2951–2962, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1981:CRS

[Muk81] N. Mukhopadhyay. Convergence rates of sequential confidence intervals and tests for the mean of a U-statistic. Communications in Statistics: Theory and Methods, 10(21):2231–2244, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1986:COF

[Muk86] Rahul Mukerjee. Construction of orthogonal factorial designs controlling interaction efficiencies. Communications in Statistics: Theory and Methods, 15(5):1535–1548, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1987:FOR

[Muk87a] Rahul Mukerjee. On fourth order rotatable designs. Communications in Statistics: Theory and Methods, 16(6):1697–1702, 1987.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1987:OMU

[Muk87b] Sunanda Bagchi Mukhopadhyay. On the optimalities of the MBGDD’s under mixed effects model. Communications in Statistics: Theory and Methods, 16(12):3565–3576, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1988:SEP

[Muk88] N. Mukhopadhyay. Sequential estimation problems for negative exponential populations. Communications in Statistics: Theory and Methods, 17(8):2471–2506, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murray:1989:NCF

[Mur89] Leigh W. Murray. A note on connectedness in fixed effects MANOVA and GMANOVA models with missing cells. Communications in Statistics: Theory and Methods, 18(7):2527–2531, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mustafi:1989:WCR

[Mus89] C. K. Mustafi. Weak convergence of residual median process for dependent or randomly censored samples. Communications in Statistics: Theory and Methods, 18(8):3087–3093, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morin-Wahhab:1985:MRT

[MW85] Danielle Morin-Wahhab. Moments of a ratio of two quadratic forms. Communications in Statistics: Theory and Methods, 14 (2):499–508, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:1985:RET

[MY85] Robert L. Mason and John C. Young. Re-examining two tests for bivariate normality. Communications in Statistics: Theory and Methods, 14(7):1531–1546, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marco:1987:NES

[MYT87] Virgil R. Marco, Dean M. Young, and Danny W. Turner. A note on the effect of simple equicorrelation in detecting a spurious multivariate observation. Communications in Statistics: Theory and Methods, 16(4):1027–1036, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:1988:DCP

[MZ88] B. Q. Miao and L. C. Zhao. Detection of change points using rank methods. Communications in Statistics: Theory and Methods, 17 (9):3207–3217, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naes:1986:DMO

[Nae86] Tormod Naes. Detection of multivariate outliers in linear mixed models. Communications in Statistics: Theory and Methods, 15(1): 33–47, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:1980:SDP

[Nag80] H. N. Nagaraja. Some distributional problems connected with order statistics. Communications in Statistics: Theory and Methods, 9(9):935–942, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagao:1985:LDJ

[Nag85] Hisao Nagao. On the limiting distributions of the jackknife statistics for eigenvalues of a sample covariance matrix. Communications in Statistics: Theory and Methods, 14(7):1547–1567, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:1988:RVR

[Nag88] H. N. Nagaraja. Record values and related statistics — a review. Communications in Statistics: Theory and Methods, 17(7):2223– 2238, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:1982:SEA

[Nai82] V. N. Nair. Some extensions of ANOVA techniques to locationscale models. Communications in Statistics: Theory and Methods, 11(14):1551–1570, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naik:1989:DOM

[Nai89] Dayanand N. Naik. Detection of outliers in the multivariate linear regression model. Communications in Statistics: Theory and Methods, 18(6):2225–2232, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandram:1989:DBC

[Nan89] Balgobin Nandram. Discrimination between the complementary log-log and logistic model for ordinal data. Communications in Statistics: Theory and Methods, 18(6):2155–2164, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nardia:1981:SHE

[Nar81] K. V. Nardia. Are statisticians helping earth scientists enough? Communications in Statistics: Theory and Methods, 10(15):v, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Narodia:1983:COO

[Nar83] C. K. Narodia. Characterizations for the original observations with linearity of regression in additive damage models. Communications in Statistics: Theory and Methods, 12(3):303–310, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Natrella:1985:SQC

[Nat85] Mary C. Natrella. Statistical and quality control standards in the U.S. Communications in Statistics: Theory and Methods, 14(11): 2753–2764, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:1985:DMB

[Nay85] Iapan Kumar Nayak. On diversity measures based on entropy functions. Communications in Statistics: Theory and Methods, 14 (1):203–215, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:1988:TEC

[Nay88] Tanan Kumar Nayak. Testing equality of conditionally independent exponential distributions. Communications in Statistics: Theory and Methods, 17(3):807–820, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagarsenker:1982:DLT

[NCLN82] B. N. Nagarsenker, J. A. Mc Culloch, C. A. Loch, and P. B. Nagarsenker. On the distribution of LRT for testing the equality of exponential distributions. Communications in Statistics: Theory and Methods, 11(10):1119–1127, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nath:1981:GRT

[ND81] R. Nath and B. S. Duran. A generalization of the rank transform in the two-sample location problem. Communications in Statistics: Theory and Methods, 10(14):1437–1455, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nel:1984:PAD

[NDM84] D. G. Nel, D. J. De Waal, and D. G. Marx. A predictive approach to the detection of additional information in a multivariate regression model. Communications in Statistics: Theory and Methods, 13 (17):2119–2132, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelson:1982:ECP

[Nel82] Peter R. Nelson. Exact critical points for the analysis of means. Communications in Statistics: Theory and Methods, 11(6):699– 709, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelsen:1986:POP

[Nel86] Roger B. Nelsen. Properties of a one-parameter family of bivariate distributions with specified marginals. Communications in Statistics: Theory and Methods, 15(11):3277–3285, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nel:1987:PAT

[Nel87] D. G. Nel. A profile analysis for two groups under the heteroscedastic assumption. Communications in Statistics: Theory and Methods, 16(12):3577–3589, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neter:1985:ASA

[Net85] John Neter. The American Statistical Association’s response to the challenges of improving quality. Communications in Statistics: Theory and Methods, 14(11):2765, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neuhaus:1988:AOR

[Neu88] Georg Neuhaus. Asymptotically optimal rank tests for the twosample problem with randomly censored data. Communications in Statistics: Theory and Methods, 17(6):2037–2058, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neudecker:1989:GRT

[Neu89] H. Neudecker. A generalization of a result of Trenkler and Trenkler. Communications in Statistics: Theory and Methods, 18(11):4303– 4304, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neyman:1980:LE

[Ney80a] J. Neyman. Letter to the editor. Communications in Statistics: Theory and Methods, 9(18):1989–1996, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neyman:1980:CSI

[Ney80b] Jerzy Neyman. Comments on the special issue of Communications in Statistics (vol. 8, no. 10) concerned with weather modification experiments. Communications in Statistics: Theory and Methods, 9(9):965–992, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See rejoinders [MSVS80, Smi80].

Neyman:1980:TRR

[Ney80c] Jerzy Neyman. Two radiation related public health studies. Communications in Statistics: Theory and Methods, 9(1):1–11, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1989:NCM

[Ng89a] Tie-Hua Ng. A new class of modified binomial distributions with applications to certain toxicological experiments. Communications in Statistics: Theory and Methods, 18(9):3477–3492, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1989:PDC

[Ng89b] Vee Ming Ng. Prediction distributions for changing sequences. Communications in Statistics: Theory and Methods, 18(2):507– 513, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nigm:1987:BPB

[NH87] A. M. Nigm and H. I. Hamdy. Bayesian prediction bounds for the Pareto lifetime model. Communications in Statistics: Theory and Methods, 16(6):1761–1772, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nichols:1989:ESM

[Nic89] D. T. Mc Nichols. The exact solution to the MLE of a decreasing density based on arbitrarily right censored data. Communications in Statistics: Theory and Methods, 18(6):2043–2062, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nigm:1988:PBB

[Nig88] A. M. Nigm. Prediction bounds for the Burr model. Communications in Statistics: Theory and Methods, 17(1):287–297, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nigm:1989:IBP

[Nig89] A. M. Nigm. An informative Bayesian prediction for the Weibull lifetime distribution. Communications in Statistics: Theory and Methods, 18(3):897–911, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishii:1988:IIE

[Nis88] Ryuei Nishii. Identities and inequalities for the expected values of random matrices and Loewner’s theory. Communications in Statistics: Theory and Methods, 17(6):2021–2035, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishii:1989:CPD

[Nis89] Ryuei Nishii. A characterization of probability densities with expected log likelihood. Communications in Statistics: Theory and Methods, 18(7):2657–2662, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neill:1984:TLF

[NJ84] James W. Neill and Dallas E. Johnson. Testing for lack of fit in regression — a review. Communications in Statistics: Theory and Methods, 13(4):485–511, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neill:1989:CSL

[NJ89] James W. Neill and Dallas E. Johnson. A comparison of some lack of fit tests based on near replicates. Communications in Statistics: Theory and Methods, 18(10):3533–3570, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noser:1988:BOT

[NM88] Barry K. Noser and Virgil R. Marco. Bayesian outlier testing using the predictive distribution for a linear model of constant intraclass form. Communications in Statistics: Theory and Methods, 17(3): 849–860, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noor:1989:SSR

[NM89] Islam Noor and J. S. Mehta. A study of some ridge-type shrinkage estimators. Communications in Statistics: Theory and Methods, 18 (12):4437–4457, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagarsenker:1982:DLA

[NN82a] B. Nagarsenker and P. B. Nagarsenker. Distributions of LRT associated with two-parameter exponential distributions. Communications in Statistics: Theory and Methods, 11(14):1583–1593, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagarsenker:1982:DIT

[NN82b] B. Nagarsenker and P. B. Nagarsenker. On the distribution of independence test for the residuals of space-time ARMA models. Communications in Statistics: Theory and Methods, 11(6):653– 661, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagarsenker:1984:AEN

[NNQ84] B. N. Nagarsenker, P. B. Nagarsenker, and D. W. Quinn. Asymptotic expansions of the non-central distribution of Wilks’ statistic in the complex Gaussian case. Communications in Statistics: Theory and Methods, 13(12):1493–1505, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nomura:1985:NCR

[NO85] Masuo Nomura and Tsuneharu Ohkubo. A note on combining ridge and principal component regression. Communications in Statistics: Theory and Methods, 14(10):2489–2493, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nordberg:1982:VSG

[Nor82] Lennart Nordberg. On variable selection in generalized linear and related regression models. Communications in Statistics: Theory and Methods, 11(21):2427–2449, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nordstrom:1989:SFA

[Nor89] Kenneth Nordström. Some further aspects of the Löwner-ordering antitonicity of the Moore–Penrose inverse. Communications in Statistics: Theory and Methods, 18(12):4471–4489, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Njui:1988:FOR

[NP88] F. Njui and M. S. Patel. Fifth order rotability. Communications in Statistics: Theory and Methods, 17(3):833–848, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1983:RPM

[NR83] Tie-Hua Ng and Ronald H. Randles. Rank procedures in many population forced discrimination problems. Communications in Statistics: Theory and Methods, 12(17):1943–1959, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nour:1983:EIV

[NS83] El-Sayed Nour Nour and C. M. Suchindran. The effect of individual variability on the life table. Communications in Statistics: Theory and Methods, 12(3):323–339, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nomakuchi:1984:UIP

[NS84] Kentaro Nomakuchi and Toshio Sakata. The union-intersection principle applied to the test of dimensionality. Communications in Statistics: Theory and Methods, 13(6):753–760, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nakanishi:1985:PLQ

[NS85] Hiroko Nakanishi and Yoshiharu Sato. The performance of the linear and quadratic discriminant functions for three types of nonnormal distribution. Communications in Statistics: Theory and Methods, 14(5):1181–1200, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nirmalan:1986:PMS

[NS86] T. Nirmalan and N. Singh. On the problem of model selection in bilinear time series analyses. Communications in Statistics: Theory and Methods, 15(11):3445–3457, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nakano:1987:EPM

[NT87] Junji Nakano and Sigani Tagami. Estimation of parameters of a multivariate moving average model from estimates of the inverse autocovariance function. Communications in Statistics: Theory and Methods, 16(1):181–192, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagelkerke:1984:HTL

[NV84] N. J. D. Nagelkerke and E. A. Van der Velde. Hypothesis tests for linear functional relationships. Communications in Statistics: Theory and Methods, 13(22):2807–2816, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nel:1986:SMB

[NvdM86] D. G. Nel and C. A. van der Merwe. A solution to the multivariate Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 15(12):3719–3735, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nyquist:1983:ONE

[Nyq83] H. Nyquist. The optimal l_{p} norm estimator in linear regression models. Communications in Statistics: Theory and Methods, 12 (21):2511–2524, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nyquist:1986:EHE

[Nyq86] H. Nyquist. An evaluation of the Huber estimator. Communications in Statistics: Theory and Methods, 15(1):205–214, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OBrien:1989:TBT

[O.’89] Ralph G. O.’Brien. On the Tiku–Balakrishnan tests for equality of covariance matrices. Communications in Statistics: Theory and Methods, 18(2):423–428, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Olejnik:1987:ASP

[OA87] Stephen F. Olejnik and James Algina. An analysis of statistical power for parametric ANCOVA and rank transform ANCOVA. Communications in Statistics: Theory and Methods, 16(7):1923– 1949, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Olejnik:1989:CAS

[OAH89] Stephen F. Olejnik, James Algina, and Kelly Hern. Contrast analysis for scale differences. Communications in Statistics: Theory and Methods, 18(8):2965–2987, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Obenchain:1984:MLR

[Obe84] Robert L. Obenchain. Maximum likelihood ridge displays. Communications in Statistics: Theory and Methods, 13(2):227–240, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Obremski:1982:ASR

[Obr82] Thomas E. Obremski. Attaining specified reliability in component design: a stochastic approximation approach. Communications in Statistics: Theory and Methods, 11(21):2475–2483, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OCinneide:1987:OTS

[O’C87a] Colm A. O’Cinneide. On optimality of a two-stage sampling method for binomial populations. Communications in Statistics: Theory and Methods, 16(5):1461–1471, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OCinneide:1987:OST

[O’C87b] Colm A. O’Cinneide. Optimality of some two-stage sampling methods. Communications in Statistics: Theory and Methods, 16(1): 253–261, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odhiambo:1986:PMG

[Odh86] J. W. Odhiambo. The performance of multistage group screening designs. Communications in Statistics: Theory and Methods, 15 (8):2467–2481, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odhiambo:1987:PTS

[Odh87] John W. Odhiambo. On the performance of three-stage group screening experiments. Communications in Statistics: Theory and Methods, 16(9):2727–2745, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1985:PDJ

[OEJ^{+}85] Donald B. Owen, Ralph A. Evans, Richard A. Johnson, Jerald F. Lawless, and Peter R. Nelson. Panel discussion by journal editors. Communications in Statistics: Theory and Methods, 14(11):2815– 2844, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Officer:1986:SUE

[Off86] B. Prasad Research Officer. Some unbiased estimators versus mean per unit and ratio estimators in finite population sample surveys. Communications in Statistics: Theory and Methods, 15(12):3647– 3657, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1984:SSP

[OH84] Kazuhiro Ohtani and Yuzo Honda. On small sample properties of the mixed regression predictor under misspecification. Communications in Statistics: Theory and Methods, 13(22):2817–2825, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1986:SSP

[Oht86] Kazuhiro Ohtani. On small sample properties of the almost unbiased generalized ridge estimator. Communications in Statistics: Theory and Methods, 15(5):1571–1578, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1987:ECM

[Oht87a] Kazuhiro Ohtani. On estimating the common mean in two normal distributions after a preliminary test for equality of variances.

Communications in Statistics: Theory and Methods, 16(7):1977– 1993, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1987:SSP

[Oht87b] Kazuhiro Ohtani. Some sampling properties of the two-stage test in a linear regression with a proxy variable. Communications in Statistics: Theory and Methods, 16(3):717–729, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1988:TDV

[Oht88] Kazuhiro Ohtani. Testing the disturbance variance after a pretest for a linear hypothesis on coefficients in a linear regression. Communications in Statistics: Theory and Methods, 17(12):4231– 4250, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ojo:1988:NAI

[Ojo88] M. O. Ojo. A new approximation to the incomplete beta function. Communications in Statistics: Theory and Methods, 17(5):1423– 1435, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oksanen:1988:PCE

[Oks88] E. H. Oksanen. Principal components in econometrics. Communications in Statistics: Theory and Methods, 17(8):2507–2532, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:1986:GNC

[OL86] S. H. Ong and P. A. Lee. On a generalized non-central negative binomial distribution. Communications in Statistics: Theory and Methods, 15(3):1065–1079, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oldford:1987:DGR

[Old87] Andrey Oldford. On the n-dimensional geometry of regression diagnostics. Communications in Statistics: Theory and Methods, 16 (9):2517–2540, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OReilly:1980:ANM

[OM80] Federico J. O’Reilly and Paul W. Mielke, Jr. Asymptotic normality of MRPP statistics from invariance principles of U-statistics. Communications in Statistics: Theory and Methods, 9(6):629–637, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odhiambo:1987:SWG

[OM87] J. W. Odhiambo and M. M. Manene. Step-wise group screening designs with errors in observations. Communications in Statistics: Theory and Methods, 16(10):3095–3115, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ottieno:1984:TSG

[OP84] J. A. M. Ottieno and M. S. Patel. Two stage group-screening designs with unequal a-priori probabilities. Communications in Statistics: Theory and Methods, 13(6):761–779, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odhiambo:1986:MGS

[OP86] J. W. Odhiambo and M. S. Patel. On multistage group screening designs. Communications in Statistics: Theory and Methods, 15 (5):1627–1645, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odhiambo:1987:TSG

[OP87] J. W. Odhiambo and M. S. Patel. Three-stage group screening with unequal group sizes and with errors in observations. Communications in Statistics: Theory and Methods, 16(10):2957–2979, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oza:1986:PUU

[OS86] Markand Oza and G. K. Shukla. Prediction under uncertainty of degree of polynomial in growth curve models. Communications in Statistics: Theory and Methods, 15(5):1653–1664, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1988:DBP

[OS88] D. B. Owen and Jyh-Jen Horng Shiau. On the duality between points and lines. Communications in Statistics: Theory and Methods, 17(1):207–228, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ord:1981:BTM

[OSG81] J. K. Ord, Erwin M. Saniga, and William V. Gehrlein. Bayesian tests for matched proportions. Communications in Statistics: Theory and Methods, 10(18):1885–1899, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oppenlander:1988:SSR

[OSH88] J. E. Oppenlander, J. Scmee, and C. J. Hahn. Some simple robust estimators of normal distribution tail percentiles and their properties. Communications in Statistics: Theory and Methods, 17(7): 2279–2301, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trenkler:1985:MSE

[öT85] ötz Trenkler. Mean square error matrix comparisons of estimators in linear regression. Communications in Statistics: Theory and Methods, 14(10):2495–2509, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Orban:1980:DFP

[OW80] John Orban and Douglas A. Wolfe. Distribution-free partially sequential placement procedures. Communications in Statistics: Theory and Methods, 9(9):883–904, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ord:1986:PIG

[OW86] J. Keith Ord and G. Alex Whitmore. The Poisson-inverse Gaussian distribution as a model for species abundance. Communications in Statistics: Theory and Methods, 15(3):853–871, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1981:Ea

[Owe81a] D. B. Owen. Editorial. Communications in Statistics: Theory and Methods, 10(1):iii, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1981:Ec

[Owe81b] D. B. Owen. Editorial. Communications in Statistics: Theory and Methods, 10(3):iii–iv, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1981:EB

[Owe81c] D. B. Owen. Editorial Board. Communications in Statistics: Theory and Methods, 10(2):iii, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1985:E

[Owe85a] D. B. Owen. Editorial. Communications in Statistics: Theory and Methods, 14(1):iii–v, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Owen:1985:ACC

[Owe85b] Roger J. Owen. Accuracy, confidence and consensus in Bayesian hypothesis inference. Communications in Statistics: Theory and Methods, 14(5):1035–1052, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padmanabhan:1980:HLE

[Pad80] A. R. Padmanabhan. Hodges–Lehmann estimators in the case or grouped data — III. Communications in Statistics: Theory and Methods, 9(4):415–425, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pagel:1981:CHK

[Pag81] Mark D. Pagel. Comment on Hoerl and Kennard’s ridge regression simulation methodology. Communications in Statistics: Theory and Methods, 10(22):2361–2367, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pal:1988:DTE

[Pal88] Nabendu Pal. Decision-theoretic estimation of generalized variance and generalized precision. Communications in Statistics: Theory and Methods, 17(12):4221–4230, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panaretos:1983:GMI

[Pan83] John Panaretos. A generating model involving Pascal and logarithmic series distributions. Communications in Statistics: Theory and Methods, 12(7):841–848, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papageorgiou:1986:BSD

[Pap86] H. Papageorgiou. Bivariate short distributions. Communications in Statistics: Theory and Methods, 15(3):893–905, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papadopoulos:1989:HAS

[Pap89] Alex G. Papadopoulos. A hierarchical approach to the study of the exponential failure model. Communications in Statistics: Theory and Methods, 18(12):4375–4392, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1981:DNU

[Par81] C. J. Park. On the distribution of the number of unobserved elements when m-samples of size n are drawn from a finite population. Communications in Statistics: Theory and Methods, 10 (4):371–383, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parr:1982:NAS

[Par82] William C. Parr. A note on adaptive L-statistics. Communications in Statistics: Theory and Methods, 11(13):1511–1518, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patefield:1981:CIS

[Pat81] W. M. Patefield. Confidence intervals for the slope of a linear functional relationship. Communications in Statistics: Theory and Methods, 10(17):1759–1764, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1983:UBM

[Pat83] H. I. Patel. Use of baseline measurements in the two-period crossover design. Communications in Statistics: Theory and Methods, 12(23):2693–2712, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1985:AEC

[Pat85] Y. C. Patel. An asymptotic expression for cumulative sum of probabilities of the Hermite distribution. Communications in Statistics: Theory and Methods, 14(9):2233–2241, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1986:TLR

[Pat86] J. K. Patel. Tolerance limits — a review. Communications in Statistics: Theory and Methods, 15(9):2719–2762, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patwardhan:1988:TE

[Pat88] Girish Patwardhan. Tests for exponentiality. Communications in Statistics: Theory and Methods, 17(11):3705–3722, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1989:PIR

[Pat89] J. K. Patel. Prediction intervals — a review. Communications in Statistics: Theory and Methods, 18(7):2393–2465, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1985:TPG

[Pau85] S. R. Paul. A three-parameter generalization of the binomial distribution. Communications in Statistics: Theory and Methods, 14 (6):1497–1506, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1988:ETS

[Pau88] S. R. Paul. Estimation of and testing significance for a common correlation coefficient. Communications in Statistics: Theory and Methods, 17(1):39–53, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1983:Ea

[PB83a] Jagdish K. Patel and Lee J. Bain. Editorial. Communications in Statistics: Theory and Methods, 12(5):vii, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1983:Eb

[PB83b] Jagdish K. Patel and Lee J. Bain. Editorial. Communications in Statistics: Theory and Methods, 12(19):vii, 1983. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1983:EBa

[PB83c] Jagdish K. Patel and Lee J. Bain. Editorial Board. Communications in Statistics: Theory and Methods, 12(19):ebi, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1983:SR

[PB83d] Jagdish K. Patel and Lee J. Bain. Statistical reviews. Communications in Statistics: Theory and Methods, 12(5):v, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1984:Ea

[PB84a] Jagdish K. Patel and Lee J. Bain. Editorial. Communications in Statistics: Theory and Methods, 13(4):vii, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1984:Eb

[PB84b] Jagdish K. Patel and Lee J. Bain. Editorial. Communications in Statistics: Theory and Methods, 13(13):7, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1984:SR

[PB84c] Jagdish K. Patel and Lee J. Bain. Statistical reviews. Communications in Statistics: Theory and Methods, 13(4):v, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pendergast:1985:REG

[PB85] Jane F. Pendergast and James D. Broffitt. Robust estimation in growth curve models. Communications in Statistics: Theory and Methods, 14(8):1919–1939, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1986:SR

[PB86] Jagdish K. Patel and Lee J. Bain. Statistical reviews. Communications in Statistics: Theory and Methods, 15(9):5–6, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1987:DOP

[PB87] S. R. Paul and R. K. Barnwal. Detection of outliers in Poisson samples. Communications in Statistics: Theory and Methods, 16 (8):2391–2403, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Poole:1982:GCM

[PC82] W. Kenneth Poole and C. Andrew Clayton. Generalizations of a contamination model for continuous type random variables. Communications in Statistics: Theory and Methods, 11(15):1733–1742, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Passi:1987:OOD

[PCP87] Ranjit M. Passi, Michael J. Carpenter, and Harsh A. Passi. Operational outlier detection. Communications in Statistics: Theory and Methods, 16(11):3379–3391, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parr:1981:MCM

[PD81] W. C. Parr and T. De Wet. On minimum Cramér–von Mises-norm parameter estimation. Communications in Statistics: Theory and Methods, 10(12):1149–1166, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Petruccelli:1984:SRU

[PD84] Joseph D. Petruccelli and Neville Davies. Some restrictions on the use of corner method hypothesis tests. Communications in Statistics: Theory and Methods, 13(5):543–551, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pederzoli:1983:RGV

[Ped83] Giorgio Pederzoli. On the ratio of generalized variances. Communications in Statistics: Theory and Methods, 12(24):2903–2909, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pederzoli:1985:BTD

[Ped85a] Giorgio Pederzoli. On beta Type-1 distributed random points in an n-ball. Communications in Statistics: Theory and Methods, 14 (1):195–202, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pederzoli:1985:UDR

[Ped85b] Giorgio Pederzoli. On uniformly distributed random points in an n-ball. Communications in Statistics: Theory and Methods, 14 (1):185–194, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pederzoli:1986:RBS

[Ped86] Giorgio Pederzoli. A representation in beta series for the exact density of some random volume contents. Communications in Statistics: Theory and Methods, 15(11):3299–3306, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peiris:1986:PTD

[Pei86a] M. Shelton Peiris. On prediction with time dependent ARMA models. Communications in Statistics: Theory and Methods, 15 (12):3659–3668, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peixoto:1986:THS

[Pei86b] Julio L. Peixoto. Testable hypotheses in singular fixed linear models. Communications in Statistics: Theory and Methods, 15(6): 1957–1973, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peiris:1988:PMA

[Pei88] M. Shelton Peiris. On the prediction of multivariate ARMA processes with a time dependent covariance structure. Communications in Statistics: Theory and Methods, 17(1):27–37, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pendleton:1985:IOAa

[Pen85a] O. J. Pendleton. Influential observations in the analysis of variance. Communications in Statistics: Theory and Methods, 14(3):551– 565, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pendleton:1985:IOAb

[Pen85b] O. J. Pendleton. Influential observations in the analysis of variance. Communications in Statistics: Theory and Methods, 14(3):567– 583, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pepple:1988:BTE

[Pep88] Patricia A. Pepple. Bayesian testing of an exponential point null hypothesis. Communications in Statistics: Theory and Methods, 17 (10):3483–3503, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pepple:1989:SEG

[Pep89] Patricia A. Pepple. Simultaneous estimation of gamma means using a hierarchical generalized linear model. Communications in Statistics: Theory and Methods, 18(3):835–852, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peres:1981:TEB

[Per81] Clovis A. Peres. Testing the effect of blocking in a randomized complete block design (RCBD). Communications in Statistics: Theory and Methods, 10(23):2447–2459, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peritz:1982:ETM

[Per82] Eric Peritz. Exact tests for matched pairs: studies with covariates. Communications in Statistics: Theory and Methods, 11(19):2157– 2167, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peritz:1985:MMH

[Per85] Eric Peritz. Modified Mantel–Haenszel procedures for matched pairs. Communications in Statistics: Theory and Methods, 14(10): 2263–2285, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perng:1987:NIE

[Per87] S. K. Perng. A note on the inverse estimator for the linear calibration problem. Communications in Statistics: Theory and Methods, 16(6):1743–1747, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peter:1982:NAG

[Pet82] D. Hill Peter. Numerical Algorithms Group Manual Mark 7. Communications in Statistics: Theory and Methods, 11(20):2343–2356, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pillai:1984:PPL

[PF84] K. C. S. Pillai and Bernhard N. Flury. Percentage points of the largest characteristic root of the multivariate beta matrix. Communications in Statistics: Theory and Methods, 13(18):2199–2237, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perng:1982:AOT

[PG82] S. K. Perng and D. S. Gill. An asymptotically optimal test for the mean and the variance of a normal distribution. Communications in Statistics: Theory and Methods, 11(16):1817–1829, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perng:1983:MOS

[PG83] S. K. Perng and Shu Geng. Modified d-optimal sequential procedure in allocation problem. Communications in Statistics: Theory and Methods, 12(22):2633–2644, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1984:REP

[PG84] Rappoport Paul and Michael L. Goetz. Ridge estimation of price sensitivities for long distance telephone demand for bell competitors. Communications in Statistics: Theory and Methods, 13(2): 241–250, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pena:1989:OCM

[PG89] Daniel Peña and Irwin Guttman. Optimal collapsing of mixture distributions in robust recursive estimation. Communications in Statistics: Theory and Methods, 18(3):817–833, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1980:MAT

[PH80] H. I. Patel and Erwin M. Hearne III. Multivariate analysis for the two-period repeated measures crossover design with application to clinical trials. Communications in Statistics: Theory and Methods, 9(18):1919–1929, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Portier:1984:TET

[PH84] C. Portier and D. Hoel. Type 1 error of trend tests in proportions and the design of cancer screens. Communications in Statistics: Theory and Methods, 13(1):1–14, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1985:NEM

[PH85] M. S. Prasad and V. R. Harischandrakar. A note on estimation in a minimum-type model with an exponential factor. Communications in Statistics: Theory and Methods, 14(12):2899–2905, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paetzold:1988:BPI

[PH88] Kamona L. Paetzold and H. Fenwick Huss. Bernoulli processes interrelated by constrain. Communications in Statistics: Theory and Methods, 17(5):1377–1383, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1989:EBP

[PH89] S. R. Paul and N. I. Ho. Estimation in the bivariate Poisson distribution and hypothesis testing concerning independence. Communications in Statistics: Theory and Methods, 18(3):1123–1133, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Phillips:1980:AAJ

[Phi80] G. D. A. Phillips. All asymptotic justification for using a jackknifed two stage least squares estimator for bias reduction in a simultaneous equation model. Communications in Statistics: Theory and Methods, 9(1):81–86, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Piegorsch:1983:QFG

[Pie83] Walter W. Piegorsch. The questions of fit in the Gregor Mendel controversy. Communications in Statistics: Theory and Methods, 12(19):2289–2304, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pillai:1985:SLD

[Pil85] R. N. Pillai. Semi-α-Laplace distributions. Communications in Statistics: Theory and Methods, 14(4):991–1000, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1981:AID

[PK81] H. I. Patel and C. G. Khatri. Analysis of incomplete data in experiments with repeated measurements in clinical trials using a stochastic model. Communications in Statistics: Theory and Methods, 10(22):2259–2277, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1985:CFO

[PK85a] M. S. Patel and J. K. Arap Koske. Conditions for fourth order rotatability in K dimensions. Communications in Statistics: Theory and Methods, 14(6):1343–1351, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1985:CPE

[PK85b] P. D. Puri and Sanpei Kageyama. Constructions of partially efficiency-balanced designs and their analysis. Communications in Statistics: Theory and Methods, 14(6):1315–1342, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peddada:1986:PNV

[PK86] Shyamal Das Peddada and Ravindra Khattree. On Pitman nearness and variance of estimators. Communications in Statistics: Theory and Methods, 15(10):3005–3017, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pinsker:1987:UCC

[PKG87] Sh I. Pinsker, Victor Kipnis, and Eugen Grechanovsky. The use of conditional cutoffs in a forward selection procedure. Communications in Statistics: Theory and Methods, 16(8):2227–2241, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perng:1989:TSB

[PKN89] S. K. Perng, K. E. Kemp, and P. I. Nelson. Testing for a separation between two normal distributions. Communications in Statistics: Theory and Methods, 18(5):1895–1912, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pavur:1982:TPA

[PL82] R. J. Pavur and T. O. Lewis. Test procedures for the analysis of experimental designs with correlated nonnormal data. Communications in Statistics: Theory and Methods, 11(20):2315–2334, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pavur:1983:UTF

[PL83] R. J. Pavur and T. O. Levis. Unbiased F tests for factorial experiments for correlated data. Communications in Statistics: Theory and Methods, 12(7):829–840, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pavur:1986:TPC

[PL86] R. J. Pavur and T. O. Lewis. Test procedures in covariance analysis assuming a general correlation pattern. Communications in Statistics: Theory and Methods, 15(2):473–491, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papageorgiou:1988:CEP

[PL88] H. Papageorgiou and S. Loukas. Conditional even point estimation for bivariate discrete distributions. Communications in Statistics: Theory and Methods, 17(10):3403–3412, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pliskin:1987:RTE

[Pli87] Jeffrey L. Pliskin. A ridge-type estimator and good prior means. Communications in Statistics: Theory and Methods, 16(12):3429– 3437, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parthasarathy:1982:SCM

[PM82] P. R. Parthasarathy and P. Mayilswami. Stochastic compartmental models with birth, death and immigration of particles. Communications in Statistics: Theory and Methods, 11(14):1625–1642, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parthasarathy:1983:STD

[PM83] P. R. Parthasarathy and P. Mayilswami. Sojourn time distribution in a stochastic compartmental system. Communications in Statistics: Theory and Methods, 12(14):1619–1635, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padgett:1984:NDE

[PM84] W. J. Padgett and Diane T. McNichols. Nonparametric density estimation from censored data. Communications in Statistics: Theory and Methods, 13(13):1581–1611, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1985:PAD

[PM85] Badri Prasad and V. V. Menon. Poisson approximation to the distribution of empty cells in randomized occupancy. Communications in Statistics: Theory and Methods, 14(1):41–48, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1988:SIE

[PM88] S. N. Pandey and H. J. Malik. Some improved estimators for a measures of dispersion of an inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 17(11):3935–3949, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1989:STS

[PM89a] B. N. Pandey and H. J. Malik. Shrinkage testimators for the shape parameter of Weibull distribution under Type II censoring. Communications in Statistics: Theory and Methods, 18(4):1175–1199, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Passi:1989:LSA

[PM89b] Ranjit M. Passi and Claude Morel. Least squares adaptive polynomials. Communications in Statistics: Theory and Methods, 18 (1):315–329, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pantula:1985:ELM

[PNA85] Sastry G. Pantula, Larry A. Nelson, and Richard L. Anderson. Estimation of linear models for field experiments. Communications in Statistics: Theory and Methods, 14(9):2199–2217, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peddada:1989:IRE

[PNS89] S. D. Peddada, A. K. Nigam, and A. K. Saxena. On the inadmissibility of ridge estimator in a linear model. Communications in Statistics: Theory and Methods, 18(10):3571–3585, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1984:OTS

[PO84a] M. S. Patel and J. A. M. Ottieno. Optimum two stage groupscreening designs. Communications in Statistics: Theory and Methods, 13(21):2649–2663, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1984:TSG

[PO84b] M. S. Patel and J. A. M. Ottieno. Two-stage group-screening designs with equal prior probabilities and no errors in decisions. Communications in Statistics: Theory and Methods, 13(9):1147–1159, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1987:OTS

[PO87] M. S. Patel and J. A. M. Ottieno. Optimum two stage groupscreening with unequal group sizes and with errors in decisions. Communications in Statistics: Theory and Methods, 16(10):2799– 2820, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pocock:1982:WRC

[Poc82] Stuart J. Pocock. When not to rely on the central limit theorem — an example from absenteeism data. Communications in Statistics: Theory and Methods, 11(19):2169–2179, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ponnapalli:1983:MPE

[Pon83] R. Ponnapalli. On multinomial parametric estimation. Communications in Statistics: Theory and Methods, 12(1):47–72, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pourahmadi:1983:EFS

[Pou83] Mohsen Pourahmadi. Exact factorization of the spectral density and its application to forecasting and time series analysis. Communications in Statistics: Theory and Methods, 12(18):2085–2094, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandian:1981:RII

[PP81] Soundara V. S. Pandian and K. N. Ponnuswamy. On the recovery of intra- and interblock information in n-ary designs. Communications in Statistics: Theory and Methods, 10(17):1779–1788, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Penas:1987:CNP

[PP87] Linda M. Penas and Jagdish K. Patel. Classifying k normal populations with respect to a control: making at most m (0 ≤ m < k) incorrect decisions. Communications in Statistics: Theory and Methods, 16(8):2287–2311, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pal:1988:NVB

[PP88a] Satyabrata Pal and Samarendranath Pal. Nonproper variance balanced designs and optimality. Communications in Statistics: Theory and Methods, 17(5):1685–1695, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paulo:1988:DGL

[PP88b] Gilberto A. Paulo and Clovis A. Peres. Diagnostics for GLMS with linear inequality parameter constraints. Communications in Statistics: Theory and Methods, 17(12):4205–4219, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandit:1984:NDA

[PR84] S. N. Narahari Pandit and C. Raghavendra Rao. On a new distribution arising in decaying inventory systems. Communications in Statistics: Theory and Methods, 13(9):1107–1120, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Popovich:1985:CTC

[PR85] E. A. Popovich and P. V. Rao. Conditional tests for censored matched pairs. Communications in Statistics: Theory and Methods, 14(9):2041–2056, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Press:1986:EBE

[PR86] S. James Press and John E. Rolph. Empirical Bayes estimation of the mean in a multivariate normal distribution. Communications in Statistics: Theory and Methods, 15(7):2201–2228, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pena:1988:EPU

[PR88] Edsel A. Pena and Vijay K. Rohatgi. Estimation of parameters under a random censorship model. Communications in Statistics: Theory and Methods, 17(8):2819–2829, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prajneshu:1983:ESG

[Pra83] Prajneshu. Environmental stochasticity of Gompertz model with time delay. Communications in Statistics: Theory and Methods, 12 (12):1359–1369, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Praagman:1989:ENTa

[Pra89a] Jaap Praagman. Efficiency of nonparametric tests for scale shift at an unknown time point. Communications in Statistics: Theory and Methods, 18(4):1325–1338, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Praagman:1989:ENTb

[Pra89b] Jaap Praagman. Efficiency of nonparametric tests for scale shift at an unknown time point. Communications in Statistics: Theory and Methods, 18(10):3665–3678, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1989:SIR

[Pra89c] B. Prasad. Some improved ratio type estimators of population mean and ratio in finite population sample surveys. Communications in Statistics: Theory and Methods, 18(1):379–392, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See erratum [Pra91].

Prasad:1991:ESI

[Pra91] B. Prasad, Research Officer. Erratum to “Some improved ratio type estimators of population mean and ratio in finite population sample survey”. Communications in Statistics: Theory and Methods, 20(7):2331–2332, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Pra89c].

Preuss:1980:CSB

[Pre80] Lucien G. Preuss. A class of statistics based on the information concept. Communications in Statistics: Theory and Methods, 9 (15):1563–1585, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Press:1989:NMM

[Pre89] S. James Press. A note on the multivariate miser for balancing experiments. Communications in Statistics: Theory and Methods, 18 (3):1099–1105, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Price:1982:CAR

[Pri82] J. Michael Price. Comparisons among regression estimators under the generalized mean square error criterion. Communications in Statistics: Theory and Methods, 11(17):1965–1984, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Proctor:1984:DCA

[Pro84] C. H. Proctor. On the detection of clustering and anisotropy using binary data from a lattice patch. Communications in Statistics: Theory and Methods, 13(5):617–638, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Provost:1986:DST

[Pro86] Serge B. Provost. On the distribution of some test statistics connected with the multivariate linear functional relationship model. Communications in Statistics: Theory and Methods, 15(4):1285– 1298, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Provost:1988:LRT

[Pro88] Serge B. Provost. Likelihood ratio test for independence with partial multivariate normal data. Communications in Statistics: Theory and Methods, 17(6):1763–1775, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1988:RTC

[PRPR88] N. G. N. Prasad, J. N. K. Rao, N. G. N. Prasad, and J. N. K. Rao. Robust tests and confidence intervals for error variance in a regression model and for functions of variance components in an unbalanced random one-way model. Communications in Statistics: Theory and Methods, 17(4):1111–1133, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patil:1986:DWD

[PRR86] G. P. Patil, C. R. Rao, and M. V. Ratnaparkhi. On discrete weighted distributions and their use in model choice for observed data. Communications in Statistics: Theory and Methods, 15(3): 907–918, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1980:SES

[PS80] B. N. Pandey and Padmakar Singh. Shrinkage estimation of scale in exponential distribution from censored samples. Communications in Statistics: Theory and Methods, 9(8):875–882, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1981:EVN

[PS81] M. Pandey and J. Singh. Estimation of variance of a normal population. Communications in Statistics: Theory and Methods, 10 (18):1879–1884, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Philip:1983:LBC

[PS83] Smith Philip and J. Sedransk. Lower bounds for confidence coefficients for confidence intervals for finite population quantiles. Communications in Statistics: Theory and Methods, 12(12):1329–1344, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patterson:1985:AFS

[PS85a] Blossom H. Patterson and Paul J. Smith. Asymptotic and finite sample behavior of the time on test estimator under random censorship when lifetimes are not exponential. Communications in Statistics: Theory and Methods, 14(7):1643–1658, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Press:1985:BMM

[PS85b] S. James Press and Kazuo Shigemasu. Bayesian MANOVA and MANOCOVA under exchangeability. Communications in Statistics: Theory and Methods, 14(5):1053–1078, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1986:AES

[PS86] M. S. Prasad and Alladi Subramanyam. Asymptotic efficiency of sequential and non-sequential procedures based on Fisher’s method of combining tests. Communications in Statistics: Theory and Methods, 15(5):1665–1680, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1987:NPS

[PS87] Dong Ho Park and K. M. Lai Saxena. Nonparametric procedure for selecting the least NBU population: large sample approach. Communications in Statistics: Theory and Methods, 16(11):3359– 3377, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pasternack:1987:GTH

[PST87] Bernard S. Pasternack, Milton Sobel, and James Thomas. Grouptesting, halving procedures, and binary search. Communications in Statistics: Theory and Methods, 16(10):2851–2871, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parkash:1986:CQQ

[PT86] Om Parkash and H. C. Taneja. Characterization of the quantitative-qualitative measure of inaccuracy for discrete generalized probability distributions. Communications in Statistics: Theory and Methods, 15(12):3763–3771, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Philippou:1989:NPI

[PTA89] Andreas N. Philippou, Gregory A. Tripsiannis, and Demetris L. Antzoulakos. New Pólya and inverse Pólya distributions of order k. Communications in Statistics: Theory and Methods, 18(6):2125– 2137, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1986:APL

[PU86] M. Pandey and S. K. Upadhyay. Approximate prediction limit for Weibull failure based on preliminary test estimator. Communications in Statistics: Theory and Methods, 15(1):241–250, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1986:BPA

[Pur86] Prem S. Puri. A branching process approach to a problem connected with a T.V. game show. Communications in Statistics: Theory and Methods, 15(3):919–933, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pendleton:1986:IFT

[PVB86] O. J. Pendleton, M. Von Tress, and R. Bremer. Interpretation of the four types of analysis of variance tables in SAS. Communications in Statistics: Theory and Methods, 15(9):2785–2808, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padgett:1981:BNE

[PW81] W. J. Padgett and L. J. Wei. A Bayesian nonparametric estimator of survival probability assuming increasing failure rate. Communications in Statistics: Theory and Methods, 10(1):49–63, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panaretos:1986:SDA

[PX86] John Panaretos and Evdokia Xekalaki. On some distributions arising from certain generalized sampling schemes. Communications in Statistics: Theory and Methods, 15(3):873–891, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Posten:1982:RTS

[PYO82] Harry O. Posten, Hon Cheng Yeh, and Donald B. Owen. Robustness of the two-sample t-test under violations of the homogeneity of variance assumption. Communications in Statistics: Theory and Methods, 11(2):109–126, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramey:1980:BSP

[RA80] James T. Ramey, Jr. and Khursheed Alam. A Bayes sequential procedure for selecting the most probable multinomial event. Communications in Statistics: Theory and Methods, 9(3):265–276, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radhakrishnan:1982:IML

[Rad82] R. Radhakrishnan. Inadmissibility of the maximum likelihood estimator for a multivariate normal distribution when some observations are missing. Communications in Statistics: Theory and Methods, 11(8):941–955, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radhakrishnan:1984:EMD

[Rad84] R. Radhakrishnan. Estimating Mahalanobis’s distance using Bayesian analysis. Communications in Statistics: Theory and Methods, 13(20):2583–2600, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radhakrishnan:1985:IFC

[Rad85] R. Radhakrishnan. Influence functions for certain parameters in discriminant analysis when a single discriminant function is not adequate. Communications in Statistics: Theory and Methods, 14 (3):535–549, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raftery:1984:CME

[Raf84] Adrian E. Raftery. A continuous multivariate exponential distribution. Communications in Statistics: Theory and Methods, 13 (8):947–965, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramirez:1984:MSP

[Ram84] Manuel Mendoza Ramírez. A modification to some proposed tests in relation to the problem of switching regression models. Communications in Statistics: Theory and Methods, 13(7):901–914, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramakrishna:1987:CPH

[Ram87] M. V. Ramakrishna. Computing the probability of hash table/urn overflow. Communications in Statistics: Theory and Methods, 16 (11):3343–3353, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1981:CBN

[Rao81a] B. Raja Rao. Correlation between the numbers of two types of children in a family with the MPSD for the family size. Communications in Statistics: Theory and Methods, 10(3):249–254, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1981:LIM

[Rao81b] C. Radhakrishna Rao. A lemma on G-inverse of a matrix and computation of correlation coefficients in the singular case. Communications in Statistics: Theory and Methods, 10(1):1–10, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1982:UPI

[Rao82] Poduri S. R. S. Rao. Use of prior information for estimating the variance components. Communications in Statistics: Theory and Methods, 11(15):1659–1669, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1986:JDS

[Rao86] B. Raja Rao. Joint distribution of simultaneous exposures to several carcinogens in a case-control study: sample size determination. Communications in Statistics: Theory and Methods, 15(10):

3035–3065, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1987:CUP

[Rao87] T. J. Rao. On certain unbiased product estimators. Communications in Statistics: Theory and Methods, 16(12):3631–3640, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rathnam:1989:ADU

[Rat89] Damaraju V. Rathnam. Analysis of a design under the competing effects model with missing values. Communications in Statistics: Theory and Methods, 18(8):2825–2833, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rauhut:1982:EME

[Rau82] Burkhard O. Rauhut. Estimation of the mean of the exponential distribution with an outlying observation. Communications in Statistics: Theory and Methods, 11(13):1439–1452, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raul:1987:BCB

[Rau87] S. R. Raul. On the beta-correlated binomial (BCB) distribution — a three parameter generalization of the binomial distribution. Communications in Statistics: Theory and Methods, 16(5):1473– 1478, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raveh:1981:NNA

[Rav81] Adi Raveh. A numerical nonmetric approach for analyzing time series data. Communications in Statistics: Theory and Methods, 10(8):809–821, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raveh:1989:GCM

[Rav89] Adi Raveh. Gini correlation as a measure of monotonicity and two of its usages. Communications in Statistics: Theory and Methods, 18(4):1415–1423, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rayner:1982:COU

[Ray82] J. C. W. Rayner. Constructing optimal univariate tests. Communications in Statistics: Theory and Methods, 11(17):1903–1917, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raynor:1983:CPM

[Ray83] William J. Raynor, Jr. Caliper pair-matching on a continuous variable in case-control studies. Communications in Statistics: Theory and Methods, 12(13):1499–1509, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rayner:1985:TVC

[Ray85] Robert K. Rayner. A test for variance–covariance parameters in normal linear models. Communications in Statistics: Theory and Methods, 14(10):2379–2392, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajagopalan:1983:BIV

[RB83a] Muthiya Rajagopalan and Lyle Broemeling. Bayesian inference for the variance components in general mixed linear models. Communications in Statistics: Theory and Methods, 12(6):701–723, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rizzuto:1983:BPB

[RB83b] G. T. Rizzuto and T. L. Bouillon. Busy period in batch queues. Communications in Statistics: Theory and Methods, 12(22):2655– 2662, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rayner:1988:STG

[RB88] J. C. W. Rayner and D. J. Best. Smooth tests of goodness of fit for regular distributions. Communications in Statistics: Theory and Methods, 17(10):3235–3267, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rudolfer:1985:LSI

[RC85] S. M. Rudolfer and M. G. Campbell. Large sample inference for diagnostic normal limits in Gaussian populations. Communications in Statistics: Theory and Methods, 14(8):1801–1835, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1988:USS

[RD88] P. V. Rao and R. R. Daley. On the use of the Scholz–Sievers method for distribution-free inference about the difference in slopes of two regression lines. Communications in Statistics: Theory and Methods, 17(5):1437–1448, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reinsel:1984:EEC

[Rei84] Gregory C. Reinsel. Effects of the estimation of covariance matrix parameters in the generalized multivariate linear model. Communications in Statistics: Theory and Methods, 13(5):639–650, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robert:1984:STC

[RF84a] W. Chen Robert and K. Hwang Frank. Some theorems, counterexamples, and conjectures in multinomial selection theory. Communications in Statistics: Theory and Methods, 13(10):1289–1298, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rust:1984:MKW

[RF84b] Steven W. Rust and Michael A. Filgner. A modification of the Kruskal–Wallis statistic for the generalized Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 13(16): 2013–2027, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ragab:1984:OSL

[RG84] Aisha Ragab and Jack Green. On order statistics from the loglogistic distribution and their properties. Communications in Statistics: Theory and Methods, 13(21):2713–2724, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rezayat:1988:UIV

[RG88] Fahimeh Rezayat and Anandalingam G. Using instrumental variables for selecting the order of ARMA models. Communications in Statistics: Theory and Methods, 17(9):3029–3065, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radaelli:1989:DCD

[RG89] Giovanni Radaelli and Giuseppe Gallus. On detection of a change in the dynamics of rare health events. Communications in Statistics: Theory and Methods, 18(2):579–590, 1989. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rattihaili:1988:OBC

[RH88] R. N. Rattihaili and M. R. Huggi. Optimal Bayesian confidence interval of fixed width. Communications in Statistics: Theory and Methods, 17(10):3477–3482, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Redner:1987:EPM

[RHB87] Richard A. Redner, Richard J. Hathaway, and James C. Bezdek. Estimating the parameters of mixture models with modal estimators. Communications in Statistics: Theory and Methods, 16(9): 2639–2660, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rieder:1980:LRC

[Rie80] Helmut Rieder. Locally robust correlation coefficients. Communications in Statistics: Theory and Methods, 9(8):803–819, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rieder:1982:GRP

[Rie82] Helmut Rieder. A general robustness property of rank correlations. Communications in Statistics: Theory and Methods, 11(3):233– 234, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rigdon:1989:TGF

[Rig89] Steven E. Rigdon. Testing goodness-of-fit for the power law process. Communications in Statistics: Theory and Methods, 18(12): 4665–4676, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Risko:1985:SBB

[Ris85] Kenneth J. Risko. Selecting the better binomial population with unequal sample sizes. Communications in Statistics: Theory and Methods, 14(1):123–158, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rivest:1988:DDU

[Riv88] Louis-Paul Rivest. A distribution for dependent unit vectors. Communications in Statistics: Theory and Methods, 17(2):461–483, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramalingam:1984:CLS

[RJ84] Shanmugam Ramalingam and Singh Jagbir. A characterization of the logarithmic series distribution and its application. Communications in Statistics: Theory and Methods, 13(7):865–875, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radhakrishnan:1981:IFC

[RK81] R. Radhakrishnan and A. M. Kshirsagar. Influence functions for certain parameters in multivariate analysis. Communications in Statistics: Theory and Methods, 10(6):515–529, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raubertas:1986:HTN

[RLN86] Richard F. Raubertas, Chu-In Charles Lee, and Erik V. Nordheim. Hypothesis tests for normal means constrained by linear inequalities. Communications in Statistics: Theory and Methods, 15(9): 2809–2833, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1988:AVF

[RM88] B. Raja Rao and Gary M. Marsh. Approximate variance formulas and asymptotic joint sampling distributions of standardized risk ratios in the presence of competing risks in cohort studies. Communications in Statistics: Theory and Methods, 17(3):745– 777, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robinson:1989:MRC

[RM89] Tony Robinson and Rana Moyeed. Making robust the crossvalidatory choice of smoothing parameter in spline smoothing regression. Communications in Statistics: Theory and Methods, 18 (2):523–539, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1986:STS

[RMW86] B. Raja Rao, Gary M. Marsh, and Joan Winwood. Sidak-type simultaneous confidence intervals for the risk measures RSMR_{1}, based on proportional mortality measures in epidemiologic studies involving several competing risks of death. Communications in Statistics: Theory and Methods, 15(2):515–536, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robert:1981:EPG

[Rob81] Gilchrist Robert. Estimation of the parameters of the gamma distribution by means of a maximal invariant. Communications in Statistics: Theory and Methods, 10(11):1095–1110, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robert:1989:LBR

[Rob89] Christian Robert. A lower bound for the risk of classes of shrinkage estimators in a general multivariate estimation problem and some deduced estimators. Communications in Statistics: Theory and Methods, 18(6):2289–2299, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rocke:1982:IRL

[Roc82] David M. Rocke. Inference for response-limited time-series models. Communications in Statistics: Theory and Methods, 11(22):2587– 2596, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez:1985:CTS

[Rod85] Robert N. Rodriguez. Computer tools for statistical quality control. Communications in Statistics: Theory and Methods, 14(11): 2767, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rohatgi:1987:DOS

[Roh87] Vijay K. Rohatgi. Distribution of order statistics with random sample size. Communications in Statistics: Theory and Methods, 16 (12):3739–3743, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roio:1987:ARL

[Roi87] Javier Roio. On the admissibility of cX̄ + d with respect to the LINEX loss function. Communications in Statistics: Theory and Methods, 16(12):3745–3748, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rojo:1985:EEL

[Roj85] Javier Rojo. Examples of the effect of the loss function on the best equivariant estimator. Communications in Statistics: Theory and Methods, 14(12):3007–3032, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rojo:1986:NSC

[Roj86] Javier Rojo. A necessary and sufficient condition for an estimator to be optimal. Communications in Statistics: Theory and Methods, 15(5):1647–1651, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roller:1988:TSP

[Rol88] William F. Roller. A two- sample partially sequential probability ratio test. Communications in Statistics: Theory and Methods, 17 (9):2907–2920, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rosenstein:1989:TSD

[Ros89] Rebecca B. Rosenstein. Tests of symmetry derived as components of Pearson’s phi-squared distance measure. Communications in Statistics: Theory and Methods, 18(5):1617–1626, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1988:VCL

[RP88] Poduri S. R. S. Rao and Kuranchie Pasmore. Variance components of the linear regression model with a random intercept. Communications in Statistics: Theory and Methods, 17(4):1011–1026, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roller:1989:TSS

[RP89] William F. Roller and Walter R. Pirie. Truncated sequential sign tests for location shift. Communications in Statistics: Theory and Methods, 18(2):721–742, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Razzaghi:1987:MML

[RR87] Mehdi Razzaghi and William S. Rayens. Modified maximum likelihood estimator for the mixing distribution in a finite mixture. Communications in Statistics: Theory and Methods, 16(9):2661– 2676, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajarshi:1988:BDR

[RR88a] Sujata Rajarshi and M. B. Rajarshi. Bathtub distributions: a review. Communications in Statistics: Theory and Methods, 17 (8):2597–2621, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ram:1988:AGL

[RR88b] C. Trioathi Ram and C. Gupta Ramesh. Another generalization of the logarithmic series and the geometric distributions. Communications in Statistics: Theory and Methods, 17(5):1541–1547, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1986:PNC

[RRM86] C. R. Rao, I. P. Reating, and Robert D. Mason. The Pitman nearness criterion and its determination. Communications in Statistics: Theory and Methods, 15(11):3173–3191, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raghavarao:1980:CDT

[RS80] D. Raghavarao and K. R. Shah. A class of d_{o} designs for twoway elimination of heterogeneity. Communications in Statistics: Theory and Methods, 9(1):75–80, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1981:APT

[RS81] C. V. Rao and K. P. Saxena. On approximation of power of a test procedure based on preliminary tests of significance. Communications in Statistics: Theory and Methods, 10(13):1305–1321, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1984:AMT

[RS84a] Poduri S. R. S. Rao and Edward A. Sylvestre. Anova and MINQUE type of estimators for the one-way random effects model. Communications in Statistics: Theory and Methods, 13(14):1667–1673, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rosenberger:1984:IST

[RS84b] James L. Rosenberger and Dennis E. Smith. Interruptible supersaturated two-level designs. Communications in Statistics: Theory and Methods, 13(5):599–609, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reeves:1987:CRP

[RS87] Jack Reeves and Marc Sobel. Complete ranking procedures with quadratic loss. Communications in Statistics: Theory and Methods, 16(3):837–854, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rohatgi:1988:CDC

[RS88] Vijay K. Rohatgi and A. K. Md. Ehsanes Saleh. A class of distributions connected to order statistics with nonintegral sample size. Communications in Statistics: Theory and Methods, 17(6): 2005–2012, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruymgaart:1989:NHT

[RS89] Frits H. Ruymgaart and William B. Smith. A note on homogeneity tests of covariance matrices. Communications in Statistics: Theory and Methods, 18(6):2301–2310, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rilstone:1989:NER

[RU89] P. Rilstone and A. Ullah. Nonparametric estimation of response coefficients. Communications in Statistics: Theory and Methods, 18 (7):2615–2627, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruberg:1986:CAN

[Rub86a] Stephen J. Ruberg. A continuously adaptive nonparametric twosample test. Communications in Statistics: Theory and Methods, 15(10):2899–2920, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruberg:1986:EST

[Rub86b] Stephen J. Ruberg. Efficiencies of some two-sample location tests for a broad class of distributions. Communications in Statistics: Theory and Methods, 15(10):2991–3004, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Russell:1980:FRC

[Rus80] K. G. Russell. Further results on the connectedness and optimality of designs of type O:XB. Communications in Statistics: Theory and Methods, 9(4):439–447, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruskeepaa:1985:CGD

[Rus85] Heikki Ruskeepää. Conditionally Gaussian distributions and an application to Kalman filtering with stochastic regressors. Communications in Statistics: Theory and Methods, 14(12):2919–2942, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruskeepaa:1988:EPG

[Rus88] Heikki Ruskeepää. Exact predictors for a generalized AR(1) process with an AR(1) parameter. Communications in Statistics: Theory and Methods, 17(3):875–885, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rai:1982:NMV

[RV82] Kamta Rai and John Van Ryzin. A note on a multivariate version of Rolle’s Theorem and uniqueness of maximum likelihood roots. Communications in Statistics: Theory and Methods, 11(13):1505– 1510, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stanish:1981:CVS

[SA81] William M. Stanish and Randy U. Allred. Categorical variable selection based on entropy reduction. Communications in Statistics: Theory and Methods, 10(17):1733–1750, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1982:RHT

[SA82] M. S. Srivastava and Hayat M. Awan. On the robustness of Hotelling’s T^{2}-test and distribution of linear and quadratic forms III sampling from a mixture of two multivariate normal populations. Communications in Statistics: Theory and Methods, 11(1): 81–107, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silver:1985:AAD

[SA85] J. Lew Silver and Mukhtar Ali. Additional approximations to the distribution of a structural coefficient estimator for the case of two included endogenous variables for the case of two included endogenous variables. Communications in Statistics: Theory and Methods, 14(6):1269–1282, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:1986:EPR

[SA86] Brajendra C. Sutradhar and Mir M. Ali. Estimation of the parameters of a regression model with a multivariate t error variable. Communications in Statistics: Theory and Methods, 15(2): 429–450, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Senyo:1989:MDS

[SA89] B. Senyo and S. K. Adjibolosoo. Measuring the degree of severity of heteroskedasticity and the choice between the OLS estimator and the 2SAE. Communications in Statistics: Theory and Methods, 18 (9):3451–3462, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadasivan:1982:TTM

[Sad82] G. Sadasivan. A Thurstone-type model for paired comparisons with unequal numbers of repetitions. Communications in Statistics: Theory and Methods, 11(7):821–833, 1982. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sager:1983:EMI

[Sag83] Thomas W. Sager. Estimating modes and isopleths. Communications in Statistics: Theory and Methods, 12(5):529–557, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Speckman:1987:SSC

[SAH87] Paul Speckman, Sharon Anderson, and John Hewett. Sample sizes for comparing means of two multivariate normal populations. Communications in Statistics: Theory and Methods, 16(4):1207–1218, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sakata:1987:LRT

[Sak87] Toshio Sakata. Likelihood ratio test for one-sided hypothesis of covariance matrices of two normal populations. Communications in Statistics: Theory and Methods, 16(11):3157–3168, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sakata:1988:DCV

[Sak88] Toshio Sakata. Detecting a change in variances. Communications in Statistics: Theory and Methods, 17(3):641–655, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samanta:1984:ELP

[Sam84] M. Samanta. Estimation of the location parameter of an exponential distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 13(11):1357–1364, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samanta:1985:GDU

[Sam85] M. Samanta. On a geometrical derivation of the upper tail of the null distribution of W -exponential. Communications in Statistics: Theory and Methods, 14(12):2947–2950, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samanta:1988:UAM

[Sam88] M. Samanta. A unified approach to minimum variance unbiased estimation of a probability function belonging to an exponential family. Communications in Statistics: Theory and Methods, 17 (10):3413–3426, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sambamoorthi:1989:ITC

[Sam89] N. Sambamoorthi. Information theoretic criterion approach to dimensionality reduction in multinomial logistic regression models. Part I: theory. Communications in Statistics: Theory and Methods, 18(6):2087–2106, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanchez:1987:MLF

[San87] Susan M. Sanchez. A modified least-failures sampling procedure for Bernoulli subset selection. Communications in Statistics: Theory and Methods, 16(12):3609–3629, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:1989:NAR

[SAN89a] S. M. Sadooghi-Alvandi and N. Nematollahi. A note on the admissibility of relative to the LINEX loss function. Communications in Statistics: Theory and Methods, 18(5):1871–1873, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanchez:1989:UEF

[San89b] Susan M. Sanchez. Unbiased estimation following selection for Bernoulli populations. Communications in Statistics: Theory and Methods, 18(11):4275–4301, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarkar:1985:BCT

[Sar85] Nityaoanda Sarkar. Box-cox transformation and the problem of heteroscedasticity. Communications in Statistics: Theory and Methods, 14(2):363–379, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1988:CCB

[Sat88] Y. S. Sathe. On the correlation coefficient between the first and the r-th smallest order statistics based on n independent exponential random variables. Communications in Statistics: Theory and Methods, 17(10):3295–3299, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saunders:1982:MLE

[Sau82] Sam C. Saunders. On maximum likelihood estimators of shape and scale parameters and their application in constructing confidence contours. Communications in Statistics: Theory and Methods, 11 (16):1769–1791, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saw:1987:MCA

[Saw87] John G. Saw. MANOVA with covariates and an application in profile analysis. Communications in Statistics: Theory and Methods, 16(6):1683–1696, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saw:1988:NND

[Saw88] John G. Saw. The non-null distributions in ANOVA. Communications in Statistics: Theory and Methods, 17(11):3833–3845, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiue:1982:APR

[SB82a] Wei Kei Shiue and Lee J. Bain. Approximate percentiles for ratios of F -variates. Communications in Statistics: Theory and Methods, 11(16):1831–1838, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stevenson:1982:BIM

[SB82b] David J. Stevenson and R. P. Bland. Bayes ian multiple comparisons of noemal populations without the assumption Op equal variances. Communications in Statistics: Theory and Methods, 11(1):49–57, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:1984:BIF

[SB84a] Samir Shaarawy and Lyle Broemeling. Bayesian inferences and forecasts with moving averages processes. Communications in Statistics: Theory and Methods, 13(15):1871–1888, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1984:BLS

[SB84b] John Shin and H. D. Brunk. Bayes least squares linear estimation of densities. Communications in Statistics: Theory and Methods, 13 (18):2253–2291, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:1985:IPA

[SB85] Samir Shaarawy and Lyle Broemeling. Inference and prediction with ARMA processes. Communications in Statistics: Theory and Methods, 14(10):2531–2548, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Senn:1989:MLE

[SB89a] Stephen Senn and Richard Brown. Maximum likelihood estimation of treatment effects for samples subject to regression to the mean. Communications in Statistics: Theory and Methods, 18(9):3389– 3406, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:1989:ADR

[SB89b] Brajendra C. Sutradhar and Roy F. Bartlett. An approximation to the distribution of the ratio of two general quadratic forms with application. Communications in Statistics: Theory and Methods, 18 (4):1563–1588, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salazar:1981:PCR

[SBC81] Diego Salazar, Lyle Broemeling, and Albert Chi. Parameter changes in a regression model with autocorrelated errors. Communications in Statistics: Theory and Methods, 10(17):1751–1758, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuel-Cahn:1980:CST

[SC80] Ester Samuel-Cahn. Comparisons of sequential two-sided tests for normal hypothesis. Communications in Statistics: Theory and Methods, 9(3):277–290, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:1985:ABD

[SC85] Arvind K. Shah and P. L. Claypool. Analysis of binary data in the randomized complete block design. Communications in Statistics: Theory and Methods, 14(5):1175–1179, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1986:ADT

[SC86] M. S. Srivastava and S. R. Chakravorti. Asymptotic distributions of two test statistics for testing independence with missing data. Communications in Statistics: Theory and Methods, 15(2):571– 588, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1989:CIR

[SC89] V. K. Srivastava and A. Chaturvedi. Comparison of improved regression estimators with and without moments. Communications in Statistics: Theory and Methods, 18(3):989–999, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1985:TMA

[SCG85] Bimal Kumar Sinha, Bernard Clement, and Narayan C. Giri. Tests for means with additional information. Communications in Statistics: Theory and Methods, 14(6):1427–1451, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schervish:1980:TRR

[Sch80] Mark J. Schervish. Tolerance regions for random vectors and best linear predictors. Communications in Statistics: Theory and Methods, 9(11):1177–1183, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schechtman:1982:NTD

[Sch82] Edna Schechtman. A nonparametric test for detecting changes in location. Communications in Statistics: Theory and Methods, 11 (13):1475–1482, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schechtman:1983:CND

[Sch83] Edna Schechtman. A conservative nonparametric distribution-free confidence bound for the shift in the changepoint problem. Communications in Statistics: Theory and Methods, 12(21):2455–2464, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schemper:1984:GWT

[Sch84a] M. Schemper. A generalized Wilcoxon test for data defined by intervals. Communications in Statistics: Theory and Methods, 13 (6):681–684, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schemper:1984:SPT

[Sch84b] M. Schemper. A survey of permutation tests for censored survival data. Communications in Statistics: Theory and Methods, 13(13): 1655–1665, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schilling:1985:RAS

[Sch85a] Edward G. Schilling. The role of acceptance sampling in modern quality control. Communications in Statistics: Theory and Methods, 14(11):2769–2783, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schott:1985:CSG

[Sch85b] James R. Schott. Comparison of some goodness of fit tests for a single non-isotropic hypothetical principal component. Communications in Statistics: Theory and Methods, 14(5):1201–1215, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schuster:1985:ISC

[Sch85c] Eugene F. Schuster. Incorporating support constraints into nonparametric estimators of densities. Communications in Statistics:

Theory and Methods, 14(5):1123–1136, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schmoyer:1986:CSM

[Sch86a] Richard L. Schmoyer. Convex sets in minimum-distance estimation. Communications in Statistics: Theory and Methods, 15(12): 3625–3635, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schott:1986:NCV

[Sch86b] James R. Schott. A note on the critical values used in stepwise tests for multiplicative components of interaction. Communications in Statistics: Theory and Methods, 15(5):1561–1570, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schott:1986:TCT

[Sch86c] James R. Schott. Tests concerning two non-isotropic principal components. Communications in Statistics: Theory and Methods, 15 (11):3307–3320, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schemper:1987:BCL

[Sch87] M. Schemper. On bootstrap confidence limits for possibly skew distributed statistics. Communications in Statistics: Theory and Methods, 16(6):1585–1590, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schneider:1989:DCR

[Sch89] Christina Schneider. Dempster’s combination rule and sufficient statistics. Communications in Statistics: Theory and Methods, 18 (10):3841–3850, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sclove:1980:OFC

[Scl80] Stanley L. Sclove. The occurrence of fingerprint characteristics as a two-dimensional Poisson process. Communications in Statistics: Theory and Methods, 9(7):675–695, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sparks:1983:M

[SCT83] R. S. Sparks, D. Coutsourides, and L. Troskie. The multivariate C_{p}. Communications in Statistics: Theory and Methods, 12(15):

1775–1793, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarmah:1983:EPR

[SD83] P. Sarmah and A. D. Dharmadhikari. Estimation of parameters of 1-out-of-2:G repairable system. Communications in Statistics: Theory and Methods, 12(14):1609–1618, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarmah:1984:EPG

[SD84a] P. Sarmah and A. D. Dharmadhikari. Estimation of parameters of 1-out-of-n: g repairable system. Communications in Statistics: Theory and Methods, 13(11):1365–1374, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scariano:1984:CTG

[SD84b] Stephen M. Scariano and James M. Davenport. Corrected F -tests in the general linear model. Communications in Statistics: Theory and Methods, 13(25):3155–3172, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steinberg:1985:DFC

[SD85] Seth M. Steinberg and C. E. Davis. Distribution-free confidence intervals for quantiles in small samples. Communications in Statistics: Theory and Methods, 14(4):979–990, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scariano:1986:FMA

[SD86] Stephen M. Scariano and James M. Davenport. A four-moment approach and other practical solutions to the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 15(5): 1467–1505, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schall:1987:AVA

[SD87a] R. Schall and T. T. Dunne. Additional variables and adjusted estimates with arbitrary known variance-covariance structure. Communications in Statistics: Theory and Methods, 16(5):1253–1263, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1987:NIB

[SD87b] M. Singh and A. Dey. A note on incomplete block designs for symmetrical parallel line assays. Communications in Statistics: Theory and Methods, 16(12):3555–3564, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steinberg:1987:CNP

[SD87c] Seth M. Steinberg and C. E. Davis. Comparison of nonparametric point estimators for interquantile differences in moderate sized samples. Communications in Statistics: Theory and Methods, 16 (6):1607–1616, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarmah:1988:SAT

[SD88] P. Sarmah and A. D. Dharmadhikari. Statistical analysis of a twounit dependent repairable system. Communications in Statistics: Theory and Methods, 17(5):1385–1400, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1989:NTA

[SdB89] Jan W. H. Swanepoel and C. F. de Beer. On a new test for autocorrelation in regression models under nonnormality. Communications in Statistics: Theory and Methods, 18(5):1793–1804, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastva:1983:FSP

[SDCB83] V. K. Srivastva, T. D. Dwivedi, Y. P. Chaubey, and S. Bhatnagar. Finite sample properties of Beale’s ratio estimator. Communications in Statistics: Theory and Methods, 12(15):1795–1805, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Searle:1980:AHL

[Sea80] S. R. Searle. Arbitrary hypotheses in linear mod fi.s with unbalanced data. Communications in Statistics: Theory and Methods, 9(2):181–200, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Searle:1988:MMU

[Sea88] Shayle R. Searle. Mixed models and unbalanced data: wherefrom, whereat and whereto? Communications in Statistics: Theory and Methods, 17(4):935–968, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sehr:1988:CCL

[Seh88] Joachim Sehr. On a conjecture concerning the least favorable configuration of a two-stage selection procedure. Communications in Statistics: Theory and Methods, 17(10):3221–3233, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Selvavel:1989:UES

[Sel89] K. Selvavel. Unbiased estimation in sampling from two onetruncation parameter families when both samples are Type II censored. Communications in Statistics: Theory and Methods, 18(9): 3519–3531, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1980:TSP

[Sen80] Pranab Kumar Sen. On time-sequential point estimation of the mean of an exponential distribution. Communications in Statistics: Theory and Methods, 9(1):27–38, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1981:SIP

[Sen81] Pranab Kumar Sen. Some invariance principles for mixed rank statistics and induced order statistics and some application. Communications in Statistics: Theory and Methods, 10(17):1691–1718, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1984:JST

[Sen84] Pranab Kumar Sen. A James–Stein type detour of U-statistics. Communications in Statistics: Theory and Methods, 13(22):2725– 2747, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1986:WJS

[Sen86] Pranab Kumar Sen. Whither jackknifing in Stein-rule estimation. Communications in Statistics: Theory and Methods, 15(7):2245– 2266, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

SenGupta:1987:GBH

[Sen87] Ashis SenGupta. Generalizations of Barlett’s and Hartley’s tests of homogeneity using “overall variability”. Communications in Statistics: Theory and Methods, 16(4):987–996, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schelchter:1985:PHS

[SF85] Mark D. Schelchter and Alan B. Forsythe. Post-hoc selection of covariates in randomized experiments. Communications in Statistics: Theory and Methods, 14(3):679–699, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwager:1984:NCS

[SFR84] S. J. Schwager, W. T. Federer, and B. L. Raktoe. Nonisomorphic complete sets of orthogonal f-squares and Hadamard matrices. Communications in Statistics: Theory and Methods, 13(11): 1391–1406, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schmidt:1982:IVG

[SG82] Peter Schmidt and David K. Gullkey. An improved version of the Geary test. Communications in Statistics: Theory and Methods, 11(4):359–374, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sielken:1984:MTS

[SG84a] Robert L. Sielken and Edward E. Gbur. Multiyear, through the season crop acreage estimation using estimated acreage in sample segments. Communications in Statistics: Theory and Methods, 13 (23):2961–2974, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1984:AGB

[SG84b] Bimal Kumar Sinha and Anirban Das Gupta. Admissibility of generalized Bayes and Pitman estimates in the non-regular family. Communications in Statistics: Theory and Methods, 13(14):1709– 1721, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1984:PBA

[SG84c] Philip J. Smith and Erdogan Gunel. Practical Bayesian approaches to the analystcs of a 2×2 contingency table with incompletely categorized data. Communications in Statistics: Theory and Methods, 13(16):1941–1963, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanubhogue:1985:NPT

[SG85a] Ashok Shanubhogue and A. P. Gore. Non-parametric tests for several sample location problem based on subsample extrema. Communications in Statistics: Theory and Methods, 14(5):1137–1150, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stephenson:1985:TSN

[SG85b] W. Robert Stephenson and Malay Ghosh. Two sample nonparametric tests based on subsamples. Communications in Statistics: Theory and Methods, 14(7):1669–1684, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soofi:1986:EBM

[SG86a] Ehsan S. Soofi and D. V. Gokhale. An extension of Bayesian measure of information to regression. Communications in Statistics: Theory and Methods, 15(12):3607–3616, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stephenson:1986:AP

[SG86b] W. R. Stephenson and Malay Ghosh. Acknowledgment of priority. Communications in Statistics: Theory and Methods, 15(8):2569, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subhash:1986:SCM

[SG86c] Kochar C. Subhash and R. P. Gupta. Some competitors of the mood test for the two-sample scale problem. Communications in Statistics: Theory and Methods, 15(1):231–239, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shan:1987:ECD

[SG87] D. N. Shan and M. R. Gupta. An efficiency comparison of dual ratio and product estimators. Communications in Statistics: Theory and Methods, 16(3):693–703, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shelty:1988:TST

[SG88a] I. D. Shelty and Z. Govindarajulu. A two-sample test for location. Communications in Statistics: Theory and Methods, 17(7):2389– 2401, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1988:BAL

[SG88b] S. K. Sinha and Irwin Guttman. Bayesian analysis of life-testing problems involving the Weibull distribution. Communications in Statistics: Theory and Methods, 17(2):343–356, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1983:NED

[SGP83] R. S. Singh, T. Gasser, and B. Prasad. Nonparametric estimates of distribution functions. Communications in Statistics: Theory and Methods, 12(18):2095–2108, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seely:1982:SDU

[SH82] Justus Seely and Robert V. Hogg. Symmetrically distributed and unbiased estimators in linear models. Communications in Statistics: Theory and Methods, 11(7):721–729, 1982. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1984:RCC

[SH84] M. S. Srivastava and Mohammad Awan Hyat. On the robustness of the correlation coefficient in sampling from a mixture of two bivariate normals. Communications in Statistics: Theory and Methods, 13(3):371–382, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1989:ECD

[SH89] Paul W. Stewart and Ronald W. Helms. Evaluating the cumulative distribution function of a Bonferroni test statistic from a Type III bivariate t distribution. Communications in Statistics: Theory and Methods, 18(2):461–477, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shapiro:1980:SAA

[Sha80] Jesse M. Shapiro. Sequential analysis applied to testing the mean of a normal population with known coefficient of variation. Communications in Statistics: Theory and Methods, 9(1):39–49, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaban:1981:CPI

[Sha81] S. A. Shaban. Computation of the Poisson-inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 10 (14):1389–1399, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sharma:1983:SSS

[Sha83] Divakar Sharma. Sufficiency of a sufficient statistic for equivariant estimation. Communications in Statistics: Theory and Methods, 12(5):611–623, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:1984:THW

[Sha84] Ramalingam Shanmugam. A test for homogeneity when the sample is a positive Lagrangian Poisson distribution. Communications in Statistics: Theory and Methods, 13(25):3243–3250, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sharma:1986:ECA

[Sha86] Subhash C. Sharma. The effects of correlation among observations on the consistency property of sample variance. Communications in Statistics: Theory and Methods, 15(4):1125–1152, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:1988:UBC

[Sha88a] K. R. Shah. On uniformly better combined estimates in rowcolumn designs with adjusted orthogonality. Communications in Statistics: Theory and Methods, 17(9):3121–3124, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:1988:CJE

[Sha88b] Jun Shao. Consistency of jackknife estimators of the variances of simple quantiles. Communications in Statistics: Theory and Methods, 17(9):3017–3028, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:1989:BGS

[Sha89] Jun Shao. Bootstrapping for generalized L-statistics. Communications in Statistics: Theory and Methods, 18(6):2005–2016, 1989.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1985:OSJ

[SHB85] A. K. Md Ehsanes Saleh, Khatab M. Hassanein, and Edward F. Brown. Optimum spacings for the joint estimation and tests of hypothesis of location and scale parameters of the Cauchy distribution. Communications in Statistics: Theory and Methods, 14 (1):247–254, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sheehan:1986:ADD

[She86] Dennis P. Sheehan. Approximating the distribution of the Durbin– Watson statistic. Communications in Statistics: Theory and Methods, 15(1):73–88, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirahata:1980:RTS

[Shi80] S. Shirahata. Rank tests for the k-sample problem with restricted alternatives. Communications in Statistics: Theory and Methods, 9 (10):1071–1086, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirahata:1982:NMIb

[Shi82a] S. Shirahata. A nonparametric measure of interclass correlation. Communications in Statistics: Theory and Methods, 11(15):1723– 1732, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirahata:1982:NMIa

[Shi82b] S. Shirahata. Nonparametric measures of intraclass correlation. Communications in Statistics: Theory and Methods, 11(15):1707– 1721, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirakura:1982:NAM

[Shi82c] Teruhiro Shirakura. Norm of alias matrices for (l + 1)-factor interactions in balanced fractional 2^{M} factorial designs of resolution 2l + 1. Communications in Statistics: Theory and Methods, 11(1): 31–47, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:1983:CBU

[Shi83a] Kunio Shimizu. Covariances between UMVU estimators for means of transformed variables. Communications in Statistics: Theory and Methods, 12(21):2525–2532, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:1983:UEC

[Shi83b] Kunio Shimizu. Umvu estimation for covariances and first product moments of transformed variables. Communications in Statistics: Theory and Methods, 12(14):1661–1674, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirahata:1985:APK

[Shi85] S. Shirahata. Asymptotic properties of Kruskal–Wallis test and Friedman test in the analysis of variance models with random effects. Communications in Statistics: Theory and Methods, 14(7): 1685–1692, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiau:1987:NMC

[Shi87] Jyh-Jen Horng Shiau. A note on MSE coverage intervals in a partial spline model. Communications in Statistics: Theory and Methods, 16(7):1851–1866, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1988:THU

[Shi88] Ning-Zhong Shi. A test of homogeneity for umbrella alternatives and tables of the level probabilities. Communications in Statistics: Theory and Methods, 17(3):657–670, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shorack:1982:BRR

[Sho82a] Galen R. Shorack. Bootstrapping robust regression. Communications in Statistics: Theory and Methods, 11(9):961–972, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoukri:1982:NET

[Sho82b] Mohamed M. Shoukri. A note on estimation and test for the Rayleigh amplitude distribution. Communications in Statistics: Theory and Methods, 11(17):1919–1934, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoukri:1982:GDP

[Sho82c] Mohamed M. Shoukri. On a generalization for the double Poisson distribution. Communications in Statistics: Theory and Methods, 11(2):151–164, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoemaker:1984:RPC

[Sho84] Lewis H. Shoemaker. Robustness properties for a class of scale estimators. Communications in Statistics: Theory and Methods, 13(1):15–28, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:1981:UMV

[SI81] K. Shimizu and K. Iwase. Uniformly minimum variance unbiased estimation in lognormal and related distributions. Communications in Statistics: Theory and Methods, 10(11):1127–1147, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seto:1985:UEM

[SI85] Noriaki Setô and Kosei Iwase. Umvu estimators of the mode and limits of an interval for the inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 14(5):1151–1161, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:1987:UEA

[SI87] Kunio Shimizu and Kōsei Iwase. Unbiased estimation of the autocovariance function in a stationary generalized lognormal process. Communications in Statistics: Theory and Methods, 16(7): 2145–2154, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sichel:1986:PEW

[Sic86] H. S. Sichel. Parameter estimation for a word frequency distribution based on occupancy theory. Communications in Statistics: Theory and Methods, 15(3):935–949, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sievers:1983:WDF

[Sie83] Gerald L. Sievers. A weighted dispersion function for estimation in linear models. Communications in Statistics: Theory and Methods, 12(10):1161–1179, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvapulle:1986:NSC

[Sil86] Mervyn J. Silvapulle. A note on some conditions for optimality of certain ridge estimators. Communications in Statistics: Theory and Methods, 15(5):1579–1585, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvapulle:1988:SEC

[Sil88] Mervyn J. Silvapulle. On the symmetry of M-estimators computed by the Huber–Dutter algorithm. Communications in Statistics: Theory and Methods, 17(12):4279–4284, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simonoff:1984:CRM

[Sim84] Jeffrey S. Simonoff. A comparison of robust methods and detection of outliers techniques when estimating a location parameter. Communications in Statistics: Theory and Methods, 13(7):813– 842, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simonoff:1987:BIP

[Sim87] Jeffrey S. Simonoff. The breakdown and influence properties of outlier rejection-plus-mean procedures. Communications in Statistics: Theory and Methods, 16(6):1749–1760, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1980:BIC

[Sin80a] Bahadur Singh. Bayesian inference for comparing several treatment means under a Poisson model. Communications in Statistics: Theory and Methods, 9(11):1137–1145, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1980:E

[Sin80b] N. Singh. On the estimation of Pr(X_{1} < Y < X_{2}). Communications in Statistics: Theory and Methods, 9(15):1551–1561, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1982:NUM

[Sin82] Murari Singh. A note on the use of multi-auxiliary information. Communications in Statistics: Theory and Methods, 11(8):933– 939, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1983:BEP

[Sin83] S. K. Sinha. Bayesian estimators of the parameters and reliability function from mixed exponentially distributed time-censored life test data. Communications in Statistics: Theory and Methods, 12 (2):141–151, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1985:SAO

[Sin85] N. Singh. A simple and asymptotically optimal test for the equality of k(≥ 2) exponential distributions based on type II censored samples. Communications in Statistics: Theory and Methods, 14 (7):1615–1625, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1986:NEV

[Sin86] Housila P. Singh. A note on the estimation of variance of sample mean using the knowledge of coefficient of variation in normal population. Communications in Statistics: Theory and Methods, 15(12):3737–3746, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1987:OGR

[Sin87] Avinash C. Singh. On the optimality and a generalization of Rao–Robson’s statistic. Communications in Statistics: Theory and Methods, 16(11):3255–3273, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1988:EOC

[Sin88] Sukhminder Singh. Estimation in overlapping clusters. Communications in Statistics: Theory and Methods, 17(2):613–621, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sirvanci:1984:ESP

[Sir84] Mete Sirvanci. An estimator for the scale parameter of the two parameter Weibull distribution for Type I singly right-censored data. Communications in Statistics: Theory and Methods, 13(14): 1759–1768, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sirvanci:1986:CTW

[Sir86] Mete Sirvanci. Comparison of two Weibull distributions under random censoring. Communications in Statistics: Theory and Methods, 15(6):1819–1836, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1985:AUR

[SIU85] H. P. Singh, R. Iachan, and L. N. Upadhyaya. Almost unbiased ratio and product estimators based on interpenetrating subsamples. Communications in Statistics: Theory and Methods, 14(4): 963–978, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sivaganesan:1988:RPM

[Siv88] S. Sivaganesan. Range of posterior measures for priors with arbitrary contaminations. Communications in Statistics: Theory and Methods, 17(5):1591–1612, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Six:1981:RMN

[Six81] F. B. Six. Representations of multivariate normal distributions with special correlation structures. Communications in Statistics: Theory and Methods, 10(13):1285–1295, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:1987:EPE

[SJ87a] B. N. Shetty and P. C. Joshi. Estimation of parameters of k exponential distributions in doubly censored samples. Communications in Statistics: Theory and Methods, 16(7):2115–2123, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stein:1987:BCP

[SJ87b] Gillian Z. Stein and June M. Juritz. Bivariate compound Poisson distributions. Communications in Statistics: Theory and Methods, 16(12):3591–3607, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stein:1988:LMI

[SJ88] Gillian Z. Stein and June M. Juritz. Linear models with an inverse Gaussian Poisson error distribution. Communications in Statistics: Theory and Methods, 17(2):557–571, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:1989:LRT

[SJ89a] B. N. Shetty and P. C. Joshi. Likelihood ratio test for testing equality of location parameters of two exponential distributions from doubly censored samples. Communications in Statistics: Theory and Methods, 18(6):2063–2072, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1989:DCS

[SJ89b] B. Singh and D. D. Joshi. On the distributions of components sums of squares in one-way unbalanced random model under heterogeneous error variances. Communications in Statistics: Theory and Methods, 18(11):4137–4144, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shultis:1986:NCP

[SJM86] J. Kenneth Shultis, D. E. Johnson, and G. A. Milliken. Nonconjugate prior distributions and their estimation for use in failure models. Communications in Statistics: Theory and Methods, 15 (9):2835–2865, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1982:NTL

[SK82] A. K. Md E. Saleh and Sen P. K. Nonparametric tests for location after a preliminary test on regression. Communications in Statistics: Theory and Methods, 11(6):639–652, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1986:TEA

[SK86] Harshinder Singh and Subhash C. Kochar. A test for exponentiality against HNBUE alternatives. Communications in Statistics: Theory and Methods, 15(8):2295–2304, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1987:TGF

[SK87a] Harshinder Singh and Amarjot Kaur. Testing goodness of fit in terms of the mean residual life function. Communications in Statistics: Theory and Methods, 16(2):489–497, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sunwoo:1987:SLR

[SK87b] Ha Sik Sunwoo and Byung Chun Kim. L_{1} for the simple linear regression model. Communications in Statistics: Theory and Methods, 16(6):1703–1715, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schucany:1988:CTMa

[SK88] W. R. Schucany and A. M. Kshirsagar. Communications in theory and methods Editor-in-chief. Communications in Statistics: Theory and Methods, 17(1):edi–edii, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1983:UMA

[SKC83] M. Singh, P. Kumar, and R. R. Chandak. Use of multi-auxiliary variables as a condensed auxiliary variable in selecting a sample. Communications in Statistics: Theory and Methods, 12(14):1685– 1697, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sposito:1980:UGP

[SKG80] V. A. Sposito, W. J. Kennedy, and J. E. Gentle. Useful generalized properties of L_{1} estimators. Communications in Statistics: Theory and Methods, 9(12):1309–1315, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sahai:1985:SBV

[SKK85] Hardeo Sahai, Andre I. Khuri, and C. H. Kapadia. A second bibliography on variance components. Communications in Statistics: Theory and Methods, 14(1):63–115, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stanley:1984:AEH

[SKN84] P. Stanley, J. W. Kennerley, and J. M. Newton. Assignment errors in histograms with equal class intervals of arbitrary width. Communications in Statistics: Theory and Methods, 13(18):2293–2306, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Semenya:1983:LMM

[SKSF83] K. A. Semenya, G. G. Koch, M. E. Stokes, and R. N. Forthofer. Linear models methods for some rank function analyses of ordinal categorical data. Communications in Statistics: Theory and Methods, 12(11):1277–1298, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1980:NIT

[SL80a] Y. S. Sathe and S. R. Lingras. A note on the inequalities for tail probability of the multivariate normal distribution. Communications in Statistics: Theory and Methods, 9(7):711–715, 1980. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1980:DEI

[SL80b] J. H. Smith and T. O. Lewis. Determining the effects of intraclass correlation on factorial experiments. Communications in Statistics: Theory and Methods, 9(13):1353–1364, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sirvanci:1982:NRC

[SL82a] Mete Sirvanci and Ibrahim Levent. Numerical results concerning the MLE of an exponential parameter used in life testing. Communications in Statistics: Theory and Methods, 11(11):1269–1278, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1982:EIC

[SL82b] J. H. Smith and T. O. Lewis. Effects of intraclass correlation on covariance analysis. Communications in Statistics: Theory and Methods, 11(1):71–80, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1983:CTC

[SL83] M. S. Srivastava and G. C. Lee. On the choice of transformations of the correlation coefficient with or without an outlier. Communications in Statistics: Theory and Methods, 12(21):2533–2547, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Slymen:1984:SDA

[SL84] D. J. Slymen and P. A. Lachenbruch. Survival distributions arising from two families and generated by transformations. Communications in Statistics: Theory and Methods, 13(10):1179–1201, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1985:RTC

[SL85] M. S. Srivastava and G. C. Lee. On the robustness of tests of correlation coefficient in the presence of an outlier. Communications in Statistics: Theory and Methods, 14(1):25–40, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stahlecker:1987:AME

[SL87] P. Stahlecker and J. Lauterbach. Approximate minimax estimation in linear regression: theoretical results. Communications in Statistics: Theory and Methods, 16(4):1101–1116, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saikkonen:1988:EME

[SL88] Pentti Saikkonen and Ritva Luukkone. An efficient method for the estimation of multivariate moving average models. Communications in Statistics: Theory and Methods, 17(12):4257–4270, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saikkonen:1989:EMA

[SL89a] Pentti Saikkonen and Ritva Luukkonen. Estimating multivariate autoregressive moving average models by fitting long autoregressions. Communications in Statistics: Theory and Methods, 18(4): 1589–1615, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stahlecker:1989:ALM

[SL89b] P. Stahlecker and J. Lauterbach. Approximate linear minimax estimation in regression analysis with ellipsoidal constraints. Communications in Statistics: Theory and Methods, 18(8):2755–2784, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saunders:1983:TIP

[SLH83] Roy Saunders, Purushottam Laud, and Tsushung A. Hua. A test of independence in the presence of discrete points. Communications in Statistics: Theory and Methods, 12(16):1831–1839, 1983. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steel:1987:RBG

[SlR87] S. J. Steel and N. J. le Roux. A reparameterisation of a bivariate gamma extension. Communications in Statistics: Theory and Methods, 16(1):293–305, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Slud:1986:IIP

[Slu86] Eric V. Slud. Inefficiency of inferences with the partial likelihood. Communications in Statistics: Theory and Methods, 15(11):3333– 3351, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1982:ECM

[SM82a] Bimal Kumar Sinha and Omar Mouqadem. Estimation of the common mean of two univariate normal populations. Communications in Statistics: Theory and Methods, 11(14):1603–1614, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sway:1982:EMR

[SM82b] P. A. V. B. Sway and J. S. Mehta. The existence of moments of ridge-like k-class and partially restricted reduced form estimators. Communications in Statistics: Theory and Methods, 11(24):2793– 2799, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1985:SSP

[SM85] V. Soundararajan and D. Muthuraj. Single sampling plans indexed by point of control and inflection point. Communications in Statistics: Theory and Methods, 14(10):2393–2410, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seth:1986:EPP

[SM86a] Anand K. Seth and Sati Mazumdar. Estimation of parameters of a polynomial model under intra class correlation structure for incomplete longitudinal data. Communications in Statistics: Theory and Methods, 15(5):1549–1559, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stefanski:1986:REL

[SM86b] Leonard A. Stefanski and Michael Meredith. Robust estimation of location in samples of size three. Communications in Statistics: Theory and Methods, 15(10):2921–2933, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:1987:CAP

[SM87] Brajendra C. Sutradhar and Ian B. MacNeill. Charts for autocorrelation and partial-autocorrelation functions for the AR(2) and MA(2) processes. Communications in Statistics: Theory and Methods, 16(9):2717–2725, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1989:DSP

[SM89] V. Soundararajan and D. Muthuraj. Double sampling plans indexed by point of control and inflection point. Communications in Statistics: Theory and Methods, 18(7):2569–2597, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1980:RCS

[Smi80] E. J. Smith. Rejoinder to “ ‘Comments on the special issue of Communications in Statistics (vol. 8, no. 10) concerned with weather modification experiments’ by Jerzy Neyman”. Communications in Statistics: Theory and Methods, 9(18):1979–1983, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Ney80b].

Smith:1982:DFT

[Smi82] Theodore M. Smith. Distribution-free test procedures for a factorial experiments. Communications in Statistics: Theory and Methods, 11(4):439–443, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1983:IBD

[Smi83] Woollcott Smith. Inequalities for bivariate distribution with x ≤ y and marginals given. Communications in Statistics: Theory and Methods, 12(12):1371–1379, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1984:EPF

[Smi84] Dennis E. Smith. Examination of a parsimonious factor screening technique: screening K = 2^{N−1}−1 factors in N runs. Communications in Statistics: Theory and Methods, 13(3):383–393, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1985:MLM

[Smi85] Richard H. Smith, Jr. Maximum likelihood mean and covariance matrix estimation constrained to general positive semi-definiteness. Communications in Statistics: Theory and Methods, 14(9):2163– 2179, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smouse:1983:NBI

[Smo83] Evan P. Smouse. Nonparametric Bayesian inference for dichotomous response models. Communications in Statistics: Theory and Methods, 12(24):2847–2859, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:1984:ADP

[SMS84] Wong Chi Song, J. F. McDonald, and M. A. Selby. The asymptotic distribution of a pivotal quantity for a capture–recapture model. Communications in Statistics: Theory and Methods, 13(1):93–98, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spector:1982:BMB

[SN82] Philip C. Spector and H. Joseph Newton. The bias of the MLE of Box’s degrees of freedom correction factor for correlation in ANOVA. Communications in Statistics: Theory and Methods, 11 (25):2937–2945, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1989:ICA

[SN89a] N. Singh and T. Nirmalan. Identification of composite (Σ + Π) ARMA models by relatively simpler models. Communications in Statistics: Theory and Methods, 18(3):1067–1083, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1989:IEL

[SN89b] Nariaki Sugiura and May Thu Naing. Improved estimators for the location of double exponential distribution. Communications in Statistics: Theory and Methods, 18(2):541–554, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scariano:1984:TRF

[SND84] Stephen M. Scariano, James W. Neill, and James M. Davenport. Testing regression function adequacy with correlation and without replication. Communications in Statistics: Theory and Methods, 13 (10):1227–1237, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuel:1984:IEL

[SNR84] Kotz Samuel, L. Johnson Norman, and N. Rodriguez Robert. Inspection errors in link sampling. Communications in Statistics:

Theory and Methods, 13(10):1203–1213, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1987:CIW

[Sob87] Mare J. Sobel. Confidence intervals which admit positively (semi) relevant sursets that are not invariant. Communications in Statistics: Theory and Methods, 16(8):2269–2276, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1989:CLC

[Sob89a] Marc Sobel. The conditional level of confidence intervals for the normal variance and applications. Communications in Statistics: Theory and Methods, 18(8):2813–2823, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1989:ENB

[Sob89b] Marc Sobel. The estimation of the number of balls put in a sequence of urns. Communications in Statistics: Theory and Methods, 18(4):1297–1307, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shumway:1981:ETH

[SOL81] R. H. Shumway, D. E. Olsen, and L. J. Levy. Estimation and tests of hypotheses for the initial mean and covariance in the Kalman filter model. Communications in Statistics: Theory and Methods, 10 (16):1625–1641, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solomon:1983:JSJ

[Sol83] Herbert Solomon. Jury size and jury verdicts. Communications in Statistics: Theory and Methods, 12(19):2179–2215, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soms:1984:NER

[Som84] Andrew P. Soms. Note on an extension of rational bounds for the t-tail area to arbitrary degrees of freedom. Communications in Statistics: Theory and Methods, 13(7):887–891, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soms:1985:PTS

[Som85] Andrew P. Soms. Permutation tests for k-sample binomial data with comparisons of exact and approximate P -levels. Communications in Statistics: Theory and Methods, 14(1):217–233, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soms:1988:LMM

[Som88] Andrew P. Soms. The Lindstrom–Madden method for series systems with repeated components. Communications in Statistics: Theory and Methods, 17(6):1845–1856, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soms:1989:AST

[Som89a] Andrew P. Soms. Applications of Schur theory to reliability inference for k of n systems. Communications in Statistics: Theory and Methods, 18(12):4459–4469, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soms:1989:SRR

[Som89b] Andrew P. Soms. Some recent results for series system reliability. Communications in Statistics: Theory and Methods, 18(11):4211– 4227, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soofi:1988:PCR

[Soo88] Ehsan S. Soofi. Principal component regression under exchangeability. Communications in Statistics: Theory and Methods, 17 (6):1717–1733, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spall:1985:IFB

[Spa85] James C. Spall. An implicit function based procedure for analyzing maximum likelihood estimates from nonidentically distributed data. Communications in Statistics: Theory and Methods, 14(7): 1719–1730, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spall:1986:AAB

[Spa86] James C. Spall. An approximation for analyzing a broad class of implicitly and explicitly defined estimators. Communications in Statistics: Theory and Methods, 15(12):3747–3762, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spall:1988:ESP

[Spa88a] James C. Spall. Effect of the sample on the posterior probability in Bayesian analysis. Communications in Statistics: Theory and Methods, 17(6):1811–1827, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sparks:1988:DFM

[Spa88b] Ross S. Sparks. A distribution free method of detecting outliers in regression. Communications in Statistics: Theory and Methods, 17(3):705–727, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spall:1989:EIK

[Spa89] James C. Spall. Effect of imprecisely known nuisance parameters on estimates of primary parameters. Communications in Statistics: Theory and Methods, 18(1):219–237, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sprott:1982:RML

[Spr82] D. A. Sprott. Robustness and maximum likelihood estimation. Communications in Statistics: Theory and Methods, 11(22):2513– 2529, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spruill:1989:ADE

[Spr89] M. C. Spruill. Accurate detection of excess over a standard based on a random sample. Communications in Statistics: Theory and Methods, 18(9):3305–3312, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1982:TST

[Spu82] John D. Spurrier. Two-stage tests of hypotheses in the general linear model. Communications in Statistics: Theory and Methods, 11 (24):2775–2791, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1983:CTI

[Spu83a] John D. Spurrier. Comparison of two independent life tests subject to type II censoring. Communications in Statistics: Theory and Methods, 12(16):1813–1829, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1983:EUC

[Spu83b] John D. Spurrier. Exact uniform confidence bands for periodic regression. Communications in Statistics: Theory and Methods, 12 (9):969–973, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1984:OTE

[Spu84] John D. Spurrier. An overview of tests for exponentiality. Communications in Statistics: Theory and Methods, 13(13):1635–1654, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silva:1980:EWR

[SQ80] Claudio Silva and Dana Quade. Evaluation of weighted rankings using expected significance level. Communications in Statistics: Theory and Methods, 9(10):1087–1096, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salama:1981:UWR

[SQ81] Ibrahim A. Salama and Dana Quade. Using weighted rankings to test against ordered alternatives in complete blocks. Communications in Statistics: Theory and Methods, 10(4):385–399, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salama:1982:NCT

[SQ82] Ibrahim Salama and Dana Quade. A nonparametric comparison of two multiple regressions by means of a weighted measure of correlation. Communications in Statistics: Theory and Methods, 11 (11):1185–1195, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1988:EPL

[SR88a] A. K. Md. Ehsanes Saleh and Vijay K. Rohatgi. Estimation of parameters of location-scale family of distributions in complete and Type II censored samples. Communications in Statistics: Theory and Methods, 17(9):2831–2858, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samara:1988:TCB

[SR88b] Basil Samara and Ronald H. Randles. A test for correlation based on Kendall’s tau. Communications in Statistics: Theory and Methods, 17(9):3191–3205, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seely:1988:WCR

[SR88c] Justus F. Seely and El-Houssainy Rady. When can random effects be treated as fixed effects for computing a test statistic for a linear hypothesis? Communications in Statistics: Theory and Methods, 17(4):1089–1109, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1983:DHM

[Sri83] M. S. Srivastava. On the distribution of Hotelling’s T^{2} and multiple correlation R^{2} when sampling from a mixture of two normals. Communications in Statistics: Theory and Methods, 12(13): 1481–1497, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1985:MDM

[Sri85] M. S. Srivastava. Multivariate data with missing observations. Communications in Statistics: Theory and Methods, 14(4):775– 792, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1987:ICO

[Sri87a] Jaya Srivastava. On the inadequacy of the customary orthogonal arrays in quality control and general scientific experimentation, and the need of probing designs of higher revealing power. Communications in Statistics: Theory and Methods, 16(10):2901–2941, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1987:BMR

[Sri87b] M. S. Srivastava. Bootstrap method in ranking and slippage problems. Communications in Statistics: Theory and Methods, 16(11): 3285–3299, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1987:PAS

[Sri87c] M. S. Srivastava. Profile analysis of several groups. Communications in Statistics: Theory and Methods, 16(3):909–926, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1987:ATP

[Sri87d] Rakesh Srivastava. Admissibility of a test procedure in a conditionally specified random model. Communications in Statistics:

Theory and Methods, 16(6):1789–1796, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1989:AST

[Sri89] Jaya Srivastava. Advances in the statistical theory of comparison of lifetimes of machines under the generalized Weibull distribution. Communications in Statistics: Theory and Methods, 18(3):1031– 1045, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sana:1982:PAF

[SRP82] G. M. Sana, B. L. Raktoe, and H. Pesotan. On the problem of augmented fractional factorial designs. Communications in Statistics: Theory and Methods, 11(23):2731–2745, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schaefer:1984:RLE

[SRW84] R. L. Schaefer, L. D. Roi, and R. A. Wolfe. A ridge logistic estimator. Communications in Statistics: Theory and Methods, 13 (1):99–113, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1981:EED

[SS81a] Y. S. Sathe and S. P. Shah. On estimating P (X > Y) for the exponential distribution. Communications in Statistics: Theory and Methods, 10(1):39–47, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

So:1981:RST

[SS81b] Ying Chiu So and Pranab Kumar Sen. Repeated significance tests based on likelihood ratio statistics. Communications in Statistics: Theory and Methods, 10(21):2149–2176, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1983:NTL

[SS83] A. K. Md. Ehsanes Saleh and P. K. Sen. Nonparametric tests of location after a preliminary test on regression in the multivariate case. Communications in Statistics: Theory and Methods, 12(16): 1855–1872, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1985:ARE

[SS85a] A. K. Md. Ehsanesh Saleh and Pranab Kumar Sen. Asymptotic relative efficiency of some joint-tests for location and scale parameters based on selected order statistics. Communications in Statistics: Theory and Methods, 14(3):621–633, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1985:SEL

[SS85b] Md A. K. Ehsanes Saleh and Pranab Kumar Sen. On shrinkage Mestimators of location parameters. Communications in Statistics: Theory and Methods, 14(10):2313–2329, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:1985:KDE

[SS85c] David W. Scott and Simon J. Sheather. Kernel density estimation with binned data. Communications in Statistics: Theory and Methods, 14(6):1353–1359, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1986:SLS

[SS86a] A. K. Md. Ehsanes Saleh and Pranab Kumar Sen. On shrinkage least squares estimation in a parallelism problem. Communications in Statistics: Theory and Methods, 15(5):1451–1466, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1986:SEM

[SS86b] Md A. K. Ehsanes Saleh and Pranab Kumar Sen. On shrinkage R-estimation in a multiple regression model. Communications in Statistics: Theory and Methods, 15(7):2229–2244, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1986:SBR

[SS86c] Y. S. Sathe and R. G. Shenoy. On sharper bounds for the risk of James–Stein estimators. Communications in Statistics: Theory and Methods, 15(2):613–617, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stumpf:1986:EDA

[SS86d] R. H. Stumpf and H. S. Steyn. Exact distributions associated with an i × j × k contingency table. Communications in Statistics: Theory and Methods, 15(6):1889–1904, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1987:RPS

[SS87a] A. K. Md Ehsanes Saleh and Pranab Kumar Sen. Relative performance of Stein-rule and preliminary test estimators in linear models least squares theory. Communications in Statistics: Theory and Methods, 16(2):461–476, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1987:WOC

[SS87b] J. N. Srivastava and M. M. Siddiqui. When A- and D-optimality conflict. Communications in Statistics: Theory and Methods, 16 (6):1675–1682, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schork:1988:SMN

[SS88a] Nicholas J. Schork and M. Anthony Schork. Skewness and mixtures of normal distributions. Communications in Statistics: Theory and Methods, 17(11):3951–3969, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1988:REW

[SS88b] Xiquan Shi and Jun Shao. Resampling estimation when observations are m-dependent. Communications in Statistics: Theory and Methods, 17(11):3923–3934, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurr:1988:GFU

[SS88c] B. D. Spurr and C. D. Sinclair. A general formula for the upper tail significance levels of empirical distribution function test statistics. Communications in Statistics: Theory and Methods, 17(12):4121– 4131, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stam:1988:ECI

[SS88d] A. J. Stam and A. G. M. Steerneman. Equality conditions in an inequality for affinities of finite measures. Communications in Statistics: Theory and Methods, 17(11):3971–3979, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:1989:HSV

[SS89] Jun Shao and Xiquan Shi. Half-sample variance estimation. Communications in Statistics: Theory and Methods, 18(11):4197–4210, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivnstava:1989:CRT

[SSK89] Rani S. Srivnstava, S. P. Srivastava, and B. B. Khare. Chain ratio type estimator for ratio of two population means using auxiliary characters. Communications in Statistics: Theory and Methods, 18 (10):3917–3926, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schucany:1988:CTMb

[SSKK88] W. R. Schucany, W. R. Schucany, A. M. Kshirsagar, and A. M. Kshirsagar. Communications in theory and methods Editor-in-chief. Communications in Statistics: Theory and Methods, 17(2): –, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1985:IBP

[SSS^{+}85] A. F. M. Smith, A. M. Skene, J. E. H. Shaw, J. C. Naylor, and M. Dransfield. The implementation of the Bayesian paradigm. Communications in Statistics: Theory and Methods, 14(5):1079– 1102, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smit:1987:WCR

[SSS87] C. F. Smit, N. G. N. Swart, and D. J. Stoker. Weighted combinations of rank tests for location-scale alternatives. Communications in Statistics: Theory and Methods, 16(12):3535–3553, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivenkataramana:1983:IRP

[ST83] T. Srivenkataramana and D. S. Tracy. Interchangeability of the ratio and product methods in sample surveys. Communications in Statistics: Theory and Methods, 12(18):2143–2150, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sposito:1986:EVM

[ST86a] V. A. Sposito and Michael D. Tveite. On the estimation of the variance of the median used in L_{1} linear inference procedures. Communications in Statistics: Theory and Methods, 15(4):1367–1375, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1986:CSE

[ST86b] Anil K. Srivastava and Derrick S. Tracy. Computation of standard errors in seemingly unrelated regression equation models. Communications in Statistics: Theory and Methods, 15(12):3583–3588, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stavros:1986:CSD

[Sta86] Kourouklis Stavros. Characterizations of some discrete distributions based on a variant of the Rao–Rubin condition. Communications in Statistics: Theory and Methods, 15(3):839–851, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steece:1986:RSO

[Ste86a] Bert M. Steece. Regressor space outliers in ridge regression. Communications in Statistics: Theory and Methods, 15(12):3599–3605, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1986:EBB

[Ste86b] Theodor J. Stewart. Experience with a Bayesian bootstrap method incorporating proper prior information. Communications in Statistics: Theory and Methods, 15(11):3205–3225, 1986. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stefanski:1989:CDM

[Ste89a] Leonard A. Stefanski. Correcting data for measurement error in generalized linear models. Communications in Statistics: Theory and Methods, 18(5):1715–1733, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stefanski:1989:UEN

[Ste89b] Leonard A. Stefanski. Unbiased estimation of a nonlinear function of a normal mean with application to measurement error models. Communications in Statistics: Theory and Methods, 18(12):4335– 4358, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1989:EHD

[Ste89c] Paul W. Stewart. Extension of the Harrell–Davis quantile estimator to finite populations. Communications in Statistics: Theory and Methods, 18(3):853–875, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singhal:1988:MRI

[STJ88] L. C. Singhal, R. K. Tuteja, and Priti Jain. On measures of relative information with preference. Communications in Statistics: Theory and Methods, 17(5):1449–1464, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sana:1982:GOR

[StR82] G. M. Sana, H. Peso tan, and B. L. Raktoe. Globally orthogonal regular fractions of the sn factorial. Communications in Statistics: Theory and Methods, 11(1):13–30, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stroud:1984:BSE

[Str84] T. W. F. Stroud. Bayesian shrinkage estimates for regression coefficients in m populations. Communications in Statistics: Theory and Methods, 13(17):2085–2109, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Street:1988:SRM

[Str88] Deborah J. Street. Some repeated measurements designs. Communications in Statistics: Theory and Methods, 17(1):87–104, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singhal:1989:ETP

[STSM89] R. A. Singhal, C. B. Tiwari, H. Sahai, and S. C. Misra. Estimation and testing procedures for non-normal samples: an introduction and a bibliography. Communications in Statistics: Theory and Methods, 18(7):2663–2741, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stufken:1988:ELT

[Stu88] John Stufken. On the existence of linear trend free block designs. Communications in Statistics: Theory and Methods, 17(11):3857– 3863, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1986:NRE

[SU86] Radhey S. Singh and Aman Ullah. Nonparametric recursive estimation of a multivariate, marginal and conditional DGP with an application to specification of econometric models. Communications in Statistics: Theory and Methods, 15(12):3489–3513, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:1988:GRM

[Sun88] Hong-Jie Sun. A general reduction method for n-variate normal orthant probability. Communications in Statistics: Theory and Methods, 17(11):3913–3921, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suzuki:1983:SAR

[Suz83] Manabu Suzuki. Some asymptotic results in the multivariate lognormal estimation theory. Communications in Statistics: Theory and Methods, 12(15):1761–1774, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1989:DSA

[SV89] V. Soundararajan and R. Vijayaraghavan. Double sampling attributes plans indexed by AQL and AOQL. Communications in Statistics: Theory and Methods, 18(8):3033–3046, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1982:FWC

[SvW82] J. W. H. Swanepoel and J. W. J. van Wyk. Fixed width confidence intervals for the location parameter of an exponential distribution. Communications in Statistics: Theory and Methods, 11(11):1279– 1289, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singpurwalla:1983:EFR

[SW83a] Nozer D. Singpurwalla and Man-Yuen Wong. Estimation of the failure rate — a survey of nonparametric methods. Part I: Non-Bayesian methods. Communications in Statistics: Theory and Methods, 12(5):559–588, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soares:1983:SRR

[SW83b] Jose F. Soares and C. F. Jeff Wu. Some restricted randomization rules in sequential designs. Communications in Statistics: Theory and Methods, 12(17):2017–2034, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoukri:1984:EIC

[SW84a] M. M. Shoukri and R. H. Ward. On the estimation of the intraclass correlation. Communications in Statistics: Theory and Methods, 13 (10):1239–1255, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spall:1984:ADT

[SW84b] James C. Spall and Kent D. Wall. Asymptotic distribution theory for the Kalman filter state estimator. Communications in Statistics: Theory and Methods, 13(16):1981–2003, 1984. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwertman:1986:NMM

[SW86a] Neil C. Schwertman and Buck Ware. A note on the multivariate multisample rank sum and median tests. Communications in Statistics: Theory and Methods, 15(11):3245–3252, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solomon:1986:RPP

[SW86b] Herbert Solomon and Howard Weiner. A review of the packing problem. Communications in Statistics: Theory and Methods, 15 (9):2571–2607, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scariano:1988:NPE

[SW88] Stephen M. Scariano and Terry A. Watkins. Nonparametric point estimators for the change-point problem. Communications in Statistics: Theory and Methods, 17(11):3645–3675, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1989:PFL

[SW89] Bahadur Singh and F. T. Wright. The power functions of the likelihood ratio tests for a simply ordered trend in normal means. Communications in Statistics: Theory and Methods, 18(7):2351– 2392, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1980:OSB

[Swa80] Meena N. Swamy. Optimum spring balance weighing designs for estimating the total weight. Communications in Statistics: Theory and Methods, 9(11):1185–1190, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1982:UBB

[Swa82] Meena N. Swamy. Use of balanced bipartite weighing designs as chemical balance designs. Communications in Statistics: Theory and Methods, 11(7):769–785, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1983:BSP

[Swa83] J. W. H. Swanepoel. Bootstrap selection procedures based on robust estimators. Communications in Statistics: Theory and Methods, 12(18):2059–2083, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1986:NPM

[Swa86a] Jan W. H. Swanepoel. A note on proving that the (modified) bootstrap works. Communications in Statistics: Theory and Methods, 15(11):3193–3203, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1986:CND

[Swa86b] Jan W. H. Swanepoel. On the construction of nonparametric density function estimators using the bootstrap. Communications in Statistics: Theory and Methods, 15(5):1399–1415, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1987:OKW

[Swa87] Jan W. H. Swanepoel. Optimal kernels when estimating nonsmooth densities. Communications in Statistics: Theory and Methods, 16(6):1835–1848, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1988:MIS

[Swa88] Jan W. H. Swanepoel. Mean integrated squared error properties and optimal kernels when estimating a distribution function. Communications in Statistics: Theory and Methods, 17(11):3785–3799, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schezhtman:1987:MAB

[SY87] Edna Schezhtman and Shlomo Yitzhaki. A measure of association based on Gin’s mean difference. Communications in Statistics: Theory and Methods, 16(1):207–231, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seaman:1989:RED

[SYT89] Samuel L. Seaman, Dean M. Young, and Danny W. Turner. On the robustness of the extreme deviate test for a single multivariate outlier against heavy-tailed distributions. Communications in Statistics: Theory and Methods, 18(9):3289–3303, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szatrowski:1985:RBT

[Sza85] Ted H. Szatrowski. The relationship between two parameterizations of linear patterned mean and covariance models. Communications in Statistics: Theory and Methods, 14(7):1709–1712, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sparks:1985:VSM

[SZC85] R. S. Sparks, W. Zucchini, and D. Coutsourides. On variable selection in multivariate regression. Communications in Statistics: Theory and Methods, 14(7):1569–1587, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taka:1983:AMC

[TA83] M. T. Taka and P. Armitage. Autoregressive models in clinical trials. Communications in Statistics: Theory and Methods, 12(8): 865–876, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taguchi:1985:QEJ

[Tag85] Genichi Taguchi. Quality engineering in Japan. Communications in Statistics: Theory and Methods, 14(11):2785–2801, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tahir:1989:AES

[Tah89] Mohamed Tahir. Asymptotic expansions for sequential confidence levels in inverse linear regression. Communications in Statistics: Theory and Methods, 18(12):4501–4509, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takahasi:1988:NHR

[Tak88] Koiti Takahasi. A note on hazard rates of order statistics. Communications in Statistics: Theory and Methods, 17(12):4133–4136, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Talwalker:1980:CNP

[Tal80] Sheela Talwalker. On characterizing the normal and Poisson distributions. Communications in Statistics: Theory and Methods, 9 (3):347–354, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Talwaiker:1986:FEC

[Tal86] Sheela Talwaiker. Functional equations in characterizations of discrete distributions by Rao–Rubin condition and its variants. Communications in Statistics: Theory and Methods, 15(3):961–979, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tamhane:1986:SLQ

[Tam86] Ajit C. Tamhane. A survey of literature on quantal response curves with a view toward application to the problem of selecting the curve with the smallest q-quantile (EDIOOq). Communications in Statistics: Theory and Methods, 15(9):2679–2718, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1982:CSS

[Tan82] W. Y. Tan. On comparing several straight lines under heteroscedasticity and robustness with respect to departure from normality. Communications in Statistics: Theory and Methods, 11(7):731– 750, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1988:SAP

[Tan88] Yukata Tanaka. Sensitivity analysis in principal component analysis: influence on the subspace spanned by principal components. Communications in Statistics: Theory and Methods, 17(9):3157– 3175, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1989:IFR

[Tan89] Yutaka Tanaka. Influence functions related to eigenvalue problems which appear in multivariate analysis. Communications in Statistics: Theory and Methods, 18(11):3991–4010, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tarone:1986:CTT

[Tar86a] Robert E. Tarone. Correctings tests for trend in proportions for skewness. Communications in Statistics: Theory and Methods, 15 (2):317–328, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tarone:1986:TCS

[Tar86b] Robert E. Tarone. Tests for comparing a sequence of Bernoulli random variables to a sequence of known probabilities. Communications in Statistics: Theory and Methods, 15(3):981–997, 1986. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tarone:1988:HST

[Tar88] Robert E. Tarone. Homogeneity score tests with nuisance parameters. Communications in Statistics: Theory and Methods, 17(5): 1549–1556, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1982:URE

[TB82] M. L. Tiku and Punara Bhasin. Usefulness of robust estimators in sample survey. Communications in Statistics: Theory and Methods, 11(22):2597–2610, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1984:RMC

[TB84a] M. L. Tiku and N. Balakrishnan. Robust multivariate classification procedures based on the MML estimators. Communications in Statistics: Theory and Methods, 13(8):967–986, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1984:TEP

[TB84b] M. L. Tiku and N. Balakrishnan. Testing equality of population variances the robust way. Communications in Statistics: Theory and Methods, 13(17):2143–2159, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1985:TEV

[TB85] M. L. Tiku and N. Balakrishnan. Testing the equality of variancecovariance matrices the robust way. Communications in Statistics: Theory and Methods, 14(12):3033–3051, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1988:BAR

[TB88a] W. Y. Tan and N. Balakrishnan. Bayesian approach to robust procedures based on censored samples. Communications in Statistics: Theory and Methods, 17(7):2403–2418, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1988:GRB

[TB88b] M. L. Tiku and N. Balakrishnan. Generalization of the robust bivariate t^{2} statistic to multivariate populations. Communications in Statistics: Theory and Methods, 17(11):3899–3911, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1988:RHT

[TB88c] M. L. Tiku and N. Balakrishnan. Robust Hotelling-type T^{2} statistics based on the modified maximum likelihood estimators. Communications in Statistics: Theory and Methods, 17(6):1789–1810, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1989:RCP

[TB89] M. L. Tiku and N. Balakrishnan. Robust classification procedures based on the MML estimators. Communications in Statistics: Theory and Methods, 18(3):1047–1066, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1981:CBF

[TC81] W. Y. Tan and S. S. Cheng. Cumulants of bilinear forms and quadratic forms and their applications. Communications in Statistics: Theory and Methods, 10(3):283–298, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1989:BEM

[TCJ89] Ram C. Tiwari, Siddhartha Chib, and S. Rao Jammalamadaka. Bayes estimation of the multiple correlation coefficient. Communications in Statistics: Theory and Methods, 18(4):1401–1413, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teather:1984:EMD

[Tea84a] D. Teather. Exchangeability and medical diagnosis. Communications in Statistics: Theory and Methods, 13(3):327–352, 1984. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teather:1984:EEB

[Tea84b] Derck Teather. The estimation of exchangeable binomial parameters. Communications in Statistics: Theory and Methods, 13(6): 671–680, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terasvirta:1981:CMM

[Ter81] Timo Teräsvirta. A comparison of mixed and minimax estimators of linear models. Communications in Statistics: Theory and Methods, 10(17):1765–1778, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terasvirta:1982:SCH

[Ter82] Timo Teräsvirta. Superiority comparisons of homogeneous linear estimators. Communications in Statistics: Theory and Methods, 11 (14):1595–1601, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terasvirta:1986:SCH

[Ter86] Timo Teräsvirta. Superiority comparisons of heterogeneous linear estimators. Communications in Statistics: Theory and Methods, 15 (4):1319–1336, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terasvirta:1988:SCB

[Ter88] Timo Teräsvirta. Superiority comparisons between mixed regression estimators. Communications in Statistics: Theory and Methods, 17(10):3537–3546, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrade:1984:MLE

[tFAH84] Dal ton F. Andrade and Ronald W. Helms. Maximum likelihood estimates in the multivariate normal with patterned mean and covariance via the em algorithm. Communications in Statistics: Theory and Methods, 13(18):2239–2251, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1983:ALC

[TG83] W. Y. Tan and R. P. Gupta. On approximating a linear combination of central Wishart matrices with positive coefficients. Communications in Statistics: Theory and Methods, 12(22):2589–2600, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1985:GLS

[TG85] Ram C. Tripathi and Ramesh C. Gupta. A generalization of the log-series distribution. Communications in Statistics: Theory and Methods, 14(8):1779–1799, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1986:UAE

[TG86] Ram C. Tripathi and John Gurland. A unified approach to estimating parameters in some generalized Poisson distributions. Communications in Statistics: Theory and Methods, 15(3):1017–1034, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripatni:1987:CBO

[TG87] Ram C. Tripatni and Ramesh C. Gupta. A comparison between the ordinary and the length-biased modified power series distributions with applications. Communications in Statistics: Theory and Methods, 16(4):1195–1206, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1989:MML

[TG89] M. L. Tiku and P. S. Gill. Modified maximum likelihood estimators for the bivariate normal based on Type II censored samples. Communications in Statistics: Theory and Methods, 18(9):3505–3518, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1989:RPT

[TGB89] M. L. Tiku, P. S. Gill, and N. Balakrishnan. A robust procedure for testing the equality of mean vectors of two bivariate populations with unequal covariance matrices. Communications in Statistics: Theory and Methods, 18(9):3249–3265, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1986:IMM

[TGG86] Ram C. Tripathi, Pushpa L. Gupta, and Ramesh C. Gupta. Incomplete moments of modified power series distributions with applications. Communications in Statistics: Theory and Methods, 15 (3):999–1015, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1982:RSTb

[Tik82] M. L. Tiku. Robust statistics for testing equality of means or variances. Communications in Statistics: Theory and Methods, 11 (22):2543–2558, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1988:MML

[Tik88a] M. L. Tiku. Modified maximum likelihood estimation for the bivariate normal. Communications in Statistics: Theory and Methods, 17(3):893–910, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1988:OSG

[Tik88b] M. L. Tiku. Order statistics in goodness of fit tests. Communications in Statistics: Theory and Methods, 17(7):2369–2387, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Titterington:1983:KBD

[Tit83] D. M. Titterington. Kernel-based density estimation using censored, truncated or grouped data. Communications in Statistics: Theory and Methods, 12(18):2151–2167, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1988:CDI

[Tiw88] Ram Chandra Tiwari. Convergence of Dirichlet invariant measures and the limits of Bayes estimates. Communications in Statistics: Theory and Methods, 17(2):357–393, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1988:BEB

[TJZ88] R. C. Tiwari, S. Rao Jammalamadaka, and Jyoti N. Zalkikar. Bayes and empirical Bayes estimation of survival function under progressive censoring. Communications in Statistics: Theory and Methods, 17(10):3591–3606, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1984:EQF

[tL84] Chun tu Lin. Extrema of quadratic forms and statistical applications. Communications in Statistics: Theory and Methods, 13(12): 1517–1520, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1989:EBE

[TL89a] Ram C. Tiwari and Parthasarathi Lahiri. Empirical Bayes estimation of the mean of a “nonhomogeneous” finite population. Communications in Statistics: Theory and Methods, 18(7):2553–2568, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1989:REB

[TL89b] Ram C. Tiwari and Parthasarathi Lahiri. On robust empirical Bayes analysis of means and variances from stratified samples. Communications in Statistics: Theory and Methods, 18(3):913– 926, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tlku:1983:EES

[Tlk83] M. L. Tlku. Exact efficiencies of some robust estimators in sample survey. Communications in Statistics: Theory and Methods, 12 (17):2043–2051, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tolley:1985:NHT

[TO85] H. D. Tolley and A. R. Olsen. A note on homogeneity tests when combining 2 × 2 tables. Communications in Statistics: Theory and Methods, 14(12):2857–2871, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1989:SAM

[TO89] Yutaka Tanaka and Yoshimasa Odaka. Sensitivity analysis in maximum likelihood factor analysis. Communications in Statistics: Theory and Methods, 18(11):4067–4084, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tomizawa:1987:DWM

[Tom87] Sadao Tomizawa. Diagonal, weighted marginal homogeneity model and decompositions for linear diagonals-parameter symmetry model. Communications in Statistics: Theory and Methods, 16 (2):477–488, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tse:1985:THD

[TP85] Y. K. Tse and C. T. Phoon. Testing for heteroscedasticity in a dynamic simultaneous equation model. Communications in Statistics:

Theory and Methods, 14(6):1283–1300, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taillie:1986:FDR

[TP86] C. Taillie and G. P. Patil. The Fibonacci distribution revisited. Communications in Statistics: Theory and Methods, 15(3):951– 959, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trader:1983:BAI

[Tra83a] Ramona L. Trader. A Bayesian analysis of imperfect inspection models. Communications in Statistics: Theory and Methods, 12 (4):397–408, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trader:1983:BPA

[Tra83b] Ramona L. Trader. A Bayesian predictive approach to the selection of variables in multiple regression. Communications in Statistics: Theory and Methods, 12(13):1553–1557, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1984:MSM

[Tra84] Derrick S. Tracy. Moments of sample moments. Communications in Statistics: Theory and Methods, 13(5):553–562, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trader:1985:BIT

[Tra85] Ramona L. Trader. Bayesian inference for truncated exponential distributions. Communications in Statistics: Theory and Methods, 14(3):585–592, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trenkler:1980:GMS

[Tre80] G. Trenkler. Generalized mean squared error comparisons of biased regression estimators. Communications in Statistics: Theory and Methods, 9(12):1247–1259, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trenkler:1983:NSC

[Tre83] G. Trenkler. A note on superiority comparisons of homogeneous linear estimators. Communications in Statistics: Theory and Methods, 12(7):799–808, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trenkler:1988:SRR

[Tre88] Gotz Trenkler. Some remarks on a ridge-type-estimator and good prior means. Communications in Statistics: Theory and Methods, 17(12):4251–4256, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tritchler:1988:ENA

[Tri88] David L. Tritchler. The exact nonparametric analysis of a class of experimental designs. Communications in Statistics: Theory and Methods, 17(5):1351–1363, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1981:RTM

[TS81a] M. L. Tiku and M. Singh. Robust test for means when population variances are unequal. Communications in Statistics: Theory and Methods, 10(20):2057–2071, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1981:TTP

[TS81b] M. L. Tiku and M. Singh. Testing the two parameter Weibull distribution. Communications in Statistics: Theory and Methods, 10(9):907–918, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1982:RSTa

[TS82] M. L. Tiku and M. Singh. Robust statistics for testing mean vectors of multivariate distributions. Communications in Statistics: Theory and Methods, 11(9):985–1001, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1988:SSR

[TS88] Derrick S. Tracy and Anil K. Srivastava. A synthesis of Steinrule and mixed regression procedures in linear regression models under non-normality. Communications in Statistics: Theory and Methods, 17(3):691–704, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tseng:1988:SMR

[Tse88] Sheng-Tsaing Tseng. Selecting more reliable Weibull populations than a control. Communications in Statistics: Theory and Methods, 17(1):169–181, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tse:1989:CPG

[Tse89] Siu-Keung Tse. A comparison of procedures for general contrasts — unequal sample size case. Communications in Statistics: Theory and Methods, 18(2):613–632, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thurman:1986:EDL

[TSM86] S. S. Thurman, P. A. V. B. Swamy, and J. S. Mehta. An examination of distributed lag model coefficients estimated with smoothness priors. Communications in Statistics: Theory and Methods, 15 (6):1723–1749, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsurumi:1985:LIB

[Tsu85] Hiroki Tsurumi. Limited information Bayesian analysis of a structural coefficient in a simultaneous equations system. Communications in Statistics: Theory and Methods, 14(5):1103–1120, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tidmore:1983:CCT

[TT83] Eugene F. Tidmore and Danny W. Turner. On clustering with Chernoff-type faces. Communications in Statistics: Theory and Methods, 12(4):381–396, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trenkler:1984:EDB

[TT84] Dietrich Trenkler and Götz Trenkler. On the Euclidean distance between biased estimators. Communications in Statistics: Theory and Methods, 13(3):273–284, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1985:EMR

[TT85] D. S. Tracy and I. H. Tajuddin. Extended moment results for improving inferences based on MRPP. Communications in Statistics: Theory and Methods, 14(6):1485–1496, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1986:EPC

[TT86] T. S. Tracy and I. H. Tajuddin. Empirical power comparisons of two MRPP rank tests. Communications in Statistics: Theory and Methods, 15(2):551–570, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1988:HMAa

[TT88] W. I. Tan and M. A. Tabatabai. Harmonic mean approach to unbalanced mixed effects models under heteroscedastic variances. Communications in Statistics: Theory and Methods, 17(4):1231– 1259, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1989:NBE

[TT89a] Ram C. Tiwari and Ram C. Tripathi. Nonparametric Bayes estimation of the probability of discovering a new species. Communications in Statistics: Theory and Methods, 18(3):877–895, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trybus:1989:DRV

[TT89b] Elzbieta Trybus and Ginter Trybus. The distance random variable and its applications. Communications in Statistics: Theory and Methods, 18(3):1085–1098, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1988:HMAb

[TTB88] W. Y. Tan, M. A. Tabatabai, and N. Balakrishnan. Harmonic mean approach to unbalanced random effects models under heteroscedasticity. Communications in Statistics: Theory and Methods, 17(4):1261–1286, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tubbs:1980:ESC

[Tub80] J. D. Tubbs. The effect of serial correlation on confidence regions for the parameters of a multivariate normal population. Communications in Statistics: Theory and Methods, 9(13):1341–1351, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tubbs:1986:MRC

[Tub86] J. D. Tubbs. Moments for a ratio of correlated gamma variates. Communications in Statistics: Theory and Methods, 15(1):251– 259, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tucker:1982:PTR

[Tuc82] W. T. Tucker. On the Padé table and its relationship to the r and s arrays and ARMA modeling. Communications in Statistics: Theory and Methods, 11(12):1335–1379, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tuncer:1985:EPI

[Tun85] Yaçin Tuncer. Equivalence of parametric identifiability and estimability. Communications in Statistics: Theory and Methods, 14 (2):393–400, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tutz:1989:CVD

[Tut89] Gerhard Tutz. On cross-validation for discrete kernel estimates in discrimination. Communications in Statistics: Theory and Methods, 18(11):4145–4162, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takahashi:1981:AEN

[TW81] Hajime Takahashi and Michael Woodroofe. Asymptotic expansions in non-linear renewal theory. Communications in Statistics: Theory and Methods, 10(21):2113–2135, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Turner:1986:NRL

[TYM86] Danny W. Turner, Dean M. Young, and Virgil R. Marc. A note on the robustness of the Lilliefors test for univariate normality with respect to equicorrelated data. Communications in Statistics: Theory and Methods, 15(8):2355–2361, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1985:EBE

[TZ85] R. C. Tiwari and J. N. Zalkikar. Empirical Bayes estimation of functionals of unknown probability measures. Communications in Statistics: Theory and Methods, 14(12):2963–2995, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Umbach:1981:IMP

[Umb81] Dale Umbach. On inference for a mixture of a Poisson and a degenerate distribution. Communications in Statistics: Theory and Methods, 10(3):299–306, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Urzua:1988:CME

[Urz88] Carlos Urzúa. A class of maximum-entropy multivariate distributions. Communications in Statistics: Theory and Methods, 17(12): 4039–4057, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Utts:1982:RTL

[Utt82] Jessica M. Utts. The rainbow test for lack of fit in regression. Communications in Statistics: Theory and Methods, 11(24):2801– 2815, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ullah:1984:IRS

[UV84] Aman Ullah and H. D. Vinod. Improvement ranges for shrinkage estimators with stochastic target. Communications in Statistics: Theory and Methods, 13(2):207–215, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vadiveloo:1983:TMR

[Vad83] Jeyaraj Vadiveloo. On the theory of modified randomization tests for nonparametric hypotheses. Communications in Statistics: Theory and Methods, 12(14):1581–1596, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Valliant:1986:MSE

[Val86] Richard Valliant. Mem squared error estimation in finite populations under nonlinear models. Communications in Statistics: Theory and Methods, 15(6):1975–1993, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vidmar:1980:OAF

[VB80] T. J. Vidmar and M. N. Brunden. Optimal allocation with fixed power in a completely randomized design with levels of subsampling. Communications in Statistics: Theory and Methods, 9(7): 757–763, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanderMerwe:1987:LGV

[vdMN87] Cay A. van der Merwe and D. G. Nel. Linear generalizations of various matric-t distributions. Communications in Statistics: Theory and Methods, 16(12):3665–3681, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Venezia:1980:OAI

[Ven80] Itzhak Venezia. Optimal aggregation of input-output coefficients — a decision theoretic approach. Communications in Statistics: Theory and Methods, 9(5):557–570, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Venter:1988:EMS

[Ven88] J. H. Venter. Estimation of the mean of the selected population. Communications in Statistics: Theory and Methods, 17(3):791– 805, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verdooren:1988:ETC

[Ver88a] L. R. Verdooren. Exact tests and confidence intervals for ratio of variance components in unbalanced two-and three-stage nested designs. Communications in Statistics: Theory and Methods, 17 (4):1197–1230, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verdooren:1988:LSE

[Ver88b] L. R. Verdooren. Least squares estimators and non-negative estimators of variance components. Communications in Statistics: Theory and Methods, 17(4):1027–1051, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vecchia:1989:EDF

[VI89] Dominic F. Vecchia and Hari K. Iyer. Exact distribution-free tests for equality of several linear models. Communications in Statistics: Theory and Methods, 18(7):2467–2488, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viana:1982:CEC

[Via82] Marios A. G. Viana. Combined estimators for the correlation coefficient. Communications in Statistics: Theory and Methods, 11 (13):1483–1504, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinod:1984:FSR

[Vin84] H. D. Vinod. Fordham symposium on ridge regression: introduction. Communications in Statistics: Theory and Methods, 13(2): v, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verrill:1988:ADQ

[VJ88] Steve Verrill and Richard A. Johnson. Asymptotic distributions for quadratic forms with applications to censored data tests of fit. Communications in Statistics: Theory and Methods, 17(12):4011– 4024, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vere-Jones:1981:SS

[VJS81] D. Vere-Jones and W. G. C. Smith. Statistics in seismology. Communications in Statistics: Theory and Methods, 10(15):1559–1585, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vogt:1986:NID

[Vog86] Arthur Vogt. The natural index and the Divisia index on the straight line as well as the Divisia index on the exponential line. Communications in Statistics: Theory and Methods, 15(12):3567– 3582, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vogt:1987:SSC

[Vog87] Arthur Vogt. Some suggestions concerning an axiom system for statistical price and quantity indices. Communications in Statistics: Theory and Methods, 16(12):3641–3663, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vogt:1989:NID

[Vog89] Arthur Vogt. Numerical integration of the Divisia price index and an application of the equation of measurement to prices. Communications in Statistics: Theory and Methods, 18(5):1943–1959, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vonesh:1983:ERM

[Von83] Edward F. Vonesh. Efficiency of repeated measures designs versus completely randomized designs based on multiple comparisons. Communications in Statistics: Theory and Methods, 12(3):289– 301, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Voss:1986:GCM

[Vos86] D. T. Voss. On generalizations of the classical method of confounding to asymmetric factorial experiments. Communications in Statistics: Theory and Methods, 15(4):1299–1314, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Voss:1988:GMR

[Vos88] Daniel T. Voss. Generalized modulus-ratio tests for analysis of factorial designs with zero degrees of freedom for error. Communications in Statistics: Theory and Methods, 17(10):3345–3359, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vardeman:1981:ERB

[VS81] Stephen Vardeman and Ashok Singh. Empirical restricted Bayes estimation in a muitivariate discrete exponential family. Communications in Statistics: Theory and Methods, 10(1):79–100, 1981. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viveros:1986:CIM

[VS86] R. Viveros and D. A. Sprott. Conditional inference and maximum likelihood in a capture–recapture model. Communications in Statistics: Theory and Methods, 15(3):1035–1046, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vellaisamy:1988:EMSa

[VS88] P. Vellaisamy and D. Sharma. Estimation of the mean of the selected gamma population. Communications in Statistics: Theory and Methods, 17(8):2797–2817, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vellaisamy:1989:NEM

[VS89] P. Vellaisamy and D. Sharma. A note on the estimation of the mean of the selected gamma population. Communications in Statistics: Theory and Methods, 18(2):555–560, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vellaisamy:1988:EMSb

[VSD88] P. Vellaisamy, Kumar Somesh, and Sharma Divakar. Estimating the mean of the selected uniform population. Communications in Statistics: Theory and Methods, 17(10):3447–3475, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Villalobos:1983:MTP

[VW83] Miguel A. Villalobos and Grace Wahba. Multivariate thin plate spline estimates for the posterior probabilities in the classification problem. Communications in Statistics: Theory and Methods, 12 (13):1449–1479, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:1981:FCT

[WA81] William W. S. Wei and Bovas Abraham. Forecasting contemporal time series aggregates. Communications in Statistics: Theory and Methods, 10(13):1335–1344, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1984:SUP

[WA84] M. C. Wang and V. Aggarwal. Stratification under a particular Pareto distribution. Communications in Statistics: Theory and Methods, 13(6):711–735, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wahba:1981:NET

[Wah81] Grace Wahba. Numerical experiments with the thin plate histospline. Communications in Statistics: Theory and Methods, 10 (24):2475–2514, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walker:1980:EPT

[Wal80] Homer F. Walker. Estimating the proportions of two populations in a mixture using linear maps. Communications in Statistics: Theory and Methods, 9(8):837–849, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Waldman:1984:MCT

[Wal84a] Donald M. Waldman. The mean of the conditional, truncated multinormal distribution. Communications in Statistics: Theory and Methods, 13(21):2679–2689, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walter:1984:NCR

[Wal84b] Shur Walter. The negative contagion reflection of the Pólya– Eggenberger distribution. Communications in Statistics: Theory and Methods, 13(7):877–885, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wallenstein:1987:CID

[Wal87] Sylvan Wallenstein. Confidence intervals for dose or treatment effect in several 2×j tables. Communications in Statistics: Theory and Methods, 16(7):2083–2096, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walker:1989:DCI

[Wal89] Esteban Walker. Detection of collinearity- influential observations. Communications in Statistics: Theory and Methods, 18(5):1675– 1690, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1983:AMR

[Wan83] C. Ming Wang. On the analysis of multivariate repeated measures designs. Communications in Statistics: Theory and Methods, 12 (14):1647–1659, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1986:PVE

[Wan86] M. C. Wang. A prior-valued estimator applied to multinomial classification. Communications in Statistics: Theory and Methods, 15(2):405–427, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1987:PLE

[Wan87a] Mel Cheng Wang. Product limit estimates: a generalized maximum likelihood study. Communications in Statistics: Theory and Methods, 16(11):3117–3132, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1987:STI

[Wan87b] P. C. Wang. Simultaneous transformations and influence in regression. Communications in Statistics: Theory and Methods, 16(11): 3417–3425, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1987:IAT

[Wan87c] Suojin Wang. Improved approximation for transformation diagnostics. Communications in Statistics: Theory and Methods, 16 (6):1797–1819, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1988:REW

[Wan88] C. Ming Wang. A robust estimator for wall following. Communications in Statistics: Theory and Methods, 17(2):411–422, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1989:EAS

[Wan89a] C. Ming Wang. Error analysis of spatial representation and estimation of mobile robots. Communications in Statistics: Theory and Methods, 18(3):1107–1122, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1989:BSB

[Wan89b] Suojin Wang. On the bootstrap and smoothed bootstrap. Communications in Statistics: Theory and Methods, 18(11):3949–3962, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2009:CEW

[Wan09] Zhu Wang. Correction to “Estimation of the Weibull Parameters with Grouped Data”. Communications in Statistics: Theory and Methods, 38(20):3755, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [CC88].

Whitten:1988:PCS

[WAS88] Betty Jones Whitten, Clifford Cohen A., and Vani Sundaraiyer. A pseudo-complete sample technique for estimation from censored samples. Communications in Statistics: Theory and Methods, 17 (7):2239–2258, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watson:1987:RGS

[Wat87] Geoffrey S. Watson. Reflections on group screening. Communications in Statistics: Theory and Methods, 16(10):2795–2798, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watson:1980:EST

[WB80] Dan A. Watson and Saul Blumenthal. Estimating the size of a truncated sample. Communications in Statistics: Theory and Methods, 9(15):1535–1550, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1987:SRC

[WC87] Song-Gui Wang and Shein-Chung Chow. Some results on canonical correlations and measures of multivariate association. Communications in Statistics: Theory and Methods, 16(2):339–351, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wheeler:1985:WCP

[WCK85] Noel C. Wheeler, Bonnie H. Corre, and Anant M. Kshirsagar. The weighted compounding of probabilities from three independent hypothesis tests. Communications in Statistics: Theory and Methods, 14(5):1163–1174, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wyckoff:1980:CRL

[WE80] James Wyckoff and Max Engelhardt. Cramér–Rao lower bounds for estimators based on censored data. Communications in Statistics: Theory and Methods, 9(13):1385–1399, 1980. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walter:1980:MLE

[WEA80] W. Stroup Walter, James W. Evans, and R. L. Anderson. Maximum likelihood estimation of variance components in a completely random bib design. Communications in Statistics: Theory and Methods, 9(7):725–756, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wechsler:1985:MTD

[Wec85] Sergio Wechsler. Mood’s test for dispersion as a counting of triplets. Communications in Statistics: Theory and Methods, 14(2):303– 306, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wefelmeyer:1987:UAL

[Wef87] W. Wefelmeyer. Uniform approximation of log-likelihood ratios in the i.i.d. case. Communications in Statistics: Theory and Methods, 16(5):1265–1280, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissman:1980:ETP

[Wei80] Ishay Weissman. Estimation of tail parameters under Type I censoring. Communications in Statistics: Theory and Methods, 9(11):

1165–1175, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:1981:ESS

[Wei81a] William W. S. Wei. Effect of systematic sampling on ARIMA models. Communications in Statistics: Theory and Methods, 10 (23):2389–2398, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weier:1981:BEB

[Wei81b] D. R. Weier. Bayes estimation for a bivariate survival model based on exponential distributions. Communications in Statistics: Theory and Methods, 10(14):1415–1427, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissman:1981:CIT

[Wei81c] Ishay Weissman. Confidence intervals for the threshold parameter. Communications in Statistics: Theory and Methods, 10(6):549– 557, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissman:1982:CIT

[Wei82] Ishay Weissman. Confidence intervals for the threshold parameter. II: unknown shape parameter. Communications in Statistics: Theory and Methods, 11(21):2451–2474, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1989:MDM

[Wei89] Lisa A. Weissfeld. A multicollinearity diagnostic for models fit to censored data. Communications in Statistics: Theory and Methods, 18(6):2073–2085, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wallenstein:1988:TPC

[WF88] Sylvan Wallenstein and Joseph L. Fleiss. The two-period crossover design with baseline measurements. Communications in Statistics: Theory and Methods, 17(10):3333–3343, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1981:CIR

[WG81] Chih-Ming Wang and Franklin A. Graybill. Confidence intervals on a ratio of variances in the two-factor nested components of variance model. Communications in Statistics: Theory and Methods, 10 (14):1357–1368, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walter:1987:EBE

[WH87] G. G. Walter and G. G. Hamedani. Empiric Bayes estimation of binomial probability. Communications in Statistics: Theory and Methods, 16(2):559–577, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wun:1988:NDM

[WH88] Lap-Ming Wun and Ken Hung. A note on the distribution of mean square error of forecasting. Communications in Statistics: Theory and Methods, 17(6):1929–1934, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Whaley:1987:OWW

[Wha87] Fredrick S. Whaley. Optimizing the Wald–Wolfowitz runs statistic using a linkage tolerance: guidelines based on computer simulation. Communications in Statistics: Theory and Methods, 16(7):2125– 2138, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Whitney:1989:AIT

[Whi89a] Paul Whitney. Approximation and information topologies. Communications in Statistics: Theory and Methods, 18(8):2909–2933, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Whittinghill:1989:NTF

[Whi89b] Dexter C. Whittinghill III. A note on trend-free block designs. Communications in Statistics: Theory and Methods, 18(1):277– 285, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wiens:1985:ECR

[Wie85] Doug Wiens. Ellipsoidal confidence regions for a normal covariance matrix. Communications in Statistics: Theory and Methods, 14 (6):1405–1411, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wijsman:1981:CSB

[Wij81] Robert A. Wijsman. Confidence sets based on sequential tests. Communications in Statistics: Theory and Methods, 10(21):2137– 2147, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

William:1981:MDE

[Wil81] C. Parr William. Minimum distance estimation: a bibliography. Communications in Statistics: Theory and Methods, 10(12):1205– 1224, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1986:ADT

[Wil86] Jeffrey R. Wilson. Approximate distribution and test of fit for the clustering effect in the Dirichlet multinomial model. Communications in Statistics: Theory and Methods, 15(4):1235–1249, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

William:1988:EMD

[Wil88a] M. Bolstad William. Estimation in the multiprocess dynamic generalized linear model. Communications in Statistics: Theory and Methods, 17(12):4179–4204, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1988:MLE

[Wil88b] P. David Wilson. Maximum likelihood estimation using differences in an autoregressive-1 process. Communications in Statistics: Theory and Methods, 17(1):17–26, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wingo:1982:UPD

[Win82] Dallas R. Wingo. Unimodality of the Pareto distribution likelihood function for multicensored samples and implications for estimation. Communications in Statistics: Theory and Methods, 11(10):1129– 1138, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Winter:1987:NEC

[Win87] B. B. Winter. Nonparametric estimation with censored data from a distribution with nonincreasing density. Communications in Statistics: Theory and Methods, 16(1):93–120, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wittig:1984:GOU

[Wit84] Timothy A. Wittig. A generalized Ornstein–Uhlenbeck process. Communications in Statistics: Theory and Methods, 13(1):29–43, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1987:BRT

[Wit87a] C. S. Withers. Bias reduction by Taylor series. Communications in Statistics: Theory and Methods, 16(8):2369–2383, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1987:UAO

[Wit87b] C. S. Withers. Unbiased asymptotically optimal estimates for the problem of the Nile. Communications in Statistics: Theory and Methods, 16(8):2361–2367, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1988:SAS

[Wit88] C. S. Withers. Some asymptotics for U-statistics. Communications in Statistics: Theory and Methods, 17(10):3269–3276, 1988. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1989:ACI

[Wit89a] C. S. Withers. Accurate confidence intervals when nuisance parameters are present. Communications in Statistics: Theory and Methods, 18(11):4229–4259, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1989:PRD

[Wit89b] C. S. Withers. The probability of a random deviation from a sample mean. Communications in Statistics: Theory and Methods, 18(1): 35–52, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wells:1989:TGF

[WJT89] Martin T. Wells, S. Rao Jammalamadaka, and Ram C. Tiwari. A test of goodness-of-fit based on extreme multinomial cell frequencies. Communications in Statistics: Theory and Methods, 18(4): 1527–1547, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wheeler:1981:UBE

[WK81] Noel C. Wheeler and Anant M. Kshirsagar. Uniformly better estimators with application in two-way designs. Communications in Statistics: Theory and Methods, 10(7):699–711, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wansbeek:1982:SWO

[WK82] Tom Wansbeek and Arie Kapteyn. A simple way to obtain the spectral decomposition of variance components models for balanced data. Communications in Statistics: Theory and Methods, 11 (18):2105–2112, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wheeler:1986:ATT

[WK86] Noel C. Wheeler and Anant M. Kshirsagar. Adjustments to the treatment total when inter-row and inter-column information is recovered. Communications in Statistics: Theory and Methods, 15(1):49–58, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woolson:1980:GCA

[WL80] Robert F. Woolson and James D. Leeper. Growth curve analysis of complete and incomplete longitudinal data. Communications in Statistics: Theory and Methods, 9(14):1491–1513, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

White:1982:GCD

[WL82] Andrew A. White and J. Richard Landis. A general categorical data methodology for evaluating medical diagnostic tests. Communications in Statistics: Theory and Methods, 11(5):567–605, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1987:CIP

[WM87] Chi Song Wong and J. F. McDonald. Confidence intervals for a population size based on a capture-recapture model. Communications in Statistics: Theory and Methods, 16(1):79–91, 1987. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wegman:1980:TSA

[WNK80] Edward J. Wegman, El-Sayed Nour, and Cris Kukuk. A time series approach to life table construction. Communications in Statistics:

Theory and Methods, 9(15):1587–1607, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wolfe:1980:CTA

[Wol80] Robert A. Wolfe. A circular test of association between a point process and a continuous time series. Communications in Statistics: Theory and Methods, 9(12):1299–1307, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wolfe:1981:TDS

[Wol81] Robert A. Wolfe. Testing disease–survival dependence. Communications in Statistics: Theory and Methods, 10(9):919–926, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1982:APB

[Won82] M. Anthony Wong. Asymptotic properties of bivariate k-means clusters. Communications in Statistics: Theory and Methods, 11 (10):1155–1171, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1988:ANP

[Won88] George Y. Wong. Asymptotic normality of poly-t densities. Communications in Statistics: Theory and Methods, 17(5):1613–1627, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1989:LMG

[Won89] Chi Song Wong. Linear models in a general parametric form. Communications in Statistics: Theory and Methods, 18(8):3095–3115, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woodroofe:1980:BRI

[Woo80] Michael Woodroofe. On the Bayes risk incurred by using asymptotic shapes. Communications in Statistics: Theory and Methods, 9 (16):1727–1748, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wood:1988:SNF

[Woo88] Andrew T. A. Wood. Some notes on the Fisher–Bingham family on the sphere. Communications in Statistics: Theory and Methods, 17 (11):3881–3897, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weideman:1987:NAH

[WR87] C. A. Weideman and D. Raghavarao. Non-adaptive hypergeometric group testing designs for identifying at most two defectives. Communications in Statistics: Theory and Methods, 16(10):2991–3006, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wright:1983:BAS

[Wri83] Tommy Wright. On Bayes allocation of the sample for estimation of the mean when each stratum has a Poisson distribution. Communications in Statistics: Theory and Methods, 12(4):409– 426, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wright:1988:NOA

[Wri88a] Tommy Wright. A note on optimal allocation and an ordering of the family of beta prior distributions. Communications in Statistics: Theory and Methods, 17(1):123–138, 1988. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wright:1988:SPV

[Wri88b] Tommy Wright. On some properties of variable size simple random sampling and a limit theorem. Communications in Statistics: Theory and Methods, 17(9):2997–3016, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williford:1982:BEC

[WS82] W. O. Williford and W. Jan Shan. Bayesian estimation of the complete sample size from an incomplete Poisson sample. Communications in Statistics: Theory and Methods, 11(7):835–846, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walker:1984:ULC

[WS84] Gail Walker and Richard Sigman. The use of LANDSAT for county estimates of crop areas: evaluation of the Huddleston–Ray and the Battese–Fuller estimators for the case of stratified sampling. Communications in Statistics: Theory and Methods, 13(23):2975– 2996, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1987:IBB

[WS87] Lisa A. Weissfeld and Helmut Schneider. Inferences based on the Buckley–James procedure. Communications in Statistics: Theory and Methods, 16(6):1773–1787, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1989:TMC

[WS89] Dong Qian Wang and David J. Scott. Testing a Markov chain for independence. Communications in Statistics: Theory and Methods, 18(11):4085–4103, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Waikar:1984:TSS

[WSR84] V. B. Waikar, F. J. Schuurmann, and T. E. Raghunathan. On a two stage shrinkage testimator of the mean of a normal distribution. Communications in Statistics: Theory and Methods, 13(15):1901– 1913, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

William:1983:BTI

[WT83] Notz I. William and Ajit C. Tamhane. Balanced treatment incomplete block (BTIB) designs for comparing treatments with a control: minimal complete sets of generator designs for k = 3, p = 3(1)10. Communications in Statistics: Theory and Methods, 12(12):1391–1412, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wells:1989:BQP

[WT89] Martin T. Wells and Ram C. Tiwari. Bayesian quantile plots and statistical inference for nonlinear models in the two sample case with incomplete data. Communications in Statistics: Theory and Methods, 18(8):2955–2964, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1982:RTI

[Wu82] Shu-Chen Wu. Rank tests for independence based on partially right-censored pairs. Communications in Statistics: Theory and Methods, 11(19):2207–2216, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1988:DAW

[Wu88] Yuehua Wu. Discrimination analysis when the variates are grouped and observed in sequential order. Communications in Statistics: Theory and Methods, 17(12):4059–4074, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1989:CMS

[Wu89] Yuehua Wu. On a class of model selection procedures. Communications in Statistics: Theory and Methods, 18(9):3267–3287, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1984:BES

[WW84] Lisa A. Weissfeld and H. Samuel Wieand. Bounds on efficiencies for some two-sample nonparametric statistics. Communications in Statistics: Theory and Methods, 13(14):1741–1757, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1987:BAB

[WYF87] Linda J. Wilson, J. H. Young, and J. Leroy Folks. A biological application of Bose–Einstein statistics. Communications in Statistics: Theory and Methods, 16(2):445–459, 1987. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:1983:HFL

[Xek83a] Evdokia Xekalaki. Hazard functions and life distributions in discrete time. Communications in Statistics: Theory and Methods, 12 (21):2503–2509, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:1983:PYD

[Xek83b] Evdokia Xekalaki. A property of the Yule distribution and its applications. Communications in Statistics: Theory and Methods, 12 (10):1181–1189, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:1986:MGW

[Xek86] Evdokia Xekalaki. The multivariate generalized Waring distribution. Communications in Statistics: Theory and Methods, 15(3): 1047–1064, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:1989:SRR

[Xie89] Min Xie. Some results on renewal equations. Communications in Statistics: Theory and Methods, 18(3):1159–1171, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xekalaki:1989:SDA

[XP89] Evdokia Xekalaki and J. Panaretos. On some distributions arising in inverse cluster sampling. Communications in Statistics: Theory and Methods, 18(1):355–366, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1989:EOS

[YAH89] David H. Young and Mohamad S. Abdel-Hameed. Efficiency of observation schemes for proportional hazard regression models with censoring. Communications in Statistics: Theory and Methods, 18 (5):1665–1674, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1986:BEE

[Yam86] Hajime Yamato. Bayes estimates of estimable parameters with a Dirichlet invariant process. Communications in Statistics: Theory and Methods, 15(8):2383–2390, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1987:NBE

[Yam87] Hajime Yamato. Nonparametric Bayes estimates of estimable parameters with a Dirichlet invariant process and invariant Ustatistics. Communications in Statistics: Theory and Methods, 16 (2):525–543, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagimoto:1988:CMT

[Yan88] Takemi Yanagimoto. The conditional MLE in the two-parameter geometric distribution and its competitors. Communications in Statistics: Theory and Methods, 17(8):2779–2787, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagimoto:1989:IBD

[Yan89a] Takemi Yanagimoto. The inverse binomial distribution as a statistical model. Communications in Statistics: Theory and Methods, 18(10):3625–3633, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1989:STL

[Yan89b] Hwa-Ming Yang. On selecting the treatment with the largest probability of survival. Communications in Statistics: Theory and Methods, 18(2):639–659, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:1986:MLE

[Yao86] Yi-Ching Yao. Maximum likelihood estimation in hazard rate models with a change-point. Communications in Statistics: Theory and Methods, 15(8):2455–2466, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yassaee:1981:IDD

[Yas81a] Hedayat Yassaee. On integrals of Dirichlet distributions and their applications. Communications in Statistics: Theory and Methods, 10(9):897–906, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yasuko:1981:RCM

[Yas81b] Chikuse Yasuko. Representations of the covariance matrix for robustness in the Gauss–Markov model. Communications in Statistics: Theory and Methods, 10(19):1997–2004, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yatchew:1986:MDI

[Yat86] Adonis Yatchew. Multivariate distributions involving ratios of normal variables. Communications in Statistics: Theory and Methods, 15(6):1905–1926, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yfantis:1982:EMD

[YB82] E. A. Yfantis and L. E. Borgman. An extension of the von Mises distribution. Communications in Statistics: Theory and Methods, 11(15):1695–1706, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:1983:TFB

[YB83] Ching-Ming Yeh and Ralph A. Bradley. Trend-free block designs: existence and construction results. Communications in Statistics: Theory and Methods, 12(1):1–24, 1983. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:1988:SOO

[Yeh88] Ching-Mirig Yeh. Schur optimality of one-way designs. Communications in Statistics: Theory and Methods, 17(8):2705–2713, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yogendra:1982:NAD

[YG82] P. Chaubey Yogendra and S. Mudholkar Govind. A new approximation for the distribution of the difference of two t-variables. Communications in Statistics: Theory and Methods, 11(20):2335– 2342, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:1988:NCN

[YI88] Y. C. Yao and H. K. Iyer. A note on the completeness of the normal family with constant coefficient of variation. Communications in Statistics: Theory and Methods, 17(6):2075–2083, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamamoto:1985:MNC

[YKY85] Sumiyasu Yamamoto, Masahide Kuwada, and Fuzhi Yuan. On the maximum number of constraints for s-symbol balanced arrays of strength t. Communications in Statistics: Theory and Methods, 14 (10):2447–2456, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Youn-Min:1985:NRD

[YM85] Chou Youn-Min. New representations for the doubly noncentral F - distribution and derived distribution. Communications in Statistics: Theory and Methods, 14(3):527–534, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1986:IS

[YM86] Hajime Yamato and Yoshihiko Maesono. Invariant U-statistics. Communications in Statistics: Theory and Methods, 15(11):3253– 3263, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1989:DSD

[YM89] Hajime Yamato and Yoshihiko Maesono. Deficiencies of Ustatistics of degree 2 under symmetric distributions. Communications in Statistics: Theory and Methods, 18(1):53–66, 1989. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagawa:1986:CPM

[YN86] Takashi Yanagawa and Kentaro Nomakuchi. Characterization of probability models by an odds ratio in prospective and retrospective studies. Communications in Statistics: Theory and Methods, 15(8):2347–2353, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1986:FSS

[YO86] Dean M. Young and Patrick L. Odell. Feature-subset selection for statistical classification problems involving unequal covariance matrices. Communications in Statistics: Theory and Methods, 15 (1):137–157, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1985:IQS

[You85] Brian G. Young, P.Eng., C.I.M. International quality system standards. Communications in Statistics: Theory and Methods, 14(11): 2803–2813, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1981:DST

[YS81] Mark C. K. Yang and Jack F. Schreckengost. Difference sign test for comovements between two time series. Communications in Statistics: Theory and Methods, 10(4):355–369, 1981. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1982:CITb

[YS82a] Robert E. Young and Arnold L. Sweet. Confidence intervals for transition probabilities in multi-state Markov chains. II. Communications in Statistics: Theory and Methods, 11(2):181–198, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1982:CITa

[YS82b] Robert E. Young and Arnold L. Sweet. Confidence intervals for transition probabilities in two-state Markov chains. Communications in Statistics: Theory and Methods, 11(2):165–179, 1982. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadava:1984:OSA

[YS84] Surendar S. Yadava and Ravindra Singh. Optimum stratification for allocation proportional to strata totals for simple random sampling scheme. Communications in Statistics: Theory and Methods, 13(22):2793–2806, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suen:1989:SRO

[yS89] Chung yi Suen. Some resolvable orthogonal arrays with two symbols. Communications in Statistics: Theory and Methods, 18(10): 3875–3881, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yoshizawa:1985:AEH

[YSD85] Carl N. Yoshizawa, Pranab K. Sen, and C. Edward Davis. Asymptotic equivalence of the Harrell–Davis median estimator and the sample median. Communications in Statistics: Theory and Methods, 14(9):2129–2136, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1988:PIC

[YSTM88] Dean M. Young, John W. Seaman, Jr., Danny W. Turner, and Virgil R. Marco. Probability inequalities for continuous unimodal random variables with finite support. Communications in Statistics: Theory and Methods, 17(10):3505–3519, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1985:AEV

[Yu85] Kai F. Yu. On the asymptotic expectation and variance of the number of boundary crossings. Communications in Statistics: Theory and Methods, 14(7):1745–1753, 1985. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1986:BRE

[Yu86] Kai F. Yu. On the bounded regret of empirical Bayes estimators. Communications in Statistics: Theory and Methods, 15(8):2391– 2403, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1988:LRU

[Yu88] Kai F. Yu. A linear regression with unobserved dependent variables. Communications in Statistics: Theory and Methods, 17(9):

3075–3087, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yum:1985:APO

[Yum85] B. J. Yum. Asymptotic properties of the OLS and GRLS estimator for the replicated functional relationship model. Communications in Statistics: Theory and Methods, 14(8):1981–1996, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:1981:BEEb

[Zac81a] S. Zacks. Bayes equivariant estimators of the variance of a finite population for exponential priors. Communications in Statistics: Theory and Methods, 10(5):427–437, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:1981:PDE

[Zac81b] S. Zacks. The probability distribution and the expected value of a stopping variable associated with one-sided cusum procedures for non-negative integer valued random variables. Communications in Statistics: Theory and Methods, 10(21):2245–2258, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zehnwirth:1985:NLB

[Zeh85] Ben Zehnwirth. Nonparametric linear Bayes estimation of survival curves from incomplete observations. Communications in Statistics: Theory and Methods, 14(8):1769–1778, 1985. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zeisel:1989:PDW

[Zei89] Helmut Zeisel. On the power of the Durbin–Watson test under high autocorrelation. Communications in Statistics: Theory and Methods, 18(10):3907–3916, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zerbe:1981:LE

[Zer81] Gary O. Zerbe. Letter to the editor. Communications in Statistics: Theory and Methods, 10(13):1355–1356, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1987:PGD

[ZF87] Nien Fan Zhang and Robert V. Foutz. Partial group delay and time-lagged relationships among multiple time series. Communications in Statistics: Theory and Methods, 16(7):2139–2144, 1987.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1988:HOR

[ZF88] Nien Fan Zhang and Robert V. Foutz. A higher-order randomparameter process for modeling and forecasting time series. Communications in Statistics: Theory and Methods, 17(2):485–498, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zografos:1989:OPP

[ZFP89] K. Zografos, K. Ferentinos, and T. Papaioannou. Order preserving property of measures of information. Communications in Statistics: Theory and Methods, 18(7):2647–2656, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zerbe:1980:CIC

[ZG80] Cary O. Zerbe and David E. Goldgar. Comparison of intraclass correlation coefficients with the ratio of two independent F -statistics. Communications in Statistics: Theory and Methods, 9(15):1641– 1655, 1980. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zerbe:1982:MCR

[ZLMK82] Gary O. Zerbe, Eugene Laska, Morris Meisner, and H. B. Kushner. On multivariate confidence regions and simultaneous confidence limits for ratios. Communications in Statistics: Theory and Methods, 11(21):2401–2425, 1982. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:1981:BEEa

[ZS81] S. Zacks and H. Solomon. Bayes and equivariant estimators of the variance of a finite population: Part I, simple random sampling. Communications in Statistics: Theory and Methods, 10(5):407– 426, 1981. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1988:RTS

[zS88] Ning zhong Shi. Rank test statistics for umbrella alternatives. Communications in Statistics: Theory and Methods, 17(6):2059– 2073, 1988. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zahedi:1989:DMS

[ZS89] Hassan Zahedi and Samuel S. Shapiro. A discrete model for the segmentation of chains and strings. Communications in Statistics: Theory and Methods, 18(1):199–218, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zalkikar:1986:BEB

[ZTJ86] Jyoti N. Zalkikar, Ram C. Tiwari, and S. Rao Jammalamadaka. Bayes and empirical Bayes estimation of the probability that Z > X + Y. Communications in Statistics: Theory and Methods, 15 (10):3079–3101, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).